### v1

# Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

Abstract.
We develop a real-analytic counterpart of the Bers embedding for the Fréchet Lie group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ of decay-controlled diffeomorphisms of the line, and establish its connection to $L^{p}$ Fisher-Rao geometry on densities. For $p\in[1,\infty)$, the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically linearizes the homogeneous $\dot{W}^{1,p}$ Finsler metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, yielding explicit geodesics and a canonical flat connection whose Eulerian geodesic equation is the generalized Hunter-Saxton equation; for $p=\infty$, logarithmic coordinates $\varphi\mapsto\log\varphi^{\prime}$ provide a global isometry and the Schwarzian derivative emerges as the projective curvature. We construct a real Bers map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ via this Schwarzian, prove it is a Fréchet-smooth injective immersion whose linearization admits a tame right inverse given by an explicit Volterra operator, and characterize its image through Sturm-Liouville spectral theory. We introduce an $L^{p}$-Schwarzian family $S_{p}$ that interpolates between affine and projective cocycles, establish full asymptotic expansions as $p\to\infty$ in Fréchet and Orlicz-Sobolev scales, and extend the Bers embedding to Orlicz diffeomorphism groups. Through the Jacobian correspondence, these structures transfer to a manifold of densities asymptotic to Lebesgue measure, where the nonlinear Eulerian transport reduces to a pointwise Riccati law and the Schwarzian becomes the score curvature governing Fisher information. The compact-manifold $L^{p}$ Fisher-Rao linearization of Bauer, Bruveris, Harms, and Michor is recalled as a guiding framework.
Key words and phrases:
Schwarzian derivative, Bers embedding, Fisher-Rao metric, diffeomorphism groups, Orlicz spaces
2020 Mathematics Subject Classification:
58D05, 58B20, 53C21, 34B24, 46E30
1. Introduction
The Schwarzian derivative
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}$
is a classical object in complex analysis and projective geometry, encoding the failure of a diffeomorphism to be a Möbius transformation. Its appearance in diverse contexts ranging from univalent function theory and Teichmüller spaces to conformal field theory and one-dimensional dynamics, which reflects a fundamental principle, that is, the Schwarzian is the unique differential invariant of third order that vanishes precisely on projective transformations and satisfies the chain rule cocycle
$S(\varphi\circ\psi)=(S(\varphi)\circ\psi)\,(\psi^{\prime})^{2}+S(\psi).$
The classical Bers embedding realizes the universal Teichmüller space as a domain in the Banach space of bounded holomorphic quadratic differentials on the lower half-plane. This embedding, fundamental to Teichmüller theory, is constructed via the Schwarzian derivative of quasiconformal extensions.
In this paper we develop a real-analytic counterpart of the Bers embedding for a Fréchet Lie group of diffeomorphisms of $\mathbb{R}$ with controlled decay at infinity and construct, around it, a unified geometric theory that interweaves four subjects:
  1. (i)
    the $L^{p}$ Finsler geometry of diffeomorphism groups and its complete isometric linearization,
  2. (ii)
    the Schwarzian derivative as a projective curvature that governs the passage from affine to projective structures,
  3. (iii)
    one-dimensional scattering theory for Schrödinger operators, which furnishes nonlinear spectral coordinates on the quotient $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$, and
  4. (iv)
    the Fisher-Rao geometry of positive densities, where the Schwarzian reappears as a score curvature and Fisher information becomes a spectral entropy.
These threads are linked by a single mechanism: the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically flattens the right-invariant Finsler geometry on diffeomorphisms and, in the limit $p\to\infty$, produces logarithmic coordinates in which the Schwarzian takes the potential form $S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$, $f=\log(\varphi^{\prime})$. The real Bers embedding, the Sturm-Liouville spectral characterization of its image, the scattering transform, the Virasoro-Bott central extension, the $L^{p}$-Schwarzian interpolation, the Orlicz-Sobolev generalization, the Jacobian correspondence to densities, the forced sign-change theorems for projective curvature, and the sharp functional inequalities all follow from this single organizing principle.
Broader programme. This is the first of a two-part work. The one-dimensional scalar theory developed here serves as both the conceptual template and the technical foundation for Part II [21], which constructs a Koopman-equivariant Orlicz-Sobolev statistical bundle over a general non-singular dynamical system. There the Schwarzian becomes a Koopman spectral potential, the Bers embedding becomes a statistical bundle projection, and the scattering transform becomes an EDMD-type spectral decomposition. The blow-up analysis and spine decomposition of Part I supply the explicit one-dimensional inputs for the WKB expansions and Hamilton-Jacobi theory on the infinite-dimensional path space in Part II. The two parts together aim to provide a geometric and spectral framework in which the interplay between diffeomorphism groups, projective cocycles, statistical estimation, and branching process asymptotics can be treated in a unified manner.
1.1. Main results
Let $W^{\infty,1}(\mathbb{R})=\bigcap_{k\geq 0}W^{k,1}(\mathbb{R})$ denote the Fréchet space of smooth functions whose derivatives of all orders lie in $L^{1}(\mathbb{R})$. Define the decay-controlled diffeomorphism group
$\operatorname{Diff}^{-\infty}(\mathbb{R})=\left\{\varphi=\operatorname{id}+f\mid f^{\prime}\in W^{\infty,1}(\mathbb{R}),\;f^{\prime}>-1,\;\lim_{x\to-\infty}f(x)=0\right\}.$
This is a Fréchet Lie group with Lie algebra $\mathfrak{g}^{-\infty}=\{u\mid u^{\prime}\in W^{\infty,1}(\mathbb{R}),\;\lim_{x\to-\infty}u(x)=0\}$, equipped with the right-invariant homogeneous $\dot{W}^{1,p}$ Finsler structure
$F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\left\|u^{\prime}\right\|_{L^{p}(\mathbb{R})},\qquad u=\delta\varphi\circ\varphi^{-1}.$
A.  Isometric linearization, flat connections, and geodesic dynamics
Our first main result establishes that this Finsler geometry admits a global linearization.
Theorem 1.1 (Isometric linearization, Theorem 2.8).
For $p\in[1,\infty)$, the map
$\Phi_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$
is an isometric embedding from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ into $(W^{\infty,1}(\mathbb{R}),\left\|\cdot\right\|_{L^{p}})$. Its image is the open set $\{f\in W^{\infty,1}(\mathbb{R})\mid f>-p\}$. For $p=\infty$, the map $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is an isometric diffeomorphism onto $W^{\infty,1}(\mathbb{R})$.
The linearization trivializes the geodesic problem: minimizing $F_{\dot{W}^{1,p}}$-geodesics become affine segments in the target space. We derive from this a canonical flat connection $\nabla^{(p)}$ on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, compute its parallel transport explicitly (which has trivial holonomy), and show that the geodesic equation in Eulerian form is the strain equation
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0,$
whose $x$-derivative gives the generalized Hunter-Saxton equation [16]. For $p=2$ this is the classical Hunter-Saxton equation. As $p\to\infty$, the quadratic nonlinearity vanishes and geodesics become solutions of the pure transport equation $\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})=0$. The connections $\nabla^{(p)}$ admit a two-parameter deformation $\nabla^{(p,\alpha)}$ by a projective tensor $\mathcal{C}(u,a)=u_{xx}a-ua_{xx}$; the resulting curvature involves the Lie bracket structure of vector fields through the commutator $\partial_{x}^{2},\mathcal{L}_{[u,v]}]$.
The variational formulation is equally clean: we prove an Euler-Lagrange characterization of geodesics (Theorem 2.20), establish the Hamiltonian duality through a Legendre transform on $L^{p}$-$L^{q}$ pairs (Theorem 2.21), and show that in the linearizing chart $\Phi_{p}$ the value function solves a Hamilton-Jacobi equation on the Banach space $L^{p}(\mathbb{R})$ with an explicit closed-form expression (Proposition 2.22). The family $\{\Phi_{p}\}_{p\in[1,\infty]}$ admits a full asymptotic expansion in Fréchet seminorms as $p\to\infty$ (Proposition 2.23).
B.  Blow-up analysis and large deviations
The Lagrangian strain $V(t,\xi)=u_{x}(t,\varphi(t,\xi))$ satisfies the autonomous Riccati ordinary differential equation
$\frac{\mathrm{d}V}{\mathrm{d}t}+\frac{1}{p}V^{2}=0$
at each label $\xi$, yielding the closed-form solution $V(t,\xi)=V_{0}(\xi)/(1+\frac{t}{p}V_{0}(\xi))$ and the Jacobian evolution $\partial_{\xi}\varphi(t,\xi)=\varphi_{0}^{\prime}(\xi)(1+\frac{t}{p}V_{0}(\xi))^{p}$. From this we derive a sharp blow-up criterion (Theorem 3.3): the geodesic loses smoothness at time $T^{*}=-p/\inf(v_{0})$ if and only if $\inf(v_{0})<0$, where the strain develops a $|x-x^{*}|^{-1}$ singularity with an explicit self-similar profile (Theorem 3.5).
Beyond blow-up, we prove global existence and uniqueness of dissipative weak solutions for the generalized Hunter-Saxton equation (Theorem 3.7) extending the framework of [9] from $p=2$ to general $p$. Geometrically, the dissipative continuation corresponds to the affine geodesic in $L^{p}(\mathbb{R})$ exiting the positivity constraint $\{f>-p\}$-the nonlinearity of the diffeomorphism group is encoded entirely in this half-space condition.
The blow-up admits a probabilistic interpretation through a branching factor and local Malthusian parameter (Definitions 3.8). We establish a small-noise large deviation principle for the catastrophe time $T^{*}$ under Gaussian perturbation of the initial strain (Theorem 3.18), with a quadratic rate function $J(T)\sim\frac{m^{4}}{2p^{2}\kappa}(T-T^{*})^{2}$ governed by the Cameron-Martin geometry. A spine decomposition (Proposition 3.20, Theorem 3.21) shows that the unique minimal-cost perturbation achieving a prescribed shift in blow-up time is proportional to the Green function $\frac{1}{2}e^{-|x-x^{*}|}$, the Riesz representer of point evaluation in $H^{1}(\mathbb{R})$.
C.  The Schwarzian as projective curvature
The Schwarzian derivative emerges naturally in the $p\to\infty$ limit through the potential identity
$S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2},\qquad f=\Phi_{\infty}(\varphi)=\log(\varphi^{\prime}).$
We develop the Schwarzian as a score curvature on densities: if $\mu=g\,dx$ is a positive density with score $s=(\log g)^{\prime}$, then $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. The mean projective curvature satisfies the identity $\int S(\mu)\,g\,dx=-\frac{3}{2}I(\mu)$ where $I(\mu)=\int s^{2}\,g\,dx$ is the Fisher information (Proposition 4.8), and the Schwarzian admits a projective decomposition inside the second-order score composite that yields a Bhattacharyya-type variance bound for unbiased estimation (Theorem 4.10).
We introduce the $L^{p}$-Schwarzian to interpolate between the affine linearizations $\Phi_{p}$ and the projective Schwarzian.
Theorem 1.2 ($L^{p}$-Schwarzian, Propositions 4.20 and 4.24).
For $p\in[1,\infty)$, the quantity
$S_{p}(\varphi)=\left(\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}+S(\varphi)\right)\varphi^{\prime 1/p}$
satisfies the cocycle identity $S_{p}(\varphi\circ\psi)=(S_{p}(\varphi)\circ\psi)\,\psi^{\prime 1/p}+S_{p}(\psi)$ and converges to $S(\varphi)$ as $p\to\infty$ with full asymptotic expansion in $W^{k,1}(\mathbb{R})$:
$S_{p}(\varphi)=S(\varphi)+\frac{1}{p}\left(uS+\frac{3}{2}(u^{\prime})^{2}\right)+O_{W^{k,1}}\!\left(\frac{1}{p^{2}}\right),\qquad u=\log(\varphi^{\prime}).$
The Sturm-Liouville equation $y^{\prime\prime}+qy=0$ provides a canonical differential-geometric framework for the Schwarzian, and we develop a semiclassical interpretation in which the WKB expansion of solutions reproduces the classical and $L^{p}$ Schwarzians as successive curvature corrections (Section 4.7).
D.  Central extensions and cocycle geometry
The Schwarzian is the fundamental obstruction to lifting the composition law on diffeomorphisms to an affine structure. We develop the cohomological content of this obstruction through several constructions. The Gelfand-Fuchs $2$-cocycle $\omega(u,v)=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx$ is shown to represent a nontrivial class in $H^{2}(\mathfrak{g}^{-\infty};\mathbb{R})$ (Proposition 5.2), and it integrates to the Virasoro-Bott central extension $\widehat{\operatorname{Diff}}^{-\infty}(\mathbb{R})$ via the Bott-Thurston group cocycle $B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})$.
We prove a rigidity theorem: the $p$-root Jacobian cocycle produces no new Lie algebra cohomology classes beyond the Gelfand-Fuchs class, up to explicit rescaling (Section 5). In contrast, the $L^{p}$-Schwarzian fails to satisfy a two-term cocycle identity; the unavoidable cross-term in its composition law is a manifestation of the fact that projective curvature cannot be captured by a simple $L^{p}$ cocycle.
A new ingredient is the cumulant transgression: although the Bott cocycle is not a group coboundary on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, it becomes an explicit groupoid coboundary on the action groupoid $\operatorname{Diff}^{-\infty}(\mathbb{R})\ltimes|\Omega|_{\infty,1}(\mathbb{R})$ once one allows dependence on a density through its score (Section 5, Subsection 3.6). This yields a canonical transgression functional whose coboundary is the Bott phase, providing a categorical absorption of the projective multiplier into a central extension of Koopman transport. We formalize this through a $U(1)$-banded gerbe on the density quotient.
E.  The real Bers embedding
Since $S(\varphi)=0$ if and only if $\varphi$ is affine, the Schwarzian descends to the quotient $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
Theorem 1.3 (Real Bers embedding, Theorem 6.7).
The map
$\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\longrightarrow W^{\infty,1}(\mathbb{R}),\qquad\beta^{-\infty}([\varphi])=\frac{1}{2}S(\varphi)$
is a Fréchet smooth injective immersion. Its differential at $\varphi]$ admits a continuous tame right inverse $R_{u}\colon W^{k,1}(\mathbb{R})\to W^{k+2,1}(\mathbb{R})$ for every $k\geq 0$, given by the explicit Volterra operator
$(R_{u}v)(x)=\int_{-\infty}^{x}\mathrm{e}^{u(t)}\int_{-\infty}^{t}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y\,\mathrm{d}t,\qquad u=\log(\varphi^{\prime}).$
The Bers potential $q=\beta^{-\infty}(\varphi)$ admits a Miura-type factorization $q=\frac{1}{2}u^{\prime\prime}-\frac{1}{4}(u^{\prime})^{2}$ (Proposition 6.6), where $u=\log(\varphi^{\prime})$ plays the role of a Miura variable. This factorization guarantees that the associated Schrödinger operator $H_{q}=-\partial_{x}^{2}-q$ is nonnegative and has no bound states: $H_{q}=A^{*}A$ where $A=\partial_{x}+\frac{1}{2}u^{\prime}$. The image of $\beta^{-\infty}$ admits a spectral characterization: a potential $q\in W^{\infty,1}(\mathbb{R})$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the Sturm-Liouville equation $y^{\prime\prime}+qy=0$ admits a positive solution $y_{2}$ with $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$, and the reconstruction is $\varphi=y_{1}/y_{2}$ where $y_{1}$ is the second solution with Wronskian $W(y_{1},y_{2})\equiv 1$ (Theorem 6.8).
We also construct local Bers coordinates centered at arbitrary basepoints, compute their transition maps as Schwarzian-shifted affine connections (Section 6.8), and establish a spectral transport formula that conjugates Schrödinger operators along paths in the Bers image (Section 6.9).
F.  Scattering coordinates and nonlinear spectral analysis
The Miura factorization implies that the Schrödinger operator $H_{q}$ for a Bers potential $q=\beta^{-\infty}(\varphi)$ has essential spectrum $0,\infty)$ with no negative eigenvalues. The scattering data therefore consist of the reflection coefficient alone, and this defines a nonlinear spectral coordinate.
Theorem 1.4 (Scattering Bers map, Theorem 7.6).
The map $\mathcal{S}(\varphi):=R_{\beta^{-\infty}(\varphi)}$, assigning to each diffeomorphism the reflection coefficient of its Bers potential, descends to an injective map
$\mathcal{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\hookrightarrow\mathscr{S}(\mathbb{R}).$
The reflection coefficient satisfies the reality condition $\overline{R(-k)}=R(k)$, the strict sub-unitarity $|R(k)|<1$ for $k\neq 0$, and the transmission denominator is the outer function (79) determined by $|R|$.
We develop the scattering theory in detail. Jost solutions and their Volterra integral equations are constructed for potentials in $W^{\infty,1}(\mathbb{R})$, yielding Schwartz-class scattering data (Section 7). We introduce a relative spectral zeta function $\zeta_{\mu}(z)$ for the Bers Schrödinger operator and prove that its residue at $z=-\frac{1}{2}$ is a multiple of the Fisher information (Proposition 7.8). The Bers-Mercer kernel $K_{q,\rho}$, obtained by integrating the Green function of $H_{q}$ against an admissible spectral weight, is shown to be positive definite, giving rise to a reproducing kernel Hilbert space $\mathcal{H}_{q,\rho}$ that provides the canonical arena for finite-rank spectral approximation (Section 7.6). The Marchenko reconstruction is developed in Section 7.7.
G.  Extension to Orlicz-Sobolev scales
The entire theory extends to Orlicz-Sobolev scales. For an admissible Young function $\Phi$ satisfying a $\Delta_{2}$ condition, we construct the Orlicz diffeomorphism group $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ in logarithmic coordinates and prove it carries a Fréchet Lie group structure (Section 8). The construction proceeds through Orlicz fiber bundles over the density manifold $|\Omega|_{\infty,1}(\mathbb{R})$, equipped with the Luxemburg norm fiberwise. We develop the Orlicz-Young duality on fibers (Section 8.5), prove that Koopman operators act as bounded bundle morphisms on the Orlicz fibers (Section 8.6), and establish that the Bers embedding extends to a Fréchet smooth injective immersion $\beta_{\Phi}\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,\Phi}(\mathbb{R})$. The convergence $S_{p}\to S$ is proved to be stable in Orlicz-Sobolev norms (Section 5.8).
H.  Density geometry: projective curvature and forced sign-change
The Jacobian map $J(\varphi)=\varphi^{\prime}\,\mathrm{d}x$ establishes a Fréchet diffeomorphism and Finsler isometry between $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the manifold $|\Omega|_{\infty,1}(\mathbb{R})$ of positive densities asymptotic to Lebesgue measure (Theorem 9.4). This correspondence intertwines all linearizations: the $p$-root map $\Phi_{p}$ on diffeomorphisms corresponds to the statistical $p$-root map $\Psi_{p}(g\,dx)=p(g^{1/p}-1)$ on densities, and the flat connection, geodesics, and Fisher-Rao metric transfer canonically.
On the density side, the Schwarzian becomes the score curvature $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$, the Bers potential is $\beta(\mu)=\frac{1}{2}S(\mu)$, and the generalized Hunter-Saxton equation becomes the pointwise Riccati law $\partial_{t}v+\frac{1}{p}v^{2}=0$ for the statistical velocity $v=\partial_{t}\log(g)$.
A central qualitative result is the forced sign-change theorem: for every nontrivial density $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$, the Bers potential $\beta(\mu)$ necessarily takes both positive and negative values (Theorem 9.28). We derive a hierarchy of structural consequences: every nontrivial density has at least two projective inflection points (Proposition 9.30), the Bers potential is negative at local maxima and positive at local minima of the density (Proposition 9.31), and log-concavity regions are always contained in the negative projective curvature region (Proposition 9.32). Perhaps most strikingly, Fisher information alone does not control the positive projective curvature (Theorem 9.34).
We develop the Koopman operator theory on density space, constructing a statistical Hilbert bundle with a unitary half-density representation and a half-density generator whose drift is the score (Section 9.5). The projective covariance of the Koopman action produces the Schwarzian cocycle (Theorem 9.19). A Feynman-Kac representation identifies $\beta(\mu)/2$ as the potential of a branching-killing diffusion (Theorem 9.38), and the forced sign-change theorem translates into the statement that every nontrivial density has both a genuine branching region and a genuine killing region. The spine change of measure identifies the score $s_{\mu}$ as (twice) the drift of the spine process (Theorem 9.40).
I.  Sharp functional inequalities
The Bers image identifies a diffeomorphism with a one-dimensional Schrödinger potential subject to a critical spectral constraint: the operator $H_{q}=-\partial_{x}^{2}-q$ is nonnegative with no bound states. We extract several precise analytic consequences.
The ground state transform $\psi=(\varphi^{\prime})^{-1/2}$ gives a sharp Hardy inequality $\int qf^{2}\,dx\leq\int(f^{\prime})^{2}\,dx$ in which $q$ is a critical Hardy weight: no positive remainder can be added (Theorem 10.1). On finite windows, the correct substitute for a spectral gap is a Hardy-Poincaré inequality after projection onto the orthogonal complement of the resonant direction, with an explicit lower bound (Theorem 10.3).
The deepest results in this section are trace identities that express the geometric integrals of the Bers potential in spectral variables. The first trace identity, which equates $\int q\,dx$ with a spectral entropy of the reflection coefficient (Theorem 10.10), is given by
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log(1-|R_{q}(k)|^{2})\,dk.$
On the density side, this becomes the identity $\int s^{2}\,dx=\frac{4}{\pi}\int\log(1-|R_{q}(k)|^{2})^{-1}\,dk$, expressing Fisher information as a spectral entropy. The second trace identity upgrades this to the quadratic level, expressing $\int q^{2}\,dx$ as a frequency-weighted spectral entropy (Theorem 10.11); on the density side this yields an identity for the mixed Sobolev/Orlicz functional $\frac{1}{4}\int|s^{\prime}|^{2}\,dx+\frac{1}{16}\int|s|^{4}\,dx$ in terms of $\int k^{2}\log(1-|R_{q}|^{2})^{-1}\,dk$.
1.2. Relation to prior work
The $L^{p}$ Fisher-Rao geometry on spaces of probability densities over compact manifolds was systematically developed by Bauer, Bruveris, Harms, and Michor [6], who proved that the $p$-root map $\mu\mapsto p(\mu/\mathrm{d}x)^{1/p}$ isometrically embeds the space of positive densities into $L^{p}$. Our treatment of $L^{p}$ geometry on densities in Section 11 recalls their framework and extends it to the noncompact setting $\mathbb{R}$ with decay conditions. The classification of diffeomorphism-invariant Riemannian metrics on densities is due to Bauer, Bruveris, and Michor [5].
The Fréchet Lie group structure on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the existence of right-invariant Sobolev metrics were established by Bauer, Bruveris, and Michor [4]. The classical Bers embedding in Teichmüller theory is treated in Lehto [22]. The Schwarzian derivative and its role in projective geometry are classical; see Ovsienko and Tabachnikov [25] for a modern treatment and Fuks [14] for the cohomological aspects. The Hunter-Saxton equation as a geodesic equation on diffeomorphism groups was identified by Khesin and Misiołek [18]; the original derivation of the equations is in [16]. The geometry of infinite-dimensional groups is treated comprehensively in Khesin and Wendt [19]. The one-dimensional inverse scattering theory on the line is due to Deift and Trubowitz [10]; see also Beals, Deift, and Tomei [7] and Faddeev and Takhtajan [12]. The Miura transformation connecting KdV and mKdV equations originates in [23]. The Bott-Thurston cocycle and its role in the theory of characteristic classes of diffeomorphism groups appears in Bott [8] and the Virasoro extension in Pressley and Segal [26]; the cohomological framework is developed in Feigin and Fuchs [13] and Fuks [14]. The integration of the Lie algebra cocycle to a group cocycle follows the general pattern of [15]. We verify the details in the present decay-controlled setting. Orlicz space theory follows Rao and Ren [27] and the Sobolev embedding theorems are from Adams and Fournier [1].
What is new in the present work, beyond the construction of the real Bers embedding itself, is the synthesis: the simultaneous development of the $L^{p}$ linearization, the $L^{p}$-Schwarzian interpolation, the scattering coordinate, the Orlicz-Sobolev extension, the density-side projective curvature theory with forced sign-change and Fisher non-control phenomena, the central extension gerbe on the density quotient with cumulant transgression, the complete blow-up theory with large deviations and spine decomposition, and the trace identities expressing Fisher energies as spectral entropies-all derived from the single organizing principle of $p$-root coordinates and their $p\to\infty$ limit.
1.3. Organization
The paper is organized as follows.
Section 2 establishes the isometric linearization of $\operatorname{Diff}^{-\infty}(\mathbb{R})$, derives the flat connection, geodesics, parallel transport, the Hamiltonian formulation with Hamilton-Jacobi equation, the two-parameter family of connections with its torsion and curvature, and the Fréchet asymptotics as $p\to\infty$.
Section 3 develops the complete blow-up theory for the generalized Hunter-Saxton equation: the Lagrangian Riccati structure, sharp blow-up criterion and rate, global existence and uniqueness of dissipative weak solutions, the probabilistic branching-factor interpretation, and the large deviation principle for the catastrophe time with its spine decomposition.
Section 4 develops the Schwarzian derivative as a projective potential in logarithmic coordinates, introduces the score curvature and its Fisher-Rao identities including the Bhattacharyya-type bound, constructs the projective connection bundle, introduces the $L^{p}$-Schwarzian family with its cocycle identity and $p\to\infty$ asymptotics, and gives the Sturm-Liouville and semiclassical interpretations.
Section 5 treats central extensions and cocycle geometry: the Gelfand-Fuchs cocycle and its nontriviality, the Virasoro-Bott group extension, $L^{p}$ cocycle rigidity, the Bott-Thurston cocycle in $p$-root coordinates, the gerbe structure on the density quotient, and the cumulant transgression producing Koopman central extensions.
Section 6 constructs the real Bers embedding, proves injectivity and the tame right inverse theorem, establishes the Miura factorization, the spectral characterization of the image, the explicit inverse construction, local Bers coordinates with their transition maps, and spectral transport along paths.
Section 7 develops the scattering theory for Bers potentials: Jost solutions, the absence of bound states, the scattering transform as a nonlinear Fourier map, spectral zeta functions and regularized determinants, Mercer kernels and reproducing kernel Hilbert spaces, and the Marchenko reconstruction.
Section 8 extends the theory to Orlicz-Sobolev scales: Orlicz-Sobolev spaces, admissible Young functions, Orlicz diffeomorphism groups, Orlicz fiber bundles and their Koopman theory, Orlicz-Young duality, the Orlicz Bers embedding, and the Orlicz stability of $S_{p}\to S$.
Section 9 develops the density-side theory: the Jacobian isomorphism, score functions and projective curvature, the linearization intertwining, the forced sign-change theorems with their inflection-point and Fisher non-control consequences, the Koopman operators and statistical Hilbert bundle with unitary half-density representation, the path space Fisher-Rao action, and the branching diffusion viewpoint with Feynman-Kac and spine constructions.
Section 10 establishes the sharp functional inequalities: critical Hardy theory and resonance rigidity, Hardy-Poincaré coercivity on finite windows, $p$-Fisher energy theory, and the trace hierarchy expressing Fisher energies as spectral entropies of the reflection coefficient.
Section 11 recalls the $L^{p}$ Fisher-Rao framework on compact manifolds and the BBM classification theorem, serving as context for the noncompact extensions developed in the body of the paper.
1.4. Relation to Part II
This paper is the first of a two-part work. Part II [21] constructs the Koopman-equivariant Orlicz-Sobolev statistical bundle over a general measure-preserving or non-singular dynamical system, where the one-dimensional structures developed here lift to an infinite-dimensional fiber bundle. Specifically, the Luxemburg-Finsler geometry from Section 8, the Koopman action on Orlicz fibers from Section 8.6, and the cumulant transgression from Section 5 become the building blocks for a bundle of Orlicz function spaces on which non-singular dynamics act as bundle automorphisms preserving the Finsler structure.
In that setting, the Bers-Mercer kernels of Section 7.6 yield invariant reproducing kernels that enable spectral reductions, and the blow-up analysis and spine decomposition of Section 3 supply the one-dimensional WKB inputs for a semiclassical Hamilton-Jacobi problem on the path space of multitype branching processes. This produces WKB expansions with Berry-phase corrections, where the phase is controlled by the Bott-Thurston cocycle (Section 5), and an explicit spine action functional whose critical points recover the large deviation rate. The trace identities of Section 10 provide the spectral entropy variables that serve as the correct action-angle coordinates for the infinite-dimensional problem.
Acknowledgments
The present paper was initiated during the author’s doctoral studies at Northwestern University and continued during his appointment at Worcester Polytechnic Institute. The geometric and microlocal analysis traditions at Northwestern shaped the analytical perspective underlying this work. The author thanks the late Steve Zelditch, Jared Wunsch, and Elton P. Hsu for many inspiring conversations. Partially supported by NSF RTG grant DMS-1502632.
2. Isometric linearization and geodesic structure of the diffeomorphism group
2.1. The group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the $\dot{W}^{1,p}$ metric
Let $W^{k,1}(\mathbb{R})$ denote the Sobolev space of functions with $k$ derivatives in $L^{1}(\mathbb{R})$, and let $W^{\infty,1}(\mathbb{R})=\bigcap_{k\geq 0}W^{k,1}(\mathbb{R})$ with the Fréchet topology induced by the seminorms $\left\|\cdot\right\|_{W^{k,1}}$.
Definition 2.1.
The decay-controlled diffeomorphism group is
$\operatorname{Diff}^{-\infty}(\mathbb{R})=\left\{\varphi=\operatorname{id}+f\ \middle|\ f^{\prime}\in W^{\infty,1}(\mathbb{R}),\ f^{\prime}>-1,\ \lim_{x\to-\infty}f(x)=0\right\}.$
Its Lie algebra is
$\mathfrak{g}^{-\infty}=\left\{u\ \middle|\ u^{\prime}\in W^{\infty,1}(\mathbb{R}),\ \lim_{x\to-\infty}u(x)=0\right\}.$
The condition $f^{\prime}\in W^{1,1}(\mathbb{R})$ implies that $f^{\prime}$ extends continuously to $\infty$$\infty$] with $f^{\prime}(\pm\infty)=0$, hence $\varphi^{\prime}(x)\to 1$ as $x\to\pm\infty$. Together with $f^{\prime}>-1$ this gives $\varphi^{\prime}>0$ on $\mathbb{R}$, so $\varphi$ is a diffeomorphism.
Proposition 2.2 ([4]).
$\operatorname{Diff}^{-\infty}(\mathbb{R})$ is a Fréchet Lie group with Lie algebra $\mathfrak{g}^{-\infty}$.
2.2. Fiberwise trivialization and proto-bundle geometry
Proposition 2.3 (Tangent bundle trivialization).
The right-trivialization
$\Psi\colon T\operatorname{Diff}^{-\infty}(\mathbb{R})\to\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty},\qquad\Psi(\varphi,\delta\varphi)=\bigl(\varphi,\,\delta\varphi\circ\varphi^{-1}\bigr)$
is a Fréchet diffeomorphism, with inverse
$\Psi^{-1}(\varphi,u)=(\varphi,u\circ\varphi).$
Under this trivialization:
  1. (i)
    for $p\in[1,\infty]$ the right-invariant $\dot{W}^{1,p}$ Finsler structure satisfies
    $F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\|u^{\prime}\|_{L^{p}(\mathbb{R})}$
    for $(\varphi,u)=\Psi(\varphi,\delta\varphi)$, and is independent of $\varphi$;
  2. (ii)
    for $p\in[1,\infty)$ one has the fiberwise identity
    $D_{\varphi}\Phi_{p}(u\circ\varphi)=(u^{\prime}\circ\varphi)\,\varphi^{\prime 1/p}.$
Proof.
Because $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is a Fréchet Lie group (Proposition 2.2), right translation $R_{\varphi}(\psi)=\psi\circ\varphi$ and inversion $\varphi\mapsto\varphi^{-1}$ are smooth maps. The differential of $R_{\varphi}$ at the identity is $T_{\operatorname{id}}R_{\varphi}(u)=u\circ\varphi$. Consequently the map
$\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty}\to T\operatorname{Diff}^{-\infty}(\mathbb{R}),\qquad(\varphi,u)\longmapsto(\varphi,u\circ\varphi)$
is smooth. Composing with inversion gives the smooth map $\Psi$, and the displayed formula $\Psi^{-1}(\varphi,u)=(\varphi,u\circ\varphi)$ is its inverse. This proves that $\Psi$ is a Fréchet diffeomorphism.
For (i), by definition $u=\delta\varphi\circ\varphi^{-1}$ is the Eulerian velocity, hence $F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\|u^{\prime}\|_{L^{p}}$.
For (ii), differentiate $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ in the direction $\delta\varphi=u\circ\varphi$. As in the computation in Theorem 2.8, one has $D_{\varphi}\Phi_{p}(\delta\varphi)=\varphi^{\prime 1/p-1}\delta\varphi^{\prime}$, and $\delta\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, which yields the stated identity. ∎
Definition 2.4.
For $p\in[1,\infty]$, the $\dot{W}^{1,p}$ right-invariant Finsler structure on $T\operatorname{Diff}^{-\infty}(\mathbb{R})$ is
$F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\bigl\|(\delta\varphi\circ\varphi^{-1})^{\prime}\bigr\|_{L^{p}(\mathbb{R})}.$
Definition 2.5 (Jacobian density manifold).
Let $|\Omega|_{\infty,1}(\mathbb{R})$ be the Fréchet manifold of positive densities $\mu=\rho\,dx$ with $\rho>0$ and $\rho-1\in W^{\infty,1}(\mathbb{R})$. We identify $|\Omega|_{\infty,1}(\mathbb{R})$ with the open subset $\{\rho\in W^{\infty,1}(\mathbb{R})\mid\rho>0\}$ via $\mu\leftrightarrow\rho$.
Proposition 2.6 (Jacobian map as a global chart).
The Jacobian map
$J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R}),\qquad J(\varphi)=\varphi^{\prime}\,dx$
is a Fréchet diffeomorphism. Its inverse is
$J^{-1}(\rho\,dx)(x)=x+\int_{-\infty}^{x}(\rho(s)-1)\,ds.$
Proof.
If $\varphi=\operatorname{id}+f\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, then $\rho=\varphi^{\prime}=1+f^{\prime}$ satisfies $\rho-1=f^{\prime}\in W^{\infty,1}(\mathbb{R})$ and $\rho>0$, hence $J(\varphi)\in|\Omega|_{\infty,1}(\mathbb{R})$.
Conversely, let $\rho\in W^{\infty,1}(\mathbb{R})$ satisfy $\rho>0$ and $\rho-1\in W^{\infty,1}(\mathbb{R})$, and define
$\varphi(x)=x+\int_{-\infty}^{x}(\rho(s)-1)\,ds.$
Then $\varphi^{\prime}=\rho$, and $\varphi-\operatorname{id}$ has derivative $\rho-1\in W^{\infty,1}(\mathbb{R})$ and satisfies $\lim_{x\to-\infty}(\varphi(x)-x)=0$. Therefore $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $J(\varphi)=\rho\,dx$. This proves bijectivity with the displayed inverse.
Smoothness follows because $\varphi\mapsto\varphi^{\prime}$ is a continuous linear map $\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ in the model topology, and the inverse is the continuous linear operator $h\mapsto\int_{-\infty}^{x}h$ composed with addition of the identity. ∎
Definition 2.7 (Proto-statistical bundle).
Define the product bundle
$\pi\colon E^{-\infty}:=|\Omega|_{\infty,1}(\mathbb{R})\times W^{\infty,1}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R})$
with fiber $E_{\mu}=W^{\infty,1}(\mathbb{R})$. The derivative map on Eulerian fields induces a bundle isomorphism
$\widetilde{J}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty}\to E^{-\infty},\qquad\widetilde{J}(\varphi,u)=\bigl(J(\varphi),\,u^{\prime}\bigr),$
covering $J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R})$.
Proof.
The map $\widetilde{J}$ is smooth because $J$ is smooth and $u\mapsto u^{\prime}$ is a continuous linear map $\mathfrak{g}^{-\infty}\to W^{\infty,1}(\mathbb{R})$.
To prove bijectivity, let $(\mu,w)\in E^{-\infty}$ with $\mu=\rho\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and $w\in W^{\infty,1}(\mathbb{R})$. Define $\varphi:=J^{-1}(\mu)$ using Proposition 2.6, and define
$u(x):=\int_{-\infty}^{x}w(s)\,ds.$
Then $u\in\mathfrak{g}^{-\infty}$ and $u^{\prime}=w$. Moreover $\widetilde{J}(\varphi,u)=(\mu,w)$ by construction. This provides a two-sided inverse, and the inverse map is smooth because it is given by the smooth chart $J^{-1}$ and a continuous linear integration operator. ∎
2.3. Isometric embedding via $p$-root coordinates
Theorem 2.8.
For $p\in[1,\infty)$, the map
$\Phi_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$
is an isometric embedding from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ into $(W^{\infty,1}(\mathbb{R}),\left\|\cdot\right\|_{L^{p}})$. Its image is $U_{p}=\{f\in W^{\infty,1}(\mathbb{R})\mid f>-p\}$, and the inverse is
$\Phi_{p}^{-1}(f)(x)=x+\int_{-\infty}^{x}\left((1+f(\tilde{x})/p)^{p}-1\right)\mathrm{d}\tilde{x}$.
Proof.
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\varphi^{\prime}=1+h^{\prime}$ where $h^{\prime}\in W^{\infty,1}(\mathbb{R})$ and $h^{\prime}>-1$. Since $W^{\infty,1}(\mathbb{R})\subset C^{\infty}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$, there exist constants $0<c\leq\varphi^{\prime}\leq C<\infty$. The function $\rho\mapsto\rho^{1/p}$ is smooth on $(0,\infty)$, so $\varphi^{\prime 1/p}$ is smooth with
$\partial_{x}(\varphi^{\prime 1/p})=\frac{1}{p}\varphi^{\prime 1/p-1}\varphi^{\prime\prime}.$
Since $\varphi^{\prime\prime}\in L^{1}(\mathbb{R})$ and $\varphi^{\prime 1/p-1}$ is bounded, we have $\partial_{x}(\varphi^{\prime 1/p})\in L^{1}(\mathbb{R})$. Higher derivatives are handled similarly using Faà di Bruno’s formula. Each $\partial_{x}^{k}(\varphi^{\prime 1/p})$ is a finite sum of products of bounded functions of $\varphi^{\prime}$ with derivatives $\varphi^{(j)}$ for $j\geq 2$, all of which lie in $L^{1}(\mathbb{R})$. Thus $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)\in W^{\infty,1}(\mathbb{R})$.
The image characterization follows from $\varphi^{\prime 1/p}=1+f/p>0$ if and only if f>pf>-p. The inverse formula is verified by differentiation. If $\varphi=\Phi_{p}^{-1}(f)$, then $\varphi^{\prime}=(1+f/p)^{p}$, so $\varphi^{\prime 1/p}=1+f/p$ and $\Phi_{p}(\varphi)=f$.
For the isometry, let $\delta\varphi\in T_{\varphi}\operatorname{Diff}^{-\infty}(\mathbb{R})$ and write $\delta\varphi=u\circ\varphi$ with $u\in\mathfrak{g}^{-\infty}$. Differentiating $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ gives
$D_{\varphi}\Phi_{p}(\delta\varphi)=\varphi^{\prime 1/p-1}\delta\varphi^{\prime}.$
Since $\delta\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, we have
$D_{\varphi}\Phi_{p}(\delta\varphi)=(u^{\prime}\circ\varphi)\varphi^{\prime 1/p}.$
Computing the $L^{p}$ norm and changing variables $y=\varphi(x)$,
$\left\|D_{\varphi}\Phi_{p}(\delta\varphi)\right\|_{L^{p}}^{p}=\int_{\mathbb{R}}|u^{\prime}(\varphi(x))|^{p}\varphi^{\prime}(x)\,\mathrm{d}x=\int_{\mathbb{R}}|u^{\prime}(y)|^{p}\,\mathrm{d}y=\left\|u^{\prime}\right\|_{L^{p}}^{p}.$
Thus $\left\|D_{\varphi}\Phi_{p}(\delta\varphi)\right\|_{L^{p}}=F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)$.
Smoothness of $\Phi_{p}$ and $\Phi_{p}^{-1}$ follows from smoothness of $\varphi\mapsto\varphi^{\prime}$ and the Nemytskii operators $\rho\mapsto\rho^{1/p}$ and $f\mapsto(1+f/p)^{p}$ on the relevant open sets. ∎
Corollary 2.9.
The map $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is an isometric diffeomorphism from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),\dot{W}^{1,\infty})$ onto $(W^{\infty,1}(\mathbb{R}),L^{\infty})$ with inverse
$\Phi_{\infty}^{-1}(f)(x)=x+\int_{-\infty}^{x}(\mathrm{e}^{f(\tilde{x})}-1)\,\mathrm{d}\tilde{x}$.
Moreover, $\Phi_{p}(\varphi)\to\Phi_{\infty}(\varphi)$ pointwise as $p\to\infty$.
Remark 2.10.
The Jacobian $\varphi^{\prime}$ is the Radon-Nikodym derivative $\mathrm{d}(\varphi^{*}\mathrm{d}x)/\mathrm{d}x$. The coordinate $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ is the centered $p$-root of this density ratio, with centering chosen to ensure membership in $W^{\infty,1}(\mathbb{R})$ given the normalization $\varphi^{\prime}(\pm\infty)=1$. If $A=a\circ\varphi\in T_{\varphi}\operatorname{Diff}^{-\infty}(\mathbb{R})$, then
$D_{\varphi}\Phi_{p}(A)=(a_{x}\circ\varphi)\,\varphi^{\prime 1/p}.$
2.4. The multiplicative cocycle
Proposition 2.11.
Define $\Theta_{p}(\varphi)=1+\Phi_{p}(\varphi)/p=\varphi^{\prime 1/p}$ for $p<\infty$. Then
$\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\,\Theta_{p}(\psi).$
For $p=\infty$, this becomes the additive cocycle $\Phi_{\infty}(\varphi\circ\psi)=(\Phi_{\infty}(\varphi)\circ\psi)+\Phi_{\infty}(\psi)$.
Proof.
The chain rule gives $(\varphi\circ\psi)^{\prime}=(\varphi^{\prime}\circ\psi)\psi^{\prime}$. Taking $p$-th roots yields $\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\Theta_{p}(\psi)$. The $p=\infty$ case follows by taking logarithms. ∎
2.5. Distance and geodesics
Corollary 2.12.
For $p\in(1,\infty)$, the induced path metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is
$d_{p}(\varphi_{0},\varphi_{1})=\left\|\Phi_{p}(\varphi_{0})-\Phi_{p}(\varphi_{1})\right\|_{L^{p}}.$
The unique minimizing geodesic from $\varphi_{0}$ to $\varphi_{1}$ is $\varphi(t)=\Phi_{p}^{-1}((1-t)\Phi_{p}(\varphi_{0})+t\Phi_{p}(\varphi_{1}))$, explicitly
$\varphi(t,x)=x+\int_{-\infty}^{x}\left(\left((1-t)\varphi_{0}^{\prime 1/p}(s)+t\varphi_{1}^{\prime 1/p}(s)\right)^{p}-1\right)\mathrm{d}s.$
Proof.
The map $\Phi_{p}$ is an isometric embedding into the strictly convex Banach space $L^{p}(\mathbb{R})$ for $p\in(1,\infty)$. The induced path distance therefore equals the ambient norm distance, and the unique minimizing curve is the pullback of the affine segment between the endpoints. ∎
2.6. The canonical flat connection and a projective deformation
Proposition 2.13.
For $p\in(1,\infty]$, there exists a unique torsion-free flat connection $\nabla^{(p)}$ on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ characterized by
$D_{\varphi(t)}\Phi_{p}\left(\nabla^{(p)}_{\dot{\varphi}}A\right)=\frac{\mathrm{d}}{\mathrm{d}t}\left(D_{\varphi(t)}\Phi_{p}(A(t))\right)$
for smooth curves $\varphi(t)$ and vector fields $A(t)$ along them. In Eulerian variables $A=a\circ\varphi$, $\dot{\varphi}=u\circ\varphi$, one has
$\nabla^{(p)}_{\dot{\varphi}}A=\left(\partial_{t}a+u\,\partial_{x}a-\frac{p-1}{p}\,\partial_{x}^{-1}(a_{x}u_{x})\right)\circ\varphi,\qquad(\partial_{x}^{-1}f)(x):=\int_{-\infty}^{x}f(s)\,ds.$
The geodesic equation $\nabla^{(p)}_{\dot{\varphi}}\dot{\varphi}=0$ is equivalent to the strain equation
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0.$
Proof.
Since $\Phi_{p}$ is a global chart from $\operatorname{Diff}^{-\infty}(\mathbb{R})$ onto the open subset $U_{p}\subset W^{\infty,1}(\mathbb{R})$, define $\nabla^{(p)}$ by requiring that for every smooth curve $\varphi(t)$ and every smooth vector field $A(t)$ along $\varphi(t)$,
$D_{\varphi(t)}\Phi_{p}\!\left(\nabla^{(p)}_{\dot{\varphi}}A\right)=\frac{d}{dt}\Bigl(D_{\varphi(t)}\Phi_{p}(A(t))\Bigr).$
This determines a unique affine connection. The connection is torsion-free and flat because it is obtained from the canonical flat connection on the affine space $W^{\infty,1}(\mathbb{R})$ by transport through the chart $\Phi_{p}$.
To obtain the Eulerian formula, write $A=a\circ\varphi$ and $\dot{\varphi}=u\circ\varphi$. By Remark 2.10,
$D_{\varphi}\Phi_{p}(A)=(a_{x}\circ\varphi)\,\varphi^{\prime 1/p}.$
Differentiate in $t$:
$\frac{d}{dt}\Bigl((a_{x}\circ\varphi)\,\varphi^{\prime 1/p}\Bigr)=\Bigl((\partial_{t}a_{x}+u\,\partial_{x}a_{x})\circ\varphi\Bigr)\varphi^{\prime 1/p}+(a_{x}\circ\varphi)\,\frac{d}{dt}\bigl(\varphi^{\prime 1/p}\bigr).$
Since $\partial_{t}\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, one has
$\frac{d}{dt}\bigl(\varphi^{\prime 1/p}\bigr)=\frac{1}{p}\,\varphi^{\prime 1/p-1}\,\partial_{t}\varphi^{\prime}=\frac{1}{p}\,(u^{\prime}\circ\varphi)\,\varphi^{\prime 1/p}.$
Therefore
$\frac{d}{dt}\Bigl(D_{\varphi}\Phi_{p}(A)\Bigr)=\Bigl(\partial_{t}a_{x}+u\,\partial_{x}a_{x}+\frac{1}{p}a_{x}u_{x}\Bigr)\circ\varphi\cdot\varphi^{\prime 1/p}.$
Write $\nabla^{(p)}_{\dot{\varphi}}A=b\circ\varphi$. Then $D_{\varphi}\Phi_{p}(b\circ\varphi)=(b_{x}\circ\varphi)\varphi^{\prime 1/p}$, so the defining identity yields
$b_{x}=\partial_{t}a_{x}+u\,\partial_{x}a_{x}+\frac{1}{p}a_{x}u_{x}=\partial_{x}(\partial_{t}a+ua_{x})-\frac{p-1}{p}a_{x}u_{x}.$
Imposing the normalization $b(-\infty)=0$ and integrating from -$\infty$ gives
$b=\partial_{t}a+ua_{x}-\frac{p-1}{p}\,\partial_{x}^{-1}(a_{x}u_{x}),$
which is the stated formula.
For the geodesic equation, take $A=\dot{\varphi}$, so $a=u$. Then $\nabla^{(p)}_{\dot{\varphi}}\dot{\varphi}=0$ is equivalent to
$u_{t}+uu_{x}-\frac{p-1}{p}\,\partial_{x}^{-1}(u_{x}^{2})=0.$
Differentiating in $x$ and using $\partial_{x}\partial_{x}^{-1}=\mathrm{Id}$ gives
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0.$
Corollary 2.14 (Generalized Hunter-Saxton equation, third-order form).
The Eulerian velocity $u$ of an $F_{\dot{W}^{1,p}}$-geodesic satisfies
$u_{txx}+uu_{xxx}+\left(1+\frac{2}{p}\right)u_{x}u_{xx}=0.$
For $p=2$, this is the classical Hunter-Saxton equation. For $p=\infty$, it reduces to $u_{txx}+uu_{xxx}+u_{x}u_{xx}=0$.
Proof.
Set $v:=u_{x}$. The strain equation in Proposition 2.13 is
$v_{t}+uv_{x}+\frac{1}{p}v^{2}=0.$
Differentiating $v_{t}+uv_{x}+\frac{1}{p}v^{2}=0$ with respect to $x$ and then substituting $v=u_{x}$ (so that $v_{tx}=u_{txx}$, $v_{xx}=u_{xxx}$, and $v_{x}=u_{xx}$) yields the equation.
Definition 2.15 (Two-parameter family of connections).
For $p\in(1,\infty]$ and $\alpha\in\mathbb{R}$, define the connection $\nabla^{(p,\alpha)}$ on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ by
$\nabla^{(p,\alpha)}_{\dot{\varphi}}A=\nabla^{(p)}_{\dot{\varphi}}A+\alpha\cdot\bigl(\mathcal{C}(u,a)\bigr)\circ\varphi,\qquad\mathcal{C}(u,a):=u_{xx}a-ua_{xx}.$
Theorem 2.16 (Torsion and curvature of $\nabla^{(p,\alpha)}$).
Fix $p\in(1,\infty]$ and $\alpha\in\mathbb{R}$. Then $\nabla^{(p,\alpha)}$ is a smooth affine connection on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ with:
  1. (i)
    $\nabla^{(p,0)}=\nabla^{(p)}$ is torsion-free and flat;
  2. (ii)
    for right-trivialized fields $U=u\circ\varphi$ and $V=v\circ\varphi$,
    $T^{(p,\alpha)}(U,V)=\bigl(2\alpha\,\mathcal{C}(u,v)\bigr)\circ\varphi;$
  3. (iii)
    for right-trivialized fields $U=u\circ\varphi$, $V=v\circ\varphi$, $W=w\circ\varphi$,
    $R^{(p,\alpha)}(U,V)W=\alpha\,\Bigl([\partial_{x}^{2},\,\mathcal{L}_{[u,v]}]\,w\Bigr)\circ\varphi\;+\;O(\alpha^{2})\qquad\text{in }W^{\infty,1}(\mathbb{R}),$
    where $\mathcal{L}_{[u,v]}w=[u,v]\,w_{x}$ and $u,v]=uv_{x}-vu_{x}$. Equivalently,
    $\partial_{x}^{2},\,\mathcal{L}_{[u,v]}]\,w=([u,v]_{xx})\,w_{x}+2([u,v]_{x})\,w_{xx}.$
Proof.
Item (i) is immediate from the definition $\nabla^{(p,0)}=\nabla^{(p)}$ and Proposition 2.13. For (ii), note that $\mathcal{C}(u,v)=-\mathcal{C}(v,u)$, so the antisymmetrization of the added term contributes $2\alpha\,\mathcal{C}(u,v)$ to the torsion on right-trivialized fields.
For (iii), write $\nabla^{(p,\alpha)}=\nabla^{(p)}+\alpha\,\mathcal{C}$, where $\mathcal{C}$ is regarded as a $(1,2)$-tensor. For a connection perturbation of this form one has the exact identity
$R^{(p,\alpha)}(U,V)W=R^{(p)}(U,V)W+\alpha\,(d^{\nabla^{(p)}}\mathcal{C})(U,V)W+\alpha^{2}\bigl(\mathcal{C}(U,\mathcal{C}(V,W))-\mathcal{C}(V,\mathcal{C}(U,W))\bigr),$ (1)
where $d^{\nabla^{(p)}}\mathcal{C}$ is the covariant exterior derivative of the $\mathrm{End}(T\operatorname{Diff}^{-\infty}(\mathbb{R}))$-valued $1$-form $U\mapsto(W\mapsto\mathcal{C}(U,W))$. Since $\nabla^{(p)}$ is flat, $R^{(p)}=0$, and the $\alpha^{2}$-term is $O(\alpha^{2})$. Therefore it suffices to compute $(d^{\nabla^{(p)}}\mathcal{C})(U,V)W$.
Introduce the weight $\theta:=\varphi^{\prime 1/p}$. For a vector field $A=a\circ\varphi$ along $\varphi$ define the strain representative
$\widehat{A}:=D_{\varphi}\Phi_{p}(A)=(a_{x}\circ\varphi)\,\theta\in W^{\infty,1}(\mathbb{R})$
(using Remark 2.10). The defining property of $\nabla^{(p)}$ implies that for any smooth curve $\varphi(t)$ with $\dot{\varphi}(t)=U_{\varphi(t)}$ and any vector field $B(t)$ along it,
$D_{\varphi(t)}\Phi_{p}\!\left(\nabla^{(p)}_{\dot{\varphi}}B\right)=\frac{d}{dt}\bigl(D_{\varphi(t)}\Phi_{p}(B(t))\bigr).$
In particular, if $U$ and $B$ are genuine vector fields on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, this identity yields
[…]
where U(B^)U($\widehat${B}) denotes the directional derivative of the map $\varphi\mapsto\widehat{B}(\varphi)$ along $U$.
We now compute the directional derivative needed below. Fix $F\in W^{\infty,1}(\mathbb{R})$ and consider the function
$G(\varphi):=(F\circ\varphi)\,\theta=(F\circ\varphi)\,\varphi^{\prime 1/p}.$
Let $\varphi(t)$ be a curve with $\varphi(0)=\varphi$ and $\dot{\varphi}(0)=U_{\varphi}=u\circ\varphi$. Then
$\frac{d}{dt}\Big|_{t=0}(F\circ\varphi(t))=(F_{x}\circ\varphi)\,\dot{\varphi}=(F_{x}\circ\varphi)\,(u\circ\varphi),$
and since $\theta(t)=\varphi^{\prime}(t)^{1/p}$, one has
$\frac{d}{dt}\Big|_{t=0}\theta(t)=\frac{1}{p}\,\varphi^{\prime 1/p-1}\,\frac{d}{dt}\Big|_{t=0}\varphi^{\prime}(t)=\frac{1}{p}\,\varphi^{\prime 1/p-1}\,(u_{x}\circ\varphi)\,\varphi^{\prime}=\frac{1}{p}\,(u_{x}\circ\varphi)\,\theta.$
Therefore
$U\bigl((F\circ\varphi)\theta\bigr)=\bigl((uF_{x}+\tfrac{1}{p}u_{x}F)\circ\varphi\bigr)\,\theta.$ (3)
Next, observe that for $u,w$ on $\mathbb{R}$,
$\mathcal{C}(u,w)=u_{xx}w-uw_{xx}=-(uw_{x}-wu_{x})_{x}=-[u,w]_{x},$
hence $(\mathcal{C}(u,w))_{x}=-[u,w]_{xx}$. For right-trivialized fields $U=u\circ\varphi$ and $W=w\circ\varphi$,
$\widehat{\mathcal{C}(U,W)}=D_{\varphi}\Phi_{p}(\mathcal{C}(u,w)\circ\varphi)=\bigl((\mathcal{C}(u,w))_{x}\circ\varphi\bigr)\theta=-\bigl([u,w]_{xx}\circ\varphi\bigr)\theta.$
Define $A(U)W:=\mathcal{C}(U,W)$, so that $U\mapsto A(U)$ is an $\mathrm{End}(T\operatorname{Diff}^{-\infty}(\mathbb{R}))$-valued $1$-form. By definition of the covariant exterior derivative,
$(d^{\nabla^{(p)}}A)(U,V)W=\nabla^{(p)}_{U}(A(V)W)-\nabla^{(p)}_{V}(A(U)W)-A([U,V])W-A(V)\nabla^{(p)}_{U}W+A(U)\nabla^{(p)}_{V}W.$
[…]
$\widehat{(d^{\nabla^{(p)}}\mathcal{C})(U,V)W}=U\!\left(\widehat{\mathcal{C}(V,W)}\right)-V\!\left(\widehat{\mathcal{C}(U,W)}\right)-\widehat{\mathcal{C}([U,V],W)}$. (4)
For right-trivialized fields one has $U,V]=-([u,v]\circ\varphi)$, hence $\widehat{\mathcal{C}([U,V],W)}=\bigl([[u,v],w]_{xx}\circ\varphi\bigr)\theta$.
Applying (3) to the functions $F=-[v,w]_{xx}$ and $F=-[u,w]_{xx}$ gives
[…]
Substituting these expressions into (4) yields
$\widehat{(d^{\nabla^{(p)}}\mathcal{C})(U,V)W}$ $=\Bigl(-u([v,w]_{xx})_{x}+v([u,w]_{xx})_{x}-\tfrac{1}{p}u_{x}[v,w]_{xx}+\tfrac{1}{p}v_{x}[u,w]_{xx}-[[u,v],w]_{xx}\Bigr)\circ\varphi\;\theta.$ (5)
It remains to simplify the scalar expression in (5). The Jacobi identity gives
[u,[v,w]][v,[u,w]]=[[u,v],w].[u,[v,w]]-[v,[u,w]]=[[u,v],w].
Differentiating twice gives
$u,[v,w]]_{xx}-[v,[u,w]]_{xx}=[[u,v],w]_{xx}.$ (6)
A direct expansion of $v,w]=vw_{x}-wv_{x}$ and its derivatives yields the identity
$u([v,w]_{xx})_{x}-u_{x}[v,w]_{xx}=[u,[v,w]]_{xx}+([u,v]_{xx})w_{x}+2([u,v]_{x})w_{xx},$ (7)
and the same identity with $(u,v)$ exchanged. Substituting (7) and its $(u,v)$-swap into (5), and using (6) to cancel the terms involving $u,v],w]_{xx}$, one obtains
$\widehat{(d^{\nabla^{(p)}}\mathcal{C})(U,V)W}=\Bigl(([u,v]_{xx})w_{x}+2([u,v]_{x})w_{xx}\Bigr)\circ\varphi\;\theta.$
Since $D_{\varphi}\Phi_{p}$ is injective on the normalized tangent space, this implies
$(d^{\nabla^{(p)}}\mathcal{C})(U,V)W=\bigl(([u,v]_{xx})w_{x}+2([u,v]_{x})w_{xx}\bigr)\circ\varphi=\bigl([\partial_{x}^{2},\mathcal{L}_{[u,v]}]w\bigr)\circ\varphi.$
Combining this with (1) proves (iii). ∎
Remark 2.17 (Schwarzian as the projective curvature density).
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ define the Schwarzian derivative
$S(\varphi):=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}.$
Because $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$, one has $\varphi^{\prime\prime}\in W^{\infty,1}(\mathbb{R})$ and $\varphi^{\prime\prime\prime}\in W^{\infty,1}(\mathbb{R})$; moreover $\varphi^{\prime}$ is bounded away from 0, hence $S(\varphi)\in W^{\infty,1}(\mathbb{R})$. The cocycle identity $S(\varphi\circ\psi)=(S(\varphi)\circ\psi)\psi^{\prime 2}+S(\psi)$ exhibits $S$ as the universal projective cocycle for one-dimensional diffeomorphisms.
2.7. Parallel transport
Theorem 2.18 (Parallel transport for $\nabla^{(p)}$).
Let $\gamma\colon[0,1]\to\operatorname{Diff}^{-\infty}(\mathbb{R})$ be a smooth curve and let $A_{0}\in T_{\gamma(0)}\operatorname{Diff}^{-\infty}(\mathbb{R})$. There exists a unique $\nabla^{(p)}$-parallel vector field $A(t)$ along $\gamma$ with $A(0)=A_{0}$. It is characterized by the requirement that $D_{\gamma(t)}\Phi_{p}(A(t))$ is independent of $t$, namely
$D_{\gamma(t)}\Phi_{p}(A(t))\equiv D_{\gamma(0)}\Phi_{p}(A_{0})\qquad\text{in }W^{\infty,1}(\mathbb{R}).$
Equivalently, writing $A_{0}=\delta\varphi_{0}$ and setting $w:=D_{\gamma(0)}\Phi_{p}(A_{0})=\gamma(0)^{\prime\,1/p-1}(\delta\varphi_{0})^{\prime}$, the transported vector at $t=1$ is
[…]
For every loop $\gamma$ based at $\varphi$, the holonomy satisfies $\mathrm{Hol}_{\nabla^{(p)}}(\gamma)=\mathrm{Id}$.
Proof.
A vector field $A(t)$ along $\gamma(t)$ is $\nabla^{(p)}$-parallel if and only if $\nabla^{(p)}_{\dot{\gamma}}A=0$. By Proposition 2.13, this is equivalent to
$\frac{d}{dt}\Bigl(D_{\gamma(t)}\Phi_{p}(A(t))\Bigr)=0,$
hence $D_{\gamma(t)}\Phi_{p}(A(t))$ is constant and equal to $w:=D_{\gamma(0)}\Phi_{p}(A_{0})$.
At $t=1$ we have $D_{\gamma(1)}\Phi_{p}(A(1))=w$. Writing $A(1)=\delta\varphi_{1}$ and using $D_{\gamma(1)}\Phi_{p}(\delta\varphi_{1})=\gamma(1)^{\prime\,1/p-1}(\delta\varphi_{1})^{\prime}$, we obtain
$(\delta\varphi_{1})^{\prime}(x)=\gamma(1)^{\prime}(x)^{\,1-1/p}\,w(x).$
With the normalization $\delta\varphi_{1}(-\infty)=0$, integration yields
$\delta\varphi_{1}(x)=\int_{-\infty}^{x}\gamma(1)^{\prime}(s)^{\,1-1/p}\,w(s)\,ds,$
which is the stated formula.
Let $\gamma$ be a loop based at $\varphi$, so $\gamma(0)=\gamma(1)=\varphi$, and let $A(t)$ be the unique $\nabla^{(p)}$-parallel field with $A(0)=A_{0}$. The parallelism implies $D_{\gamma(t)}\Phi_{p}(A(t))\equiv D_{\varphi}\Phi_{p}(A_{0})$, hence $D_{\varphi}\Phi_{p}(A(1))=D_{\varphi}\Phi_{p}(A_{0})$. If $D_{\varphi}\Phi_{p}(\delta\varphi)=0$, then $\varphi^{\prime\,1/p-1}\delta\varphi^{\prime}=0$, so $\delta\varphi^{\prime}$ vanishes and $\delta\varphi$ is constant; the normalization at -$\infty$ forces $\delta\varphi\equiv 0$. Thus $D_{\varphi}\Phi_{p}$ is injective on the normalized tangent space, and $A(1)=A_{0}$. Therefore the holonomy is the identity. ∎
2.8. Geodesic action functional and Hamiltonian formulation
Definition 2.19 (Geodesic action functional).
For $p\in(1,\infty)$ and a $C^{1}$ path $\gamma\colon[0,T]\to\operatorname{Diff}^{-\infty}(\mathbb{R})$, define the $p$-action
$\mathcal{A}_{p}[\gamma]=\int_{0}^{T}L_{p}(\gamma(t),\dot{\gamma}(t))\,dt,\qquad L_{p}(\varphi,\delta\varphi)=\frac{1}{p}\bigl\|\bigl(\delta\varphi\circ\varphi^{-1}\bigr)^{\prime}\bigr\|_{L^{p}}^{p}.$
In Eulerian variables $\dot{\gamma}=u\circ\gamma$, this reads $L_{p}=\frac{1}{p}\int_{\mathbb{R}}|u^{\prime}(x)|^{p}\,dx.$
Theorem 2.20 (Euler-Lagrange equation).
Critical points of $\mathcal{A}_{p}$ among paths with fixed endpoints are precisely the $F_{\dot{W}^{1,p}}$-geodesics. In Eulerian variables $u=\dot{\gamma}\circ\gamma^{-1}$ they satisfy the generalized Hunter-Saxton equation of Corollary 2.14.
Proof.
Let $\gamma_{\varepsilon}$ be a smooth variation with fixed endpoints $\gamma_{\varepsilon}(0)=\gamma(0)$ and $\gamma_{\varepsilon}(T)=\gamma(T)$. Define
$u_{\varepsilon}:=\dot{\gamma}_{\varepsilon}\circ\gamma_{\varepsilon}^{-1},\qquad\eta:=\left.\partial_{\varepsilon}\gamma_{\varepsilon}\right|_{\varepsilon=0}\circ\gamma^{-1}.$
Then $\delta\gamma:=\left.\partial_{\varepsilon}\gamma_{\varepsilon}\right|_{\varepsilon=0}=\eta\circ\gamma$. Differentiating the identity $\gamma_{\varepsilon}\circ\gamma_{\varepsilon}^{-1}=\operatorname{id}$ at $\varepsilon=0$ gives
$\delta(\gamma_{\varepsilon}^{-1})=-(\eta/\gamma^{\prime})\circ\gamma^{-1}.$
Differentiate $u_{\varepsilon}=\dot{\gamma}_{\varepsilon}\circ\gamma_{\varepsilon}^{-1}$ at $\varepsilon=0$:
$\delta u=(\delta\dot{\gamma})\circ\gamma^{-1}+(\dot{\gamma}_{x}\circ\gamma^{-1})\,\delta(\gamma^{-1}).$
Since $\delta\dot{\gamma}=\partial_{t}(\delta\gamma)=\partial_{t}(\eta\circ\gamma)=(\eta_{t}+u\eta_{x})\circ\gamma$, one has $(\delta\dot{\gamma})\circ\gamma^{-1}=\eta_{t}+u\eta_{x}$. Also $\dot{\gamma}_{x}=(u_{x}\circ\gamma)\gamma^{\prime}$, so $(\dot{\gamma}_{x}\circ\gamma^{-1})=u_{x}\,(\gamma^{\prime}\circ\gamma^{-1})$ and $\delta(\gamma^{-1})=-(\eta/(\gamma^{\prime}\circ\gamma^{-1}))$, which gives $(\dot{\gamma}_{x}\circ\gamma^{-1})\,\delta(\gamma^{-1})=-u_{x}\eta$. Therefore
$\delta u=\eta_{t}+u\eta_{x}-u_{x}\eta=\eta_{t}+[u,\eta],\qquad[u,\eta]=u\eta_{x}-\eta u_{x}.$ (8)
Differentiating in $x$ yields
$\delta u_{x}=\partial_{t}\eta_{x}+u\,\eta_{xx}-\eta\,u_{xx}.$ (9)
Set $m:=|u_{x}|^{p-2}u_{x}$. Since
$\mathcal{A}_{p}[\gamma]=\frac{1}{p}\int_{0}^{T}\int_{\mathbb{R}}|u_{x}|^{p}\,dx\,dt,$
the first variation is
$\delta\mathcal{A}_{p}=\int_{0}^{T}\int_{\mathbb{R}}m\,\delta u_{x}\,dx\,dt.$
Insert (9). Using that $\eta(0,\cdot)=\eta(T,\cdot)=0$ and the decay at $\pm\infty$, integration by parts gives
$\int_{0}^{T}\int_{\mathbb{R}}m\,\partial_{t}\eta_{x}\,dx\,dt=-\int_{0}^{T}\int_{\mathbb{R}}m_{t}\,\eta_{x}\,dx\,dt=\int_{0}^{T}\int_{\mathbb{R}}(m_{t})_{x}\,\eta\,dx\,dt,$
$\int_{0}^{T}\int_{\mathbb{R}}m\,u\,\eta_{xx}\,dx\,dt=-\int_{0}^{T}\int_{\mathbb{R}}(mu)_{x}\,\eta_{x}\,dx\,dt=\int_{0}^{T}\int_{\mathbb{R}}(mu)_{xx}\,\eta\,dx\,dt,$
and
$\int_{0}^{T}\int_{\mathbb{R}}(-m\,\eta\,u_{xx})\,dx\,dt=-\int_{0}^{T}\int_{\mathbb{R}}(m\,u_{xx})\,\eta\,dx\,dt.$
Hence
$\delta\mathcal{A}_{p}=\int_{0}^{T}\int_{\mathbb{R}}\Bigl((m_{t})_{x}+(mu)_{xx}-m\,u_{xx}\Bigr)\eta\,dx\,dt.$
Since $\eta$ is arbitrary, stationarity is equivalent to
$(m_{t})_{x}+(mu)_{xx}-m\,u_{xx}=0,\qquad m=|u_{x}|^{p-2}u_{x}.$ (10)
Expanding $(mu)_{xx}=u\,m_{xx}+2u_{x}m_{x}+mu_{xx}$ shows (10) is equivalent to
$(m_{t})_{x}+u\,m_{xx}+2u_{x}m_{x}=0.$
This is equivalent to the generalized Hunter-Saxton equation in Corollary 2.14 after differentiating the relation $m=|u_{x}|^{p-2}u_{x}$ and eliminating $m$. ∎
Theorem 2.21 (Hamiltonian and Legendre duality).
Let q=p/(p1)q=p/(p-1). The fiberwise Legendre transform of $L_{p}$ in Eulerian strain variables yields the Hamiltonian
$H_{p}(m)=\sup_{v\in L^{p}}\left\{\int_{\mathbb{R}}m(x)v(x)\,dx-\frac{1}{p}\int_{\mathbb{R}}|v(x)|^{p}\,dx\right\}=\frac{1}{q}\int_{\mathbb{R}}|m(x)|^{q}\,dx=\frac{p-1}{p}\int_{\mathbb{R}}|m(x)|^{p/(p-1)}\,dx.$
Proof.
Let q=p/(p1)q=p/(p-1). For real numbers $m,v$ one has Young’s inequality
$mv\leq\frac{1}{q}|m|^{q}+\frac{1}{p}|v|^{p},$
with equality when $v=|m|^{q-2}m$. Integrating over $\mathbb{R}$ gives, for any $v\in L^{p}(\mathbb{R})$,
$\int_{\mathbb{R}}mv\,dx-\frac{1}{p}\int_{\mathbb{R}}|v|^{p}\,dx\leq\frac{1}{q}\int_{\mathbb{R}}|m|^{q}\,dx.$
Equality is attained by $v=|m|^{q-2}m\in L^{p}(\mathbb{R})$. Therefore the displayed supremum equals $\frac{1}{q}\int_{\mathbb{R}}|m|^{q}\,dx$. ∎
Proposition 2.22 (Hamilton-Jacobi on the linearizing chart).
Fix $p\in(1,\infty)$ and let $f(t):=\Phi_{p}(\gamma(t))\in U_{p}\subset L^{p}(\mathbb{R})$. Then
$\mathcal{A}_{p}[\gamma]=\frac{1}{p}\int_{0}^{T}\|\dot{f}(t)\|_{L^{p}}^{p}\,dt,$
and the value function
$S_{T}(f_{0},f_{1}):=\inf\{\mathcal{A}_{p}[\gamma]\mid\Phi_{p}(\gamma(0))=f_{0},\ \Phi_{p}(\gamma(T))=f_{1}\}$
is explicitly
$S_{T}(f_{0},f_{1})=\frac{T^{1-p}}{p}\,\|f_{1}-f_{0}\|_{L^{p}}^{p}.$
Consequently the dynamic value function $S(t,f)$ for fixed $f_{0}$ solves the Hamilton-Jacobi equation on $L^{p}(\mathbb{R})$
$\partial_{t}S(t,f)+H_{p}\bigl(\nabla S(t,f)\bigr)=0$
in the sense of the duality between $L^{p}$ and $L^{q}$.
Proof.
Let $\gamma(t)$ be a $C^{1}$ path in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $f(t):=\Phi_{p}(\gamma(t))$. By Theorem 2.8, $\Phi_{p}$ is a fiberwise isometry, hence
$F_{\dot{W}^{1,p}}(\gamma(t),\dot{\gamma}(t))=\|\dot{f}(t)\|_{L^{p}}.$
Therefore
$\mathcal{A}_{p}[\gamma]=\int_{0}^{T}\frac{1}{p}\,F_{\dot{W}^{1,p}}(\gamma(t),\dot{\gamma}(t))^{p}\,dt=\frac{1}{p}\int_{0}^{T}\|\dot{f}(t)\|_{L^{p}}^{p}\,dt.$
The endpoint minimization problem is thus the $p$-energy minimization for curves in $L^{p}(\mathbb{R})$ between $f_{0}$ and $f_{1}$. The unique minimizer is the constant-velocity segment
$f(t)=\Bigl(1-\frac{t}{T}\Bigr)f_{0}+\frac{t}{T}f_{1},\qquad\dot{f}(t)=\frac{f_{1}-f_{0}}{T}.$
Evaluating its action gives
$S_{T}(f_{0},f_{1})=\frac{1}{p}\int_{0}^{T}\left\|\frac{f_{1}-f_{0}}{T}\right\|_{L^{p}}^{p}\,dt=\frac{T^{1-p}}{p}\,\|f_{1}-f_{0}\|_{L^{p}}^{p}.$
The Hamilton-Jacobi equation is the standard dynamic programming identity for the Lagrangian $L(v)=\frac{1}{p}\|v\|_{L^{p}}^{p}$ and its Legendre dual Hamiltonian $H(\xi)=\frac{1}{q}\|\xi\|_{L^{q}}^{q}$ (Theorem 2.21), with $\nabla S$ interpreted via the $L^{p}$-$L^{q}$ dual pairing. ∎
2.9. Asymptotics as $p\to\infty$
Proposition 2.23.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $f_{\infty}=\log(\varphi^{\prime})\in W^{\infty,1}(\mathbb{R})$,
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+O_{W^{k,1}}\!\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
Set $\rho:=\varphi^{\prime}>0$ and $f_{\infty}:=\log\rho\in W^{\infty,1}(\mathbb{R})$. Then for each $p\geq 1$,
$\Phi_{p}(\varphi)=p\bigl(\rho^{1/p}-1\bigr)=p\bigl(e^{f_{\infty}/p}-1\bigr).$
Fix $k\in\mathbb{N}$. Let $\mathcal{B}\subset\mathrm{Diff}^{-\infty}(\mathbb{R})$ be a bounded set. By definition of the Fréchet structure, boundedness of $\mathcal{B}$ implies that for every $m\leq k+1$ there is a constant $M_{m}(\mathcal{B})$ such that
$\sup_{\varphi\in\mathcal{B}}\|f_{\infty}\|_{W^{m,1}(\mathbb{R})}\leq M_{m}(\mathcal{B}).$
In particular, for each $j\leq k$ the derivative $f_{\infty}^{(j)}$ belongs to $W^{1,1}(\mathbb{R})$ uniformly on $\mathcal{B}$, hence it is bounded uniformly in $L^{\infty}(\mathbb{R})$. Indeed, if $g\in W^{1,1}(\mathbb{R})$ and $g\in L^{1}(\mathbb{R})$, then $g$ has limits $g(\pm\infty)=0$ and for every $x\in\mathbb{R}$,
$|g(x)|=\left|\int_{x}^{\infty}g^{\prime}(t)\,dt\right|\leq\|g^{\prime}\|_{L^{1}(\mathbb{R})},$
so $\|g\|_{L^{\infty}}\leq\|g^{\prime}\|_{L^{1}}\leq\|g\|_{W^{1,1}}$. Applying this with $g=f_{\infty}^{(j)}$ yields uniform bounds
$\sup_{\varphi\in\mathcal{B}}\|f_{\infty}^{(j)}\|_{L^{\infty}(\mathbb{R})}\leq M_{j+1}(\mathcal{B}),\qquad 0\leq j\leq k.$
We now expand the exponential with an integral remainder. For every real $z$,
$e^{z}=1+z+\frac{z^{2}}{2}+\frac{z^{3}}{6}+\frac{z^{4}}{6}\int_{0}^{1}(1-t)^{3}e^{tz}\,dt,$
which follows by applying Taylor’s theorem with integral remainder to the $C^{\infty}$ function $e^{z}$. With $z=f_{\infty}/p$ this gives the pointwise identity
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+R_{p},\qquad R_{p}(x)=\frac{1}{6p^{3}}f_{\infty}(x)^{4}E_{p}(x),$
where
$E_{p}(x):=\int_{0}^{1}(1-t)^{3}e^{tf_{\infty}(x)/p}\,dt.$
Since $0\leq(1-t)^{3}\leq 1$ and $e^{tf_{\infty}/p}\leq e^{\|f_{\infty}\|_{L^{\infty}}/p}\leq e^{\|f_{\infty}\|_{L^{\infty}}}$, we have the uniform bound
$\|E_{p}\|_{L^{\infty}(\mathbb{R})}\leq e^{\|f_{\infty}\|_{L^{\infty}(\mathbb{R})}}.$
In particular, for $m=0$,
$\|R_{p}\|_{L^{1}(\mathbb{R})}\leq\frac{1}{6p^{3}}\|E_{p}\|_{L^{\infty}}\|f_{\infty}^{4}\|_{L^{1}}\leq\frac{C(\mathcal{B})}{p^{3}},$
because $f_{\infty}\in L^{\infty}\cap L^{1}$ uniformly on $\mathcal{B}$, hence $\|f_{\infty}^{4}\|_{L^{1}}\leq\|f_{\infty}\|_{L^{\infty}}^{3}\|f_{\infty}\|_{L^{1}}$ is uniformly bounded on $\mathcal{B}$.
To control $\partial_{x}^{m}R_{p}$ in $L^{1}$ for $1\leq m\leq k$, we estimate $\partial_{x}^{m}(f_{\infty}^{4}E_{p})$ and then divide by $p^{3}$. By Leibniz,
$\partial_{x}^{m}(f_{\infty}^{4}E_{p})=\sum_{a=0}^{m}\binom{m}{a}\partial_{x}^{a}(f_{\infty}^{4})\,\partial_{x}^{m-a}E_{p}.$
The derivatives $\partial_{x}^{a}(f_{\infty}^{4})$ are finite sums of products of $f_{\infty}^{(j)}$ with total differentiation order $a$ and with at most four factors. For each such product, one factor can be placed in $L^{1}$ (choose a factor carrying at least one derivative when $a\geq 1$; when $a=0$ use $f_{\infty}\in L^{1}$), and all remaining factors are placed in $L^{\infty}$. Using the uniform $L^{\infty}$ bounds for $f_{\infty}^{(j)}$ recorded above, it follows that for each $a\leq k$,
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{a}(f_{\infty}^{4})\|_{L^{1}(\mathbb{R})}\leq C_{a}(\mathcal{B}).$
It remains to bound $\partial_{x}^{r}E_{p}$ in $L^{\infty}$ for $0\leq r\leq k$. Write $E_{p}=F_{p}\circ f_{\infty}$, where
$F_{p}(s):=\int_{0}^{1}(1-t)^{3}e^{ts/p}\,dt.$
Then $F_{p}^{(n)}(s)=\int_{0}^{1}(1-t)^{3}(t/p)^{n}e^{ts/p}\,dt$, hence for every $n\geq 0$ and $s\in\mathbb{R}$,
$|F_{p}^{(n)}(s)|\leq\frac{1}{p^{n}}e^{|s|/p}\leq e^{|s|},$
using $p\geq 1$. Applying the Faà di Bruno formula to the composition $F_{p}\circ f_{\infty}$ yields, for each $r\geq 1$,
$\partial_{x}^{r}E_{p}(x)=\sum_{n=1}^{r}F_{p}^{(n)}(f_{\infty}(x))\sum_{\begin{subarray}{c}\alpha_{1}+\cdots+\alpha_{r}=n\\ \alpha_{1}+2\alpha_{2}+\cdots+r\alpha_{r}=r\end{subarray}}c_{\alpha}\prod_{j=1}^{r}\bigl(f_{\infty}^{(j)}(x)\bigr)^{\alpha_{j}},$
with combinatorial coefficients $c_{\alpha}$ depending only on $r$. Each product on the right involves derivatives $f_{\infty}^{(j)}$ with $1\leq j\leq r$, hence every factor is bounded in $L^{\infty}$ uniformly on $\mathcal{B}$. Since $F_{p}^{(n)}(f_{\infty})$ is also bounded in $L^{\infty}$ uniformly on $\mathcal{B}$ by the previous inequality and the uniform $L^{\infty}$ bound for $f_{\infty}$, we obtain
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{r}E_{p}\|_{L^{\infty}(\mathbb{R})}\leq C_{r}(\mathcal{B}),\qquad 0\leq r\leq k,$
where $r=0$ is the earlier bound on $\|E_{p}\|_{L^{\infty}}$.
Combining these bounds with the Leibniz expansion gives, for each $m\leq k$,
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m}(f_{\infty}^{4}E_{p})\|_{L^{1}(\mathbb{R})}\leq\sum_{a=0}^{m}\binom{m}{a}\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{a}(f_{\infty}^{4})\|_{L^{1}}\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m-a}E_{p}\|_{L^{\infty}}\leq C_{m}(\mathcal{B}).$
Therefore
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m}R_{p}\|_{L^{1}(\mathbb{R})}\leq\frac{1}{6p^{3}}C_{m}(\mathcal{B}),\qquad 0\leq m\leq k,$
and hence
$\sup_{\varphi\in\mathcal{B}}\|R_{p}\|_{W^{k,1}(\mathbb{R})}\leq\frac{C_{k}(\mathcal{B})}{p^{3}}.$
And so
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+O_{W^{k,1}}\!\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly for $\varphi$ ranging in bounded subsets of $\mathrm{Diff}^{-\infty}(\mathbb{R})$.
3. Blow-up Analysis and Large Deviations for the Generalized Hunter-Saxton Equation
The generalized Hunter-Saxton equation (see [16] for the original $p=2$ case)
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0$ (11)
is the Eulerian geodesic strain equation for the $\dot{W}^{1,p}$ metric. We develop a complete blow-up theory.
3.1. The strain variable and Riccati structure
Definition 3.1 (Strain variable).
For a smooth curve $\varphi(t)\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with Eulerian velocity $u=\dot{\varphi}\circ\varphi^{-1}$, define the strain $v=u_{x}$ and the Lagrangian strain $V(t,\xi)=v(t,\varphi(t,\xi))$.
Proposition 3.2 (Lagrangian Riccati equation).
Along any $F_{\dot{W}^{1,p}}$-geodesic.
  1. (i)
    The Lagrangian strain satisfies the autonomous Riccati ODE
    $\frac{\mathrm{d}V}{\mathrm{d}t}+\frac{1}{p}V^{2}=0.$
  2. (ii)
    The explicit solution with initial data $V_{0}(\xi)=v_{0}(\xi)$ is
    $V(t,\xi)=\frac{V_{0}(\xi)}{1+\frac{t}{p}V_{0}(\xi)}.$
  3. (iii)
    The Jacobian evolves as $\partial_{\xi}\varphi(t,\xi)=\varphi^{\prime}_{0}(\xi)\left(1+\frac{t}{p}V_{0}(\xi)\right)^{p}$.
Proof.
Let $\varphi(t,\cdot)$ be an $F_{\dot{W}^{1,p}}$-geodesic and write $u(t,x)=\partial_{t}\varphi(t,\varphi^{-1}(t,x))$ for its Eulerian velocity. Denote by
$V(t,\xi):=(u_{x}\circ\varphi)(t,\xi)$
the Lagrangian strain, i.e. the spatial derivative of the Eulerian velocity pulled back along the flow.
The straightening theorem asserts that along an $F_{\dot{W}^{1,p}}$-geodesic the quantity
$w(t,x):=\Phi_{p}(\varphi(t))(x)=p\bigl(\varphi^{\prime}(t,x)^{1/p}-1\bigr)$
evolves as an affine function of $t$ in Eulerian coordinates, equivalently that the Eulerian strain $u_{x}$ satisfies the scalar evolution
$(\partial_{t}+u\,\partial_{x})u_{x}+\frac{1}{p}u_{x}^{2}=0$ (12)
pointwise wherever the geodesic is defined. Composing (12) with the Lagrangian map $\varphi$ and using that $\partial_{t}(f\circ\varphi)=(\partial_{t}f+u\,\partial_{x}f)\circ\varphi$ for any smooth Eulerian field $f$, we obtain
$\partial_{t}V=\partial_{t}(u_{x}\circ\varphi)=\bigl(\partial_{t}u_{x}+u\,\partial_{x}u_{x}\bigr)\circ\varphi.$
Fix an $F_{\dot{W}^{1,p}}$-geodesic $\varphi(t,\cdot)$ and let $u(t,x)$ be its Eulerian velocity, so that the flow equation is
$\partial_{t}\varphi(t,\xi)=u\bigl(t,\varphi(t,\xi)\bigr).$
For any sufficiently regular Eulerian field $f(t,x)$, define its pullback to Lagrangian variables by
$F(t,\xi):=(f\circ\varphi)(t,\xi)=f\bigl(t,\varphi(t,\xi)\bigr).$
Differentiating in $t$ and using the chain rule gives
$\partial_{t}F(t,\xi)=\partial_{t}f\bigl(t,\varphi(t,\xi)\bigr)+\partial_{x}f\bigl(t,\varphi(t,\xi)\bigr)\,\partial_{t}\varphi(t,\xi)=\Bigl(\partial_{t}f+u\,\partial_{x}f\Bigr)\bigl(t,\varphi(t,\xi)\bigr).$
In other words,
$\frac{d}{dt}\bigl(f\circ\varphi\bigr)=\bigl((\partial_{t}+u\partial_{x})f\bigr)\circ\varphi,$
so pullback by the flow converts the Eulerian material derivative $\partial_{t}+u\partial_{x}$ into the ordinary time derivative $\frac{d}{dt}$ at fixed label $\xi$.
Applying this with $f=u_{x}$ yields, for the Lagrangian strain $V:=u_{x}\circ\varphi$,
$\partial_{t}V=\partial_{t}(u_{x}\circ\varphi)=\bigl(\partial_{t}u_{x}+u\,\partial_{x}u_{x}\bigr)\circ\varphi.$
Thus the term $u\,\partial_{x}u_{x}$ does not disappear by itself; rather, it is exactly the extra contribution coming from the time-dependence of the footpoint $x=\varphi(t,\xi)$, and it is absorbed into $\partial_{t}V$, which is the pullback of the Eulerian material derivative of $u_{x}$.
Thus (12) transports to the autonomous Riccati ODE
$\frac{d}{dt}V(t,\xi)+\frac{1}{p}V(t,\xi)^{2}=0,$
for each fixed label $\xi\in\mathbb{R}$, which is (i).
To solve it, fix $\xi$ and write $V(t)=V(t,\xi)$. On any time interval where $1+\frac{t}{p}V(0)\neq 0$, the equation is separable, that is,
$\frac{dV}{V^{2}}=-\frac{1}{p}\,dt.$
Integrating gives 1/V=t/p+C-1/V=-t/p+C, and imposing $V(0)=V_{0}(\xi)$ yields $C=-1/V_{0}(\xi)$. Solving for $V$ gives
$V(t,\xi)=\frac{V_{0}(\xi)}{1+\frac{t}{p}V_{0}(\xi)},$
which is (ii).
For (iii), differentiate the flow equation $\partial_{t}\varphi(t,\xi)=u(t,\varphi(t,\xi))$ with respect to $\xi$ to obtain
$\partial_{t}(\partial_{\xi}\varphi)(t,\xi)=u_{x}(t,\varphi(t,\xi))\,\partial_{\xi}\varphi(t,\xi)=V(t,\xi)\,\partial_{\xi}\varphi(t,\xi).$
Hence $\partial_{\xi}\varphi$ satisfies the linear ODE
$\partial_{t}\log\bigl(\partial_{\xi}\varphi(t,\xi)\bigr)=V(t,\xi).$
Integrating from 0 to $t$ and using $\partial_{\xi}\varphi(0,\xi)=\varphi_{0}^{\prime}(\xi)$ gives
$\partial_{\xi}\varphi(t,\xi)=\varphi_{0}^{\prime}(\xi)\exp\!\left(\int_{0}^{t}V(s,\xi)\,ds\right).$
Substituting the explicit expression from (ii) yields
$\int_{0}^{t}V(s,\xi)\,ds=\int_{0}^{t}\frac{V_{0}(\xi)}{1+\frac{s}{p}V_{0}(\xi)}\,ds=p\log\!\left(1+\frac{t}{p}V_{0}(\xi)\right),$
and therefore
$\partial_{\xi}\varphi(t,\xi)=\varphi_{0}^{\prime}(\xi)\left(1+\frac{t}{p}V_{0}(\xi)\right)^{p},$
which is (iii).
3.2. Precise blow-up criterion
Theorem 3.3 (Sharp blow-up criterion).
Let $\varphi(t)$ be the $F_{\dot{W}^{1,p}}$-geodesic with initial data $(\varphi_{0},u_{0})\in T\operatorname{Diff}^{-\infty}(\mathbb{R})$ and let $v_{0}=\partial_{x}u_{0}$. Define
$T^{*}=\begin{cases}-p/\inf_{\mathbb{R}}v_{0}&\text{if }\inf_{\mathbb{R}}v_{0}<0,\\ +\infty&\text{if }\inf_{\mathbb{R}}v_{0}\geq 0.\end{cases}$
Then.
  1. (i)
    The geodesic exists smoothly for $t\in[0,T^{*})$.
  2. (ii)
    If $T^{*}<\infty$, then $\lim_{t\nearrow T^{*}}\inf_{x}u_{x}(t,x)=$-$\infty$.
  3. (iii)
    The blow-up is of gradient catastrophe type. $u(t,\cdot)$ remains bounded as $t\nearrow T^{*}$, but $u_{x}$ blows up at precisely the point(s) where $v_{0}$ achieves its infimum.
Proof.
Let $v=u_{x}$ and $V(t,\xi)=v(t,\varphi(t,\xi))$ be the Lagrangian strain. Since $\varphi_{0}$ is an increasing diffeomorphism, composition preserves infima, hence
$\inf_{\xi\in\mathbb{R}}V_{0}(\xi)=\inf_{x\in\mathbb{R}}v_{0}(x)=:m.$
Proposition 3.2 gives
$V(t,\xi)=\frac{V_{0}(\xi)}{1+\frac{t}{p}V_{0}(\xi)},\qquad\partial_{\xi}\varphi(t,\xi)=\varphi_{0}^{\prime}(\xi)\left(1+\frac{t}{p}V_{0}(\xi)\right)^{p},$ (13)
as long as $1+\frac{t}{p}V_{0}(\xi)>0$. With $T^{*}$ as in the statement, for every $t\in[0,T^{*})$ we have $1+\frac{t}{p}V_{0}(\xi)\geq 1+\frac{t}{p}m>0$ for all $\xi$, hence $\partial_{\xi}\varphi(t,\xi)>0$ everywhere. Therefore $\varphi(t,\cdot)$ remains an increasing diffeomorphism on $0,T^{*})$, and the explicit formulas show $\varphi(t)\in\mathrm{Diff}^{-\infty}(\mathbb{R})$ and depends smoothly on $t$ on this interval. This proves (i).
Assume $T^{*}<\infty$, so $m<0$. For fixed $t<T^{*}$ the map $s\mapsto s/(1+\frac{t}{p}s)$ is strictly increasing on $\{1+\frac{t}{p}s>0\}$, hence
$\inf_{\xi}V(t,\xi)=\frac{m}{1+\frac{t}{p}m}.$
Since $v(t,x)=V(t,\varphi^{-1}(t,x))$ and $\varphi(t,\cdot)$ is a diffeomorphism for $t<T^{*}$, we also have $\inf_{x}u_{x}(t,x)=\inf_{x}v(t,x)=\inf_{\xi}V(t,\xi)$, so
$\inf_{x}u_{x}(t,x)=\frac{m}{1+\frac{t}{p}m}\xrightarrow[t\nearrow T^{*}$$\infty$,
which is (ii).
For (iii), the explicit formula for $V(t,\xi)$ shows that $V(t,\xi)\to$-$\infty$ as $t\nearrow T^{*}$ precisely for those $\xi$ with $V_{0}(\xi)=m$, and $V(t,\xi)$ remains bounded for $\xi$ with $V_{0}(\xi)>m$. Translating to Eulerian variables using $u_{x}(t,\varphi(t,\xi))=V(t,\xi)$ shows that the blow-up of $u_{x}$ occurs exactly along the trajectories issuing from the minimizers of $v_{0}$.
Finally, since $u(t,x)\to 0$ as $x\to$-$\infty$, we have $u(t,x)=\int_{-\infty}^{x}v(t,y)\,dy$ and therefore $\|u(t,\cdot)\|_{L^{\infty}}\leq\|v(t,\cdot)\|_{L^{1}}$. Changing variables $y=\varphi(t,\xi)$ and using (13) gives
$\|v(t,\cdot)\|_{L^{1}}=\int_{\mathbb{R}}|V(t,\xi)|\,\partial_{\xi}\varphi(t,\xi)\,d\xi=\int_{\mathbb{R}}|V_{0}(\xi)|\,\varphi_{0}^{\prime}(\xi)\left(1+\frac{t}{p}V_{0}(\xi)\right)^{p-1}d\xi.$
Because $V_{0}\in W^{\infty,1}(\mathbb{R})\subset L^{\infty}(\mathbb{R})$ and $t<T^{*}$, the factor $\bigl(1+\frac{t}{p}V_{0}(\xi)\bigr)^{p-1}$ is uniformly bounded in $\xi$, and $\varphi_{0}^{\prime}$ is bounded. Since $V_{0}\in L^{1}$, the right-hand side stays finite (indeed uniformly bounded) for $t\in[0,T^{*})$, hence $\|u(t,\cdot)\|_{L^{\infty}}$ remains bounded as $t\nearrow T^{*}$.
Corollary 3.4 (Characterization of complete geodesics).
The geodesic $\varphi(t)$ exists for all $t\geq 0$ if and only if $u_{0}$ is non-decreasing, i.e., $\partial_{x}u_{0}\geq 0$ everywhere. In this case, $\varphi(t)$ converges as $t\to\infty$ to a limiting diffeomorphism $\varphi_{\infty}\in\operatorname{Diff}^{-\infty}(\mathbb{R})$.
3.3. Blow-up rate and profile
Theorem 3.5 (Blow-up rate).
Assume $v_{0}$ achieves a unique non-degenerate minimum at $\xi^{*}$ with $v_{0}(\xi^{*})=m<0$ and $v_{0}^{\prime\prime}(\xi^{*})>0$. Then as $t\nearrow T^{*}=-p/m$,
  1. (i)
    the blow-up point is $x^{*}(t)=\varphi(t,\xi^{*})$ with $x^{*}(t)\to x^{*}_{\infty}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\varphi_{T^{*}}(\xi^{*})$;
  2. (ii)
    the strain satisfies
    $\sup_{x}|u_{x}(t,x)|=|V(t,\xi^{*})|=\frac{p}{T^{*}-t};$
  3. (iii)
    near the blow-up point, the profile is asymptotically
    $u_{x}(t,x)\sim-\frac{p}{T^{*}-t}\cdot\frac{1}{1+c(T^{*}-t)^{-2(p+1)}(x-x^{*}(t))^{2}}$
    where $c=\frac{v_{0}^{\prime\prime}(\xi^{*})p^{2p+1}}{2|m|^{2p+2}}$.
Proof.
Part (i) follows from continuity of $\varphi(t,\cdot)$ up to $t=T^{*}$. Part (ii) is immediate from the Riccati solution at the minimum.
For (iii), set $\alpha:=v_{0}^{\prime\prime}(\xi^{*})>0$. The non-degeneracy assumption gives the Taylor expansion
$V_{0}(\xi)=m+\frac{\alpha}{2}(\xi-\xi^{*})^{2}+O\bigl((\xi-\xi^{*})^{3}\bigr),\qquad\xi\to\xi^{*}.$
Substituting this into $V(t,\xi)=V_{0}(\xi)/(1+\frac{t}{p}V_{0}(\xi))$ and using $1+\frac{t}{p}m=1-\frac{t}{T^{*}}=\frac{T^{*}-t}{T^{*}}$ yields
$V(t,\xi)=\frac{m+\frac{\alpha}{2}(\xi-\xi^{*})^{2}+O((\xi-\xi^{*})^{3})}{\frac{T^{*}-t}{T^{*}}+\frac{t\alpha}{2p}(\xi-\xi^{*})^{2}+O((\xi-\xi^{*})^{3})}.$
The nontrivial blow-up profile comes from the regime where the two leading contributions in the denominator are comparable, i.e. $(\xi-\xi^{*})^{2}\sim T^{*}-t$, so we restrict to $|\xi-\xi^{*}|=O(\sqrt{T^{*}-t})$; then the cubic remainder is $o(T^{*}-t)$ and does not affect the leading order. Factoring $(T^{*}-t)/T^{*}$ out of the denominator and using $T^{*}m=-p$ gives the Lagrangian asymptotic
$V(t,\xi)\sim-\frac{p}{T^{*}-t}\,\frac{1}{1+\kappa\,\frac{(\xi-\xi^{*})^{2}}{T^{*}-t}},\qquad\kappa:=\frac{t\alpha T^{*}}{2p}\xrightarrow[t\nearrow T^{*}]{}\frac{\alpha p}{2|m|^{2}}.$ (14)
To pass to Eulerian variables, set $x=\varphi(t,\xi)$ and $x^{*}(t)=\varphi(t,\xi^{*})$. A first-order Taylor expansion in $\xi$ gives, as $\xi\to\xi^{*}$,
$x-x^{*}(t)=\partial_{\xi}\varphi(t,\xi^{*})\,(\xi-\xi^{*})+O\bigl((\xi-\xi^{*})^{2}\bigr).$
By Proposition 3.2 and $V_{0}(\xi^{*})=m$,
$\partial_{\xi}\varphi(t,\xi^{*})=\varphi_{0}^{\prime}(\xi^{*})\left(1+\frac{t}{p}m\right)^{p}=\varphi_{0}^{\prime}(\xi^{*})\left(\frac{T^{*}-t}{T^{*}}\right)^{p},$
and hence, for $\xi$ sufficiently close to $\xi^{*}$ and $t$ close to $T^{*}$,
$\xi-\xi^{*}=\frac{x-x^{*}(t)}{\partial_{\xi}\varphi(t,\xi^{*})}+o\!\left(\frac{|x-x^{*}(t)|}{\partial_{\xi}\varphi(t,\xi^{*})}\right)=\frac{x-x^{*}(t)}{\varphi_{0}^{\prime}(\xi^{*})}\left(\frac{T^{*}}{T^{*}-t}\right)^{p}+o(\cdots).$
Since the Lagrangian balance is $(\xi-\xi^{*})^{2}/(T^{*}-t)\sim 1$, we have
$|x-x^{*}(t)|\sim\partial_{\xi}\varphi(t,\xi^{*})\,|\xi-\xi^{*}|\sim(T^{*}-t)^{p}\,(T^{*}-t)^{1/2}=(T^{*}-t)^{p+\frac{1}{2}},$
so equivalently
$\frac{(\xi-\xi^{*})^{2}}{T^{*}-t}\sim\frac{T^{*\,2p}}{\varphi_{0}^{\prime}(\xi^{*})^{2}}\,\frac{(x-x^{*}(t))^{2}}{(T^{*}-t)^{2p+1}}\qquad(t\nearrow T^{*},\ x\to x^{*}(t)).$
Substituting this into the Lagrangian asymptotic for $V$ (14) and using
$u_{x}(t,x)=v(t,x)=V\bigl(t,\varphi^{-1}(t,x)\bigr)$
gives the Eulerian profile
$u_{x}(t,x)\sim-\frac{p}{T^{*}-t}\cdot\frac{1}{1+c\,(T^{*}-t)^{-(2p+1)}(x-x^{*}(t))^{2}},\qquad c:=\kappa\,\frac{T^{*\,2p}}{\varphi_{0}^{\prime}(\xi^{*})^{2}}.$
In particular, if $\varphi_{0}=\mathrm{Id}$ so that $\varphi_{0}^{\prime}(\xi^{*})=1$, then
$c=\frac{\alpha p}{2|m|^{2}}\,T^{*\,2p}=\frac{v_{0}^{\prime\prime}(\xi^{*})\,p^{2p+1}}{2|m|^{2p+2}}.$
3.4. Weak solutions and entropy conditions
To continue solutions of (11) past the first gradient catastrophe time $T^{*}$, we work with the standard dissipative (entropy-admissible) class [9], in which the $L^{p}$ strain energy is allowed to drop at singular times.
Definition 3.6 (Dissipative weak solution).
Fix $p\in(1,\infty)$ and initial data $v_{0}\in L^{p}(\mathbb{R})$ such that the primitive
$u_{0}(x)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\int_{-\infty}^{x}v_{0}(y)\,\mathrm{d}y$
is well-defined (equivalently $v_{0}\in L^{1}((-\infty,0])$), and fix this normalization of primitives. A measurable function $v\colon[0,\infty)\times\mathbb{R}\to\mathbb{R}$ is a dissipative weak solution of (11) with initial datum $v_{0}$ if
  • for a.e. $t\geq 0$, $v(t,\cdot)\in L^{p}(\mathbb{R})$ and the map $t\mapsto\|v(t,\cdot)\|_{L^{p}}^{p}$ is non-increasing;
  • letting $u(t,\cdot)$ denote the primitive of $v(t,\cdot)$ with the same normalization as $u_{0}$ (so $u_{x}=v$ a.e.), one has for every $\phi\in C_{c}^{\infty}([0,\infty)\times\mathbb{R})$,
    $\int_{0}^{\infty}\!\!\int_{\mathbb{R}}\Bigl(v\,\partial_{t}\phi+u\,v\,\partial_{x}\phi+\frac{p-1}{p}\,v^{2}\,\phi\Bigr)\,\mathrm{d}x\,\mathrm{d}t+\int_{\mathbb{R}}v_{0}(x)\,\phi(0,x)\,\mathrm{d}x=0.$ (15)
Theorem 3.7 (Global existence of dissipative solutions).
Let $v_{0}\in L^{p}(\mathbb{R})$ with $u_{0}(x)=\int_{-\infty}^{x}v_{0}$ well-defined. Then there exists a unique dissipative weak solution $v$ of (11) with $v(0,\cdot)=v_{0}$. The solution satisfies.
  1. (i)
    (Energy dissipation) For all $t\geq 0$,
    $\|v(t,\cdot)\|_{L^{p}}^{p}\leq\|v_{0}\|_{L^{p}}^{p},$
    and equality holds on every time interval on which $v$ is classical.
  2. (ii)
    If $v_{0}\in W^{\infty,1}(\mathbb{R})$ with $\inf v_{0}\geq 0$, then $v$ is smooth for all $t\geq 0$.
  3. (iii)
    If $v_{0}\in W^{\infty,1}(\mathbb{R})$ with $\inf v_{0}<0$, then $v$ is smooth for $t<T^{*}$ and its dissipative continuation loses regularity at $t=T^{*}$ by creating a shock discontinuity in $v$.
Proof.
Write $V_{0}=v_{0}$ and set
$a(t,\xi):=1+\frac{t}{p}V_{0}(\xi),\qquad a_{+}(t,\xi):=\max\{a(t,\xi),0\}.$
Define a (possibly non-strictly) increasing Lagrangian map $\varphi(t,\cdot)$ by prescribing its Jacobian
$\partial_{\xi}\varphi(t,\xi)=a_{+}(t,\xi)^{p}$ (16)
together with the normalization $\varphi(t,\xi)-\xi\to 0$ as $\xi\to$-$\infty$, namely
$\varphi(t,\xi)=\xi+\int_{-\infty}^{\xi}\bigl(a_{+}(t,\eta)^{p}-1\bigr)\,d\eta.$ (17)
Define the Lagrangian strain by
$V(t,\xi):=\begin{cases}\dfrac{V_{0}(\xi)}{a(t,\xi)},&a(t,\xi)>0,\\[8.00003pt] 0,&a(t,\xi)\leq 0,\end{cases}$ (18)
and define $u$ and $v$ in Eulerian variables via
$u\bigl(t,\varphi(t,\xi)\bigr):=\int_{-\infty}^{\xi}V(t,\eta)\,\partial_{\eta}\varphi(t,\eta)\,d\eta,\qquad v\bigl(t,\varphi(t,\xi)\bigr):=V(t,\xi).$ (19)
On $\{a(t,\xi)>0\}$ these are exactly the classical Lagrangian formulas from the Riccati proposition. On $\{a(t,\xi)\leq 0\}$ one has $\partial_{\xi}\varphi=0$ by (16) and $V=0$ by (18), so $u$ and $v$ are constant on any label interval where $\varphi(t,\cdot)$ is flat. Hence (19) defines single-valued functions of the Eulerian variable $x$.
By the change of variables $x=\varphi(t,\xi)$ and (16)-(18),
$\|v(t,\cdot)\|_{L^{p}}^{p}$ $=\int_{\mathbb{R}}|V(t,\xi)|^{p}\,\partial_{\xi}\varphi(t,\xi)\,d\xi=\int_{\{a>0\}}\left|\frac{V_{0}(\xi)}{a(t,\xi)}\right|^{p}a(t,\xi)^{p}\,d\xi$
$=\int_{\{a>0\}}|V_{0}(\xi)|^{p}\,d\xi\leq\|V_{0}\|_{L^{p}}^{p},$ (20)
and since $\{a(t,\xi)>0\}=\{\xi:V_{0}(\xi)>-p/t\}$ decreases as $t$ increases, the map $t\mapsto\|v(t,\cdot)\|_{L^{p}}^{p}$ is non-increasing.
To verify the weak formulation, fix $\phi\in C_{c}^{\infty}([0,\infty)\times\mathbb{R})$ and set $\Phi(t,\xi)=\phi(t,\varphi(t,\xi))$. The chain rule gives
$\partial_{t}\Phi(t,\xi)=(\partial_{t}\phi)(t,\varphi)+(\partial_{x}\phi)(t,\varphi)\,\partial_{t}\varphi(t,\xi).$
Differentiating (17) in $t$ and using $\partial_{t}(a_{+}^{p})=a_{+}^{p-1}V_{0}\,\mathbf{1}_{\{a>0\}}$, we obtain
$\partial_{t}\varphi(t,\xi)=\int_{-\infty}^{\xi}V_{0}(\eta)\,a_{+}^{p-1}\mathbf{1}_{\{a>0\}}\,d\eta=\int_{-\infty}^{\xi}V(t,\eta)\,\partial_{\eta}\varphi\,d\eta=u(t,\varphi(t,\xi)),$
where the second equality uses (16)-(18) and the last is (19). Define $M(t,\xi)=V(t,\xi)\,\partial_{\xi}\varphi(t,\xi)$. On $\{a>0\}$ one has $M=V_{0}\,a^{p-1}$; on $\{a\leq 0\}$ one has $M=0$. Differentiating,
$\partial_{t}M(t,\xi)=V_{0}(p-1)a^{p-2}\cdot\frac{V_{0}}{p}\cdot\mathbf{1}_{\{a>0\}}=\frac{p-1}{p}V^{2}\,\partial_{\xi}\varphi.$
Integrating by parts in $t$ and using that $\Phi$ is compactly supported,
$\int_{0}^{\infty}\!\!\int_{\mathbb{R}}M\,\partial_{t}\Phi\,d\xi\,dt=-\int_{0}^{\infty}\!\!\int_{\mathbb{R}}\partial_{t}M\,\Phi\,d\xi\,dt-\int_{\mathbb{R}}M(0,\xi)\,\Phi(0,\xi)\,d\xi.$
At $t=0$ we have $\varphi(0,\xi)=\xi$, $\partial_{\xi}\varphi=1$, $V=V_{0}$, so $M(0,\xi)=V_{0}(\xi)$ and $\Phi(0,\xi)=\phi(0,\xi)$. Changing variables back to Eulerian coordinates yields the weak formulation. This proves existence.
For uniqueness, let $v$ be any dissipative weak solution and $(\varphi,V)$ its monotone Lagrangian representation with $\varphi(0,\xi)=\xi$ and $V(t,\xi)=v(t,\varphi(t,\xi))$. The characteristic relation
$\partial_{t}\varphi(t,\xi)=u(t,\varphi(t,\xi))$ (21)
holds for a.e. $(t,\xi)$, and if $\varphi(t,\cdot)$ is constant on an interval at some time, it remains so for all later times (sticking property). Write $J=\partial_{\xi}\varphi$. On $\{J>0\}$, the weak equation reduces to the Riccati law
$\partial_{t}V+\frac{1}{p}V^{2}=0,$ (22)
and differentiating (21) in $\xi$ gives
$\partial_{t}J=V\,J.$ (23)
Set $a=J^{1/p}$ on $\{J>0\}$. Then (23) gives $\partial_{t}a=\frac{1}{p}Va$. Combining with (22),
$\partial_{t}(aV)=a\,\partial_{t}V+(\partial_{t}a)V=a\left(\partial_{t}V+\frac{1}{p}V^{2}\right)=0.$
Since $J(0,\xi)=1$ and $V(0,\xi)=v_{0}(\xi)$, we have $a(t,\xi)V(t,\xi)=v_{0}(\xi)$ as long as $a>0$. Substituting $V=v_{0}/a$ into $\partial_{t}a=\frac{1}{p}Va$ yields $\partial_{t}a=\frac{1}{p}v_{0}$, hence $a(t,\xi)=1+\frac{t}{p}v_{0}(\xi)$ up to the first time it reaches zero. If $a(t_{0},\xi)=0$, then $J(t_{0},\xi)=0$ and the sticking property forces $J(t,\xi)=0$ for all $t\geq t_{0}$. The dissipativity requirement then forces $V(t,\xi)=0$ whenever $J(t,\xi)=0$. Therefore
$a(t,\xi)=\max\left\{1+\frac{t}{p}v_{0}(\xi),0\right\},\quad J(t,\xi)=a(t,\xi)^{p},\quad V(t,\xi)=\begin{cases}v_{0}(\xi)/a(t,\xi),&a>0,\\ 0,&a=0,\end{cases}$
which coincide with (16)-(18). Since $\varphi$ is determined by integrating $J=\partial_{\xi}\varphi$, the Eulerian field $v$ is uniquely determined by $v_{0}$.
Assume $v_{0}\in W^{\infty,1}(\mathbb{R})$. If $\inf v_{0}\geq 0$, then $a(t,\xi)\geq 1$ for all $t\geq 0$ and all $\xi$, so the truncation never activates, $\varphi(t,\cdot)$ remains a diffeomorphism, and $v$ is smooth for all time, proving (ii). If $\inf v_{0}<0$, let $T^{*}=-p/\inf v_{0}$. For $t<T^{*}$ one has $a>0$ everywhere and $v$ is smooth. At $t=T^{*}$ the set $\{a(T^{*},\xi)=0\}$ is nonempty and $J(T^{*},\xi)=0$ there, so $\varphi(T^{*},\cdot)$ ceases to be strictly increasing. For $t>T^{*}$, on any interval $I$ where $a(t,\xi)=0$ one has $J=0$ and $\varphi(t,\cdot)$ is constant with image a single point $x_{s}(t)$. Since $V=0$ on $I$ but $V=v_{0}/a\neq 0$ approaching $I$ from outside, $v(t,\cdot)$ has a jump discontinuity at $x_{s}(t)$. Thus a shock forms at $T^{*}$ and persists, and the $L^{p}$ energy drops strictly because $\{a>0\}$ shrinks, proving (iii). ∎
3.5. Probabilistic interpretation of the strain blow-up
The explicit Lagrangian formulas of Proposition 3.2 admit a canonical probabilistic reading in terms of time-inhomogeneous linear birth-death processes. This interpretation is auxiliary but conceptually sharp: it assigns a stochastic meaning to the Jacobian factor and identifies the dissipative continuation of Theorem 3.7 as an absorbed evolution.
Definition 3.8 (Branching factor and local Malthusian parameter).
Let $\varphi(t,\cdot)$ be a (classical or monotone) Lagrangian flow on a time interval $[0,T)$, with Jacobian $J(t,\xi)=\partial_{\xi}\varphi(t,\xi)$ and Lagrangian strain $V(t,\xi)=u_{x}(t,\varphi(t,\xi))$ whenever defined. Define the branching factor
$g(t,\xi):=\frac{J(t,\xi)}{J(0,\xi)}=\frac{\partial_{\xi}\varphi(t,\xi)}{\varphi_{0}^{\prime}(\xi)}$ (24)
whenever the ratio is meaningful, and define the local Malthusian parameter
$b(t,\xi):=\partial_{t}\log g(t,\xi)$ (25)
on the set $\{g(t,\xi)>0\}$.
Lemma 3.9 (Strain as log-growth).
On any time interval where the geodesic is smooth and $J(t,\xi)>0$, one has
$b(t,\xi)=V(t,\xi)\qquad\text{for all }\xi\in\mathbb{R}.$ (26)
In particular, in the classical regime
$g(t,\xi)=\left(1+\frac{t}{p}V_{0}(\xi)\right)^{p},\qquad\int_{0}^{t}V(s,\xi)\,ds=p\log\!\left(1+\frac{t}{p}V_{0}(\xi)\right).$ (27)
For the dissipative weak solution of Theorem 3.7 (constructed with $\varphi(0,\xi)=\xi$), one has $g(t,\xi)=a_{+}(t,\xi)^{p}$ with $a_{+}(t,\xi)=\max\{1+\frac{t}{p}v_{0}(\xi),0\}$ and $b(t,\xi)=V(t,\xi)$ on $\{a(t,\xi)>0\}$.
Proof.
In the smooth regime, $J>0$ and Proposition 3.2(iii) gives $\partial_{t}J=VJ$. Dividing by $J$ yields $\partial_{t}\log J=V$. Since $g=J/J(0)$, we have $\partial_{t}\log g=\partial_{t}\log J$, hence (26). Integrating in time gives $g(t,\xi)=\exp(\int_{0}^{t}V)$, and substituting the explicit Riccati solution from Proposition 3.2(ii) yields (27).
In the dissipative regime, $J=a_{+}^{p}$ by (16) and $V=v_{0}/a$ on $\{a>0\}$ by (18). Therefore
$\partial_{t}\log J=p\,\partial_{t}\log a_{+}=\mathbf{1}_{\{a>0\}}\frac{v_{0}}{a}=V,$
which is the stated identity on $\{a>0\}$. ∎
Proposition 3.10 (Birth-death representation of the Jacobian).
Fix $\xi\in\mathbb{R}$ and assume $b(\cdot,\xi)$ is locally bounded on $[0,T)$. Define time-dependent per-capita birth and death rates by
$\lambda(t,\xi):=b(t,\xi)_{+},\qquad\mu(t,\xi):=(-b(t,\xi))_{+},$ (28)
where $s_{+}=\max\{s,0\}$. Then there exists a time-inhomogeneous non-explosive linear birth-death process $(Z_{t}^{\xi})_{t\in[0,T)}$ on $\mathbb{N}_{0}$ with $Z_{0}^{\xi}=1$ and rates $\lambda(t,\xi),\mu(t,\xi)$ such that
$\frac{d}{dt}\mathbb{E}[Z_{t}^{\xi}]=b(t,\xi)\,\mathbb{E}[Z_{t}^{\xi}],\qquad\mathbb{E}[Z_{0}^{\xi}]=1,$ (29)
and therefore
$\mathbb{E}[Z_{t}^{\xi}]=\exp\!\left(\int_{0}^{t}b(s,\xi)\,ds\right)=g(t,\xi).$ (30)
Proof.
For fixed $\xi$, define a time-inhomogeneous Markov chain on $\mathbb{N}_{0}$ with generator
$(\mathcal{L}_{t}f)(n)=n\lambda(t,\xi)\bigl(f(n+1)-f(n)\bigr)+n\mu(t,\xi)\bigl(f(n-1)-f(n)\bigr).$
Fix any $T_{0}<T$ and write $C_{T_{0}}:=\sup_{t\in[0,T_{0}]}(\lambda(t,\xi)+\mu(t,\xi))<\infty$. Then the total jump rate out of state $n$ on $0,T_{0}]$ is at most $nC_{T_{0}}$. Comparing with the pure-birth chain with rates $nC_{T_{0}}$ and using $\sum_{n\geq 1}(nC_{T_{0}})^{-1}=+\infty$ gives the standard non-explosion criterion: the process cannot make infinitely many jumps on $0,T_{0}]$. Since $T_{0}<T$ was arbitrary, this yields a non-explosive process on $[0,T)$.
Taking $f(n)=n$ gives
$(\mathcal{L}_{t}f)(n)=n(\lambda(t,\xi)-\mu(t,\xi))=n\,b(t,\xi).$
Dynkin’s formula yields
$\frac{d}{dt}\mathbb{E}[Z_{t}^{\xi}]=\mathbb{E}[(\mathcal{L}_{t}f)(Z_{t}^{\xi})]=b(t,\xi)\,\mathbb{E}[Z_{t}^{\xi}],$
which is (29). Solving the scalar ODE with $\mathbb{E}[Z_{0}^{\xi}]=1$ gives
$\mathbb{E}[Z_{t}^{\xi}]=\exp\!\left(\int_{0}^{t}b(s,\xi)\,ds\right),$
and by Definition 3.8 this equals $g(t,\xi)$, proving (30). ∎
Theorem 3.11 (Extinction and supercritical growth along labels).
Assume $v_{0}\in W^{\infty,1}(\mathbb{R})$ and let $V(t,\xi)$ be the classical Lagrangian strain on $t<T^{*}$ given by Proposition 3.2. For each $\xi$ set
$\nu(\xi):=V_{0}(\xi)$
and define the local critical time
$\tau(\xi):=\begin{cases}-p/\nu(\xi),&\nu(\xi)<0,\\ +\infty,&\nu(\xi)\geq 0.\end{cases}$
Then the birth-death process of Proposition 3.10 reduces to:
  1. (i)
    If $\nu(\xi)<0$, then $\lambda(\cdot,\xi)\equiv 0$ and $Z_{t}^{\xi}$ is a pure-death process with hazard
    $\mu(t,\xi)=-V(t,\xi)=\frac{-\nu(\xi)}{1+\frac{t}{p}\nu(\xi)}.$
    Its survival probability equals the branching factor:
    $\mathbb{P}(Z_{t}^{\xi}=1)=g(t,\xi)=\left(1+\frac{t}{p}\nu(\xi)\right)^{p},$
    and $\mathbb{P}(Z_{\tau(\xi)}^{\xi}=0)=1$. In particular, extinction occurs almost surely by time $\tau(\xi)$.
  2. (ii)
    If $\nu(\xi)=0$, then $V(t,\xi)\equiv 0$, $g(t,\xi)\equiv 1$, and $Z_{t}^{\xi}\equiv 1$.
  3. (iii)
    If $\nu(\xi)>0$, then $\mu(\cdot,\xi)\equiv 0$ and $Z_{t}^{\xi}$ is a pure-birth process with time-dependent rate $\lambda(t,\xi)=V(t,\xi)$. Its mean grows polynomially: $\mathbb{E}[Z_{t}^{\xi}]=g(t,\xi)\sim(t\,\nu(\xi)/p)^{p}$ as $t\to\infty$.
Moreover, the global existence time $T^{*}$ from Theorem 3.3 coincides with the earliest extinction time across labels:
$T^{*}=\inf_{\xi\in\mathbb{R}}\tau(\xi).$
Proof.
By Lemma 3.9, $b(t,\xi)=V(t,\xi)$ in the classical regime. Since $t\mapsto V(t,\xi)$ preserves the sign of $\nu(\xi)$ for as long as it exists, the decomposition (28) yields a pure-death, critical, or pure-birth process.
If $\nu(\xi)<0$, then $V(t,\xi)<0$ for $t\in[0,\tau(\xi))$ and $\lambda\equiv 0$. A pure-death process started from one dies at a (random) death time $T_{\mathrm{d}}$ whose survival function is
$\mathbb{P}(T_{\mathrm{d}}>t)=\exp\!\left(-\int_{0}^{t}\mu(s,\xi)\,ds\right).$
Here $\mu=-V$, and Lemma 3.9 gives $\exp(-\int_{0}^{t}\mu)=\exp(\int_{0}^{t}V)=g(t,\xi)$, proving the identity. Since $\int_{0}^{\tau(\xi)}\mu(s,\xi)\,ds=+\infty$ (equivalently $g(t,\xi)\downarrow 0$ as $t\uparrow\tau(\xi)$), we have $\mathbb{P}(T_{\mathrm{d}}\geq\tau(\xi))=0$, hence $\mathbb{P}(Z_{\tau(\xi)}^{\xi}=0)=1$.
If $\nu(\xi)=0$, then $V\equiv 0$ and rates vanish, so the chain is constant. If $\nu(\xi)>0$, then $\mu\equiv 0$ and (30) gives the growth of the mean; the asymptotic follows from expanding $(1+\frac{t}{p}\nu(\xi))^{p}$.
Finally, $T^{*}=\inf_{\xi}\tau(\xi)$ is exactly the relationship between the explicit Riccati law and the definition of $T^{*}$ in Theorem 3.3. ∎
Remark 3.12 (Polynomial growth versus exponential growth).
In the supercritical regime $\nu(\xi)>0$ one has $V(t,\xi)\sim p/t$ as $t\to\infty$, so $\int_{0}^{t}V(s,\xi)\,ds\sim p\log t$ and therefore $\mathbb{E}[Z_{t}^{\xi}]=\exp(\int_{0}^{t}V)=g(t,\xi)\sim(t\,\nu(\xi)/p)^{p}$. Thus the model exhibits polynomial mean growth, a genuine consequence of the Riccati decay and in sharp contrast with time-homogeneous branching where the mean typically grows exponentially.
Corollary 3.13 (Deterministic many-to-one identity [^1] In classical branching terminology, a many-to-one lemma relates expectations over all particles to a single-particle expectation under a size-biased (spine) change of measure. Here there is no random genealogical tree. The reweighting $M_{t}(d\xi)=g(t,\xi)\pi_{0}(d\xi)$ plays the role of the spine Radon-Nikodym factor, and (31) is the corresponding deterministic covariance identity.).
Let $\pi_{0}$ be any probability measure on the Lagrangian labels $\mathbb{R}$. Define the (possibly subprobability) measure $M_{t}$ by
$M_{t}(d\xi):=g(t,\xi)\,\pi_{0}(d\xi)$
in the classical regime, and $M_{t}(d\xi):=a_{+}(t,\xi)^{p}\,\pi_{0}(d\xi)$ in the dissipative regime of Theorem 3.7. Then for every bounded measurable $f$, the map $t\mapsto\int f(\xi)\,M_{t}(d\xi)$ is absolutely continuous and satisfies
$\frac{d}{dt}\int_{\mathbb{R}}f(\xi)\,M_{t}(d\xi)=\int_{\mathbb{R}}f(\xi)\,V(t,\xi)\,M_{t}(d\xi)$ (31)
for a.e. $t$. If $M_{t}(\mathbb{R})>0$ and $\widehat{\pi}_{t}:=M_{t}/M_{t}(\mathbb{R})$ is the normalized law, then
$\frac{d}{dt}\mathbb{E}_{\widehat{\pi}_{t}}[f]=\mathrm{Cov}_{\widehat{\pi}_{t}}(f,V).$
Proof.
In the classical regime, $M_{t}$ has density $g(t,\xi)$ with respect to $\pi_{0}$, and Definition 3.8 gives $\partial_{t}g=Vg$. Thus
$\frac{d}{dt}\int f\,dM_{t}=\int f\,\partial_{t}g\,d\pi_{0}=\int f\,Vg\,d\pi_{0}=\int f\,V\,dM_{t},$
which is (31). In the dissipative regime, the same computation holds on $\{a>0\}$ since $\partial_{t}(a_{+}^{p})=Va_{+}^{p}$ there, and both sides vanish on $\{a\leq 0\}$. The covariance identity follows by differentiating $M_{t}(f)/M_{t}(1)$. ∎
Remark 3.14 (Price equation and Fisher-type covariance identities).
The identity (31) is the deterministic form of the Price equation: the derivative of an average under a time-dependent reweighting equals the covariance with the “fitness” $V$. In particular, taking $f=V$ gives $\frac{d}{dt}\mathbb{E}_{\widehat{\pi}_{t}}[V]=\mathrm{Var}_{\widehat{\pi}_{t}}(V)$, a Fisher-type monotonicity statement in this strain/weighting setting.
3.6. Large deviations for the catastrophe time
While the identity $T^{*}=-p/\inf v_{0}$ is exact, it hides a delicate instability, that is, the infimum is not a smooth functional. Under a mild nondegeneracy hypothesis (unique quadratic minimum), the blow-up time becomes differentiable and admits a sharp small-noise large deviation theory.
Definition 3.15 (Blow-up time as a functional).
For a bounded function $f:\mathbb{R}\to\mathbb{R}$ with $\inf_{\mathbb{R}}f<0$ define
$\mathcal{T}(f):=-\frac{p}{\inf_{\mathbb{R}}f}.$
If $\inf_{\mathbb{R}}f\geq 0$, set $\mathcal{T}(f)=+\infty$. For smooth initial strain $f=v_{0}$, this coincides with the catastrophe time $T^{*}$ of Theorem 3.3.
Lemma 3.16 (Stability of the minimizer under perturbations).
Let $v_{0}\in C^{3}(\mathbb{R})$ and assume it has a unique nondegenerate global minimum at $x^{*}$:
$v_{0}(x^{*})=m<0,\qquad v_{0}^{\prime}(x^{*})=0,\qquad v_{0}^{\prime\prime}(x^{*})=\alpha>0.$
Then there exists $\delta>0$ and a $C^{1}$ map $h\mapsto x_{h}$ from the ball $\{h\in C^{2}(\mathbb{R}):\|h\|_{C^{2}}\leq\delta\}$ to $\mathbb{R}$ such that $x_{h}$ is the unique global minimizer of $v_{0}+h$ and
$x_{h}$ $=x^{*}-\frac{h^{\prime}(x^{*})}{\alpha}+O\bigl(\|h\|_{C^{2}}^{2}\bigr),$ (32)
$\inf_{\mathbb{R}}(v_{0}+h)$ $=m+h(x^{*})-\frac{1}{2\alpha}h^{\prime}(x^{*})^{2}+O\bigl(\|h\|_{C^{2}}^{3}\bigr).$ (33)
Proof.
Define $F(x,h):=v_{0}^{\prime}(x)+h^{\prime}(x)$. One has $F(x^{*},0)=0$ and $\partial_{x}F(x^{*},0)=v_{0}^{\prime\prime}(x^{*})=\alpha>0$. By the implicit function theorem there exist $\delta>0$ and a $C^{1}$ map $h\mapsto x_{h}$ (for $\|h\|_{C^{2}}\leq\delta$) such that $F(x_{h},h)=0$. Shrinking $\delta$ if needed, strict convexity of $v_{0}$ near $x^{*}$ and smallness of $\|h^{\prime\prime}\|_{\infty}$ imply that $x_{h}$ is the unique global minimizer.
Expanding $0=F(x_{h},h)$ around $(x^{*},0)$ gives
$0=\alpha(x_{h}-x^{*})+h^{\prime}(x^{*})+O\bigl(|x_{h}-x^{*}|^{2}+\|h\|_{C^{2}}|x_{h}-x^{*}|\bigr),$
hence $x_{h}-x^{*}=-h^{\prime}(x^{*})/\alpha+O(\|h\|_{C^{2}}^{2})$, proving (32).
Writing $m_{h}:=\inf(v_{0}+h)=(v_{0}+h)(x_{h})$ and Taylor expanding at $x^{*}$,
$v_{0}(x_{h})=m+\frac{\alpha}{2}(x_{h}-x^{*})^{2}+O(|x_{h}-x^{*}|^{3}),\quad h(x_{h})=h(x^{*})+h^{\prime}(x^{*})(x_{h}-x^{*})+O(\|h\|_{C^{2}}|x_{h}-x^{*}|^{2}).$
Substituting (32) yields (33). ∎
Proposition 3.17 (First variation of the blow-up time).
Under the assumptions of Lemma 3.16, the map $h\mapsto\mathcal{T}(v_{0}+h)$ is differentiable at $h=0$ (as a map $C^{2}(\mathbb{R})\to\mathbb{R}$), with
$\mathcal{T}(v_{0}+h)=T^{*}+\frac{p}{m^{2}}\,h(x^{*})+O\bigl(\|h\|_{C^{1}}^{2}\bigr),\qquad T^{*}=-\frac{p}{m}.$ (34)
Proof.
Lemma 3.16 gives $\inf(v_{0}+h)=m+h(x^{*})+O(\|h\|_{C^{1}}^{2})$. Writing $m_{h}=m+\delta m$ and expanding $p/m_{h}$ around $m$ gives
$\frac{p}{m+\delta m}=-\frac{p}{m}+\frac{p}{m^{2}}\delta m+O(\delta m^{2}).$
Substituting $\delta m=h(x^{*})+O(\|h\|_{C^{1}}^{2})$ yields (34). ∎
Theorem 3.18 (Small-noise LDP for the blow-up time).
Assume $v_{0}\in C^{3}(\mathbb{R})$ satisfies Lemma 3.16. Let $\eta$ be a centered Gaussian random element in a separable Banach space $B$ continuously embedded in $C^{1}(\mathbb{R})$, with Cameron-Martin space $H\subset B$. For $\varepsilon>0$ set $v_{0}^{\varepsilon}:=v_{0}+\varepsilon\eta$ and $T^{\varepsilon}:=\mathcal{T}(v_{0}^{\varepsilon})$. Then $(T^{\varepsilon})_{\varepsilon\downarrow 0}$ satisfies a large deviation principle on $\mathbb{R}$ with speed $\varepsilon^{-2}$ and good rate function
$J(T)=\inf\left\{\frac{1}{2}\|h\|_{H}^{2}\;:\;h\in H,\ \mathcal{T}(v_{0}+h)=T\right\}.$ (35)
If the evaluation functional $\ell(h):=h(x^{*})$ is continuous on $H$, then $J$ has the quadratic expansion
$J(T)=\frac{m^{4}}{2p^{2}\,\kappa}\,(T-T^{*})^{2}+o\bigl((T-T^{*})^{2}\bigr)\qquad(T\to T^{*}),$ (36)
where $\kappa:=\|\ell\|_{H^{*}}^{2}$.
Proof.
Since $B\hookrightarrow C^{1}(\mathbb{R})$, the map $f\mapsto\inf f$ is continuous from $B$ to $\mathbb{R}$. Because $m=\inf v_{0}<0$ stays separated from 0 under sufficiently small perturbations in $B$, the map $h\mapsto\mathcal{T}(v_{0}+h)$ is continuous at $h=0$ as a map $B\to\mathbb{R}$. The family $\varepsilon\eta$ satisfies an LDP on $B$ with speed $\varepsilon^{-2}$ and rate function $I(h)=\frac{1}{2}\|h\|_{H}^{2}$ on $H$ (and $I=+\infty$ otherwise). The contraction principle yields the LDP for $T^{\varepsilon}$ with rate function (35).
For the local expansion, let $\delta T:=T-T^{*}$. Proposition 3.17 gives
$\mathcal{T}(v_{0}+h)=T^{*}+\frac{p}{m^{2}}\,\ell(h)+O(\|h\|_{H}^{2}),$
using the continuous embedding $H\hookrightarrow C^{1}$. Thus, for $\delta T$ small, the constraint $\mathcal{T}(v_{0}+h)=T$ enforces $\ell(h)=\frac{m^{2}}{p}\delta T$ at leading order. Minimizing $\frac{1}{2}\|h\|_{H}^{2}$ under the affine constraint $\ell(h)=a$ yields the unique minimizer $h=a\,\psi/\kappa$, where $\psi\in H$ is the Riesz representer of $\ell$ and $\kappa=\|\psi\|_{H}^{2}=\|\ell\|_{H^{*}}^{2}$, and the minimal cost equals $a^{2}/(2\kappa)$. Taking $a=(m^{2}/p)\delta T$ gives (36). ∎
3.7. Spine decomposition and an explicit action functional
The LDP rate function from Theorem 3.18 is realized by a unique “spine” direction in the Cameron-Martin space: to leading order, only one rank-one perturbation direction moves the catastrophe time at minimal energetic cost.
Definition 3.19 (Spine direction in the Cameron-Martin space).
Assume $\ell(h)=h(x^{*})$ is continuous on $H$. Let $\psi_{x^{*}}\in H$ denote the Riesz representer of $\ell$, i.e.
$\ell(h)=\langle h,\psi_{x^{*}}\rangle_{H},\qquad h\in H,$
and set $\kappa:=\|\psi_{x^{*}}\|_{H}^{2}=\|\ell\|_{H^{*}}^{2}$.
Proposition 3.20 (Spine decomposition).
Every $h\in H$ admits the orthogonal decomposition
$h=\frac{\ell(h)}{\kappa}\,\psi_{x^{*}}+h^{\perp},\qquad\ell(h^{\perp})=0,\qquad\langle h^{\perp},\psi_{x^{*}}\rangle_{H}=0.$
Moreover, the first-order shift of the catastrophe time depends only on the spine component:
$\mathcal{T}(v_{0}+h)=T^{*}+\frac{p}{m^{2}}\,\ell(h)+O(\|h\|_{H}^{2}).$
Consequently, among perturbations producing a prescribed first-order shift $\delta T$, the unique minimal-cost perturbation is proportional to $\psi_{x^{*}}$.
Proof.
The decomposition is the orthogonal projection of $h$ onto $\mathrm{span}\{\psi_{x^{*}}\}$ in the Hilbert space $H$. Since $\ell(h)=\langle h,\psi_{x^{*}}\rangle_{H}$ and $\ell(h^{\perp})=0$, Proposition 3.17 yields the stated dependence. The minimality follows from the Hilbert space fact that under the affine constraint $\ell(h)=a$, the unique minimizer of $\|h\|_{H}$ is $h=a\,\psi_{x^{*}}/\kappa$. ∎
Theorem 3.21 (Explicit spine for $H=H^{1}(\mathbb{R})$).
Assume the perturbations have Cameron-Martin space $H=H^{1}(\mathbb{R})$ with inner product
$\langle h_{1},h_{2}\rangle_{H^{1}}:=\int_{\mathbb{R}}(h_{1}h_{2}+h_{1}^{\prime}h_{2}^{\prime})\,dx.$
Then the Riesz representer of evaluation at $x^{*}$ is the Green function
$\psi_{x^{*}}(x)=\frac{1}{2}e^{-|x-x^{*}|},$
so that $\kappa=\|\psi_{x^{*}}\|_{H^{1}}^{2}=\frac{1}{2}$. In particular, the quadratic expansion (36) becomes
$J(T)=\frac{m^{4}}{p^{2}}\,(T-T^{*})^{2}+o((T-T^{*})^{2}),$
and the optimal perturbation realizing the shift $\delta T=T-T^{*}$ at leading order is
$h_{T}(x)=\frac{m^{2}}{p}\,\delta T\,e^{-|x-x^{*}|}+o(\delta T)\quad\text{in }H^{1}.$
Proof.
The representer $\psi$ solves $(1-\partial_{x}^{2})\psi=\delta_{x^{*}}$ in distributions and is the unique $H^{1}$ solution decaying at infinity, hence $\psi(x)=\frac{1}{2}e^{-|x-x^{*}|}$. For any $h\in H^{1}$,
$\langle h,\psi\rangle_{H^{1}}=\int(h\psi+h^{\prime}\psi^{\prime})\,dx=\int h(1-\partial_{x}^{2})\psi\,dx=h(x^{*}),$
so $\psi=\psi_{x^{*}}$. Moreover, $\kappa=\|\psi\|_{H^{1}}^{2}=\langle\psi,\psi\rangle_{H^{1}}=\psi(x^{*})=1/2$. Substituting into Proposition 3.20 yields the stated formulas. ∎
4. The Schwarzian derivative: projective curvature and $L^{p}$ deformation
4.1. Definition and basic properties
Definition 4.1.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the Schwarzian derivative is
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}.$
Proposition 4.2.
Let $f=\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$. Then
$S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}.$
The map $S\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth with differential
$D_{\varphi}S(\delta\varphi)=\delta f^{\prime\prime}-f^{\prime}\delta f^{\prime},\qquad\delta f=\frac{\delta\varphi^{\prime}}{\varphi^{\prime}}.$
Proof.
From $f=\log(\varphi^{\prime})$ we have $f^{\prime}=\varphi^{\prime\prime}/\varphi^{\prime}$ and $f^{\prime\prime}=\varphi^{\prime\prime\prime}/\varphi^{\prime}-(\varphi^{\prime\prime}/\varphi^{\prime})^{2}$, which implies
$S(\varphi)=f^{\prime\prime}+(f^{\prime})^{2}-\frac{3}{2}(f^{\prime})^{2}=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}.$
Smoothness follows from $\varphi\mapsto f$ being smooth and $W^{\infty,1}$ being closed under differentiation and multiplication. The differential is computed by varying $f$. ∎
Proposition 4.3 (Cocycle identity).
For $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$.
$S(\varphi\circ\psi)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi).$
In particular, $S(\varphi)=0$ if and only if $\varphi\in\operatorname{Aff}(\mathbb{R})$.
Proof.
Direct computation using the chain rule. The kernel statement follows since $S(ax+b)=0$ for $a>0$, $b\in\mathbb{R}$, and if $S(\varphi)=0$ then $f=\log(\varphi^{\prime})$ satisfies $f^{\prime\prime}=\frac{1}{2}(f^{\prime})^{2}$. The general solution is $f^{\prime}=-2/(x-x_{0})$ or $f^{\prime}\equiv 0$. Only $f^{\prime}\equiv 0$ gives $\varphi^{\prime}-1\in W^{1,1}(\mathbb{R})$, hence $\varphi^{\prime}=c$ constant, which with $\varphi^{\prime}(\pm\infty)=1$ forces $\varphi^{\prime}=1$, so $\varphi(x)=x+b$. ∎
Remark 4.4.
The additive cocycle $\Phi_{\infty}(\varphi\circ\psi)=(\Phi_{\infty}(\varphi)\circ\psi)+\Phi_{\infty}(\psi)$ is the linear shadow of the projective cocycle for $S$. The operator $f\mapsto f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ is the one-dimensional projective curvature.
4.2. Score curvature and Fisher-Rao identities
The potential form of Proposition 4.2 extends verbatim from Jacobians $\varphi^{\prime}$ to arbitrary positive densities. This recasts the Schwarzian as an intrinsic “score curvature” and connects it to Fisher information via identities that are invisible in the purely group-theoretic formulation.
Definition 4.5 (Score, Fisher information, and density Schwarzian).
Let $\mu=g\,dx$ be a probability measure on $\mathbb{R}$ with $g>0$, $\log g\in W^{\infty,1}(\mathbb{R})$, and sufficient decay so that all integrations by parts below are legitimate. Its score is
$s_{\mu}:=(\log g)^{\prime}\in W^{\infty,1}(\mathbb{R}).$
The (location) Fisher information is
$I(\mu):=\int_{\mathbb{R}}s_{\mu}(x)^{2}\,g(x)\,dx=\|s_{\mu}\|_{L^{2}(\mu)}^{2}.$
The Schwarzian curvature of $\mu$ is the projective curvature of its score,
$S(\mu):=s_{\mu}^{\prime}-\frac{1}{2}s_{\mu}^{2}\in W^{\infty,1}(\mathbb{R}).$
Remark 4.6.
If $g=\varphi^{\prime}$ for some $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, then $S(\mu)=S(\varphi)$ by Proposition 4.2 applied to $f=\log g$. Thus $S(\mu)$ is the density-side incarnation of the group Schwarzian.
Proposition 4.7 (Square-root representation).
Let $\rho:=\sqrt{g}$. Then
$I(\mu)=4\int_{\mathbb{R}}|\rho^{\prime}(x)|^{2}\,dx.$
Proof.
Since $\rho^{\prime}/\rho=\frac{1}{2}(\log g)^{\prime}=\frac{1}{2}s_{\mu}$, we have $s_{\mu}=2\rho^{\prime}/\rho$ and hence
$I(\mu)=\int_{\mathbb{R}}s_{\mu}^{2}g\,dx=\int_{\mathbb{R}}\frac{4(\rho^{\prime})^{2}}{\rho^{2}}\,\rho^{2}\,dx=4\int_{\mathbb{R}}(\rho^{\prime})^{2}\,dx.$
Proposition 4.8 (Mean projective curvature and Fisher information).
Assume additionally that $g^{\prime}\in L^{1}(\mathbb{R})$ and $g^{\prime}(x)\to 0$ as $x\to\pm\infty$. Then
$\int_{\mathbb{R}}S(\mu)(x)\,g(x)\,dx\;=\;-\frac{3}{2}\,I(\mu).$ (37)
Proof.
Write $s=s_{\mu}$ and note $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. Using $(\log g)^{\prime\prime}=g^{\prime\prime}/g-(g^{\prime}/g)^{2}$, we obtain $s^{\prime}g=g^{\prime\prime}-(g^{\prime})^{2}/g$ and therefore
$\int_{\mathbb{R}}s^{\prime}g\,dx=\int_{\mathbb{R}}g^{\prime\prime}\,dx-\int_{\mathbb{R}}\frac{(g^{\prime})^{2}}{g}\,dx=-\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu),$
since $\int g^{\prime\prime}=g^{\prime}(+\infty)-g^{\prime}(-\infty)=0$ by hypothesis and $\frac{(g^{\prime})^{2}}{g}=s^{2}g$. Consequently,
$\int_{\mathbb{R}}S(\mu)\,g\,dx=\int_{\mathbb{R}}s^{\prime}g\,dx-\frac{1}{2}\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu)-\frac{1}{2}I(\mu)=-\frac{3}{2}I(\mu).$
Corollary 4.9.
Under the hypotheses of Proposition 4.8,
$\|S(\mu)\|_{L^{2}(\mu)}\geq\frac{3}{2}\,I(\mu).$
Proof.
By Cauchy-Schwarz and (37),
$\frac{3}{2}I(\mu)=\left|\int_{\mathbb{R}}S(\mu)\,d\mu\right|\leq\|S(\mu)\|_{L^{2}(\mu)}\cdot\|1\|_{L^{2}(\mu)}=\|S(\mu)\|_{L^{2}(\mu)}.$
Theorem 4.10 (Second-order score constraint and Bhattacharyya-type bound).
[…]
$u_{1}:=\partial_{\theta}\log g(X-\theta)\big|_{\theta=0}=-s_{\mu}(X),\qquad v_{2}:=\partial_{\theta}^{2}\log g(X-\theta)\big|_{\theta=0}+u_{1}(X)^{2}=s_{\mu}^{\prime}(X)+s_{\mu}(X)^{2}.$
Then the unbiasedness conditions imply the linear constraints in $L^{2}(\mu)$:
$\langle T,u_{1}\rangle_{\mu}=1,\qquad\langle T,v_{2}\rangle_{\mu}=0,$ (38)
and the variance satisfies
$\mathrm{Var}_{0}(\hat{\theta})=\|T\|_{L^{2}(\mu)}^{2}\;\geq\;\frac{\|v_{2}\|_{L^{2}(\mu)}^{2}}{\|u_{1}\|_{L^{2}(\mu)}^{2}\,\|v_{2}\|_{L^{2}(\mu)}^{2}-\langle u_{1},v_{2}\rangle_{\mu}^{2}}.$ (39)
Moreover,
$v_{2}\;=\;S(\mu)+\frac{3}{2}\,s_{\mu}^{2},$ (40)
so the only projectively invariant contribution to the second score composite is $S(\mu)$.
Proof.
The first constraint in (38) is the standard Cramér-Rao identity. Indeed $E_{\theta}[T]=0$, hence $0=\partial_{\theta}E_{\theta}[T]=E_{\theta}[\partial_{\theta}T]+E_{\theta}[T\,\partial_{\theta}\log p_{\theta}]$, and since $\partial_{\theta}T=-1$ and $\partial_{\theta}\log p_{\theta}=u_{1}$ this gives $E_{\theta}[Tu_{1}]=1$ at $\theta=0$.
For the second constraint, differentiate once more: $0=\partial_{\theta}^{2}E_{\theta}[T]=\partial_{\theta}E_{\theta}[Tu_{1}]$. Using $\partial_{\theta}E_{\theta}[F]=E_{\theta}[Fu_{1}]$ for $\theta$-independent $F$ and the identity $\partial_{\theta}u_{1}=\partial_{\theta}^{2}\log p_{\theta}=u_{2}$, one obtains
$0=\partial_{\theta}E_{\theta}[Tu_{1}]=E_{\theta}[Tu_{2}]+E_{\theta}[Tu_{1}^{2}]+E_{\theta}[(\partial_{\theta}T)u_{1}].$
Since $\partial_{\theta}T=-1$ and $E_{\theta}[u_{1}]=0$, the last term vanishes and we obtain $E_{0}[T(u_{2}+u_{1}^{2})]=0$, i.e. $\langle T,v_{2}\rangle_{\mu}=0$.
Now consider the Hilbert space $H=L^{2}(\mu)$ with inner product $\langle\cdot,\cdot\rangle_{\mu}$. Among all $T\in H$ satisfying (38), the minimizer of $\|T\|_{H}$ is the unique element $T_{0}\in\mathrm{span}\{u_{1},v_{2}\}$ satisfying the same constraints. Writing $T_{0}=\alpha u_{1}+\beta v_{2}$ and imposing (38) yields the $2\times 2$ linear system
$\alpha\|u_{1}\|^{2}+\beta\langle v_{2},u_{1}\rangle=1,\qquad\alpha\langle u_{1},v_{2}\rangle+\beta\|v_{2}\|^{2}=0,$
whose determinant is $\|u_{1}\|^{2}\|v_{2}\|^{2}-\langle u_{1},v_{2}\rangle^{2}>0$. A direct elimination gives
$\|T_{0}\|^{2}=\frac{\|v_{2}\|^{2}}{\|u_{1}\|^{2}\|v_{2}\|^{2}-\langle u_{1},v_{2}\rangle^{2}}.$
For any admissible $T$, the difference $T-T_{0}$ is orthogonal to $\mathrm{span}\{u_{1},v_{2}\}$ by the normal equations for constrained minimization, hence by Pythagoras $\|T\|^{2}=\|T_{0}\|^{2}+\|T-T_{0}\|^{2}\geq\|T_{0}\|^{2}$, which is exactly (39).
Finally, (40) is the algebraic identity $s^{\prime}+s^{2}=(s^{\prime}-\frac{1}{2}s^{2})+\frac{3}{2}s^{2}$ with $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. ∎
Remark 4.11.
The numerical lower bound (39) coincides with the classical second-order Bhattacharyya bound for the location model. The point of the present formulation is the projective decomposition (40): it isolates $S(\mu)$ as the projectively invariant component inside the second-order score composite.
4.3. Projective connections and cocycle lifting
Definition 4.12 (Quadratic differential pullback).
Let $\mathcal{Q}^{\infty,1}(\mathbb{R}):=W^{\infty,1}(\mathbb{R})$, viewed as coefficients of quadratic differentials $q(x)\,dx^{2}$. For $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ define the pullback
$\psi^{*}q:=(q\circ\psi)(\psi^{\prime})^{2}.$
Then pullback is contravariant:
$(\psi_{1}\circ\psi_{2})^{*}=\psi_{2}^{*}\circ\psi_{1}^{*},$
so $q\mapsto q\cdot\psi:=\psi^{*}q$ is a right action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $\mathcal{Q}^{\infty,1}(\mathbb{R})$.
Proposition 4.13 (Affine action on projective connections).
Define, for $q\in\mathcal{Q}^{\infty,1}(\mathbb{R})$ and $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$,
$q\star\psi:=\psi^{*}q+S(\psi).$
Then $\star$ is a right action:
$(q\star\psi_{1})\star\psi_{2}=q\star(\psi_{1}\circ\psi_{2}).$
Proof.
By contravariance of pullback and Proposition 4.3,
$(q\star\psi_{1})\star\psi_{2}$ $=\psi_{2}^{*}(\psi_{1}^{*}q+S(\psi_{1}))+S(\psi_{2})$
$=(\psi_{1}\circ\psi_{2})^{*}q+(S(\psi_{1})\circ\psi_{2})(\psi_{2}^{\prime})^{2}+S(\psi_{2})$
$=(\psi_{1}\circ\psi_{2})^{*}q+S(\psi_{1}\circ\psi_{2})=q\star(\psi_{1}\circ\psi_{2}).$
Definition 4.14 (Schwarzian bundle over the affine quotient).
Let $\operatorname{Aff}(\mathbb{R})\subset\operatorname{Diff}^{-\infty}(\mathbb{R})$ be the affine subgroup (in the decay class, only translations survive, but we keep the notation). Define the associated bundle
$\mathcal{Q}_{\mathrm{proj}}:=\operatorname{Diff}^{-\infty}(\mathbb{R})\times_{\operatorname{Aff}(\mathbb{R})}\mathcal{Q}^{\infty,1}(\mathbb{R}),$
where $(\varphi,q)\sim(\varphi\circ A,A^{*}q)$ for $A\in\operatorname{Aff}(\mathbb{R})$, with $A^{*}q=(q\circ A)(A^{\prime})^{2}$. We write elements as $\varphi,q]$ and the base point as $\varphi]\in\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
Theorem 4.15 (Cocycle lift as canonical section).
The Schwarzian derivative defines a well-defined section
$\sigma_{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to\mathcal{Q}_{\mathrm{proj}},\qquad\sigma_{S}([\varphi])=[\varphi,S(\varphi)].$
Moreover, the left action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on the base by left composition,
$\psi\cdot[\varphi]=[\psi\circ\varphi],$
lifts to an action on $\mathcal{Q}_{\mathrm{proj}}$ given by
$\psi\cdot[\varphi,q]:=[\psi\circ\varphi,\;q+\varphi^{*}S(\psi)],$ (41)
and $\sigma_{S}$ is equivariant:
$\psi\cdot\sigma_{S}([\varphi])=\sigma_{S}(\psi\cdot[\varphi]).$
Proof.
Well-definedness of $\sigma_{S}$. If $\varphi$ is replaced by $\varphi\circ A$ with $A\in\operatorname{Aff}(\mathbb{R})$, then by Proposition 4.3 and $S(A)=0$,
$S(\varphi\circ A)=(S(\varphi)\circ A)(A^{\prime})^{2}=A^{*}S(\varphi),$
so $\varphi\circ A,S(\varphi\circ A)]=[\varphi,A^{*}S(\varphi)]=[\varphi,S(\varphi)]$ in $\mathcal{Q}_{\mathrm{proj}}$.
Well-definedness of the lifted action. If $\varphi,q]=[\varphi\circ A,A^{*}q]$, then using contravariance of pullback,
$(\varphi\circ A)^{*}S(\psi)=A^{*}(\varphi^{*}S(\psi)),$
hence
$\psi\cdot[\varphi\circ A,A^{*}q]=[\psi\circ\varphi\circ A,\;A^{*}q+(\varphi\circ A)^{*}S(\psi)]=[\psi\circ\varphi\circ A,\;A^{*}(q+\varphi^{*}S(\psi))]=\psi\cdot[\varphi,q],$
so (41) descends to the quotient.
Equivariance. By Proposition 4.3,
$S(\psi\circ\varphi)=\varphi^{*}S(\psi)+S(\varphi),$
therefore
$\psi\cdot\sigma_{S}([\varphi])=\psi\cdot[\varphi,S(\varphi)]=[\psi\circ\varphi,\;S(\varphi)+\varphi^{*}S(\psi)]=[\psi\circ\varphi,\;S(\psi\circ\varphi)]=\sigma_{S}([\psi\circ\varphi]).$
Proposition 4.16 (Infinitesimal Schwarzian cocycle).
Let $u$ be a vector field on $\mathbb{R}$ such that $u^{\prime},u^{\prime\prime},u^{\prime\prime\prime}\in L^{1}(\mathbb{R})$ and $u^{\prime},u^{\prime\prime}\to 0$ at $\pm\infty$, and let $\varphi_{t}$ be the corresponding local flow with $\varphi_{0}=\operatorname{id}$. Then
$S(\varphi_{t})=t\,u^{\prime\prime\prime}+O_{W^{k,1}}(t^{2})\qquad(t\to 0),$
for each $k\geq 0$. In particular, $D_{\operatorname{id}}S(u)=u^{\prime\prime\prime}$.
Proof.
Write $\varphi_{t}(x)=x+tu(x)+O(t^{2})$ in $W^{k,1}$ for every $k$. Then $\varphi_{t}^{\prime}=1+tu^{\prime}+O(t^{2})$, $\varphi_{t}^{\prime\prime}=tu^{\prime\prime}+O(t^{2})$, $\varphi_{t}^{\prime\prime\prime}=tu^{\prime\prime\prime}+O(t^{2})$. Substituting into Definition 4.1 and expanding to first order yields $S(\varphi_{t})=tu^{\prime\prime\prime}+O(t^{2})$ in $W^{k,1}$. ∎
Definition 4.17 (Gelfand-Fuchs cocycle).
On the Lie algebra $\mathfrak{X}^{-\infty}(\mathbb{R})$ of such vector fields, define
$\omega_{\mathrm{GF}}(u,v):=\frac{1}{2}\int_{\mathbb{R}}\bigl(u^{\prime}v^{\prime\prime}-u^{\prime\prime}v^{\prime}\bigr)\,dx.$
Proposition 4.18 (Normalization and cocycle property).
For $u,v\in\mathfrak{X}^{-\infty}(\mathbb{R})$,
$\omega_{\mathrm{GF}}(u,v)=\int_{\mathbb{R}}u^{\prime\prime\prime}(x)\,v(x)\,dx=-\int_{\mathbb{R}}u(x)\,v^{\prime\prime\prime}(x)\,dx.$ (42)
In particular,
$\omega_{\mathrm{GF}}(u,v)=\int_{\mathbb{R}}\bigl(D_{\operatorname{id}}S(u)\bigr)(x)\,v(x)\,dx.$ (43)
Moreover, $\omega_{\mathrm{GF}}$ is a Lie algebra $2$-cocycle:
$\omega_{\mathrm{GF}}([u,v],w)+\omega_{\mathrm{GF}}([v,w],u)+\omega_{\mathrm{GF}}([w,u],v)=0.$
Proof.
Since $u^{\prime},v^{\prime}\to 0$ at $\pm\infty$ and $u^{\prime\prime},v^{\prime\prime}\in L^{1}$, integration by parts gives
$\int_{\mathbb{R}}u^{\prime\prime}v^{\prime}\,dx=-\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx,$
so
$\omega_{\mathrm{GF}}(u,v)=\frac{1}{2}\int_{\mathbb{R}}(u^{\prime}v^{\prime\prime}-u^{\prime\prime}v^{\prime})\,dx=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx.$
A second integration by parts yields $\int u^{\prime}v^{\prime\prime}=\int u^{\prime\prime\prime}v=-\int uv^{\prime\prime\prime}$, proving (42). Then (43) follows from Proposition 4.16.
For the cocycle identity, write $\omega_{\mathrm{GF}}(a,b)=\int a\,b^{\prime\prime\prime}$ using (42) and the bracket $u,v]=uv^{\prime}-vu^{\prime}$. The cyclic sum of the integrands is
$(uv^{\prime}-vu^{\prime})w^{\prime\prime\prime}+(vw^{\prime}-wv^{\prime})u^{\prime\prime\prime}+(wu^{\prime}-uw^{\prime})v^{\prime\prime\prime}=\partial_{x}\Bigl(uv^{\prime}w^{\prime\prime}-uv^{\prime\prime}w^{\prime}+vw^{\prime}u^{\prime\prime}-vw^{\prime\prime}u^{\prime}+wu^{\prime}v^{\prime\prime}-wu^{\prime\prime}v^{\prime}\Bigr),$
an exact derivative. The boundary term vanishes by the decay hypotheses, hence the cyclic integral is zero. ∎
4.4. The $L^{p}$-Schwarzian
We introduce a family interpolating between the affine cocycle structure of $\Phi_{p}$ and the projective cocycle of $S$.
Definition 4.19.
For $p\in[1,\infty)$ and $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the $L^{p}$-Schwarzian is
$S_{p}(\varphi)=\left(\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}+S(\varphi)\right)\varphi^{\prime 1/p}.$
Proposition 4.20.
The $L^{p}$-Schwarzian satisfies
  • (i)
    $S_{p}(A)=0$ for affine $A$.
  • (ii)
    $S_{p}(\varphi\circ\psi)=(S_{p}(\varphi)\circ\psi)(\psi^{\prime})^{2+1/p}+S_{p}(\psi)(\varphi^{\prime}\circ\psi)^{1/p}+\frac{3}{p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}\,((\varphi\circ\psi)^{\prime})^{1/p}$.
  • (iii)
    $S_{p}(\varphi)(x)\to S(\varphi)(x)$ pointwise as $p\to\infty$.
Proof.
We use the pointwise identity
$S_{p}(\theta)=\left(S(\theta)+\frac{3}{2p}\left(\frac{\theta^{\prime\prime}}{\theta^{\prime}}\right)^{2}\right)(\theta^{\prime})^{1/p},$ (44)
valid for any orientation-preserving $C^{3}$ diffeomorphism $\theta$. For (i), if $A(x)=ax+b$ with $a>0$, then $S(A)=0$ and $A^{\prime\prime}=0$, hence (44) gives $S_{p}(A)=0$. For (ii), set $\Theta=\varphi\circ\psi$. Then $\Theta^{\prime}=(\varphi^{\prime}\circ\psi)\psi^{\prime}$ and
$\frac{\Theta^{\prime\prime}}{\Theta^{\prime}}=\partial_{x}\log\Theta^{\prime}=\partial_{x}\bigl(\log(\varphi^{\prime}\circ\psi)+\log\psi^{\prime}\bigr)=\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime}+\frac{\psi^{\prime\prime}}{\psi^{\prime}}.$
Squaring yields
$\left(\frac{\Theta^{\prime\prime}}{\Theta^{\prime}}\right)^{2}=\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)^{2}(\psi^{\prime})^{2}+2\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}+\left(\frac{\psi^{\prime\prime}}{\psi^{\prime}}\right)^{2}.$
The classical Schwarzian satisfies $S(\Theta)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi)$. Inserting these identities into (44) for $\theta=\Theta$ and using $\Theta^{\prime 1/p}=(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}$, we obtain
$S_{p}(\Theta)$ $=\left(\left(S(\varphi)+\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}\right)\circ\psi\right)(\psi^{\prime})^{2+1/p}(\varphi^{\prime}\circ\psi)^{1/p}$
$\quad+\left(S(\psi)+\frac{3}{2p}\left(\frac{\psi^{\prime\prime}}{\psi^{\prime}}\right)^{2}\right)(\psi^{\prime})^{1/p}(\varphi^{\prime}\circ\psi)^{1/p}$
$\quad+\frac{3}{p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}\,(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}.$
The first two terms are exactly $(S_{p}(\varphi)\circ\psi)(\psi^{\prime})^{2+1/p}$ and $S_{p}(\psi)(\varphi^{\prime}\circ\psi)^{1/p}$ by (44), and the last term equals the stated formula since $(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}=\Theta^{\prime 1/p}$. For (iii), (44) gives
$S_{p}(\varphi)=S(\varphi)\,(\varphi^{\prime})^{1/p}+\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}(\varphi^{\prime})^{1/p},$
and pointwise $(\varphi^{\prime})^{1/p}\to 1$ while $\frac{1}{p}(\varphi^{\prime\prime}/\varphi^{\prime})^{2}(\varphi^{\prime})^{1/p}\to 0$, hence $S_{p}(\varphi)\to S(\varphi)$ pointwise. ∎
Remark 4.21 (Cocycle defect at finite $p$).
The third term in Proposition 4.20(ii) shows that $S_{p}$ is not a $1$-cocycle for finite $p$. It is precisely the obstruction to functoriality of the nonlinear curvature map $\varphi\mapsto S_{p}(\varphi)$ under composition. Two features justify $S_{p}$ as the correct interpolation between affine and projective geometry: (i) the defect term vanishes in the classical limit $p\to\infty$, recovering the Schwarzian cocycle; (ii) in $p$-root coordinates (Proposition 4.22) the expression for $S_{p}$ becomes a clean second-order curvature functional.
Proposition 4.22.
In $p$-root coordinates $\theta=\varphi^{\prime 1/p}$.
$S_{p}(\varphi)=p\theta^{\prime\prime}-\frac{p(p-1)}{2}\frac{(\theta^{\prime})^{2}}{\theta}.$
Equivalently, with $f=\Phi_{p}(\varphi)=p(\theta-1)$.
$S_{p}(\varphi)=f^{\prime\prime}-\frac{p-1}{2p}\frac{(f^{\prime})^{2}}{1+f/p}.$
Proof.
Let $u=\log(\varphi^{\prime})=p\log\theta$. Then $u^{\prime}=p\theta^{\prime}/\theta$ and $u^{\prime\prime}=p\theta^{\prime\prime}/\theta-p(\theta^{\prime})^{2}/\theta^{2}$. Substituting into the definition and simplifying yields both formulas. ∎
Corollary 4.23.
The map $S_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth. On bounded subsets, it is locally Lipschitz as a map into $W^{k,1}(\mathbb{R})$.
4.5. Asymptotics of $S_{p}$ as $p\to\infty$
Proposition 4.24.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $u=\log(\varphi^{\prime})$ and $S=S(\varphi)$.
$S_{p}(\varphi)=S+\frac{1}{p}\left(uS+\frac{3}{2}(u^{\prime})^{2}\right)+\frac{1}{p^{2}}\left(\frac{u^{2}S}{2}+\frac{3u(u^{\prime})^{2}}{2}\right)+O_{W^{k,1}}\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
From Definition 4.19, $S_{p}(\varphi)=(S+\frac{3}{2p}(u^{\prime})^{2})\mathrm{e}^{u/p}$. Expanding $\mathrm{e}^{u/p}=1+u/p+u^{2}/(2p^{2})+O(p^{-3})$ and collecting terms yields the expansion. The $W^{k,1}$ remainder estimate uses Fa i . Bruno and Leibniz as in Proposition 2.23. ∎
4.6. Sturm-Liouville interpretation
The Schwarzian governs a natural Sturm-Liouville problem.
Proposition 4.25.
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $S=S(\varphi)$. Define $y_{2}=(\varphi^{\prime})^{-1/2}$ and $y_{1}=\varphi(\varphi^{\prime})^{-1/2}$. Then $y_{1},y_{2}$ are linearly independent solutions of
$y^{\prime\prime}+\frac{1}{2}Sy=0$
with Wronskian $W(y_{1},y_{2})\equiv 1$. Conversely, if $y_{1},y_{2}$ solve $y^{\prime\prime}+qy=0$ with $y_{2}>0$ and $W(y_{1},y_{2})\equiv 1$, then $\varphi=y_{1}/y_{2}$ satisfies $\varphi^{\prime}>0$ and $S(\varphi)=2q$.
Proof.
Let $u=\log(\varphi^{\prime})$. For $y_{2}=\mathrm{e}^{-u/2}$. $y_{2}^{\prime}=-\frac{1}{2}u^{\prime}y_{2}$ and $y_{2}^{\prime\prime}=(\frac{1}{4}(u^{\prime})^{2}-\frac{1}{2}u^{\prime\prime})y_{2}$. Then
$y_{2}^{\prime\prime}+\frac{S}{2}y_{2}=\left(\frac{(u^{\prime})^{2}}{4}-\frac{u^{\prime\prime}}{2}+\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}\right)y_{2}=0.$
For $y_{1}=\varphi y_{2}$. $y_{1}^{\prime\prime}=\varphi y_{2}^{\prime\prime}$ after simplification, so $y_{1}^{\prime\prime}+\frac{S}{2}y_{1}=0$. The Wronskian is $W(y_{1},y_{2})=y_{1}^{\prime}y_{2}-y_{1}y_{2}^{\prime}=\varphi^{\prime}y_{2}^{2}=\varphi^{\prime}(\varphi^{\prime})^{-1}=1$. Conversely, $\varphi^{\prime}=W(y_{1},y_{2})/y_{2}^{2}=1/y_{2}^{2}>0$. Computing $S(\varphi)$ from $\log(\varphi^{\prime})=-2\log y_{2}$ gives $S(\varphi)=2q$. ∎
Theorem 4.26 (Liouville conjugation).
Define $(\mathcal{L}_{\varphi}u)(x)=(\varphi^{\prime}(x))^{-1/2}u(\varphi(x))$. For any $q\in C^{\infty}(\mathbb{R})$.
$\mathcal{L}_{\varphi}\circ(\partial_{y}^{2}+q(y))\circ\mathcal{L}_{\varphi}^{-1}=\partial_{x}^{2}+q(\varphi(x))\varphi^{\prime}(x)^{2}+\frac{1}{2}S(\varphi)(x).$
Proof.
Set $v=\mathcal{L}_{\varphi}^{-1}u$, so $v(\varphi(x))=\varphi^{\prime}(x)^{1/2}u(x)$. Differentiate twice and simplify using $\varphi^{\prime\prime}=\varphi^{\prime}(\log\varphi^{\prime})^{\prime}$. ∎
Remark 4.27.
The Schwarzian is the unique scalar correction ensuring functoriality of Liouville conjugation for second-order operators. In semiclassical analysis, it is the subprincipal correction when conjugating $h^{2}\partial_{x}^{2}$ by the Fourier integral operator implementing $\varphi$.
4.7. Semiclassical interpretation
Definition 4.28 (Semiclassical parameter).
Identify $p=1/\hbar$ where $\hbar>0$ is the semiclassical parameter. The $L^{p}$-Schwarzian becomes the $\hbar$-Schwarzian
$S_{\hbar}(\varphi):=S_{1/\hbar}(\varphi)=\left(S(\varphi)+\frac{3\hbar}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}\right)\varphi^{\prime\hbar}.$
Corollary 4.29 (WKB expansion of $S_{\hbar}$).
As $\hbar\to 0$ (equivalently $p\to\infty$), Proposition 4.24 yields the full asymptotic expansion
$S_{\hbar}(\varphi)=S(\varphi)+\hbar S^{(1)}(\varphi)+\hbar^{2}S^{(2)}(\varphi)+O_{W^{k,1}}(\hbar^{3}),$
where, with $u=\log(\varphi^{\prime})$,
$S^{(1)}(\varphi)$ $=u\,S(\varphi)+\frac{3}{2}(u^{\prime})^{2},$
$S^{(2)}(\varphi)$ $=\frac{u^{2}}{2}\,S(\varphi)+\frac{3u(u^{\prime})^{2}}{2}.$
Each coefficient $S^{(n)}$ is a universal polynomial in $u$ and its derivatives, involving derivatives up to order $n+2$.
Proposition 4.30 (Semiclassical Liouville correction).
Apply Theorem 4.26 to the semiclassical Schrödinger operator $\hbar^{2}\partial_{y}^{2}+V(y)$ by writing it as $\hbar^{2}(\partial_{y}^{2}-q(y))$ with $q(y)=V(y)/\hbar^{2}$. Then
$\mathcal{L}_{\varphi}\circ\bigl(-\hbar^{2}\partial_{y}^{2}+V(y)\bigr)\circ\mathcal{L}_{\varphi}^{-1}=-\hbar^{2}\partial_{x}^{2}+V(\varphi(x))\varphi^{\prime}(x)^{2}-\frac{\hbar^{2}}{2}\,S(\varphi)(x).$
In particular, the Schwarzian contributes the $\hbar^{2}$-order correction term dictated by Liouville functoriality.
Proof.
Set $q(y)=-V(y)/\hbar^{2}$ in Theorem 4.26 and multiply the resulting identity by $\hbar^{2}$. ∎
5. Central Extensions and Cocycle Geometry
The Schwarzian derivative is the fundamental obstruction to lifting the composition law on diffeomorphisms to an affine structure. This obstruction is measured by Lie algebra cohomology and central extensions. The results in the preceding sections show that $L^{p}$-based coordinates linearize the right-invariant Finsler geometry on diffeomorphism and density spaces, yielding explicit distances and geodesics. In contrast, the projective and cohomological features governed by the Schwarzian and by central extensions do not admit an equally clean interpolation. On the one hand, the Bott-Thurston extension is rigid in the sense that the $p$-root Jacobian cocycle produces no new real-valued Lie algebra cohomology classes beyond the Gelfand-Fuchs class, up to an explicit scalar. On the other hand, the $L^{p}$-Schwarzian fails to satisfy a two-term cocycle identity: the unavoidable cross-term in its composition law shows that the projective curvature cannot be encoded by a simple $L^{p}$ cocycle.
A complementary phenomenon - and the main genuinely new ingredient of this section - is that although the Bott cocycle is not a group coboundary on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, it becomes an explicit groupoid coboundary once one allows dependence on a density through its score. This produces a canonical transgression functional whose coboundary is exactly the Bott phase. The resulting mechanism provides a clean categorical “absorption” of the projective multiplier into a central extension of Koopman transport.
5.1. The Gelfand-Fuchs cocycle
Definition 5.1 (Gelfand-Fuchs $2$-cocycle).
On the Lie algebra $\mathfrak{g}^{-\infty}$, define the bilinear form
$\omega(u,v):=\int_{\mathbb{R}}u^{\prime}(x)\,v^{\prime\prime}(x)\,\mathrm{d}x.$
Proposition 5.2 (Nontriviality of the Gelfand-Fuchs class).
The form $\omega$ is an alternating Lie algebra $2$-cocycle on $\mathfrak{g}^{-\infty}$; we refer to the verification of antisymmetry and the cocycle identity in §4 (projective cocycle lifting). Moreover, $\omega$ is not a coboundary: there is no continuous linear functional $\Lambda\colon\mathfrak{g}^{-\infty}\to\mathbb{R}$ such that $\omega(u,v)=\Lambda([u,v])$ for all $u,v$. Hence $\omega]\neq 0$ in $H^{2}(\mathfrak{g}^{-\infty};\mathbb{R})$.
Proof.
Assume for contradiction that $\omega$ is a coboundary, that is, there exists a continuous linear functional $\Lambda:\mathfrak{g}^{-\infty}\to\mathbb{R}$ such that
$\omega(u,v)=\Lambda([u,v])\qquad\text{for all }u,v\in\mathfrak{g}^{-\infty}.$
Restricting to $C_{c}^{\infty}(\mathbb{R})\subset\mathfrak{g}^{-\infty}$, fix arbitrary $v\in C_{c}^{\infty}(\mathbb{R})$ and choose $u\in C_{c}^{\infty}(\mathbb{R})$ with $u\equiv 1$ on a neighborhood of $\mathrm{supp}(v)$. Then $u^{\prime}=0$ on $\mathrm{spt}(v)$, hence $u^{\prime}v^{\prime\prime}\equiv 0$ and $\omega(u,v)=0$. On the other hand, on $\mathrm{supp}(v)$ one has $u,v]=uv^{\prime}-u^{\prime}v=v^{\prime}$, so
$0=\omega(u,v)=\Lambda([u,v])=\Lambda(v^{\prime}).$
Therefore,
$\Lambda(\partial_{x}f)=0\qquad\text{for all }f\in C_{c}^{\infty}(\mathbb{R}).$ (45)
If $g\in C_{c}^{\infty}(\mathbb{R})$ satisfies $\int_{\mathbb{R}}g\,dx=0$, then $g=f^{\prime}$ for $f(x):=\int_{-\infty}^{x}g(t)\,dt\in C_{c}^{\infty}(\mathbb{R})$, so (45) gives $\Lambda(g)=0$. Hence $\Lambda$ vanishes on the codimension-one subspace $\{g\in C_{c}^{\infty}(\mathbb{R}):\int g=0\}$, and there exists $c\in\mathbb{R}$ with
$\Lambda(g)=c\int_{\mathbb{R}}g(x)\,dx\qquad\text{for all }g\in C_{c}^{\infty}(\mathbb{R}).$ (46)
Consequently, for $u,v\in C_{c}^{\infty}(\mathbb{R})$,
$\omega(u,v)=\Lambda([u,v])=c\int_{\mathbb{R}}(uv^{\prime}-u^{\prime}v)\,dx=-2c\int_{\mathbb{R}}u^{\prime}v\,dx,$
where $\int uv^{\prime}=-\int u^{\prime}v$ by compact support.
On the other hand, integrating by parts twice gives, for $u,v\in C_{c}^{\infty}(\mathbb{R})$,
$\omega(u,v)=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx=\int_{\mathbb{R}}u^{\prime\prime\prime}v\,dx.$
Combining the two identities yields
$\int_{\mathbb{R}}(u^{\prime\prime\prime}+2c\,u^{\prime})\,v\,dx=0\qquad\text{for all }u,v\in C_{c}^{\infty}(\mathbb{R}).$
Fixing $u$ and using the arbitrariness of $v$ implies $u^{\prime\prime\prime}+2c\,u^{\prime}=0$ as a distribution, hence as a smooth identity. Writing $w:=u^{\prime}$, we obtain
$w^{\prime\prime}+2c\,w=0\quad\text{on }\mathbb{R}.$
Every $C^{\infty}$ solution of this ODE is real-analytic, so if $w$ has compact support then $w\equiv 0$. Thus $u^{\prime}\equiv 0$, so $u$ is constant; since $u\in C_{c}^{\infty}(\mathbb{R})$, it follows that $u\equiv 0$. This contradiction shows that no such $\Lambda$ exists. ∎
Remark 5.3.
We will use the already established linearization $D_{\operatorname{id}}S(u)=u^{\prime\prime\prime}$ and the identity $\omega(u,v)=\int_{\mathbb{R}}u^{\prime\prime\prime}(x)v(x)\,dx$ (proved in §4) to identify $\omega$ as the transgression of the infinitesimal Schwarzian.
5.2. The Virasoro-Bott extension
Definition 5.4 (Central extension).
The Virasoro-Bott central extension of $\mathfrak{g}^{-\infty}$ is the Lie algebra
$\widehat{\mathfrak{g}}^{-\infty}=\mathfrak{g}^{-\infty}\oplus\mathbb{R}c$
with bracket
$(u,\alpha),(v,\beta)]=\left([u,v],\omega(u,v)\right).$
The element $c=(0,1)$ is central. $c,\cdot]=0$.
Proposition 5.5 (Integrability).
The central extension $\widehat{\mathfrak{g}}^{-\infty}$ integrates to a central extension of Fréchet Lie groups
$1\to\mathbb{R}\to\widehat{\operatorname{Diff}}^{-\infty}(\mathbb{R})\to\operatorname{Diff}^{-\infty}(\mathbb{R})\to 1$
where the group $2$-cocycle is the Bott cocycle
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,\mathrm{d}\log(\psi^{\prime}).$
Proof.
We first verify that the Bott functional
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)(x)\,\frac{\psi^{\prime\prime}(x)}{\psi^{\prime}(x)}\,dx$
is well-defined on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ under the standing decay assumptions. By definition, $\psi^{\prime}-1$ is integrable together with all derivatives, hence $\log(\psi^{\prime})$ and $(\log\psi^{\prime})^{\prime}=\psi^{\prime\prime}/\psi^{\prime}$ are integrable; similarly $\log(\varphi^{\prime}\circ\psi)$ is bounded with integrable derivatives, so the product is integrable and the integral converges. Normalization is immediate: $B(\operatorname{id},\psi)=0$ and $B(\varphi,\operatorname{id})=0$ since $\log(\operatorname{id}^{\prime})=0$ and $d\log(\operatorname{id}^{\prime})=0$.
Group cocycle identity. For $\varphi,\psi,\chi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, write the group coboundary
$(\delta B)(\varphi,\psi,\chi):=B(\psi,\chi)-B(\varphi\circ\psi,\chi)+B(\varphi,\psi\circ\chi)-B(\varphi,\psi).$
Using the chain rule and logarithms,
$\log\bigl((\psi\circ\chi)^{\prime}\bigr)=\log(\psi^{\prime}\circ\chi)+\log(\chi^{\prime}),\qquad d\log\bigl((\psi\circ\chi)^{\prime}\bigr)=d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)+d\log(\chi^{\prime}),$
we expand
$B(\varphi,\psi\circ\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log\bigl((\psi\circ\chi)^{\prime}\bigr)$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)+\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime}).$
Similarly, since $(\varphi\circ\psi)^{\prime}\circ\chi=(\varphi^{\prime}\circ\psi\circ\chi)\cdot(\psi^{\prime}\circ\chi)$,
$\log\bigl((\varphi\circ\psi)^{\prime}\circ\chi\bigr)=\log(\varphi^{\prime}\circ\psi\circ\chi)+\log(\psi^{\prime}\circ\chi),$
hence
$B(\varphi\circ\psi,\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log\bigl((\varphi\circ\psi)^{\prime}\circ\chi\bigr)\,d\log(\chi^{\prime})$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime})+\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime}).$
Substituting these two expressions into $(\delta B)(\varphi,\psi,\chi)$, the terms $\frac{1}{2}\int\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime})$ cancel, and we obtain
$(\delta B)(\varphi,\psi,\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime})-\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime})$
$\qquad+\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)-\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)-\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}).$
Now use the pullback/change-of-variables identity for exact one-forms:
$d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)=\chi^{*}(d\log(\psi^{\prime})),\qquad\int_{\mathbb{R}}F\circ\chi\;\chi^{*}\alpha=\int_{\mathbb{R}}F\,\alpha,$
valid for orientation-preserving $\chi$ and integrable forms. With $F=\log(\varphi^{\prime}\circ\psi)$ and $\alpha=d\log(\psi^{\prime})$, we get
$\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)=\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}),$
so $(\delta B)(\varphi,\psi,\chi)=0$. Hence $B$ is a group $2$-cocycle.
Infinitesimal cocycle. Let $\varphi_{t},\psi_{s}$ be the flows of $u,v\in\mathfrak{g}^{-\infty}$ with $\varphi_{0}=\psi_{0}=\operatorname{id}$. Then
$\varphi_{t}^{\prime}(x)=1+t\,u^{\prime}(x)+O(t^{2}),\qquad\psi_{s}^{\prime}(x)=1+s\,v^{\prime}(x)+O(s^{2}),\qquad\frac{\psi_{s}^{\prime\prime}(x)}{\psi_{s}^{\prime}(x)}=s\,v^{\prime\prime}(x)+O(s^{2}).$
Also $\log(\varphi_{t}^{\prime}\circ\psi_{s})=t\,u^{\prime}(x)+O(t^{2})+O(ts^{2})$, so inserting into the defining integral gives
$B(\varphi_{t},\psi_{s})=\frac{1}{2}\int_{\mathbb{R}}\Bigl(tu^{\prime}(x)+o(t)\Bigr)\Bigl(sv^{\prime\prime}(x)+o(s)\Bigr)\,dx=\frac{ts}{2}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx+o(ts).$
Thus the bilinear form obtained by differentiating $B$ at $(\operatorname{id},\operatorname{id})$ is $\frac{1}{2}\,\omega(u,v)$. Equivalently, the cocycle $2B$ differentiates to $\omega$.
The standard construction of a central extension from a normalized group $2$-cocycle therefore yields a central extension of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ whose Lie algebra is the central extension defined by $\omega$ (using the cocycle $2B$) or by $\frac{1}{2}\omega$ (using $B$). ∎
5.3. $L^{p}$ cocycles and rigidity
The quantity $\varphi^{\prime 1/p}$ is the Jacobian cocycle for the natural action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $p$th-root densities. Variants of this Jacobian cocycle, and Bott-Thurston type constructions for general $p$, appear in the literature on $L^{p}$ diffeomorphism groups and density geometry; see in particular [24] for extensions of the Virasoro algebra by modules of $\lambda$-densities. The point here is that, in the present decay-controlled setting, one can work in explicit $p$-root coordinates and obtain a clean rigidity statement, namely the induced real-valued Lie algebra $2$-cocycle is always the Gelfand-Fuchs cocycle, up to the explicit scalar factor $p^{-2}$. In particular, there is no new real-valued Lie algebra $2$-cocycle coming from the symmetric pairing $\int u^{\prime}v^{\prime}\,\mathrm{d}x$, which is not alternating and therefore does not define a Chevalley-Eilenberg $2$-cochain.
Definition 5.6 ($p$-root Jacobian cocycle and Bott-Thurston cocycle).
Fix $p\in[1,\infty)$. Define the multiplicative group $1$-cocycle
$\Theta_{p}(\varphi):=(\varphi^{\prime})^{1/p},$
and its logarithm
$\sigma_{p}(\varphi):=\log\Theta_{p}(\varphi)=\frac{1}{p}\log\varphi^{\prime}.$
Define the real-valued group $2$-cochain
$B_{p}(\varphi,\psi):=\int_{\mathbb{R}}\bigl(\sigma_{p}(\varphi)\circ\psi\bigr)\,(\sigma_{p}(\psi))^{\prime}\,dx=\int_{\mathbb{R}}\bigl(\sigma_{p}(\varphi)\circ\psi\bigr)\,d(\sigma_{p}(\psi)).$ (47)
Its infinitesimal cocycle is the alternating bilinear form
$\omega_{p}(u,v):=\frac{1}{p^{2}}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx,\qquad u,v\in\mathfrak{g}^{-\infty}.$ (48)
Theorem 5.7 (Rigidity of the Bott-Thurston class).
Let $p\in[1,\infty)$. Then the following hold.
  • (i)
    The maps $\Theta_{p}$ and $\sigma_{p}$ satisfy the cocycle identities
    $\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\,\Theta_{p}(\psi),\qquad\sigma_{p}(\varphi\circ\psi)=\sigma_{p}(\varphi)\circ\psi+\sigma_{p}(\psi).$
  • (ii)
    The cochain $B_{p}$ defined in (47) is a group $2$-cocycle. Its Lie algebra cocycle obtained by differentiating at the identity is $\omega_{p}$ in (48).
  • (iii)
    The form $\omega_{p}$ is a non-trivial Lie algebra $2$-cocycle and
    $\omega_{p}]=\frac{1}{p^{2}}[\omega]\in H^{2}(\mathfrak{g}^{-\infty};\mathbb{R}),\qquad\omega(u,v):=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx.$
    In particular the one-dimensional subspace spanned by $\omega_{p}]$ is independent of $p$.
Proof.
Part (i) is the Jacobian cocycle identity already established earlier in the paper (it is an immediate consequence of the chain rule), and the logarithmic identity follows by taking logarithms.
To show that $B_{p}$ is a group $2$-cocycle, write the coboundary as
$(\delta B_{p})(\varphi,\psi,\eta)=B_{p}(\psi,\eta)-B_{p}(\varphi\circ\psi,\eta)+B_{p}(\varphi,\psi\circ\eta)-B_{p}(\varphi,\psi).$
Using $\sigma_{p}(\psi\circ\eta)=\sigma_{p}(\psi)\circ\eta+\sigma_{p}(\eta)$ and bilinearity of the integral,
$B_{p}(\varphi,\psi\circ\eta)$ $=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)+\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\eta)),$
$B_{p}(\varphi\circ\psi,\eta)$ $=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\eta))+\int_{\mathbb{R}}(\sigma_{p}(\psi)\circ\eta)\,d(\sigma_{p}(\eta)).$
Substituting into $\delta B_{p}$, the terms involving $d(\sigma_{p}(\eta))$ cancel, as do $B_{p}(\psi,\eta)$ with the last term above, leaving
$(\delta B_{p})(\varphi,\psi,\eta)=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)-B_{p}(\varphi,\psi).$
Since $d(\sigma_{p}(\psi)\circ\eta)=(\sigma_{p}(\psi)^{\prime}\circ\eta)\eta^{\prime}\,dx$, a change of variables $y=\eta(x)$ gives
$\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi)(y)\,\sigma_{p}(\psi)^{\prime}(y)\,dy=B_{p}(\varphi,\psi),$
so $\delta B_{p}=0$.
For the infinitesimal cocycle, let $\varphi_{t}$ and $\psi_{s}$ be the flows of $u$ and $v$ with $\varphi_{0}=\psi_{0}=\mathrm{Id}$. Then $\varphi_{t}^{\prime}(x)=1+tu^{\prime}(x)+O(t^{2})$ and $\sigma_{p}(\varphi_{t})(x)=\frac{t}{p}u^{\prime}(x)+O(t^{2})$, while $(\sigma_{p}(\psi_{s}))^{\prime}(x)=\frac{s}{p}v^{\prime\prime}(x)+O(s^{2})$. Inserting into (47) yields
$B_{p}(\varphi_{t},\psi_{s})=\frac{ts}{p^{2}}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx+o(ts),$
hence $\partial_{t}\partial_{s}|_{0}B_{p}(\varphi_{t},\psi_{s})=\omega_{p}(u,v)$.
Finally, $\omega_{p}$ is alternating, which is immediate from integration by parts with vanishing boundary terms. Since $\omega$ is a non-trivial cocycle, $\omega_{p}=\frac{1}{p^{2}}\omega$ is non-trivial and $\omega_{p}]=\frac{1}{p^{2}}[\omega]$ in cohomology. ∎
Remark 5.8.
The $L^{p}$-Schwarzian does not satisfy a two-term cocycle law. Instead, its composition formula is the three-term identity in Proposition 4.20, and the cross-term is structural, in the sense that it disappears only in the affine case or in the projective limit $p\to\infty$. This highlights a qualitative difference with the metric results of the previous sections in which $L^{p}$-methods linearize the Finsler geometry, but they do not simplify the projective and cohomological structure encoded by Schwarzian-type cocycles.
5.4. The Bott-Thurston cocycle in $p$-root coordinates
Although the Bott-Thurston cocycle itself is classical (see for instance [8], [19], [26]), its expression in the $p$-root coordinates that arise from $L^{p}$ linearization seems not to be available in the literature in an explicit, coordinate-level form. The following formula is tailored to the present framework and will be used to make the dependence on $p$ transparent.
Proposition 5.9 (Explicit Bott cocycle formula in $p$-root coordinates).
Let $p\in[1,\infty)$ and write the $p$-root coordinate map $f=\Phi_{p}(\varphi):=p(\varphi^{\prime 1/p}-1)$. Then the normalized Bott cocycle $B$ of Proposition 5.5 satisfies
$B(\varphi,\psi)=\frac{p}{2}\int_{\mathbb{R}}\log\!\left(1+\frac{f_{\varphi}\circ\Phi_{p}^{-1}(f_{\psi})}{p}\right)\cdot\frac{(f_{\psi})^{\prime}}{1+f_{\psi}/p}\,\mathrm{d}x,$
where $f_{\varphi}=\Phi_{p}(\varphi)$ and $f_{\psi}=\Phi_{p}(\psi)$. Equivalently, the rescaled Bott-Thurston cocycle $B_{p}$ in (47) is given by
$B_{p}(\varphi,\psi)=\frac{1}{p}\int_{\mathbb{R}}\log\!\left(1+\frac{f_{\varphi}\circ\Phi_{p}^{-1}(f_{\psi})}{p}\right)\cdot\frac{(f_{\psi})^{\prime}}{1+f_{\psi}/p}\,\mathrm{d}x,$
using the identity $B_{p}=\frac{2}{p^{2}}B$.
Proof.
Let $\varphi,\psi$ be orientation-preserving diffeomorphisms in the class under consideration, so that $\varphi^{\prime},\psi^{\prime}>0$ and hence $\log\varphi^{\prime}$ and $\log\psi^{\prime}$ are well-defined. Moreover, by the decay/normalization assumptions on the group (e.g. $\psi^{\prime}-1$ rapidly decaying or compactly supported), the Bott-Thurston integral below converges. By definition,
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d(\log\psi^{\prime})=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,\frac{\psi^{\prime\prime}}{\psi^{\prime}}\,dx.$
Now set $f_{\varphi}=\Phi_{p}(\varphi)$ and $f_{\psi}=\Phi_{p}(\psi)$, so that
$f_{\varphi}=p(\varphi^{\prime 1/p}-1),\qquad f_{\psi}=p(\psi^{\prime 1/p}-1),$
and therefore
$\varphi^{\prime}=\left(1+\frac{f_{\varphi}}{p}\right)^{p},\qquad\psi^{\prime}=\left(1+\frac{f_{\psi}}{p}\right)^{p}.$
Composing the first identity with $\psi$ gives
$\log(\varphi^{\prime}\circ\psi)=\log\!\left(\left(1+\frac{f_{\varphi}\circ\psi}{p}\right)^{p}\right)=p\log\!\left(1+\frac{f_{\varphi}\circ\psi}{p}\right).$
Likewise,
$\frac{\psi^{\prime\prime}}{\psi^{\prime}}=\frac{d}{dx}\log\psi^{\prime}=\frac{d}{dx}\left(p\log\!\left(1+\frac{f_{\psi}}{p}\right)\right)=\frac{(f_{\psi})^{\prime}}{1+f_{\psi}/p}.$
Substituting these expressions into the defining formula for $B$ yields
$B(\varphi,\psi)=\frac{p}{2}\int_{\mathbb{R}}\log\!\left(1+\frac{f_{\varphi}\circ\psi}{p}\right)\frac{(f_{\psi})^{\prime}}{1+f_{\psi}/p}\,dx.$
Finally, since $f_{\psi}=\Phi_{p}(\psi)$ implies $\psi=\Phi_{p}^{-1}(f_{\psi})$, we may rewrite $f_{\varphi}\circ\psi$ as $f_{\varphi}\circ\Phi_{p}^{-1}(f_{\psi})$, giving the first displayed formula. The second formula follows from $B_{p}=\frac{2}{p^{2}}B$. ∎
Remark 5.10 (Normalization and notation).
The paper uses two related cocycles: the normalized Bott cocycle $B$ (with the prefactor $1/2$) integrating the Virasoro-Bott extension in Proposition 5.5, and the $p$-rescaled Bott-Thurston cocycle $B_{p}$ in (47) adapted to the $L^{p}$-root logarithmic cocycle $\sigma_{p}$. Proposition 5.9 records both formulas in $\Phi_{p}$-coordinates in a way that makes their scaling relation explicit.
5.5. Gerbes and categorical lifting
The central extension $\widehat{\operatorname{Diff}}^{-\infty}(\mathbb{R})\to\operatorname{Diff}^{-\infty}(\mathbb{R})$ can be encoded categorically as a banded gerbe: the cocycle $B$ appears as the obstruction to choosing globally compatible lifts of transition data. We only need the concrete Čech presentation.
Definition 5.11 (Smooth $A$-gerbe via Čech data).
Let $M$ be a (Fréchet) manifold and $A$ an abelian Lie group (in this section $A=\mathbb{R}$ or $A=U(1)$). Fix an open cover $\{U_{\alpha}\}$ of $M$. A smooth $A$-gerbe on $M$ is represented by smooth maps
$g_{\alpha\beta\gamma}\colon U_{\alpha\beta\gamma}:=U_{\alpha}\cap U_{\beta}\cap U_{\gamma}\to A$
satisfying the Čech $2$-cocycle identity on quadruple overlaps:
$g_{\beta\gamma\delta}\,g_{\alpha\gamma\delta}^{-1}\,g_{\alpha\beta\delta}\,g_{\alpha\beta\gamma}^{-1}=1\qquad\text{on }U_{\alpha\beta\gamma\delta}.$
Two cocycles $(g_{\alpha\beta\gamma})$ and $(g^{\prime}_{\alpha\beta\gamma})$ define equivalent gerbes if there exist smooth $A$-valued 11-cochains $h_{\alpha\beta}\colon U_{\alpha\beta}\to A$ such that $g^{\prime}_{\alpha\beta\gamma}=g_{\alpha\beta\gamma}\,h_{\beta\gamma}\,h_{\alpha\gamma}^{-1}\,h_{\alpha\beta}$.
Definition 5.12 (Lifting gerbe of a central extension).
Let $1\to A\to\widehat{G}\xrightarrow{\pi}G\to 1$ be a central extension of Lie groups and let $P\to X$ be a principal $G$-bundle with transition functions $g_{\alpha\beta}\colon U_{\alpha\beta}\to G$ on a cover $\{U_{\alpha}\}$. Choose local lifts $\widehat{g}_{\alpha\beta}\colon U_{\alpha\beta}\to\widehat{G}$ with $\pi(\widehat{g}_{\alpha\beta})=g_{\alpha\beta}$. On triple overlaps $U_{\alpha\beta\gamma}$, the product
$\widehat{g}_{\alpha\beta}\,\widehat{g}_{\beta\gamma}\,\widehat{g}_{\gamma\alpha}$
[…]
Proposition 5.13 (Čech cocycle induced by a group $2$-cocycle).
Assume $\widehat{G}=G\times A$ as a manifold and the multiplication is
$(g,a)\cdot(h,b)=(gh,ab\,\kappa(g,h))$
for a normalized smooth group $2$-cocycle $\kappa\colon G\times G\to A$. Let $P\to X$ be a principal $G$-bundle with transition functions $g_{\alpha\beta}$ and choose the canonical lifts $\widehat{g}_{\alpha\beta}:=(g_{\alpha\beta},1)$. Then the lifting gerbe cocycle is
$g_{\alpha\beta\gamma}=\kappa(g_{\alpha\beta},g_{\beta\gamma})\qquad\text{on }U_{\alpha\beta\gamma}.$
Proof of Proposition 5.13.
Let $g_{\alpha\beta}\colon U_{\alpha\beta}\to G$ be transition functions of a principal $G$-bundle, so $g_{\alpha\beta}g_{\beta\gamma}=g_{\alpha\gamma}$ on triple overlaps. Assume $\widehat{G}=G\times A$ with multiplication
$(g,a)\cdot(h,b)=(gh,ab\,\kappa(g,h)),$
where $\kappa\colon G\times G\to A$ is a normalized group $2$-cocycle, i.e.
$\kappa(g,\operatorname{id})=\kappa(\operatorname{id},g)=1,\qquad\kappa(h,k)\,\kappa(g,hk)=\kappa(gh,k)\,\kappa(g,h).$
Choose the canonical lifts $\widehat{g}_{\alpha\beta}:=(g_{\alpha\beta},1)$. Then on $U_{\alpha\beta\gamma}$,
$\widehat{g}_{\alpha\beta}\widehat{g}_{\beta\gamma}=(g_{\alpha\beta}g_{\beta\gamma},\,\kappa(g_{\alpha\beta},g_{\beta\gamma}))=(g_{\alpha\gamma},\,\kappa(g_{\alpha\beta},g_{\beta\gamma}))=\widehat{g}_{\alpha\gamma}\cdot(\operatorname{id},\kappa(g_{\alpha\beta},g_{\beta\gamma})).$
Thus the associated Čech $2$-cochain is
$g_{\alpha\beta\gamma}=\kappa(g_{\alpha\beta},g_{\beta\gamma}).$
To verify the Čech cocycle condition on $U_{\alpha\beta\gamma\delta}$, compute using $g_{\beta\delta}=g_{\beta\gamma}g_{\gamma\delta}$ and $g_{\alpha\delta}=g_{\alpha\beta}g_{\beta\delta}$:
$g_{\beta\gamma\delta}\,g_{\alpha\gamma\delta}^{-1}\,g_{\alpha\beta\delta}\,g_{\alpha\beta\gamma}^{-1}$ $=\kappa(g_{\beta\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\gamma},g_{\gamma\delta})^{-1}\,\kappa(g_{\alpha\beta},g_{\beta\delta})\,\kappa(g_{\alpha\beta},g_{\beta\gamma})^{-1}.$
Apply the group cocycle identity to the triple $(g_{\alpha\beta},g_{\beta\gamma},g_{\gamma\delta})$:
$\kappa(g_{\beta\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\beta},g_{\beta\gamma}g_{\gamma\delta})=\kappa(g_{\alpha\beta}g_{\beta\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\beta},g_{\beta\gamma}).$
Using $g_{\beta\gamma}g_{\gamma\delta}=g_{\beta\delta}$ and $g_{\alpha\beta}g_{\beta\gamma}=g_{\alpha\gamma}$, this becomes
$\kappa(g_{\beta\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\beta},g_{\beta\delta})=\kappa(g_{\alpha\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\beta},g_{\beta\gamma}),$
and rearranging yields precisely
$\kappa(g_{\beta\gamma},g_{\gamma\delta})\,\kappa(g_{\alpha\gamma},g_{\gamma\delta})^{-1}\,\kappa(g_{\alpha\beta},g_{\beta\delta})\,\kappa(g_{\alpha\beta},g_{\beta\gamma})^{-1}=1.$
Hence $(g_{\alpha\beta\gamma})$ is a Čech $2$-cocycle, as claimed. ∎
Theorem 5.14 (Gerbe class and $p$-rigidity).
Let $G=\operatorname{Diff}^{-\infty}(\mathbb{R})$ and let $B$ be the Bott cocycle in Proposition 5.5. Fix a real parameter $c$ and consider the $U(1)$-valued group cocycle
$\kappa_{c}(\varphi,\psi):=\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr).$
Then:
  1. (i)
    The cocycle $\kappa_{c}$ defines a central $U(1)$-extension of $G$ and hence, for any principal $G$-bundle $P\to X$, a lifting gerbe $\mathcal{G}_{c}(P)$ on $X$ in the sense of Definition 5.12.
  2. (ii)
    Replacing $B$ by the $p$-rescaled cocycle $B_{p}$ replaces $\kappa_{c}$ by $\kappa_{c,p}(\varphi,\psi):=\exp\!\bigl(2\pi i\,c\,B_{p}(\varphi,\psi)\bigr)$ and hence scales the resulting lifting gerbe class by the same factor in continuous cohomology.
  3. (iii)
    Infinitesimally, the corresponding Lie algebra class scales as $\omega_{p}]=p^{-2}[\omega]$ as in Theorem 5.7(iii).
Proof.
(i) Since $B$ is a normalized real-valued group $2$-cocycle on $G$ (Proposition 5.5), the map $\kappa_{c}(\varphi,\psi)=\exp(2\pi ic\,B(\varphi,\psi))$ is a normalized $U(1)$-valued group $2$-cocycle:
$\kappa_{c}(\operatorname{id},\psi)=\kappa_{c}(\varphi,\operatorname{id})=1,\qquad\delta\kappa_{c}=1,$
because $\delta\kappa_{c}=\exp(2\pi ic\,\delta B)$ and $\delta B=0$.
Define $\widehat{G}_{c}:=G\times U(1)$ as a manifold with multiplication
$(\varphi,z)\cdot(\psi,w):=(\varphi\circ\psi,\;zw\,\kappa_{c}(\varphi,\psi)).$ (49)
Associativity is equivalent to the cocycle identity for $\kappa_{c}$. Indeed, expanding $((\varphi,z)\cdot(\psi,w))\cdot(\eta,r)$ and $(\varphi,z)\cdot((\psi,w)\cdot(\eta,r))$ using (49) shows that the $G$-components coincide and the $U(1)$-components agree precisely when $\kappa_{c}(\varphi,\psi)\kappa_{c}(\varphi\psi,\eta)=\kappa_{c}(\psi,\eta)\kappa_{c}(\varphi,\psi\eta)$. Normalization implies that $(\operatorname{id},1)$ is the identity element. The subset $\{(\operatorname{id},z):z\in U(1)\}$ is central since the factor $\kappa_{c}(\operatorname{id},\cdot)=\kappa_{c}(\cdot,\operatorname{id})=1$, and the projection $\pi(\varphi,z)=\varphi$ is a surjective group homomorphism with kernel identified with $U(1)$. Thus we obtain a central extension
$1\longrightarrow U(1)\longrightarrow\widehat{G}_{c}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}G\longrightarrow 1.$
Given a principal $G$-bundle $P\to X$ with transition functions $g_{\alpha\beta}$ on a cover $\{U_{\alpha}\}$, choose the canonical lifts $\widehat{g}_{\alpha\beta}:=(g_{\alpha\beta},1)$. On triple overlaps $U_{\alpha\beta\gamma}$ one has $g_{\alpha\beta}g_{\beta\gamma}=g_{\alpha\gamma}$, hence
$\widehat{g}_{\alpha\beta}\widehat{g}_{\beta\gamma}=(g_{\alpha\gamma},\kappa_{c}(g_{\alpha\beta},g_{\beta\gamma}))=\widehat{g}_{\alpha\gamma}\cdot(\operatorname{id},\kappa_{c}(g_{\alpha\beta},g_{\beta\gamma})).$
Therefore the induced Čech $2$-cocycle is $g_{\alpha\beta\gamma}=\kappa_{c}(g_{\alpha\beta},g_{\beta\gamma})$, which satisfies the Čech cocycle condition exactly because $\delta\kappa_{c}=1$. This Čech class is the lifting gerbe $\mathcal{G}_{c}(P)$ of Definition 5.12.
(ii) By definition, we expand
$B_{p}(\varphi,\psi)=\frac{1}{p^{2}}\int\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})=\frac{2}{p^{2}}\,B(\varphi,\psi),$
since $B(\varphi,\psi)=\frac{1}{2}\int\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})$. Fix $c\in\mathbb{R}$ and set
$\kappa_{c}(\varphi,\psi):=\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr),\qquad\kappa_{c,p}(\varphi,\psi):=\exp\!\bigl(2\pi i\,c\,B_{p}(\varphi,\psi)\bigr).$
Using $B_{p}=\frac{2}{p^{2}}B$ we obtain the pointwise identity
$\kappa_{c,p}(\varphi,\psi)=\exp\!\Bigl(2\pi i\,c\,\frac{2}{p^{2}}B(\varphi,\psi)\Bigr)=\kappa_{\,2c/p^{2}}(\varphi,\psi).$
Therefore, the lifting gerbe defined using $\kappa_{c,p}$ is exactly the same gerbe as the one defined using $\kappa_{2c/p^{2}}$, i.e. the class scales by $2/p^{2}$ in the $c$-parameter. In particular, in the notation $\mathcal{G}_{c}(P)$ of part (i),
$\mathcal{G}_{c,p}(P)=\mathcal{G}_{2c/p^{2}}(P).$
(iii) is exactly Theorem 5.7(iii). ∎
5.6. Statistical lifting gerbe on the density quotient
Let $\operatorname{Aff}(\mathbb{R})$ denote the affine subgroup of $\operatorname{Diff}^{-\infty}(\mathbb{R})$. In the decay-controlled setting used throughout the paper, $\operatorname{Aff}(\mathbb{R})$ is precisely the vanishing locus of the Schwarzian and therefore the natural subgroup for passing from diffeomorphisms to densities.
Proposition 5.15 (Vanishing of the Bott cocycle on affine elements).
For every $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and every $\alpha\in\operatorname{Aff}(\mathbb{R})$,
$B(\varphi,\alpha)=0\qquad\text{and}\qquad B(\alpha,\varphi)=0.$
Proof.
Write $\alpha(x)=ax+b$ with $a>0$. Then $\alpha^{\prime}\equiv a$ is constant, hence $d\log(\alpha^{\prime})=0$ and therefore
$B(\varphi,\alpha)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\alpha)\,d\log(\alpha^{\prime})=0.$
For $B(\alpha,\varphi)$ we use that $\log(\alpha^{\prime}\circ\varphi)\equiv\log a$ is constant:
$B(\alpha,\varphi)=\frac{1}{2}\int_{\mathbb{R}}\log a\;d\log(\varphi^{\prime})=\frac{\log a}{2}\Bigl[\log(\varphi^{\prime})\Bigr]_{-\infty}^{+\infty}=0,$
since $\varphi^{\prime}-1$ decays and hence $\log(\varphi^{\prime})\to 0$ at $\pm\infty$. ∎
Let $|\Omega|_{\infty,1}(\mathbb{R})=\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$ denote the density quotient used in the paper.
Definition 5.16 (Transformation groupoid).
Let $G=\operatorname{Diff}^{-\infty}(\mathbb{R})$ act on $|\Omega|_{\infty,1}(\mathbb{R})$ by left composition. The associated action groupoid $G\ltimes|\Omega|_{\infty,1}(\mathbb{R})$ has object space $|\Omega|_{\infty,1}(\mathbb{R})$ and morphisms
$\mathrm{Mor}(G\ltimes|\Omega|_{\infty,1}(\mathbb{R}))=\{(\varphi,\mu):\ \varphi\in G,\ \mu\in|\Omega|_{\infty,1}(\mathbb{R})\},$
with source and target maps $s(\varphi,\mu)=\mu$ and $t(\varphi,\mu)=\varphi\cdot\mu$, and composition
$(\varphi,\psi\cdot\mu)\circ(\psi,\mu)=(\varphi\circ\psi,\mu).$
Definition 5.17 (Bott gerbe on the action groupoid).
Fix $c\in\mathbb{R}$ and define a $U(1)$-valued function on composable pairs of morphisms by
$\kappa_{c}\bigl((\varphi,\psi\cdot\mu),(\psi,\mu)\bigr):=\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr).$
The corresponding $U(1)$-banded gerbe on the groupoid $G\ltimes|\Omega|_{\infty,1}(\mathbb{R})$ is called the Bott gerbe at charge $c$.
Proposition 5.18 (Groupoid cocycle identity).
The function $\kappa_{c}$ in Definition 5.17 satisfies the groupoid $2$-cocycle identity on composable triples: for all $\varphi,\psi,\eta\in G$ and $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$,
$\kappa_{c}\bigl((\varphi,\psi\eta\cdot\mu),(\psi\eta,\mu)\bigr)\,\kappa_{c}\bigl((\psi,\eta\cdot\mu),(\eta,\mu)\bigr)=\kappa_{c}\bigl((\varphi\psi,\eta\cdot\mu),(\eta,\mu)\bigr)\,\kappa_{c}\bigl((\varphi,\psi\eta\cdot\mu),(\psi,\eta\cdot\mu)\bigr).$
Proof.
By definition, the groupoid cocycle $\kappa_{c}$ on the action groupoid $G\ltimes|\Omega|_{\infty,1}(\mathbb{R})$ is
$\kappa_{c}\bigl((\varphi,\psi\cdot\mu),(\psi,\mu)\bigr)=\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr),$
which depends only on the pair $(\varphi,\psi)$ and not on $\mu$.
For a composable triple
$(\varphi,\psi\eta\cdot\mu),\quad(\psi,\eta\cdot\mu),\quad(\eta,\mu),$
the groupoid $2$-cocycle condition is
$\kappa_{c}\bigl((\varphi,\psi\eta\cdot\mu),(\psi\eta,\mu)\bigr)\,\kappa_{c}\bigl((\psi,\eta\cdot\mu),(\eta,\mu)\bigr)$
$\qquad=\kappa_{c}\bigl((\varphi\psi,\eta\cdot\mu),(\eta,\mu)\bigr)\,\kappa_{c}\bigl((\varphi,\psi\eta\cdot\mu),(\psi,\eta\cdot\mu)\bigr).$
Substituting the definition of $\kappa_{c}$ reduces this to
$\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi\eta)\bigr)\,\exp\!\bigl(2\pi i\,c\,B(\psi,\eta)\bigr)=\exp\!\bigl(2\pi i\,c\,B(\varphi\psi,\eta)\bigr)\,\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr),$
which is exactly the exponential of the Bott cocycle identity
$B(\varphi,\psi)+B(\varphi\circ\psi,\eta)=B(\psi,\eta)+B(\varphi,\psi\circ\eta).$
Hence $\kappa_{c}$ satisfies the groupoid cocycle condition. ∎
5.7. Cumulant transgression and Koopman central extensions
We now exhibit an explicit $1$-cochain on the action groupoid whose groupoid coboundary is the Bott cocycle. This provides a canonical transgression functional built from the score of a density.
Definition 5.19 (Score).
Let $\mu$ be a smooth positive density on $\mathbb{R}$ in the class defining $|\Omega|_{\infty,1}(\mathbb{R})$, and write $\mu=\rho\,dx$ in the standard coordinate. The score of $\mu$ is
$s_{\mu}:=\partial_{x}\log\rho.$
Lemma 5.20 (Score transformation law).
Let $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and let $\mu=\rho\,dx$ be as above. Then $\psi_{*}\mu=\rho_{\psi_{*}\mu}\,dx$ satisfies
$s_{\psi_{*}\mu}(x)=\bigl(s_{\mu}\circ\psi^{-1}\bigr)(x)\,(\psi^{-1})^{\prime}(x)+\frac{(\psi^{-1})^{\prime\prime}(x)}{(\psi^{-1})^{\prime}(x)}.$
Proof.
By the change-of-variables formula,
$\rho_{\psi_{*}\mu}(x)=\rho(\psi^{-1}(x))\,(\psi^{-1})^{\prime}(x).$
Taking logarithms and differentiating gives
$s_{\psi_{*}\mu}(x)$ $=\partial_{x}\log\rho(\psi^{-1}(x))+\partial_{x}\log(\psi^{-1})^{\prime}(x)$
$=\bigl(\partial_{x}\log\rho\bigr)(\psi^{-1}(x))\,(\psi^{-1})^{\prime}(x)+\frac{(\psi^{-1})^{\prime\prime}(x)}{(\psi^{-1})^{\prime}(x)},$
which is the claimed identity. ∎
Definition 5.21 (Cumulant transgression functional).
Fix $c\in\mathbb{R}$. For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and a density $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s_{\mu}$, define
$\mathcal{A}_{c}(\varphi;\mu):=\frac{c}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}(x))\,s_{\mu}(x)\,dx.$ (50)
Theorem 5.22 (Bott cocycle as a groupoid coboundary).
For all $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and all $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$,
$\mathcal{A}_{c}(\varphi;\psi_{*}\mu)\;+\;\mathcal{A}_{c}(\psi;\mu)\;-\;\mathcal{A}_{c}(\varphi\circ\psi;\mu)=-c\,B(\varphi,\psi).$ (51)
Equivalently, $\exp(2\pi i\,\mathcal{A}_{c})$ is a groupoid $1$-cochain whose coboundary is the Bott cocycle at charge $c$.
Proof.
Write $\mu=\rho\,dx$ and use Lemma 5.20:
$\mathcal{A}_{c}(\varphi;\psi_{*}\mu)$ $=\frac{c}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}(x))\,s_{\psi_{*}\mu}(x)\,dx$
$=\frac{c}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}(x))\Bigl[(s_{\mu}\circ\psi^{-1})(x)\,(\psi^{-1})^{\prime}(x)\Bigr]\,dx\;+\;\frac{c}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}(x))\,\frac{(\psi^{-1})^{\prime\prime}(x)}{(\psi^{-1})^{\prime}(x)}\,dx$
$=:\frac{c}{2}\,(I_{1}+I_{2}).$
For $I_{1}$, change variables $x=\psi(y)$ so that $dx=\psi^{\prime}(y)\,dy$ and $(\psi^{-1})^{\prime}(\psi(y))=1/\psi^{\prime}(y)$. Then
$I_{1}=\int_{\mathbb{R}}\log(\varphi^{\prime}(\psi(y)))\,s_{\mu}(y)\,dy=\int_{\mathbb{R}}\log((\varphi^{\prime}\circ\psi)(y))\,s_{\mu}(y)\,dy.$
Hence
$\frac{c}{2}I_{1}+\mathcal{A}_{c}(\psi;\mu)=\frac{c}{2}\int_{\mathbb{R}}\bigl(\log(\varphi^{\prime}\circ\psi)+\log(\psi^{\prime})\bigr)\,s_{\mu}\,dy=\mathcal{A}_{c}(\varphi\circ\psi;\mu),$
using $\log((\varphi\circ\psi)^{\prime})=\log(\varphi^{\prime}\circ\psi)+\log(\psi^{\prime})$. Therefore the left-hand side of (51) reduces to $\frac{c}{2}I_{2}$.
For $I_{2}$, again set $x=\psi(y)$. Using $(\psi^{-1})^{\prime\prime}(\psi(y))/(\psi^{-1})^{\prime}(\psi(y))=-\psi^{\prime\prime}(y)/\psi^{\prime}(y)^{2}$, we obtain
$I_{2}$ $=\int_{\mathbb{R}}\log(\varphi^{\prime}(\psi(y)))\left(-\frac{\psi^{\prime\prime}(y)}{\psi^{\prime}(y)^{2}}\right)\psi^{\prime}(y)\,dy=-\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)(y)\,\frac{\psi^{\prime\prime}(y)}{\psi^{\prime}(y)}\,dy$
$=-\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}).$
By definition of the normalized Bott cocycle,
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}),$
so $I_{2}=-2B(\varphi,\psi)$ and therefore $\frac{c}{2}I_{2}=-c\,B(\varphi,\psi)$, which is (51). ∎
Definition 5.23 (Cumulant-Koopman central extension).
Fix $c\in\mathbb{R}$. Define the (smooth) central extension $\widehat{G}_{c}$ of $G=\operatorname{Diff}^{-\infty}(\mathbb{R})$ by $\mathbb{R}$ with multiplication
$(\varphi,a)\cdot(\psi,b):=\bigl(\varphi\circ\psi,\;a+b+c\,B(\varphi,\psi)\bigr),\qquad a,b\in\mathbb{R}.$
Definition 5.24 (Koopman field).
For $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ and $\varphi\in G$, define the Koopman transport
$\mathcal{K}_{\varphi}^{\mu}\colon L^{2}(\mu)\to L^{2}(\varphi_{*}\mu),\qquad(\mathcal{K}_{\varphi}^{\mu}f)(x):=f(\varphi^{-1}(x)).$
This is the standard pullback isometry between $L^{2}$-fibers; analytic properties and the corresponding operators on the Orlicz-Sobolev fibers are treated later in the manuscript (see the Koopman discussion in §9.5).
Theorem 5.25 (Central extension absorbs the Bott multiplier).
Fix $c\in\mathbb{R}$ and define the cumulant-gauged Koopman field by
$\mathcal{K}_{\varphi,c}^{\mu}:=\exp\!\bigl(-2\pi i\,\mathcal{A}_{c}(\varphi;\mu)\bigr)\,\mathcal{K}_{\varphi}^{\mu}\;:\;L^{2}(\mu)\to L^{2}(\varphi_{*}\mu).$ (52)
Then:
  1. (i)
    The family $\mathcal{K}_{\varphi,c}^{\mu}$ is projectively multiplicative with multiplier $\exp(2\pi i\,c\,B)$ in the sense that
    $\mathcal{K}_{\varphi,c}^{\psi_{*}\mu}\circ\mathcal{K}_{\psi,c}^{\mu}=\exp\!\bigl(2\pi i\,c\,B(\varphi,\psi)\bigr)\,\mathcal{K}_{\varphi\circ\psi,c}^{\mu}.$ (53)
  2. (ii)
    Defining, for $(\varphi,a)\in\widehat{G}_{c}$,
    $\widehat{\mathcal{K}}_{(\varphi,a)}^{\mu}:=\exp(2\pi i\,a)\,\mathcal{K}_{\varphi,c}^{\mu},$
    yields a strictly multiplicative action of $\widehat{G}_{c}$ on the $L^{2}$-field: for all composable pairs,
    $\widehat{\mathcal{K}}_{(\varphi,a)}^{\psi_{*}\mu}\circ\widehat{\mathcal{K}}_{(\psi,b)}^{\mu}=\widehat{\mathcal{K}}_{(\varphi,a)\cdot(\psi,b)}^{\mu}.$
Proof.
Since $\mathcal{K}_{\varphi}^{\mu}$ is functorial on the action groupoid,
$\mathcal{K}_{\varphi}^{\psi_{*}\mu}\circ\mathcal{K}_{\psi}^{\mu}=\mathcal{K}_{\varphi\circ\psi}^{\mu}.$
Using (52), we compute
$\mathcal{K}_{\varphi,c}^{\psi_{*}\mu}\circ\mathcal{K}_{\psi,c}^{\mu}$ $=\exp\!\bigl(-2\pi i\,\mathcal{A}_{c}(\varphi;\psi_{*}\mu)\bigr)\,\exp\!\bigl(-2\pi i\,\mathcal{A}_{c}(\psi;\mu)\bigr)\,\mathcal{K}_{\varphi\circ\psi}^{\mu}$
$=\exp\!\Bigl(-2\pi i\bigl[\mathcal{A}_{c}(\varphi;\psi_{*}\mu)+\mathcal{A}_{c}(\psi;\mu)-\mathcal{A}_{c}(\varphi\circ\psi;\mu)\bigr]\Bigr)\,\mathcal{K}_{\varphi\circ\psi,c}^{\mu}.$
By Theorem 5.22, the bracket equals $c\,B(\varphi,\psi)$, giving (53).
For (ii), multiply (53) by $\exp(2\pi i(a+b))$ and use Definition 5.23:
$\exp(2\pi i(a+b))\,\exp(2\pi icB(\varphi,\psi))=\exp(2\pi i(a+b+cB(\varphi,\psi))).$
This is exactly the phase appearing in $\widehat{\mathcal{K}}_{(\varphi,a)\cdot(\psi,b)}^{\mu}$, proving strict multiplicativity. ∎
Remark 5.26 ($L^{p}$-scaled transgression).
In the $L^{p}$ framework, it is natural to rescale the transgression functional to match $B_{p}$. Define
$\mathcal{A}_{c,p}(\varphi;\mu):=\frac{2}{p^{2}}\,\mathcal{A}_{c}(\varphi;\mu)=\frac{c}{p^{2}}\int_{\mathbb{R}}\log(\varphi^{\prime})\,s_{\mu}\,dx=\frac{c}{p}\int_{\mathbb{R}}\sigma_{p}(\varphi)\,s_{\mu}\,dx.$
Then Theorem 5.22 scales to
$\mathcal{A}_{c,p}(\varphi;\psi_{*}\mu)+\mathcal{A}_{c,p}(\psi;\mu)-\mathcal{A}_{c,p}(\varphi\circ\psi;\mu)=-c\,B_{p}(\varphi,\psi),$
since $B_{p}=\frac{2}{p^{2}}B$ and the groupoid coboundary is linear.
6. The real Bers embedding
The Schwarzian cocycle assigns to each $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ a potential $\frac{1}{2}S(\varphi)\in W^{\infty,1}(\mathbb{R})$. This section studies the resulting map $\beta^{-\infty}$, the real Bers map, from three perspectives: (i) as an injective embedding into a Fréchet space of potentials, with the image characterized by the existence of a distinguished positive solution of a Sturm-Liouville equation; (ii) as the curvature correction governing Liouville transport of second-order operators; and (iii) as a local coordinate system on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ near the identity, constructed via the tame inverse function theorem. A canonical inverse reconstruction, energy and factorization identities for the associated Schrödinger operator, and a spectral transport theorem relating Bers potentials to the free Laplacian are also established.
6.1. Definition and injectivity
[…]
Definition 6.1.
The real Bers map is
$\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\beta^{-\infty}(\varphi)=Q(\varphi)=\frac{1}{2}S(\varphi).$
Theorem 6.2.
[…]
$\widetilde{\operatorname{Diff}}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\longrightarrow W^{\infty,1}(\mathbb{R}).$
Its image is
$\beta^{-\infty}\bigl(\mathrm{Diff}^{-\infty}(\mathbb{R})\bigr)=\left\{\frac{1}{2}\left(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}\right):f\in W^{\infty,1}(\mathbb{R})\right\}.$
Equivalently, $q$ lies in the image if and only if there exists $f\in W^{\infty,1}(\mathbb{R})$ such that $2q=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$.
Proof.
[…]
To prove injectivity on the quotient, suppose $\beta^{-\infty}(\varphi)=\beta^{-\infty}(\psi)$, i.e. $S(\varphi)=S(\psi)$. Apply the Schwarzian cocycle identity to $\varphi=(\varphi\circ\psi^{-1})\circ\psi$:
$S(\varphi)=\bigl(S(\varphi\circ\psi^{-1})\circ\psi\bigr)(\psi^{\prime})^{2}+S(\psi).$
Since $S(\varphi)=S(\psi)$, we obtain $\bigl(S(\varphi\circ\psi^{-1})\circ\psi\bigr)(\psi^{\prime})^{2}=0$, hence $S(\varphi\circ\psi^{-1})=0$. By the classical characterization of the kernel of the Schwarzian, this implies $\varphi\circ\psi^{-1}\in\mathrm{Aff}(\mathbb{R})$, equivalently $\varphi=A\circ\psi$ for some affine $A$. Thus $\beta^{-\infty}$ is injective on $\widetilde{\operatorname{Diff}}^{-\infty}(\mathbb{R})/\mathrm{Aff}(\mathbb{R})$.
For the image formula, let $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$ and set $f:=\log(\varphi^{\prime})\in W^{\infty,1}(\mathbb{R})$. A direct computation gives
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}=\bigl(f^{\prime\prime}+(f^{\prime})^{2}\bigr)-\frac{3}{2}(f^{\prime})^{2}=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2},$
hence
$\beta^{-\infty}(\varphi)=\frac{1}{2}S(\varphi)=\frac{1}{2}\left(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}\right),$
so $\beta^{-\infty}(\mathrm{Diff}^{-\infty}(\mathbb{R}))$ is contained in the displayed set.
Conversely, given any $f\in W^{\infty,1}(\mathbb{R})$, let $\varphi:=\Phi_{\infty}^{-1}(f)$ (so that $\log(\varphi^{\prime})=f$). Then the same computation yields $\beta^{-\infty}(\varphi)=\frac{1}{2}(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2})$, proving that every element of the displayed set is attained. This identifies the image. ∎
Lemma 6.3 (Asymptotic normalization removes the affine gauge).
If $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $\varphi=A\circ\psi$ for some $A\in\operatorname{Aff}(\mathbb{R})$, then $A=\mathrm{Id}$ and hence $\varphi=\psi$. Consequently, $\beta^{-\infty}$ is injective already on $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
Write $A(x)=ax+b$ with $a>0$. Then $(A\circ\psi)^{\prime}=a\,\psi^{\prime}$ so $a=(A\circ\psi)^{\prime}(+\infty)=\varphi^{\prime}(+\infty)$. By the defining asymptotics of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ we have $\varphi^{\prime}(+\infty)=\psi^{\prime}(+\infty)=1$, hence $a=1$. With $a=1$ we have $A(x)=x+b$ and $\varphi(x)-\psi(x)\equiv b$. Since $\varphi-\psi\in W^{\infty,1}(\mathbb{R})$ and $W^{\infty,1}(\mathbb{R})\subset L^{1}(\mathbb{R})$, this forces $b=0$. ∎
Remark 6.4 (Connection to $L^{p}$ linearization).
The proof shows that the Bers map factors as
$\beta^{-\infty}=\frac{1}{2}K_{\infty}\circ\Phi_{\infty},$
where $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is the $p=\infty$ isometric embedding from Corollary 2.9 and $K_{\infty}(f)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ is the one-dimensional projective curvature operator.
6.2. Affine action on potentials
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ define the affine action on potentials
$(\mathcal{A}_{\varphi}q)(x)=q(\varphi(x))\varphi^{\prime}(x)^{2}+Q(\varphi)(x).$
Introduce also the Liouville transport on functions
$(\mathcal{L}_{\varphi}f)(x):=(\varphi^{\prime}(x))^{-1/2}f(\varphi(x)).$ (54)
Proposition 6.5.
For $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and potentials $q$ the following hold.
  1. (i)
    $\mathcal{A}_{\varphi\circ\psi}=\mathcal{A}_{\psi}\circ\mathcal{A}_{\varphi}$.
  2. (ii)
    The Liouville transport intertwines the Sturm-Liouville operators by
    $L_{\mathcal{A}_{\varphi}q}\,\mathcal{L}_{\varphi}=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}\,L_{q},\qquad L_{q}=\partial_{x}^{2}+q.$ (55)
    Equivalently,
    $\mathcal{L}_{\varphi}\,L_{q}\,\mathcal{L}_{\varphi}^{-1}=(\varphi^{\prime})^{-2}\,L_{\mathcal{A}_{\varphi}q}.$ (56)
  3. (iii)
[…]
The orbit of 0 under $\mathcal{A}$ is precisely $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$.
Proof.
(i) Let $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and let $q$ be a smooth potential. By definition,
$(\mathcal{A}_{\varphi\circ\psi}q)(x)=q(\varphi(\psi(x)))\,(\varphi\circ\psi)^{\prime}(x)^{2}+Q(\varphi\circ\psi)(x).$
Since $(\varphi\circ\psi)^{\prime}=(\varphi^{\prime}\circ\psi)\,\psi^{\prime}$, the first term equals $q(\varphi(\psi(x)))\,(\varphi^{\prime}(\psi(x))\psi^{\prime}(x))^{2}$. On the other hand,
$(\mathcal{A}_{\psi}(\mathcal{A}_{\varphi}q))(x)=(\mathcal{A}_{\varphi}q)(\psi(x))\,\psi^{\prime}(x)^{2}+Q(\psi)(x),$
and expanding $(\mathcal{A}_{\varphi}q)(\psi(x))$ gives
$(\mathcal{A}_{\psi}(\mathcal{A}_{\varphi}q))(x)=q(\varphi(\psi(x)))\,(\varphi^{\prime}(\psi(x))\psi^{\prime}(x))^{2}+(Q(\varphi)\circ\psi)(x)\,\psi^{\prime}(x)^{2}+Q(\psi)(x).$
Thus $\mathcal{A}_{\varphi\circ\psi}=\mathcal{A}_{\psi}\circ\mathcal{A}_{\varphi}$ is equivalent to
$Q(\varphi\circ\psi)=(Q(\varphi)\circ\psi)\,(\psi^{\prime})^{2}+Q(\psi),$
which holds because $Q=\frac{1}{2}S$ and the Schwarzian cocycle identity gives $S(\varphi\circ\psi)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi)$.
For (ii), it suffices to verify (55) on $f\in C_{c}^{\infty}(\mathbb{R})$. Write
$g(x):=(\mathcal{L}_{\varphi}f)(x)=(\varphi^{\prime}(x))^{-1/2}f(\varphi(x))=\mathrm{e}^{-u(x)/2}(f\circ\varphi)(x),\qquad u=\log\varphi^{\prime}.$
Then
$\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)g=\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)\bigl(\mathrm{e}^{-u/2}(f\circ\varphi)\bigr)=\mathrm{e}^{-u/2}(f^{\prime}\circ\varphi)\,\varphi^{\prime}=\mathrm{e}^{u/2}(f^{\prime}\circ\varphi),$
and therefore
$\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)g=\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\bigl(\mathrm{e}^{u/2}(f^{\prime}\circ\varphi)\bigr)=\mathrm{e}^{u/2}(f^{\prime\prime}\circ\varphi)\,\varphi^{\prime}=(\varphi^{\prime})^{3/2}(f^{\prime\prime}\circ\varphi).$
Since $Q(\varphi)=\frac{1}{2}S(\varphi)=\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}$, the left-hand side equals $(\partial_{x}^{2}+Q(\varphi))g$. Adding the potential term yields
$L_{\mathcal{A}_{\varphi}q}g=\bigl(\partial_{x}^{2}+q(\varphi)\varphi^{\prime 2}+Q(\varphi)\bigr)g=(\varphi^{\prime})^{3/2}\bigl((f^{\prime\prime}+qf)\circ\varphi\bigr)=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}(L_{q}f),$
which is (55). The conjugation form (56) follows by right-multiplying by $\mathcal{L}_{\varphi}^{-1}$.
[…]
6.3. Miura factorization
Proposition 6.6.
Let $u=\log(\varphi^{\prime})$. Then
$L_{\beta^{-\infty}(\varphi)}=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\left(\partial_{x}+\frac{u^{\prime}}{2}\right).$
The function $y_{2}=\mathrm{e}^{-u/2}=(\varphi^{\prime})^{-1/2}$ is a positive solution of $L_{\beta^{-\infty}(\varphi)}y=0$ annihilated by $(\partial_{x}+\frac{u^{\prime}}{2})$.
Proof.
The proof is a direct computation. ∎
6.4. Fréchet structure and tame right inverse
Theorem 6.7.
The map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth. For $\varphi$ with $u=\log(\varphi^{\prime})$, the differential is
$D_{\varphi}\beta^{-\infty}(\delta\varphi)=\frac{1}{2}(\delta u^{\prime\prime}-u^{\prime}\delta u^{\prime}),\qquad\delta u=\frac{\delta\varphi^{\prime}}{\varphi^{\prime}}.$
The operator $\mathcal{D}_{u}(\delta u)=\delta u^{\prime\prime}-u^{\prime}\delta u^{\prime}$ admits a continuous tame right inverse $R_{u}\colon W^{k,1}(\mathbb{R})\to W^{k+2,1}(\mathbb{R})$ for every $k\geq 0$, depending smoothly on $u$ on bounded subsets.
Proof.
Smoothness of $\beta^{-\infty}$ follows from Proposition 4.2. The differential formula is immediate.
For the right inverse, note $\mathcal{D}_{u}(\delta u)=(\partial_{x}-u^{\prime})\partial_{x}\delta u$. Setting $w=\partial_{x}\delta u$ with $\delta u(-\infty)=0$, we have $(\partial_{x}-u^{\prime})w=v$. Multiplying by $\mathrm{e}^{-u}$ gives $\partial_{x}(\mathrm{e}^{-u}w)=\mathrm{e}^{-u}v$, hence
$w(x)=\mathrm{e}^{u(x)}\int_{-\infty}^{x}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y.$
Define
$(R_{u}v)(x)=\int_{-\infty}^{x}\mathrm{e}^{u(t)}\int_{-\infty}^{t}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y\,\mathrm{d}t.$
Then $\partial_{x}(R_{u}v)=w$ and $\mathcal{D}_{u}(R_{u}v)=v$.
For tame bounds, on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ we have uniform bounds $0<c\leq\mathrm{e}^{u}\leq C$. By Faà di Bruno, $\partial_{x}^{m}(\mathrm{e}^{u(t)-u(y)})$ is a sum of terms $\mathrm{e}^{u(t)-u(y)}P_{m}(u^{\prime},\ldots,u^{(m)})$ with $P_{m}$ polynomial. Since $u^{(j)}\in W^{1,1}\hookrightarrow L^{\infty}$ with $u^{(j)}(\pm\infty)=0$, all coefficients are bounded.
For $m\leq k+2$, the structure of $\partial_{x}^{m}(R_{u}v)$ is
$\partial_{x}^{m}(R_{u}v)=\sum_{j=0}^{m-2}a_{m,j}\partial_{x}^{j}v+b_{m}\partial_{x}(R_{u}v)$
where $a_{m,j},b_{m}$ are bounded by finitely many $\left\|u^{(\ell)}\right\|_{L^{\infty}}$. The term $\partial_{x}(R_{u}v)=w$ satisfies $\left\|w\right\|_{L^{1}}\leq C\left\|v\right\|_{L^{1}}$ by Grönwall. Thus
$\left\|\partial_{x}^{m}(R_{u}v)\right\|_{L^{1}}\leq C_{m}\left(\sum_{j=0}^{m-2}\left\|\partial_{x}^{j}v\right\|_{L^{1}}+\left\|v\right\|_{L^{1}}\right)\leq C_{m}^{\prime}\left\|v\right\|_{W^{k,1}},$
giving $\left\|R_{u}v\right\|_{W^{k+2,1}}\leq C\left\|v\right\|_{W^{k,1}}$.
Smooth dependence on $u$ follows from smoothness of $u\mapsto\mathrm{e}^{\pm u}$ on bounded sets and the linear structure of $R_{u}$. ∎
6.5. Spectral characterization of the image
Theorem 6.8.
A potential $q\in W^{\infty,1}(\mathbb{R})$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if $y^{\prime\prime}+qy=0$ admits linearly independent solutions $y_{1},y_{2}$ with $y_{2}>0$, $W(y_{1},y_{2})\equiv 1$, and $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. The diffeomorphism is $\varphi=y_{1}/y_{2}$.
Proof.
If $q=\beta^{-\infty}(\varphi)$, Proposition 4.25 gives $y_{2}=(\varphi^{\prime})^{-1/2}$, $y_{1}=\varphi(\varphi^{\prime})^{-1/2}$ satisfying the required properties, with $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$.
Conversely, given such $y_{1},y_{2}$, set $\varphi=y_{1}/y_{2}$. Then $\varphi^{\prime}=W(y_{1},y_{2})/y_{2}^{2}=y_{2}^{-2}>0$ and $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. The normalization at -$\infty$ is determined by integrating $\varphi^{\prime}-1$, giving $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. By Proposition 4.25, $S(\varphi)=2q$, so $\beta^{-\infty}(\varphi)=q$. ∎
6.6. Explicit inverse construction
The spectral characterization in Theorem 6.8 becomes a canonical inverse once the asymptotic normalization built into $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is used to remove any residual affine ambiguity.
Lemma 6.9 (Distinguished positive solution).
Let $q\in W^{\infty,1}(\mathbb{R})$ lie in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$. Then there exists a unique positive solution $y_{q}$ of
$y^{\prime\prime}+qy=0$
such that $y_{q}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Moreover $y_{q}(\pm\infty)=1$ and $y_{q}^{\prime}(\pm\infty)=0$.
Proof.
Existence follows from Theorem 6.8 by taking $y_{q}=y_{2}$.
To prove the limits, note that $y_{q}^{-2}-1\in W^{1,1}(\mathbb{R})$ implies $y_{q}^{-2}(x)\to 1$ as $x\to\pm\infty$. Since $y_{q}>0$, this gives $y_{q}(x)\to 1$ as $x\to\pm\infty$. Differentiating $y_{q}^{-2}$ gives $(y_{q}^{-2})^{\prime}=-2y_{q}^{\prime}/y_{q}^{3}\in L^{1}(\mathbb{R})$, and on $W^{1,1}$-bounded subsets we have uniform bounds $0<c\leq y_{q}\leq C$, hence $y_{q}^{\prime}\in L^{1}(\mathbb{R})$ and therefore $y_{q}^{\prime}(\pm\infty)=0$.
For uniqueness, let $y,\tilde{y}>0$ be solutions with $y^{-2}-1,\tilde{y}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Their Wronskian W(y,y~)W($\tilde{y}$) is constant. Using $y(\pm\infty)=\tilde{y}(\pm\infty)=1$ and $y^{\prime}(\pm\infty)=\tilde{y}^{\prime}(\pm\infty)=0$ we obtain $W(y,\tilde{y})=\lim_{x\to\pm\infty}(y\tilde{y}^{\prime}-y^{\prime}\tilde{y})=0$, hence $W\equiv 0$ and $y=c\,\tilde{y}$. Taking $x\to\pm\infty$ yields $c=1$. ∎
Theorem 6.10 (Canonical reconstruction formula).
Let $q\in W^{\infty,1}(\mathbb{R})$. Then $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the distinguished positive solution $y_{q}$ from Lemma 6.9 exists. In that case, the unique $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\beta^{-\infty}(\varphi)=q$ is given by
$\varphi^{\prime}(x)=y_{q}(x)^{-2},\qquad\varphi(x)=x+\int_{-\infty}^{x}\bigl(y_{q}(t)^{-2}-1\bigr)\,dt.$ (57)
Equivalently, with $u=\log\varphi^{\prime}$ one has $u=-2\log y_{q}\in W^{\infty,1}(\mathbb{R})$ and $\beta^{-\infty}(\varphi)=\frac{1}{2}\bigl(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\bigr)=q$.
Proof.
Assume first $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$. Then Lemma 6.9 provides $y_{q}>0$ with $y_{q}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Define $\varphi$ by (57). Then $\varphi^{\prime}>0$ and $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$, hence $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ by the defining asymptotics. Moreover, setting $u=\log\varphi^{\prime}=-2\log y_{q}$ and using $y_{q}^{\prime\prime}=-qy_{q}$, a direct computation yields
$\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)=\frac{1}{2}\left(-2\frac{y_{q}^{\prime\prime}}{y_{q}}+2\left(\frac{y_{q}^{\prime}}{y_{q}}\right)^{2}-\frac{1}{2}\cdot 4\left(\frac{y_{q}^{\prime}}{y_{q}}\right)^{2}\right)=-\frac{y_{q}^{\prime\prime}}{y_{q}}=q.$
Thus $\beta^{-\infty}(\varphi)=q$.
Conversely, if $y_{q}$ exists, then the construction above produces $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\beta^{-\infty}(\varphi)=q$, hence $q$ lies in the image. Uniqueness of $\varphi$ follows from Lemma 6.3. ∎
Remark 6.11 (Volterra form of the distinguished solution).
If $y_{q}$ is the distinguished solution, then integrating $y_{q}^{\prime\prime}=-qy_{q}$ twice from -$\infty$ gives the Volterra equation
$y_{q}(x)=1-\int_{-\infty}^{x}(x-t)\,q(t)\,y_{q}(t)\,dt,$ (58)
which uniquely characterizes $y_{q}$ among solutions with $y_{q}(-\infty)=1$, $y_{q}^{\prime}(-\infty)=0$.
6.7. Riccati-Miura characterization and energy identity
Proposition 6.12 (Riccati-Miura criterion with integral constraint).
Let $q\in W^{\infty,1}(\mathbb{R})$. The following are equivalent:
  1. (i)
    $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$.
  2. (ii)
    There exists $w\in W^{\infty,1}(\mathbb{R})$ with $w(\pm\infty)=0$ and
    $\int_{\mathbb{R}}w(x)\,dx=0$ (59)
    such that
    $q=w^{\prime}-w^{2}.$ (60)
  3. (iii)
    There exists a positive solution $y$ of $y^{\prime\prime}+qy=0$ with $y^{-2}-1\in W^{\infty,1}(\mathbb{R})$, in which case $w=-y^{\prime}/y$ satisfies (59) and (60).
Proof.
[…]
$q=\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)=\left(\frac{u^{\prime}}{2}\right)^{\prime}-\left(\frac{u^{\prime}}{2}\right)^{2}.$
Set $w=u^{\prime}/2$. Since $u\in W^{1,1}(\mathbb{R})$ we have $u(\pm\infty)=0$ and hence $\int_{\mathbb{R}}w=\frac{1}{2}(u(\infty)-u(-\infty))=0$.
[…]
$y(x):=\exp\left(-\int_{-\infty}^{x}w(t)\,dt\right).$
Then $y>0$, $y^{\prime}=-wy$, and $y^{\prime\prime}=(w^{2}-w^{\prime})y=-qy$, so $y^{\prime\prime}+qy=0$ and $w=-y^{\prime}/y$. Moreover,
$y^{-2}(x)=\exp\left(2\int_{-\infty}^{x}w(t)\,dt\right)=\exp(u(x)),\qquad u(x):=2\int_{-\infty}^{x}w(t)\,dt.$
Since $w\in W^{\infty,1}$, we have $u\in W^{\infty,1}$. The integral constraint (59) implies $u(\infty)=0$, hence $y^{-2}-1=\mathrm{e}^{u}-1\in W^{\infty,1}(\mathbb{R})$.
[…]
Proposition 6.13 (Energy identity and correct factorization).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $u=\log\varphi^{\prime}$ and $q=\beta^{-\infty}(\varphi)$. Then, with $w=u^{\prime}/2$, one has the energy identity
$\int_{\mathbb{R}}q(x)\,dx=-\int_{\mathbb{R}}w(x)^{2}\,dx=-\frac{1}{4}\int_{\mathbb{R}}(u^{\prime}(x))^{2}\,dx\leq 0.$ (61)
Moreover the Schrödinger operator
$H_{q}:=-\partial_{x}^{2}-q$ (62)
admits the exact nonnegative factorization
$H_{q}=\left(-\partial_{x}+w\right)\left(\partial_{x}+w\right)=A^{\ast}A,\qquad A:=\partial_{x}+w,$ (63)
and therefore $(H_{q}f,f)_{L^{2}}\geq 0$ on $C_{c}^{\infty}(\mathbb{R})$ (and on the natural self-adjoint domain).
Proof.
By Proposition 6.12 we have $q=w^{\prime}-w^{2}$ with $w(\pm\infty)=0$. Integrating gives $\int_{\mathbb{R}}q=\int_{\mathbb{R}}w^{\prime}-\int_{\mathbb{R}}w^{2}=-\int_{\mathbb{R}}w^{2}$, which is (61).
For the factorization, compute for $f\in C_{c}^{\infty}(\mathbb{R})$:
$\left(-\partial_{x}+w\right)\left(\partial_{x}+w\right)f=-f^{\prime\prime}-w^{\prime}f-wf^{\prime}+wf^{\prime}+w^{2}f=-f^{\prime\prime}-(w^{\prime}-w^{2})f=-f^{\prime\prime}-qf.$
Since $A^{\ast}=-\partial_{x}+w$ on $C_{c}^{\infty}(\mathbb{R})$, we obtain $H_{q}=A^{\ast}A$ and hence $(H_{q}f,f)=\|Af\|_{L^{2}}^{2}\geq 0$. ∎
6.8. Local Bers coordinates and transition maps
Theorem 6.14 (Local Bers coordinates).
There exists a neighborhood $U$ of $\mathrm{Id}$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and a neighborhood $V$ of 0 in $W^{\infty,1}(\mathbb{R})$ such that
$\beta^{-\infty}\colon U\longrightarrow V$
is a Fréchet diffeomorphism onto its image, with Fréchet smooth inverse. Equivalently, in the linearizing coordinate $u=\log\varphi^{\prime}$ the map
$\mathcal{B}(u):=\beta^{-\infty}(\Phi_{\infty}^{-1}(u))=\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)$
is a tame Fréchet local diffeomorphism near $u=0$.
Proof.
Work in the global coordinate $u=\log\varphi^{\prime}\in W^{\infty,1}(\mathbb{R})$ via the global chart $\Phi_{\infty}$. Then $\mathcal{B}\colon W^{\infty,1}\to W^{\infty,1}$ is Fréchet smooth and its differential at $u$ is
$D_{u}\mathcal{B}(h)=\frac{1}{2}\bigl(h^{\prime\prime}-u^{\prime}h^{\prime}\bigr)=\frac{1}{2}\,\mathcal{D}_{u}(h).$
By Theorem 6.7, $\mathcal{D}_{u}$ admits a continuous tame right inverse $R_{u}$ on bounded sets, hence $D_{u}\mathcal{B}$ has a continuous tame right inverse $\frac{1}{2}R_{u}$.
It remains to show that $D_{u}\mathcal{B}$ is injective on $W^{\infty,1}(\mathbb{R})$. If $D_{u}\mathcal{B}(h)=0$, then $\mathcal{D}_{u}(h)=h^{\prime\prime}-u^{\prime}h^{\prime}=0$, hence $(\mathrm{e}^{-u}h^{\prime})^{\prime}=0$. Thus $\mathrm{e}^{-u}h^{\prime}=c$ is constant. Since $h\in W^{\infty,1}(\mathbb{R})$ we have $h^{\prime}\in W^{\infty,1}(\mathbb{R})\subset L^{1}(\mathbb{R})$ and $h^{\prime}(\pm\infty)=0$, so $c=0$ and $h^{\prime}\equiv 0$. Because $h\in L^{1}(\mathbb{R})$, this forces $h\equiv 0$.
Therefore $D_{u}\mathcal{B}$ is a tame linear isomorphism between the model Fréchet spaces at each $u$ in a sufficiently small neighborhood of 0. Applying the Nash-Moser inverse function theorem in tame Fréchet spaces yields a neighborhood of 0 on which $\mathcal{B}$ is a tame local diffeomorphism, and transporting back by $\Phi_{\infty}$ gives the claimed local diffeomorphism property of $\beta^{-\infty}$ near $\mathrm{Id}$. ∎
Corollary 6.15 (Quantitative stability on the Bers chart).
For each $k\geq 0$ there exist neighborhoods $U_{k}\subset\operatorname{Diff}^{-\infty}(\mathbb{R})$ of $\mathrm{Id}$ and $V_{k}\subset W^{k,1}(\mathbb{R})$ of 0 and a constant $C_{k}>0$ such that if $\varphi,\psi\in U_{k}$ and $q_{\varphi}=\beta^{-\infty}(\varphi)$, $q_{\psi}=\beta^{-\infty}(\psi)\in V_{k}$, then
$\|\log\varphi^{\prime}-\log\psi^{\prime}\|_{W^{k+2,1}}\leq C_{k}\,\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}.$
In particular, $\|\varphi^{\prime}-\psi^{\prime}\|_{W^{k+2,1}}\leq C_{k}^{\prime}\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}$ and $\|\varphi-\psi\|_{W^{k+3,1}}\leq C_{k}^{\prime\prime}\|q_{\varphi}-q_{\psi}\|_{W^{k,1}}$.
Proof.
This is an immediate consequence of Theorem 6.14 together with the tame estimates in Theorem 6.7 for the right inverse of the linearization and the smoothness of $u\mapsto\Phi_{\infty}^{-1}(u)$ on bounded sets. ∎
Proposition 6.16 (Transition maps between Bers charts).
Fix $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Define the Bers chart at $\psi$ by
$\beta^{-\infty}_{\psi}(\varphi):=\mathcal{A}_{\psi^{-1}}\bigl(\beta^{-\infty}(\varphi)\bigr).$
Then $\beta^{-\infty}_{\psi}(\psi)=0$ and, on neighborhoods where these charts are defined, the transition map between charts at $\psi$ and $\chi$ is
$\beta^{-\infty}_{\chi}(\varphi)=\mathcal{A}_{\chi^{-1}\circ\psi}\bigl(\beta^{-\infty}_{\psi}(\varphi)\bigr).$
Proof.
Since $\beta^{-\infty}(\psi)=\mathcal{A}_{\psi}0$ and $\mathcal{A}_{\psi^{-1}}\circ\mathcal{A}_{\psi}=\mathcal{A}_{\psi\circ\psi^{-1}}=\mathcal{A}_{\mathrm{Id}}$, we have $\beta_{\psi}^{-\infty}(\psi)=0$. For the transition, use $\mathcal{A}_{\chi^{-1}}=\mathcal{A}_{\chi^{-1}\circ\psi}\circ\mathcal{A}_{\psi^{-1}}$ from Proposition 6.5(i). ∎
6.9. Spectral transport and corrected Liouville conjugation
The next result replaces the incorrect operator-level statement that omits the $(\varphi^{\prime})^{-2}$ prefactor and uses the wrong half-density weight. The correct transport is naturally formulated using the Liouville map (54).
Theorem 6.17 (Spectral transport).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q=\beta^{-\infty}(\varphi)=Q(\varphi)$. Define $\mathcal{L}_{\varphi}$ by (54). Then, on $C_{c}^{\infty}(\mathbb{R})$, one has the exact conjugation identity
$\mathcal{L}_{\varphi}\,\partial_{y}^{2}\,\mathcal{L}_{\varphi}^{-1}=(\varphi^{\prime})^{-2}\,\bigl(\partial_{x}^{2}+q(x)\bigr)=(\varphi^{\prime})^{-2}\,L_{q}.$ (64)
Equivalently,
$L_{q}\,\mathcal{L}_{\varphi}=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}\,\partial_{y}^{2}.$ (65)
Moreover, $\mathcal{L}_{\varphi}$ extends to a unitary map
$\mathcal{L}_{\varphi}\colon L^{2}(\mathbb{R},dy)\longrightarrow L^{2}(\mathbb{R},\varphi^{\prime}(x)^{2}\,dx),$
and therefore the self-adjoint realization of $(\varphi^{\prime})^{-2}L_{q}$ on $L^{2}(\mathbb{R},\varphi^{\prime 2}dx)$ is unitarily equivalent to the free Laplacian $\partial_{y}^{2}$ on $L^{2}(\mathbb{R},dy)$.
In particular, for each $\xi\in\mathbb{R}$ the distorted plane waves
$\psi_{\xi}^{\pm}(x):=(\varphi^{\prime}(x))^{-1/2}\,\mathrm{e}^{\pm i\xi\,\varphi(x)}$ (66)
satisfy the exact eigenvalue equation
$(\varphi^{\prime})^{-2}L_{q}\,\psi_{\xi}^{\pm}=-\xi^{2}\,\psi_{\xi}^{\pm},\qquad\text{equivalently}\qquad L_{q}\,\psi_{\xi}^{\pm}=-\xi^{2}\,(\varphi^{\prime})^{2}\,\psi_{\xi}^{\pm}.$
Proof.
Using Proposition 6.6 we write $L_{q}=(\partial_{x}-\frac{u^{\prime}}{2})(\partial_{x}+\frac{u^{\prime}}{2})$ with $u=\log\varphi^{\prime}$. Let $f\in C_{c}^{\infty}(\mathbb{R})$ and set $g=\mathcal{L}_{\varphi}f=\mathrm{e}^{-u/2}(f\circ\varphi)$. A direct computation gives
$\left(\partial_{x}+\frac{u^{\prime}}{2}\right)g=\left(\partial_{x}+\frac{u^{\prime}}{2}\right)\left(\mathrm{e}^{-u/2}f(\varphi(x))\right)=\mathrm{e}^{-u/2}\,f^{\prime}(\varphi(x))\,\varphi^{\prime}(x)=\mathrm{e}^{u/2}\,(f^{\prime}\circ\varphi).$
Applying the second factor yields
$L_{q}g=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\Bigl(\mathrm{e}^{u/2}(f^{\prime}\circ\varphi)\Bigr)=\mathrm{e}^{u/2}\,\varphi^{\prime}\,(f^{\prime\prime}\circ\varphi)=(\varphi^{\prime})^{3/2}\,(f^{\prime\prime}\circ\varphi).$
Since $\mathcal{L}_{\varphi}(\partial_{y}^{2}f)=(\varphi^{\prime})^{-1/2}(f^{\prime\prime}\circ\varphi)$, this is exactly
$L_{q}\,\mathcal{L}_{\varphi}f=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}(\partial_{y}^{2}f),$
which is (65). Conjugation form (64) follows by right-multiplying by $\mathcal{L}_{\varphi}^{-1}$.
For unitarity, compute for $f\in C_{c}^{\infty}(\mathbb{R})$:
$\|\mathcal{L}_{\varphi}f\|_{L^{2}(\varphi^{\prime 2}dx)}^{2}=\int_{\mathbb{R}}\left|(\varphi^{\prime})^{-1/2}f(\varphi(x))\right|^{2}\,\varphi^{\prime}(x)^{2}\,dx=\int_{\mathbb{R}}|f(\varphi(x))|^{2}\,\varphi^{\prime}(x)\,dx=\int_{\mathbb{R}}|f(y)|^{2}\,dy,$
by the change of variables $y=\varphi(x)$. The eigenfunction statement (66) is obtained by taking $f(y)=\mathrm{e}^{\pm i\xi y}$. ∎
Corollary 6.18 (Trivial scattering for the transported Laplacian).
Let $q=\beta^{-\infty}(\varphi)$. The operator $(\varphi^{\prime})^{-2}L_{q}$ on $L^{2}(\mathbb{R},\varphi^{\prime 2}dx)$ is unitarily equivalent to the free Laplacian on $L^{2}(\mathbb{R},dy)$ and therefore has purely absolutely continuous spectrum $(-\infty,0]$. In the distorted Fourier representation determined by (66), the operator is diagonal with no mode mixing which is equivalent to a trivial monodromy.
Proof.
This is immediate from Theorem 6.17 and the spectral theorem for $\partial_{y}^{2}$ on $L^{2}(\mathbb{R},dy)$. ∎
Lemma 6.19 (Unitary Koopman operator on half-densities).
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the operator
$(\mathcal{K}_{\varphi}f)(x):=\sqrt{\varphi^{\prime}(x)}\,f(\varphi(x))$
is unitary on $L^{2}(\mathbb{R},dx)$. Moreover $\mathcal{L}_{\varphi}=M_{\varphi^{\prime-1}}\mathcal{K}_{\varphi}$, where $M_{\varphi^{\prime-1}}$ denotes multiplication by $\varphi^{\prime-1}$.
Proof.
Unitarity is the change of variables $y=\varphi(x)$:
$\|\mathcal{K}_{\varphi}f\|_{L^{2}(dx)}^{2}=\int_{\mathbb{R}}\varphi^{\prime}(x)|f(\varphi(x))|^{2}\,dx=\int_{\mathbb{R}}|f(y)|^{2}\,dy.$
The factorization $\mathcal{L}_{\varphi}=M_{\varphi^{\prime-1}}\mathcal{K}_{\varphi}$ is immediate from definitions. ∎
7. Scattering theory and the nonlinear Fourier coordinate
The Sturm-Liouville characterization of the real Bers image in Theorem 6.8 admits a refinement in terms of one-dimensional scattering. For a real potential $q\in W^{\infty,1}(\mathbb{R})$ we consider the Sturm-Liouville operator $\mathcal{L}_{q}=\partial_{x}^{2}+q$ from Proposition 6.5 and the associated Schrödinger operator
$H_{q}:=-\mathcal{L}_{q}=-\partial_{x}^{2}-q$ (67)
on $L^{2}(\mathbb{R})$ with domain $H^{2}(\mathbb{R})$. Since $q\in L^{1}(\mathbb{R})$, the operator $H_{q}$ is short-range: its essential spectrum is $0,\infty)$ and its negative spectrum, if present, consists of finitely many eigenvalues. For Bers-Miura potentials $q=\beta^{-\infty}(\varphi)$, the factorization in Proposition 6.6 implies $H_{q}\geq 0$, hence the scattering data contain no discrete bound state part. In this situation the reflection coefficient provides a nonlinear spectral coordinate on $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
7.1. Jost solutions and scattering data
Definition 7.1 (Jost solutions).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $k\in\mathbb{C}$ with $\Im k\geq 0$. A Jost solution for $H_{q}$ is a solution $y(x,k)$ of
$y^{\prime\prime}(x,k)-q(x)y(x,k)=k^{2}y(x,k)$ (68)
with prescribed plane wave asymptotics at one end. The right and left Jost solutions $f_{\pm}(x,k)$ are uniquely characterized by
[…]
For $k\in\mathbb{R}\setminus\{0\}$ the scattering coefficients $a(k),b(k)$ are defined by the identity
$f_{-}(x,k)=a(k)f_{+}(x,-k)+b(k)f_{+}(x,k),\qquad k\in\mathbb{R}\setminus\{0\}.$ (70)
The transmission and reflection coefficients are $T(k)=a(k)^{-1}$ and $R(k)=b(k)/a(k)$.
Proposition 7.2 (Existence and basic properties of scattering data).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real.
  1. (i)
    For each $k$ with $\Im k\geq 0$ and $k\neq 0$, the Jost solutions exist, are $C^{\infty}$ in $x$, and depend holomorphically on $k$ for $\Im k>0$ and continuously on $k$ for $\Im k\geq 0$. They solve the Volterra equations
    $f_{+}(x,k)=\mathrm{e}^{\mathrm{i}kx}+\int_{x}^{\infty}\frac{\sin(k(t-x))}{k}\,q(t)f_{+}(t,k)\,\mathrm{d}t,\qquad\Im k\geq 0,$ (71)
    and the analogous equation for $f_{-}(x,k)$ with integration from -$\infty$ to $x$.
  2. (ii)
    The coefficients $a(k),b(k)$ are $C^{\infty}$ on $\mathbb{R}\setminus\{0\}$ and satisfy the symmetry
    $a(-k)=\overline{a(k)},\qquad b(-k)=\overline{b(k)},\qquad k\in\mathbb{R}\setminus\{0\}.$ (72)
  3. (iii)
    One has the flux identities
    $|a(k)|^{2}=1+|b(k)|^{2},\qquad|T(k)|^{2}+|R(k)|^{2}=1,\qquad k\in\mathbb{R}\setminus\{0\}.$ (73)
    In particular $|a(k)|\geq 1$, with equality if and only if $b(k)=0$.
  4. (iv)
    For every $N\in\mathbb{N}$ there exists $C_{N}$ such that, for $|k|\geq 1$,
    $|a(k)-1|+|b(k)|\leq C_{N}|k|^{-N}.$
    Consequently a(k)1a(k)-1 and $b(k)$ extend to Schwartz functions of $k$ on $\mathbb{R}$.
Proof.
Fix $k$ with $\Im k\geq 0$ and $k\neq 0$. The Green kernel for $\partial_{x}^{2}-k^{2}$ with outgoing normalization at $\infty$ is $G_{k}(x,t)=\frac{\sin(k(t-x))}{k}\mathbf{1}_{t\geq x}$, so variation of constants gives (71). Let $X=L^{\infty}([x_{0},\infty))$ with norm $\|u\|_{X}=\sup_{t\geq x_{0}}|u(t)|$. For $u\in X$ the integral operator
$(K_{k}u)(x):=\int_{x}^{\infty}\frac{\sin(k(t-x))}{k}\,q(t)u(t)\,\mathrm{d}t$
satisfies, via the bound $|\sin(k(t-x))/k|\leq|t-x|$ valid for all $k\neq 0$,
$\|K_{k}u\|_{X}\leq\left(\int_{x_{0}}^{\infty}(t-x_{0})\,|q(t)|\,\mathrm{d}t\right)\|u\|_{X}.$
Since $q\in W^{\infty,1}(\mathbb{R})$ implies $\int_{\mathbb{R}}(1+|t|)|q(t)|\,\mathrm{d}t<\infty$, choosing $x_{0}$ large enough that $\int_{x_{0}}^{\infty}(t-x_{0})|q(t)|\,\mathrm{d}t<\frac{1}{2}$ makes the map $u\mapsto\mathrm{e}^{\mathrm{i}k\cdot}+K_{k}u$ a contraction on $X$, hence it has a unique fixed point. Patching the resulting solution as $x_{0}\to$-$\infty$ gives a globally defined $f_{+}(x,k)$ satisfying (71) for all $x\in\mathbb{R}$. Differentiating (71) in $x$ shows $f_{+}$ is $C^{\infty}$ in $x$ and solves (68). Holomorphy in $k$ for $\Im k>0$ follows by differentiating the Volterra equation in $k$: each $k$-derivative falls on $\sin(k(t-x))/k$ and produces factors bounded by $|t-x|^{m}$ times $\mathrm{e}^{-(\Im k)(t-x)}$, which remain integrable against $|q(t)|$ for $\Im k>0$. Continuity on $\Im k=0$ follows by dominated convergence, using $|\sin(k(t-x))|\leq\mathrm{e}^{(\Im k)(t-x)}$.
For real $k\neq 0$, the coefficients $a(k),b(k)$ are determined by Wronskians. Since the Wronskian $W(u,v)=uv^{\prime}-u^{\prime}v$ is constant in $x$ for any pair of solutions of (68), applying $W(f_{+}(k),\cdot)$ to (70) yields
$a(k)=\frac{1}{2\mathrm{i}k}W(f_{-}(\cdot,k),f_{+}(\cdot,k)),$ (74)
because $W(f_{+}(\cdot,k),f_{+}(\cdot,-k))=-2\mathrm{i}k$ by the asymptotics (69). Similarly, applying $W(f_{+}(\cdot,-k),\cdot)$ to (70) gives
$b(k)=\frac{1}{2\mathrm{i}k}W(f_{+}(\cdot,-k),f_{-}(\cdot,k)).$ (75)
The dependence asserted in (ii) follows from the smooth dependence of $f_{\pm}$ on $k$ on $\mathbb{R}\setminus\{0\}$ and the identities (74)-(75). The symmetry (72) follows from reality of $q$ and uniqueness of Jost solutions: for $k\in\mathbb{R}\setminus\{0\}$ the function $\overline{f_{+}(x,k)}$ solves (68) and has asymptotic $\mathrm{e}^{-\mathrm{i}kx}$ as $x\to+\infty$, hence $\overline{f_{+}(x,k)}=f_{+}(x,-k)$, and similarly $\overline{f_{-}(x,k)}=f_{-}(x,-k)$.
To prove (73), consider the conserved current for real solutions of the stationary equation: if $y$ solves (68) with real $k$, then
$J[y]:=\frac{1}{2\mathrm{i}}\bigl(\overline{y}\,y^{\prime}-y\,\overline{y}^{\prime}\bigr)$
is constant in $x$ because $J^{\prime}[y]=\Im(\overline{y}y^{\prime\prime})$ and $y^{\prime\prime}=-(k^{2}+q)y$ with $k^{2}+q$ real. Evaluate $J[f_{-}(\cdot,k)]$ as $x\to$-$\infty$ using $f_{-}(x,k)\sim\mathrm{e}^{-\mathrm{i}kx}$ to obtain $J[f_{-}]=-k$. Evaluate the same current as $x\to+\infty$ using (70) and $f_{+}(x,\pm k)\sim\mathrm{e}^{\pm\mathrm{i}kx}$:
$f_{-}(x,k)\sim a(k)\mathrm{e}^{-\mathrm{i}kx}+b(k)\mathrm{e}^{\mathrm{i}kx}\quad\text{as }x\to+\infty,$
hence $J[f_{-}]=-k|a(k)|^{2}+k|b(k)|^{2}$. Equating the two values gives $|a(k)|^{2}=1+|b(k)|^{2}$. Dividing by $|a(k)|^{2}$ yields $|T(k)|^{2}+|R(k)|^{2}=1$ and $|a(k)|\geq 1$, with equality exactly when $b(k)=0$.
For (iv), introduce $m_{+}(x,k):=\mathrm{e}^{-\mathrm{i}kx}f_{+}(x,k)$, so $m_{+}(x,k)\to 1$ as $x\to+\infty$ and $m_{+}$ satisfies the Volterra equation obtained by substituting $f_{+}=\mathrm{e}^{\mathrm{i}k\cdot}m_{+}$ into (71),
$m_{+}(x,k)=1+\int_{x}^{\infty}\frac{\mathrm{e}^{2\mathrm{i}k(t-x)}-1}{2\mathrm{i}k}\,q(t)m_{+}(t,k)\,\mathrm{d}t.$
Iterating this equation shows $m_{+}(x,k)$ is uniformly bounded in $x$ for $|k|\geq 1$ and $\Im k\geq 0$, with bounds controlled by $\|q\|_{L^{1}}$. Using (75) and the asymptotics of $f_{-}$ at -$\infty$ one obtains the standard representation
$b(k)=-\frac{1}{2\mathrm{i}k}\int_{\mathbb{R}}q(t)m_{+}(t,k)\mathrm{e}^{2\mathrm{i}kt}\,\mathrm{d}t,\qquad k\in\mathbb{R}\setminus\{0\},$
and similarly a(k)1a(k)-1 is given by an oscillatory integral with integrand built from $q$ and $m_{+}$. Since $q^{(N)}\in L^{1}(\mathbb{R})$ for every $N$ and $m_{+}(\cdot,k)$ is bounded with bounded $t$-derivatives expressible by the differential equation, repeated integration by parts in $t$ gives $|b(k)|\leq C_{N}|k|^{-N}$ for $|k|\geq 1$. The same argument applies to a(k)1a(k)-1. This proves rapid decay and hence the Schwartz property on $\mathbb{R}$. ∎
7.2. Miura factorization and absence of bound states
Lemma 7.3 (No bound states for Bers-Miura potentials).
Let $q=\beta^{-\infty}(\varphi)$ with $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and write $u=\log(\varphi^{\prime})$. Set $A:=\partial_{x}+\frac{u^{\prime}}{2}$. Then
$H_{q}=A^{*}A\geq 0$ (76)
as a self-adjoint operator on $L^{2}(\mathbb{R})$. In particular $H_{q}$ has no negative eigenvalues, hence no bound states. Moreover $H_{q}y=0$ admits a positive solution $y=\mathrm{e}^{-u/2}=(\varphi^{\prime})^{-1/2}$, which is bounded and satisfies $y(x)\to 1$ as $x\to\pm\infty$.
Proof.
Proposition 6.6 gives
$\mathcal{L}_{q}=\partial_{x}^{2}+q=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\left(\partial_{x}+\frac{u^{\prime}}{2}\right),$
hence $H_{q}=-\mathcal{L}_{q}=(-\partial_{x}+\frac{u^{\prime}}{2})(\partial_{x}+\frac{u^{\prime}}{2})$. Since the adjoint of $\partial_{x}$ on $L^{2}(\mathbb{R})$ is $\partial_{x}$ and $u^{\prime}/2$ acts by multiplication, this is exactly $A^{*}A$, proving (76). For $f\in H^{1}(\mathbb{R})$ one has $\langle f,H_{q}f\rangle=\|Af\|_{L^{2}}^{2}\geq 0$, so $H_{q}\geq 0$ and there are no negative eigenvalues. Finally $Ay=0$ has the positive solution $y=\mathrm{e}^{-u/2}$ and $u\in W^{\infty,1}(\mathbb{R})$ implies $u(x)\to 0$ as $x\to\pm\infty$, hence $y(x)\to 1$ at both ends. ∎
7.3. Spectral characterization at zero energy
Theorem 7.4 (Spectral characterization of the Bers image).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $f_{+}(x,k)$ be the right Jost solution of $H_{q}$ as in Definition 7.1. Then $q$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the zero-energy Jost solution $f_{+}(x,0)$ satisfies
$f_{+}(x,0)>0\qquad\text{and}\qquad f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R}).$ (77)
In this case the associated diffeomorphism is
$\varphi(x)=x+\int_{-\infty}^{x}\bigl(f_{+}(t,0)^{-2}-1\bigr)\,\mathrm{d}t,$ (78)
so that $\varphi^{\prime}=f_{+}(\cdot,0)^{-2}$ and $q=\beta^{-\infty}(\varphi)$. Moreover, with $w:=-f_{+}^{\prime}(\cdot,0)/f_{+}(\cdot,0)$ one has $w\in W^{\infty,1}(\mathbb{R})$ and the Miura factorization
$\mathcal{L}_{q}=(\partial_{x}-w)(\partial_{x}+w)$
holds.
Proof.
Assume first that $q=\beta^{-\infty}(\varphi)$ for some $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Proposition 6.6 gives a positive solution $y_{2}=(\varphi^{\prime})^{-1/2}$ of $\mathcal{L}_{q}y=0$, equivalently $H_{q}y=0$, and $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. Since $\varphi^{\prime}\to 1$ as $x\to+\infty$, we have $y_{2}(x)\to 1$ as $x\to+\infty$. The Volterra equation (71) at $k=0$ becomes
$y(x)=1+\int_{x}^{\infty}(t-x)q(t)y(t)\,\mathrm{d}t,$
and it has a unique bounded solution with $y(x)\to 1$ as $x\to+\infty$. By the defining property (69) at $k=0$, this solution is exactly $f_{+}(x,0)$. Thus $f_{+}(x,0)=y_{2}(x)>0$ and $f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R})$, proving (77).
Conversely, assume (77). Define $\varphi$ by (78). Since $f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R})$, we have $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$ and $\varphi^{\prime}>0$, hence $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Set $y_{2}:=f_{+}(\cdot,0)$ and $y_{1}:=\varphi y_{2}$. Then $y_{2}$ solves $y_{2}^{\prime\prime}+qy_{2}=0$. Since $\varphi^{\prime}=y_{2}^{-2}$, the quotient rule gives
$\left(\frac{y_{1}}{y_{2}}\right)^{\prime}=\varphi^{\prime}=y_{2}^{-2},$
hence $y_{1}^{\prime}y_{2}-y_{1}y_{2}^{\prime}=\varphi^{\prime}y_{2}^{2}=1$. Thus $y_{1},y_{2}$ are linearly independent solutions with unit Wronskian in the convention of Theorem 6.8. Moreover $y_{2}>0$ and $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. By Theorem 6.8, this implies $q=\beta^{-\infty}(\varphi)$.
Finally, put $w:=-y_{2}^{\prime}/y_{2}$. Since $y_{2}>0$ and $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$, the function $\log y_{2}$ has derivative w-w in $W^{\infty,1}(\mathbb{R})$, hence $w\in W^{\infty,1}(\mathbb{R})$. Writing $y_{2}=\mathrm{e}^{-\int w}$ and substituting into $y_{2}^{\prime\prime}+qy_{2}=0$ gives $q=w^{\prime}-w^{2}$ and therefore $\mathcal{L}_{q}=(\partial_{x}-w)(\partial_{x}+w)$. ∎
7.4. The scattering transform as a nonlinear Fourier map
For $q\in W^{\infty,1}(\mathbb{R})$ the scattering data consist of the reflection coefficient together with possible discrete bound state data. Lemma 7.3 shows that for $q=\beta^{-\infty}(\varphi)$ no bound states occur, so the reflection coefficient alone captures the scattering data. This yields a nonlinear spectral coordinate on the real Bers image.
Definition 7.5 (Scattering Bers map).
Define
$\mathcal{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to\mathscr{S}(\mathbb{R}),\qquad\mathcal{S}(\varphi):=R_{\beta^{-\infty}(\varphi)},$
where $R_{q}$ is the reflection coefficient of the Schrödinger operator $H_{q}$ in (67).
Theorem 7.6 (Injectivity of the scattering Bers map).
The map $\mathcal{S}$ is constant on left $\operatorname{Aff}(\mathbb{R})$-cosets and descends to an injective map on $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$. Moreover, for every $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ the function $R=\mathcal{S}(\varphi)$ satisfies:
  1. (i)
    $\overline{R(-k)}=R(k)$ for all $k\in\mathbb{R}$,
  2. (ii)
    $|R(k)|<1$ for all $k\in\mathbb{R}\setminus\{0\}$,
  3. (iii)
    the boundary modulus determines an outer function $a_{\mathrm{out}}$ on the upper half-plane,
    $a_{\mathrm{out}}(z)=\exp\!\left(-\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{R}}\frac{\log(1-|R(\xi)|^{2})}{\xi-z}\,\mathrm{d}\xi\right),\qquad\Im z>0,$ (79)
    and the transmission denominator $a(z)$ for $H_{\beta^{-\infty}(\varphi)}$ coincides with $a_{\mathrm{out}}(z)$ and has no zeros in $\Im z>0$.
Proof.
If $\psi\in\operatorname{Aff}(\mathbb{R})$ then $\beta^{-\infty}(\varphi\circ\psi)=\beta^{-\infty}(\varphi)$ by Theorem 6.2, hence $\mathcal{S}(\varphi\circ\psi)=\mathcal{S}(\varphi)$, so $\mathcal{S}$ factors through $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
For injectivity, let $\varphi_{1},\varphi_{2}\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q_{j}=\beta^{-\infty}(\varphi_{j})$. Assume $\mathcal{S}(\varphi_{1})=\mathcal{S}(\varphi_{2})$, so $R_{q_{1}}(k)=R_{q_{2}}(k)$ for all $k\in\mathbb{R}\setminus\{0\}$. By Lemma 7.3, the operators $H_{q_{j}}$ have no negative spectrum, hence their scattering data have no discrete part. The Gel’fand-Levitan-Marchenko reconstruction for short-range one-dimensional Schrödinger operators then implies $q_{1}=q_{2}$ from equality of the reflection coefficients in the no-bound-state case, see [10, Theorem 1]. The injectivity of $\beta^{-\infty}$ modulo $\operatorname{Aff}(\mathbb{R})$ from Theorem 6.2 then gives $\varphi_{1}]=[\varphi_{2}]$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
For the additional properties, (i) follows from reality of $q$ and the symmetry $b(-k)=\overline{b(k)}$, $a(-k)=\overline{a(k)}$ in (72). For (ii), Proposition 7.2 gives $|T(k)|^{2}+|R(k)|^{2}=1$ for $k\neq 0$, and $T(k)\neq 0$ for real $k\neq 0$ since $T(k)=0$ would produce a nontrivial solution of (68) that is purely outgoing at both ends at energy $k^{2}>0$, which is impossible. Thus $|R(k)|<1$ for $k\neq 0$.
For (iii), the function $a(z)$ extends holomorphically to $\Im z>0$ and has no zeros there exactly because zeros in $\Im z>0$ correspond to bound states of $H_{q}$ and these are excluded by Lemma 7.3. On the real axis one has $|a(k)|^{-2}=|T(k)|^{2}=1-|R(k)|^{2}$ for $k\in\mathbb{R}\setminus\{0\}$ by (73). Since $R\in\mathscr{S}(\mathbb{R})$, the function $\log(1-|R|^{2})$ is integrable and the outer function $a_{\mathrm{out}}$ defined by (79) lies in the Hardy class on $\Im z>0$ and has boundary modulus $|a_{\mathrm{out}}(k)|=|a(k)|$. The quotient $a/a_{\mathrm{out}}$ is then a bounded holomorphic function on $\Im z>0$ with unimodular boundary values, hence constant. The normalization $a(z)\to 1$ as $|z|\to\infty$ forces that constant to be $1$, so $a=a_{\mathrm{out}}$. ∎
7.5. Spectral zeta functions and regularized determinants
The heat semigroup for $H_{q}$ is trace class on compact sets but not on $L^{2}(\mathbb{R})$ (see [28] for background on trace ideals). The correct global object is the relative trace obtained from the difference of heat kernels.
Definition 7.7 (Relative spectral zeta function).
Let $\mu=g\,\mathrm{d}x\in|\Omega|_{\infty,1}(\mathbb{R})$ and put $q=\beta(\mu)$. Define
$H_{\mu}:=-\partial_{x}^{2}-q,\qquad H_{0}:=-\partial_{x}^{2}.$
Let $K_{\mu}(t,x,y)$ and $K_{0}(t,x,y)$ be the heat kernels of $\mathrm{e}^{-tH_{\mu}}$ and $\mathrm{e}^{-tH_{0}}$. For $t>0$ define the relative heat trace
$\operatorname{Tr}_{\mathrm{rel}}(\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}}):=\int_{\mathbb{R}}\bigl(K_{\mu}(t,x,x)-K_{0}(t,x,x)\bigr)\,\mathrm{d}x,$ (80)
which is finite for $q\in W^{\infty,1}(\mathbb{R})$. For $\Re z$ large, the relative zeta function is
$\zeta_{\mu}(z)=\frac{1}{\Gamma(z)}\int_{0}^{\infty}t^{z-1}\operatorname{Tr}_{\mathrm{rel}}(\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}})\,\mathrm{d}t.$ (81)
Proposition 7.8 (Zeta pole and score energy).
Let $\mu=g\,\mathrm{d}x\in|\Omega|_{\infty,1}(\mathbb{R})$ and $q=\beta(\mu)$. The function $\zeta_{\mu}(z)$ admits a meromorphic continuation to $\mathbb{C}$ with at most simple poles at $z=-\frac{1}{2},-\frac{3}{2},-\frac{5}{2},\ldots$. Moreover,
$\operatorname{Res}_{z=-\frac{1}{2}}\zeta_{\mu}(z)=-\frac{1}{4\pi}\int_{\mathbb{R}}\beta(\mu)\,\mathrm{d}x=\frac{1}{16\pi}\int_{\mathbb{R}}s^{2}\,\mathrm{d}x,$ (82)
where $s=(\log g)^{\prime}$ is the score.
Proof.
Write $q=\beta(\mu)\in W^{\infty,1}(\mathbb{R})$. For $t>0$, the Duhamel identity gives
$\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}}=\int_{0}^{t}\mathrm{e}^{-(t-\tau)H_{0}}\,q\,\mathrm{e}^{-\tau H_{\mu}}\,\mathrm{d}\tau,$ (83)
which is obtained by differentiating $\tau\mapsto\mathrm{e}^{-(t-\tau)H_{0}}\mathrm{e}^{-\tau H_{\mu}}$ and integrating from $\tau=0$ to $\tau=t$.
To isolate the leading small-time behavior, we substitute once more into the factor $\mathrm{e}^{-\tau H_{\mu}}$ the identity
$\mathrm{e}^{-\tau H_{\mu}}=\mathrm{e}^{-\tau H_{0}}+\int_{0}^{\tau}\mathrm{e}^{-(\tau-\sigma)H_{0}}\,q\,\mathrm{e}^{-\sigma H_{\mu}}\,\mathrm{d}\sigma,$ (84)
which is (83) with $t$ replaced by $\tau$. Inserting (84) into (83) yields the decomposition
$\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}}=\int_{0}^{t}\mathrm{e}^{-(t-\tau)H_{0}}\,q\,\mathrm{e}^{-\tau H_{0}}\,\mathrm{d}\tau+\int_{0<\sigma<\tau<t}\mathrm{e}^{-(t-\tau)H_{0}}\,q\,\mathrm{e}^{-(\tau-\sigma)H_{0}}\,q\,\mathrm{e}^{-\sigma H_{\mu}}\,\mathrm{d}\sigma\,\mathrm{d}\tau.$ (85)
Let $K_{0}(t,x,y)=(4\pi t)^{-1/2}\exp(-(x-y)^{2}/(4t))$ be the free heat kernel and let $K_{\mu}$ denote the heat kernel of $H_{\mu}$. Passing to Schwartz kernels in (85), restricting to the diagonal $y=x$, and integrating in $x$ gives
$\operatorname{Tr}_{\mathrm{rel}}(\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}})=I_{1}(t)+I_{2}(t),$ (86)
where
$I_{1}(t):=\int_{0}^{t}\int_{\mathbb{R}}\int_{\mathbb{R}}K_{0}(t-\tau,x,y)\,q(y)\,K_{0}(\tau,y,x)\,\mathrm{d}y\,\mathrm{d}x\,\mathrm{d}\tau$ (87)
and $I_{2}(t)$ is the contribution of the double integral term in (85). We first evaluate $I_{1}(t)$ exactly. By the semigroup property of the free kernel,
$\int_{\mathbb{R}}K_{0}(t-\tau,x,y)\,K_{0}(\tau,y,x)\,\mathrm{d}x=K_{0}(t,y,y),$ (88)
and therefore (87) reduces to
$I_{1}(t)=\int_{0}^{t}\int_{\mathbb{R}}q(y)\,K_{0}(t,y,y)\,\mathrm{d}y\,\mathrm{d}\tau=t\,(4\pi t)^{-1/2}\int_{\mathbb{R}}q(y)\,\mathrm{d}y=\frac{t^{1/2}}{\sqrt{4\pi}}\int_{\mathbb{R}}q(y)\,\mathrm{d}y.$ (89)
We now estimate $I_{2}(t)$ and show it is $O(t^{3/2})$ as $t\downarrow 0$. The kernel of the second term in (85) is
$\int_{0<\sigma<\tau<t}\int_{\mathbb{R}^{2}}K_{0}(t-\tau,x,y_{1})\,q(y_{1})\,K_{0}(\tau-\sigma,y_{1},y_{2})\,q(y_{2})\,K_{\mu}(\sigma,y_{2},x)\,\mathrm{d}y_{1}\,\mathrm{d}y_{2}\,\mathrm{d}\sigma\,\mathrm{d}\tau.$ (90)
Integrating (90) over $x$ and using $\int_{\mathbb{R}}K_{0}(s,x,y_{1})\,\mathrm{d}x=1$ for $s>0$, we obtain the bound
$|I_{2}(t)|\leq\int_{0<\sigma<\tau<t}\int_{\mathbb{R}^{2}}|q(y_{1})|\,K_{0}(\tau-\sigma,y_{1},y_{2})\,|q(y_{2})|\,\left(\int_{\mathbb{R}}K_{\mu}(\sigma,y_{2},x)\,\mathrm{d}x\right)\mathrm{d}y_{1}\,\mathrm{d}y_{2}\,\mathrm{d}\sigma\,\mathrm{d}\tau.$ (91)
Since $H_{\mu}$ is self-adjoint and semibounded, the heat kernel $K_{\mu}(\sigma,\cdot,\cdot)$ is nonnegative and
$\int_{\mathbb{R}}K_{\mu}(\sigma,y_{2},x)\,\mathrm{d}x\leq C$
uniformly for $\sigma\in(0,1]$, with $C$ depending only on a lower bound for $H_{\mu}$; for $q\in W^{\infty,1}$ such a bound holds and $C$ may be taken finite. Using this in (91) and the fact that $K_{0}(\tau-\sigma,\cdot,\cdot)$ has unit mass in each variable gives
$|I_{2}(t)|$ $\leq C\int_{0<\sigma<\tau<t}\int_{\mathbb{R}}|q(y_{2})|\left(\int_{\mathbb{R}}|q(y_{1})|\,K_{0}(\tau-\sigma,y_{1},y_{2})\,\mathrm{d}y_{1}\right)\mathrm{d}y_{2}\,\mathrm{d}\sigma\,\mathrm{d}\tau$
$\leq C\|q\|_{L^{1}}^{2}\int_{0<\sigma<\tau<t}\mathrm{d}\sigma\,\mathrm{d}\tau=\frac{C}{2}\|q\|_{L^{1}}^{2}\,t^{2}.$ (92)
In particular $I_{2}(t)=O(t^{2})$ as $t\downarrow 0$. Combining (86), (89), and (92) yields the small-time expansion
$\operatorname{Tr}_{\mathrm{rel}}(\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}})=\frac{t^{1/2}}{\sqrt{4\pi}}\int_{\mathbb{R}}q(x)\,\mathrm{d}x+O(t^{3/2}),\qquad t\downarrow 0.$ (93)
We now insert (93) into the Mellin representation (81). The contribution of the leading term is
$\frac{1}{\Gamma(z)}\int_{0}^{1}t^{z-1}\frac{t^{1/2}}{\sqrt{4\pi}}\left(\int_{\mathbb{R}}q\right)\mathrm{d}t=\frac{1}{\Gamma(z)}\frac{1}{\sqrt{4\pi}}\left(\int_{\mathbb{R}}q\right)\frac{1}{z+\frac{1}{2}},$
which has a simple pole at $z=-\frac{1}{2}$ with residue $\frac{1}{\Gamma(-\frac{1}{2})\sqrt{4\pi}}\int_{\mathbb{R}}q=-\frac{1}{4\pi}\int_{\mathbb{R}}q$, since $\Gamma(-\frac{1}{2})=-2\sqrt{\pi}$. The error term $O(t^{3/2})$ produces a function holomorphic in a neighborhood of $z=-\frac{1}{2}$. Thus
$\operatorname{Res}_{z=-\frac{1}{2}}\zeta_{\mu}(z)=-\frac{1}{4\pi}\int_{\mathbb{R}}q(x)\,\mathrm{d}x.$ (94)
Since $q=\beta(\mu)$, Proposition 9.11 gives
$\int_{\mathbb{R}}q\,\mathrm{d}x=\int_{\mathbb{R}}\beta(\mu)\,\mathrm{d}x=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,\mathrm{d}x,$
and substituting this into (94) yields (82).
Iterating (83) produces, for each $N\geq 1$, a finite Dyson-Duhamel expansion
$\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}}=\sum_{m=1}^{N}\int_{\Delta_{m}(t)}\mathrm{e}^{-(t-\tau_{1})H_{0}}q\,\mathrm{e}^{-(\tau_{1}-\tau_{2})H_{0}}q\cdots q\,\mathrm{e}^{-\tau_{m}H_{0}}\,\mathrm{d}\tau+\mathcal{R}_{N+1}(t),$ (95)
whose remainder $\mathcal{R}_{N+1}(t)$ contributes $O(t^{N+\frac{1}{2}})$ to the relative trace. Passing to kernels and expanding the Gaussian moments in the main terms yields an asymptotic expansion
$\operatorname{Tr}_{\mathrm{rel}}(\mathrm{e}^{-tH_{\mu}}-\mathrm{e}^{-tH_{0}})=\frac{1}{\sqrt{4\pi t}}\sum_{n=1}^{N}c_{n}[q]\,t^{n}+O\!\left(t^{N+\frac{1}{2}}\right),\qquad t\downarrow 0,$ (96)
with coefficients $c_{n}[q]$ given by integrals of universal polynomials in $q$ and its derivatives. Substituting (96) into (81) shows that each term $(4\pi t)^{-1/2}c_{n}[q]t^{n}$ produces a simple pole at $z=\frac{1}{2}-n$, hence the remaining poles can only occur at $z=-\frac{1}{2},-\frac{3}{2},-\frac{5}{2},\ldots$. ∎
Corollary 7.9 (Relative zeta determinant).
For $\mu=g\,\mathrm{d}x\in|\Omega|_{\infty,1}(\mathbb{R})$ define the relative zeta determinant by
$\log\det_{\zeta}(H_{\mu}/H_{0}):=-\zeta_{\mu}^{\prime}(0),$ (97)
where $\zeta_{\mu}$ is the relative zeta function in Definition 7.7.
7.6. Mercer kernels and reproducing Hilbert spaces
The scattering coordinate $R(k)$ packages the absolutely continuous spectrum of $H_{q}$. To pass from scattering data to canonical finite-rank approximations it is convenient to encode the spectrum into a positive definite kernel by functional calculus.
Definition 7.10 (Admissible spectral weights).
A nonnegative Borel measure $\rho$ on $(0,\infty)$ is called admissible if
$\int_{0}^{\infty}\lambda^{-1/2}\,\mathrm{d}\rho(\lambda)<\infty.$ (98)
For admissible $\rho$ define the completely monotone function $\Phi_{\rho}:[0,\infty)\to(0,\infty)$ by
$\Phi_{\rho}(\sigma):=\int_{0}^{\infty}\frac{1}{\sigma+\lambda}\,\mathrm{d}\rho(\lambda).$ (99)
Definition 7.11 (Bers-Mercer kernel).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $\rho$ be admissible. For $\lambda>0$ denote by $G_{q}(\cdot,\cdot;\lambda)$ the Green kernel of the resolvent $(H_{q}+\lambda)^{-1}$, i.e.
$((H_{q}+\lambda)^{-1}f)(x)=\int_{\mathbb{R}}G_{q}(x,y;\lambda)f(y)\,\mathrm{d}y.$
The $\rho$-Bers-Mercer kernel of $H_{q}$ is
$K_{q,\rho}(x,y):=\int_{0}^{\infty}G_{q}(x,y;\lambda)\,\mathrm{d}\rho(\lambda),\qquad x,y\in\mathbb{R}.$ (100)
Definition 7.12 (Stieltjes transform of an admissible weight).
Let $\rho$ be an admissible finite positive Borel measure on $0,\infty)$. Its Stieltjes transform is the function
$\Phi_{\rho}(z):=\int_{0}^{\infty}\frac{1}{z+\lambda}\,d\rho(\lambda),\qquad z\in\mathbb{C}\setminus(-\infty,0].$
In particular $\Phi_{\rho}$ is holomorphic on $\mathbb{C}\setminus(-\infty,0]$ and satisfies $\Phi_{\rho}(t)\geq 0$ for $t>0$.
Proposition 7.13 (Basic properties of $K_{q,\rho}$).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and assume $H_{q}\geq 0$. Let $\rho$ be admissible. Then the kernel $K_{q,\rho}$ defined by (100) is well-defined and continuous on $\mathbb{R}\times\mathbb{R}$. Moreover,
  1. (i)
    $K_{q,\rho}(x,y)=K_{q,\rho}(y,x)$ for all $x,y\in\mathbb{R}$,
  2. (ii)
    $K_{q,\rho}$ is positive definite, i.e. for any finite family $x_{1},\ldots,x_{N}\in\mathbb{R}$ and $c_{1},\ldots,c_{N}\in\mathbb{C}$,
    $\sum_{i,j=1}^{N}c_{i}\overline{c_{j}}\,K_{q,\rho}(x_{i},x_{j})\geq 0.$
Consequently there exists a unique reproducing kernel Hilbert space $\mathcal{H}_{q,\rho}$ of functions on $\mathbb{R}$ whose reproducing kernel is $K_{q,\rho}$.
Proof.
Fix $\lambda>0$ and write $\kappa=\sqrt{\lambda}$. Since $H_{q}$ is self-adjoint and $H_{q}\geq 0$, the operator $H_{q}+\lambda$ is strictly positive, hence invertible, and its resolvent $(H_{q}+\lambda)^{-1}$ is positive and self-adjoint. In particular its integral kernel $G_{q}(\cdot,\cdot;\lambda)$ is nonnegative and symmetric. In one dimension $G_{q}$ is continuous and admits the Jost representation of Proposition 7.14:
$G_{q}(x,y;\lambda)=\frac{f_{-}(x_{<},i\kappa)\,f_{+}(x_{>},i\kappa)}{W\!\bigl(f_{-}(\cdot,i\kappa),f_{+}(\cdot,i\kappa)\bigr)},\qquad x_{<}:=\min\{x,y\},\quad x_{>}:=\max\{x,y\}.$
The individual functions $f_{\pm}(\cdot,i\kappa)$ do not decay at both ends; what matters is the product appearing above. Introduce the renormalizations
$m_{+}(x,\kappa):=e^{\kappa x}f_{+}(x,i\kappa),\qquad m_{-}(x,\kappa):=e^{-\kappa x}f_{-}(x,i\kappa).$
Substituting $k=i\kappa$ into the Volterra equations (71) and its left analogue and factoring out the exponentials yields Volterra equations for $m_{\pm}$ whose kernels are nonnegative and bounded by $\frac{1}{2\kappa}|q(t)|$. Standard Volterra theory gives that for each compact $J\Subset\mathbb{R}$ there is a constant $M_{J}$ such that
$\sup_{\kappa\in(0,1]}\sup_{x\in J}\bigl(|m_{+}(x,\kappa)|+|m_{-}(x,\kappa)|\bigr)\leq M_{J},$ (101)
and for $\kappa\geq 1$ one has a uniform bound depending only on $\|q\|_{L^{1}}$. Consequently, for $(x,y)$ ranging in a fixed compact set $J\times J$,
$|f_{-}(x_{<},i\kappa)\,f_{+}(x_{>},i\kappa)|=e^{\kappa x_{<}}e^{-\kappa x_{>}}\,|m_{-}(x_{<},\kappa)m_{+}(x_{>},\kappa)|\leq M_{J}^{2}\,e^{-\kappa|x-y|}.$ (102)
Next, the Wronskian can be identified in terms of the transmission denominator. Indeed, the identity (74),
$a(k)=\frac{1}{2ik}\,W\!\bigl(f_{-}(\cdot,k),f_{+}(\cdot,k)\bigr),\qquad k\in\mathbb{R}\setminus\{0\},$
extends to $\Im k>0$ by holomorphy of $f_{\pm}(\cdot,k)$ in $k$ and constancy of the Wronskian in $x$. Evaluating at $k=i\kappa$ gives
$W\!\bigl(f_{-}(\cdot,i\kappa),f_{+}(\cdot,i\kappa)\bigr)=-2\kappa\,a(i\kappa).$ (103)
In the no-bound-state regime (in particular for Bers-Miura potentials), $a$ has no zeros in $\Im z>0$ and, moreover, the outer function formula of Theorem 7.6(iii) implies $|a(z)|\geq 1$ for $\Im z>0$. Hence
$\bigl|W(f_{-}(\cdot,i\kappa),f_{+}(\cdot,i\kappa))\bigr|=2\kappa\,|a(i\kappa)|\geq 2\kappa.$ (104)
Combining (102) with (104) yields, for $(x,y)\in J\times J$,
$0\leq G_{q}(x,y;\lambda)\leq\frac{M_{J}^{2}}{2\kappa}\,e^{-\kappa|x-y|}=\frac{M_{J}^{2}}{2\sqrt{\lambda}}\,e^{-\sqrt{\lambda}|x-y|}.$ (105)
Since $\rho$ is admissible, $\int_{0}^{\infty}\lambda^{-1/2}\,d\rho(\lambda)<\infty$, and therefore (105) implies that
$\int_{0}^{\infty}G_{q}(x,y;\lambda)\,d\rho(\lambda)$
converges absolutely and uniformly on $J\times J$. This shows $K_{q,\rho}$ is well-defined and, by dominated convergence in $(x,y)$, continuous on $\mathbb{R}\times\mathbb{R}$. Symmetry follows from symmetry of each $G_{q}(\cdot,\cdot;\lambda)$.
For positive definiteness, let $\nu=\sum_{j=1}^{N}c_{j}\delta_{x_{j}}$. Using Fubini and positivity of the resolvent (which holds because $H_{q}\geq 0$), one finds
$\sum_{i,j=1}^{N}c_{i}\overline{c_{j}}\,K_{q,\rho}(x_{i},x_{j})=\int_{0}^{\infty}\Bigl\langle\nu,(H_{q}+\lambda)^{-1}\nu\Bigr\rangle\,d\rho(\lambda)\geq 0.$
The RKHS statement is the Moore-Aronszajn theorem [2]. ∎
Proposition 7.14 (Jost formula for the resolvent kernel).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $\lambda>0$. Write $\kappa=\sqrt{\lambda}$ and let $f_{\pm}(\cdot,i\kappa)$ be the Jost solutions from Definition 7.1. Then the resolvent kernel of $H_{q}+\lambda$ is
$G_{q}(x,y;\lambda)=\frac{1}{W\bigl(f_{-}(\cdot,i\kappa),f_{+}(\cdot,i\kappa)\bigr)}\begin{cases}f_{-}(x,i\kappa)\,f_{+}(y,i\kappa),&x\leq y,\\[3.00003pt] f_{-}(y,i\kappa)\,f_{+}(x,i\kappa),&x>y.\end{cases}$ (106)
The denominator is nonzero, and $G_{q}(\cdot,\cdot;\lambda)$ is positive and continuous.
Proof.
For $\kappa>0$, the functions $f_{+}(\cdot,i\kappa)$ and $f_{-}(\cdot,i\kappa)$ solve $(H_{q}+\kappa^{2})u=0$ and decay exponentially at $\infty$ and -$\infty$, respectively, by (69). They are linearly independent, hence their Wronskian is a nonzero constant. Define $G_{q}$ by (106). For fixed $y$, the function $x\mapsto G_{q}(x,y;\lambda)$ solves the homogeneous equation away from $x=y$, is continuous at $x=y$, and has the unit jump in the first derivative
$\partial_{x}G_{q}(x,y;\lambda)\big|_{x=y+}-\partial_{x}G_{q}(x,y;\lambda)\big|_{x=y-}=1,$
which is an immediate computation from constancy of the Wronskian. Exponential decay of $f_{\pm}(\cdot,i\kappa)$ gives $G_{q}(\cdot,y;\lambda)\in L^{2}(\mathbb{R})$ and makes boundary terms vanish when testing against compactly supported smooth functions, so $(H_{q}+\lambda)G_{q}(\cdot,y;\lambda)=\delta_{y}$ in the distributional sense. Thus $G_{q}$ is the resolvent kernel. Positivity follows from positivity preservation of the resolvent (or from the maximum principle for $(H_{q}+\lambda)u=0$). ∎
Theorem 7.15 (Scattering representation and Mercer expansion).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and assume $H_{q}\geq 0$ (in particular, $q=\beta^{-\infty}(\varphi)$ implies this by Lemma 7.3). Let $\rho$ be admissible and let $K_{q,\rho}$ be the Bers-Mercer kernel and $\Phi_{\rho}$ be as in Definition 7.12. Then $K_{q,\rho}$ admits the spectral representation
$K_{q,\rho}(x,y)=\frac{1}{2\pi}\int_{\mathbb{R}}\Phi_{\rho}(k^{2})\,\frac{f_{+}(x,k)\,f_{+}(y,-k)}{|a(k)|^{2}}\,\mathrm{d}k,$ (107)
where $a(k)$ are the transmission denominator from Definition  7.1. Moreover, for every bounded interval $I\subset\mathbb{R}$, the restriction $K_{q,\rho}|_{I\times I}$ is a continuous positive definite kernel on the compact set $\overline{I}$, hence there exist eigenvalues $\{\lambda_{n}^{(I)}\}_{n\geq 1}\subset(0,\infty)$ and an $L^{2}(I)$-orthonormal basis $\{\psi_{n}^{(I)}\}_{n\geq 1}$ such that
$K_{q,\rho}(x,y)=\sum_{n=1}^{\infty}\lambda_{n}^{(I)}\,\psi_{n}^{(I)}(x)\,\psi_{n}^{(I)}(y),\qquad(x,y)\in I\times I,$ (108)
with uniform convergence on $I\times I$.
Proof.
By Definition 7.11 one has, for all $x,y\in\mathbb{R}$,
$K_{q,\rho}(x,y)=\int_{0}^{\infty}G_{q}(x,y;\lambda)\,d\rho(\lambda),$
where $G_{q}(\cdot,\cdot;\lambda)$ is the Green kernel of $(H_{q}+\lambda)^{-1}$ for $\lambda>0$.
Fix $\lambda>0$. Since $q\in W^{\infty,1}(\mathbb{R})$ is short-range and $H_{q}\geq 0$, the operator $H_{q}$ has no negative eigenvalues; hence its spectrum is purely absolutely continuous on $0,\infty)$. Let $\psi(\cdot,k)$ be the left-incident scattering eigenfunction
$\psi(x,k):=\frac{1}{a(k)}\,f_{+}(x,k),\qquad k\in\mathbb{R}\setminus\{0\},$
so that $\psi(\cdot,k)$ solves $H_{q}\psi(\cdot,k)=k^{2}\psi(\cdot,k)$ and is normalized at $\infty$. The scattering Fourier transform associated to $\psi$ diagonalizes $H_{q}$ on $L^{2}(\mathbb{R})$, with spectral measure $dk/(2\pi)$; equivalently, for the bounded Borel function $m_{\lambda}(t)=(t+\lambda)^{-1}$ one has the spectral resolution of the resolvent
$(H_{q}+\lambda)^{-1}=m_{\lambda}(H_{q}),\qquad m_{\lambda}(H_{q})(x,y)=\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\psi(x,k)\,\overline{\psi(y,k)}}{k^{2}+\lambda}\,dk,$
and the integral is absolutely convergent since $(k^{2}+\lambda)^{-1}\in L^{1}(\mathbb{R})$ and $\psi(\cdot,k)$ is bounded in $x$ for each fixed $k$. In other words,
$G_{q}(x,y;\lambda)=\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\psi(x,k)\,\overline{\psi(y,k)}}{k^{2}+\lambda}\,dk.$ (109)
For real $q$, uniqueness of Jost solutions and (72) give $\overline{f_{+}(y,k)}=f_{+}(y,-k)$ and $\overline{a(k)}=a(-k)$ for $k\in\mathbb{R}\setminus\{0\}$, hence
$\psi(x,k)\,\overline{\psi(y,k)}=\frac{f_{+}(x,k)}{a(k)}\,\frac{f_{+}(y,-k)}{\overline{a(k)}}=\frac{f_{+}(x,k)\,f_{+}(y,-k)}{|a(k)|^{2}}.$
Substituting into (109) yields
$G_{q}(x,y;\lambda)=\frac{1}{2\pi}\int_{\mathbb{R}}\frac{f_{+}(x,k)\,f_{+}(y,-k)}{|a(k)|^{2}}\,\frac{dk}{k^{2}+\lambda}.$ (110)
We now integrate (110) against $d\rho(\lambda)$ and justify exchanging the $\lambda$- and $k$-integrals using only admissibility (98). Fix $x,y\in\mathbb{R}$. By Proposition 7.2, $a(k)$ is continuous on $\mathbb{R}\setminus\{0\}$, satisfies $|a(k)|\geq 1$ by (73), and $a(k)\to 1$ rapidly as $|k|\to\infty$. Moreover, for each fixed $x$ the map $k\mapsto f_{+}(x,k)$ is continuous on $\mathbb{R}\setminus\{0\}$ and has at most polynomial growth as $|k|\to\infty$; in particular, there is a constant $C_{x,y}$ such that
$\sup_{k\in\mathbb{R}\setminus\{0\}}\frac{|f_{+}(x,k)\,f_{+}(y,-k)|}{|a(k)|^{2}}\leq C_{x,y}.$
Therefore the integrand in (110) is dominated by $C_{x,y}(k^{2}+\lambda)^{-1}$, and Tonelli’s theorem gives
$\int_{0}^{\infty}\int_{\mathbb{R}}\frac{C_{x,y}}{k^{2}+\lambda}\,dk\,d\rho(\lambda)=C_{x,y}\int_{0}^{\infty}\left(\int_{\mathbb{R}}\frac{dk}{k^{2}+\lambda}\right)d\rho(\lambda)=C_{x,y}\pi\int_{0}^{\infty}\lambda^{-1/2}\,d\rho(\lambda)<\infty$
by (98). Hence Fubini applies and we may exchange the order of integration:
$K_{q,\rho}(x,y)=\int_{0}^{\infty}G_{q}(x,y;\lambda)\,d\rho(\lambda)=\frac{1}{2\pi}\int_{\mathbb{R}}\left(\int_{0}^{\infty}\frac{1}{k^{2}+\lambda}\,d\rho(\lambda)\right)\frac{f_{+}(x,k)\,f_{+}(y,-k)}{|a(k)|^{2}}\,dk.$
The inner $\lambda$-integral is exactly $\Phi_{\rho}(k^{2})$ by (99), proving (107).
For the Mercer expansion, fix a bounded interval $I\subset\mathbb{R}$. For each $\lambda>0$, the kernel $G_{q}(\cdot,\cdot;\lambda)$ is symmetric and positive definite on $\mathbb{R}$ because $(H_{q}+\lambda)^{-1}$ is a positive self-adjoint operator. Integrating against the positive measure $\rho$ preserves symmetry and positive definiteness, so $K_{q,\rho}|_{\overline{I}\times\overline{I}}$ is a continuous positive definite kernel on the compact space $\overline{I}$. Let $T_{I}\colon L^{2}(I)\to L^{2}(I)$ be the integral operator with kernel $K_{q,\rho}|_{I\times I}$. Continuity on $\overline{I}\times\overline{I}$ implies $K_{q,\rho}|_{I\times I}\in L^{2}(I\times I)$, hence $T_{I}$ is Hilbert-Schmidt and therefore compact [28]. Positivity and symmetry of the kernel imply $T_{I}$ is positive and self-adjoint. The spectral theorem provides eigenvalues $\lambda_{n}^{(I)}\geq 0$, $\lambda_{n}^{(I)}\downarrow 0$, and an $L^{2}(I)$-orthonormal eigenbasis $\{\psi_{n}^{(I)}\}_{n\geq 1}$. Since the kernel is continuous, $T_{I}(L^{2}(I))\subset C(\overline{I})$, so eigenfunctions with $\lambda_{n}^{(I)}>0$ admit continuous representatives. Mercer’s theorem then yields (108) with uniform convergence on $I\times I$, after reindexing the strictly positive eigenvalues so that $\lambda_{n}^{(I)}\in(0,\infty)$. ∎
Corollary 7.16 (Kernel sections as a canonical dictionary).
Let $q=\beta^{-\infty}(\varphi)$ and fix an admissible weight $\rho$. Then each section $K_{q,\rho}(\cdot,x)$ belongs to $\mathcal{H}_{q,\rho}$ and satisfies
$\|K_{q,\rho}(\cdot,x)\|_{\mathcal{H}_{q,\rho}}^{2}=K_{q,\rho}(x,x).$
If $(x_{j})_{j\geq 1}\subset\mathbb{R}$ is a sequence whose intersection with every bounded interval is dense, then the linear span of $\{K_{q,\rho}(\cdot,x_{j})\}_{j\geq 1}$ is dense in $\mathcal{H}_{q,\rho}$.
Proof.
The first statement is the reproducing property in $\mathcal{H}_{q,\rho}$ together with evaluation at $x$ applied to $K_{q,\rho}(\cdot,x)$. For density, let $x_{j}\to x$. The RKHS norm satisfies
$\|K_{q,\rho}(\cdot,x_{j})-K_{q,\rho}(\cdot,x)\|_{\mathcal{H}_{q,\rho}}^{2}=K_{q,\rho}(x_{j},x_{j})+K_{q,\rho}(x,x)-2\Re K_{q,\rho}(x_{j},x),$
which tends to 0 by continuity of $K_{q,\rho}$. Thus every kernel section is in the closure of the span of sections at a dense set of points, and finite linear combinations of kernel sections are dense by construction of the RKHS. ∎
Corollary 7.17 (Galerkin convergence for kernel dictionaries).
Let $q=\beta^{-\infty}(\varphi)$ and fix an admissible weight $\rho$, with RKHS $\mathcal{H}_{q,\rho}$. Let $T:\mathbb{R}\to\mathbb{R}$ be a map such that the composition operator $\mathcal{K}_{T}f:=f\circ T$ is bounded on $\mathcal{H}_{q,\rho}$. For $x_{1},\ldots,x_{N}\in\mathbb{R}$ set $V_{N}:=\operatorname{span}\{K_{q,\rho}(\cdot,x_{j})\}_{j=1}^{N}$ and let $P_{N}$ be the orthogonal projection of $\mathcal{H}_{q,\rho}$ onto $V_{N}$. If $\bigcup_{N}V_{N}$ is dense in $\mathcal{H}_{q,\rho}$, then the finite-dimensional operators $\mathcal{K}_{T}^{(N)}:=P_{N}\mathcal{K}_{T}P_{N}$ converge to $\mathcal{K}_{T}$ in the strong operator topology:
$\|\mathcal{K}_{T}f-\mathcal{K}_{T}^{(N)}f\|_{\mathcal{H}_{q,\rho}}\to 0\qquad\text{for every }f\in\mathcal{H}_{q,\rho}.$
Proof.
Density of $\bigcup_{N}V_{N}$ implies $P_{N}\to I$ strongly on $\mathcal{H}_{q,\rho}$. Thus $P_{N}f\to f$ for every $f$, and boundedness of $\mathcal{K}_{T}$ gives $\mathcal{K}_{T}(P_{N}f)\to\mathcal{K}_{T}f$. Applying $P_{N}$ again and using $P_{N}\to I$ strongly yields $P_{N}\mathcal{K}_{T}(P_{N}f)\to\mathcal{K}_{T}f$, i.e. $\mathcal{K}_{T}^{(N)}f\to\mathcal{K}_{T}f$. ∎
7.7. Marchenko reconstruction and the inverse scattering Bers map
Theorem 7.6 used the fact that the reflection coefficient determines the potential when no bound states are present. In the present sign convention $H_{q}=-\partial_{x}^{2}-q$ the reconstruction is governed by the Marchenko equation.
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and assume $H_{q}\geq 0$, so there are no bound states. Let $R(k)$ be its reflection coefficient. Since $R\in\mathscr{S}(\mathbb{R})$ on the Bers image, we may define the Marchenko kernel
$F(x):=\frac{1}{2\pi}\int_{\mathbb{R}}R(k)\mathrm{e}^{\mathrm{i}kx}\,\mathrm{d}k,\qquad x\in\mathbb{R},$ (111)
which is a Schwartz function. For each $x\in\mathbb{R}$, define the Hankel operator $\mathcal{H}_{x}$ on $L^{2}([x,\infty))$ by
$(\mathcal{H}_{x}h)(y):=\int_{x}^{\infty}F(y+z)\,h(z)\,\mathrm{d}z.$
This operator is Hilbert-Schmidt and hence compact.
Theorem 7.18 (Marchenko equation and reconstruction).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and assume $H_{q}\geq 0$. Let $R$ be its reflection coefficient and let $F$ be defined by (111). Then for each $x\in\mathbb{R}$ the integral equation
$K(x,y)+F(x+y)+\int_{x}^{\infty}K(x,z)F(z+y)\,\mathrm{d}z=0,\qquad y\geq x,$ (112)
has a unique solution $K(x,\cdot)\in L^{2}([x,\infty))$, and this solution is $C^{\infty}$ in $(x,y)$. The potential is recovered by
$q(x)=2\frac{\mathrm{d}}{\mathrm{d}x}K(x,x),$ (113)
and the right Jost solution is obtained from $K$ by the transformation formula
$f_{+}(x,k)=\mathrm{e}^{\mathrm{i}kx}+\int_{x}^{\infty}K(x,y)\mathrm{e}^{\mathrm{i}ky}\,\mathrm{d}y,\qquad k\in\mathbb{R}.$ (114)
In particular, $f_{+}$ solves (68) for the potential $q$ in (113), has the asymptotic (69) at $\infty$, and its reflection coefficient is $R$.
Proof.
For short-range one-dimensional Schrödinger operators, the Gel’fand-Levitan-Marchenko equation reconstructs the unique potential from its scattering data. In the no-bound-state case, the scattering data consist only of the reflection coefficient, and in the present convention $H_{q}=-\partial_{x}^{2}-q$ this yields (112)-(113); see [10, Theorem 1]. We record the points that are needed in our setting.
Since $F$ is Schwartz, $\mathcal{H}_{x}$ is Hilbert-Schmidt on $L^{2}([x,\infty))$. For reflection coefficients arising from a self-adjoint $H_{q}$ with no bound states, the Fredholm operator $I+\mathcal{H}_{x}$ is invertible for every $x$; equivalently, (112) has a unique $L^{2}$ solution $K(x,\cdot)$. Differentiating the equation under the integral sign shows that $K$ is smooth in $x$ and $y$, with all derivatives decaying rapidly as $y\to\infty$.
Given such a $K$, define $f_{+}$ by (114). Smoothness of $K$ allows differentiation under the integral sign, and a standard calculation gives that $f_{+}$ satisfies the differential equation (68) provided $q$ is defined by (113); the derivative in (113) is precisely the boundary term that cancels the singular contribution at $y=x$. The asymptotic $f_{+}(x,k)\sim\mathrm{e}^{\mathrm{i}kx}$ as $x\to+\infty$ follows from $K(x,\cdot)\in L^{1}([x,\infty))$ and decay of $K$. Finally, the scattering relation between $f_{+}$ at -$\infty$ and the free plane waves is encoded in the Marchenko equation and recovers the prescribed reflection coefficient $R$; this is the standard identification in the proof of the GLM theorem [10]. ∎
Corollary 7.19 (Constructive inversion of the scattering Bers map).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $R=\mathcal{S}(\varphi)$. Let $K$ be the Marchenko kernel from Theorem 7.18. Then the zero-energy Jost solution is
$f_{+}(x,0)=1+\int_{x}^{\infty}K(x,y)\,\mathrm{d}y,$ (115)
and the diffeomorphism class $\varphi]\in\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$ is recovered by
$\varphi(x)=x+\int_{-\infty}^{x}\bigl(f_{+}(t,0)^{-2}-1\bigr)\,\mathrm{d}t.$ (116)
Proof.
By definition of $\mathcal{S}$, the reflection coefficient $R$ arises from $q=\beta^{-\infty}(\varphi)$, which satisfies $H_{q}\geq 0$ by Lemma 7.3. The Marchenko reconstruction therefore yields the same potential $q$ and the corresponding right Jost family $f_{+}(x,k)$. Evaluating (114) at $k=0$ gives (115). Theorem 7.4 then identifies $\varphi$ by (78), which is exactly (116). ∎
8. Extension to Orlicz-Sobolev scales
This section extends the constructions of Sections 2 and 6 from the $W^{\infty,1}$ scale to Orlicz-Sobolev scales. We fix an admissible Young function $\Phi$ and work on $\mathbb{R}$ with Luxemburg norms. The Schwarzian $S$ and the $L^{p}$-Schwarzian family $S_{p}$ are as in Section 4.
A structural theme that will be used repeatedly is that the logarithmic coordinate
$\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$
simultaneously linearizes the diffeomorphism group and the associated Jacobian (density) geometry. In particular, the Orlicz setting admits a canonical statistical bundle over the Orlicz density manifold, together with a Luxemburg-Finsler structure and a Koopman-equivariant pullback action. These fiberwise constructions are intrinsic and do not reduce to a rephrasing of the $L^{p}$ case.
8.1. Orlicz-Sobolev spaces
Definition 8.1.
A Young function $\Psi\colon[0,\infty)\to[0,\infty)$ is a convex, increasing function such that $\Psi(0)=0$ and
$\lim_{t\to 0^{+}}\frac{\Psi(t)}{t}=0,\qquad\lim_{t\to\infty}\frac{\Psi(t)}{t}=\infty.$
Let $(X,\mu)$ be a $\sigma$-finite measure space. The Orlicz space $L^{\Psi}(\mu)$ consists of measurable functions $g$ such that $\int_{X}\Psi(|g|/\lambda)\,d\mu<\infty$ for some $\lambda>0$, equipped with the Luxemburg norm
$\|g\|_{L^{\Psi}(\mu)}=\inf\left\{\lambda>0\;\Big|\;\int_{X}\Psi\Big(\frac{|g|}{\lambda}\Big)\,d\mu\leq 1\right\}.$
On $\mathbb{R}$ we abbreviate $L^{\Psi}(\mathbb{R})=L^{\Psi}(dx)$. For $k\in\mathbb{N}$, the Orlicz-Sobolev space is
$W^{k,\Psi}(\mathbb{R})=\{g\in L^{\Psi}(\mathbb{R})\mid\partial_{x}^{j}g\in L^{\Psi}(\mathbb{R}),\;1\leq j\leq k\},\qquad\|g\|_{W^{k,\Psi}}=\sum_{j=0}^{k}\|\partial_{x}^{j}g\|_{L^{\Psi}}.$
We also set $W^{\infty,\Psi}(\mathbb{R})=\bigcap_{k\geq 0}W^{k,\Psi}(\mathbb{R})$ with the Fréchet topology induced by the seminorms $\|\cdot\|_{W^{k,\Psi}}$.
We will use the elementary Luxemburg estimate
$\|fg\|_{L^{\Psi}(\mu)}\leq\|f\|_{L^{\Psi}(\mu)}\|g\|_{L^{\infty}(\mu)}\qquad(f\in L^{\Psi}(\mu),\ g\in L^{\infty}(\mu)),$ (117)
which follows immediately from monotonicity of $\Psi$ in the definition of $\|\cdot\|_{L^{\Psi}(\mu)}$.
Lemma 8.2 (Algebra property on the smooth scale).
Assume $\Phi$ is such that the embedding $W^{1,\Phi}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$ is continuous. Then for every $k\geq 1$ the Banach space $W^{k,\Phi}(\mathbb{R})$ is a Banach algebra under pointwise multiplication, and the bilinear map $(f,g)\mapsto fg$ is continuous $W^{k,\Phi}\times W^{k,\Phi}\to W^{k,\Phi}$.
Proof.
Fix $k\geq 1$. If $f\in W^{k,\Phi}$ then $f^{(j)}\in W^{1,\Phi}$ for $0\leq j\leq k-1$, hence each $f^{(j)}$ is bounded. Leibniz’ rule gives $(fg)^{(j)}=\sum_{m=0}^{j}\binom{j}{m}f^{(m)}g^{(j-m)}$. For each term, place one factor in $L^{\Phi}$ and the other in $L^{\infty}$ using (117). Summing over $0\leq j\leq k$ yields $\|fg\|_{W^{k,\Phi}}\leq C_{k}\|f\|_{W^{k,\Phi}}\|g\|_{W^{k,\Phi}}$ for some $C_{k}$ depending only on $k$ and the embedding constant. ∎
8.2. Admissible Young functions
Definition 8.3.
A Young function $\Phi$ is called admissible if the following hold.
  • (i)
    $\Phi$ satisfies a $\Delta_{2}$-condition.
  • (ii)
    The embedding $W^{1,\Phi}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$ is continuous.
  • (iii)
    For every bounded set $B\subset W^{\infty,\Phi}(\mathbb{R})$, the Nemytskii maps $u\mapsto e^{\pm u}$ are Fréchet smooth as maps $B\to W^{\infty,\Phi}(\mathbb{R})$.
  • (iv)
    $W^{\infty,\Phi}(\mathbb{R})$ is closed under differentiation and pointwise multiplication, and the bilinear map $(f,g)\mapsto fg$ is continuous $W^{\infty,\Phi}(\mathbb{R})\times W^{\infty,\Phi}(\mathbb{R})\to W^{\infty,\Phi}(\mathbb{R})$.
By Lemma 8.2, condition (ii) already implies continuity of multiplication on each $W^{k,\Phi}$ with $k\geq 1$; condition (iv) enforces the corresponding closedness and continuity on the Fréchet intersection $W^{\infty,\Phi}$.
Remark 8.4 (Examples and non-examples of admissible Young functions).
Definition 8.3 is designed to capture the class of Orlicz scales for which one-dimensional Sobolev embedding and the usual composition and multiplication calculus hold. A concrete family beyond pure powers is
$\Phi_{p,\beta}(t)=t^{p}\bigl(\log(e+t)\bigr)^{\beta},\qquad p>1,\ \ \beta\in\mathbb{R},$
after modifying $\Phi_{p,\beta}$ near $t=0$ so that $\Phi_{p,\beta}(t)/t\to 0$ as $t\to 0^{+}$. These Young functions satisfy the global $\Delta_{2}$ condition and generate Orlicz spaces with the standard complementary-function duality theory, and in particular are compatible with the one-dimensional embedding $W^{1,\Phi_{p,\beta}}(\mathbb{R})\hookrightarrow L^{\infty}(\mathbb{R})$; see [27, 1, 20]. In this setting the Nemytskii maps $u\mapsto e^{\pm u}$ are Fréchet smooth on bounded subsets of $W^{\infty,\Phi_{p,\beta}}(\mathbb{R})$ and the space $W^{\infty,\Phi_{p,\beta}}(\mathbb{R})$ is stable under differentiation and pointwise multiplication, so $\Phi_{p,\beta}$ is admissible.
By contrast, the exponential-type Young functions $\Phi(t)=e^{t^{\alpha}}-1$ fail the global $\Delta_{2}$ condition for every $\alpha>0$, hence are excluded by item (i) in Definition 8.3; see [27].
8.3. Orlicz diffeomorphism groups
Definition 8.5.
Let $\Phi$ be admissible and define the model Fréchet space
$E_{\Phi}=W^{\infty,\Phi}(\mathbb{R})\cap L^{1}(\mathbb{R}),\qquad\|u\|_{E_{\Phi},k}=\|u\|_{L^{1}}+\sum_{j=0}^{k}\|\partial_{x}^{j}u\|_{L^{\Phi}},\qquad k\geq 0.$
The Orlicz diffeomorphism group $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ is the set of $C^{\infty}$ diffeomorphisms $\varphi\colon\mathbb{R}\to\mathbb{R}$ such that $\varphi^{\prime}>0$,
$u_{\varphi}=\log(\varphi^{\prime})\in E_{\Phi},$
and $\varphi$ is normalized at -$\infty$ by
$\lim_{x\to-\infty}(\varphi(x)-x)=0.$ (118)
For $\varphi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$ we write $\Phi_{\infty}(\varphi)=u_{\varphi}$.
Proposition 8.6.
Let $\Phi$ be admissible.
  • (i)
    For every $u\in E_{\Phi}$ the formula
    $\varphi_{u}(x)=x+\int_{-\infty}^{x}\bigl(e^{u(t)}-1\bigr)\,dt$ (119)
    defines an element $\varphi_{u}\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$ with $\Phi_{\infty}(\varphi_{u})=u$.
  • (ii)
    The map $\Phi_{\infty}\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})\to E_{\Phi}$ is bijective with inverse $u\mapsto\varphi_{u}$ given by (119). Transporting the Fréchet structure of $E_{\Phi}$ by $\Phi_{\infty}$ makes $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ into a Fréchet manifold.
  • (iii)
    In the global chart $\Phi_{\infty}$, the group law and inversion are
    $u\odot v=u\circ\varphi_{v}+v,\qquad u^{-1}=-\,u\circ\varphi_{u}^{-1}.$ (120)
    These maps are Fréchet smooth on bounded subsets of $E_{\Phi}$, hence $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ is a Fréchet Lie group with global chart $\Phi_{\infty}$.
Proof.
Let $u\in E_{\Phi}$. By admissibility (ii), $u\in L^{\infty}(\mathbb{R})$. Since $u\in L^{1}(\mathbb{R})$, the bound
$|e^{u}-1|\leq e^{\|u\|_{L^{\infty}}}\,|u|$
implies $e^{u}-1\in L^{1}(\mathbb{R})$, so (119) defines a $C^{1}$ function with $\varphi_{u}^{\prime}(x)=e^{u(x)}>0$. Differentiating (119) repeatedly and using admissibility (iii) shows that $\varphi_{u}\in C^{\infty}(\mathbb{R})$ and that $\varphi_{u}^{(j)}\in L^{\Phi}(\mathbb{R})$ for all $j\geq 2$. The normalization (118) follows from $e^{u}-1\in L^{1}$ by construction, and $\varphi_{u}$ is strictly increasing with $\varphi_{u}(x)\to\pm\infty$ as $x\to\pm\infty$, hence $\varphi_{u}$ is a diffeomorphism. This proves (i).
If $\varphi\in\mathrm{Diff}^{-\Phi}(\mathbb{R})$ and $u=\log(\varphi^{\prime})$, then $\varphi^{\prime}=e^{u}$ and (118) forces $\varphi$ to satisfy (119). Thus $\Phi_{\infty}$ is bijective with inverse $u\mapsto\varphi_{u}$, proving (ii).
For (iii), (120) is immediate from $(\varphi_{u}\circ\varphi_{v})^{\prime}=(\varphi_{u}^{\prime}\circ\varphi_{v})\varphi_{v}^{\prime}$ and $(\varphi_{u}^{-1})^{\prime}=(\varphi_{u}^{\prime})^{-1}\circ\varphi_{u}^{-1}$.
It remains to justify that the group operations are Fréchet smooth in the logarithmic coordinate. In view of
$u\odot v=u\circ\varphi_{v}+v,\qquad u^{-1}=-\,u\circ\varphi_{u}^{-1},$ (121)
it is enough to prove the Fréchet smoothness of the maps
$(u,v)\longmapsto u\circ\varphi_{v},\qquad v\longmapsto\varphi_{v}^{-1},$ (122)
as maps on bounded subsets of $E_{\Phi}$, with values in $E_{\Phi}$ for the first map and in the diffeomorphism class for the second map. Throughout we fix a bounded set $B\subset E_{\Phi}$ and work with constants depending only on $B$ and on finitely many seminorms. For $v\in B$ we have $v\in L^{\infty}(\mathbb{R})$ by admissibility, hence
$0<m_{B}\leq\varphi_{v}^{\prime}(x)=e^{v(x)}\leq M_{B}<\infty,\qquad\bigl\|(\varphi_{v}^{-1})^{\prime}\bigr\|_{L^{\infty}}\leq m_{B}^{-1}.$ (123)
In particular $\varphi_{v}$ and $\varphi_{v}^{-1}$ are globally bi-Lipschitz, and the change of variables $y=\varphi_{v}(x)$ is legitimate in Orlicz modular estimates. We first record the basic Orlicz pullback estimate in the following. Let $\psi$ be any $C^{1}$ diffeomorphism with $\psi^{\prime}>0$ and $\|(\psi^{-1})^{\prime}\|_{L^{\infty}}\leq C$. If $\lambda>0$ satisfies $\int_{\mathbb{R}}\Phi(|f|/\lambda)\,dx\leq 1$, then
$\int_{\mathbb{R}}\Phi\left(\frac{|f(\psi(x))|}{\lambda}\right)dx=\int_{\mathbb{R}}\Phi\left(\frac{|f(y)|}{\lambda}\right)(\psi^{-1})^{\prime}(y)\,dy\leq C\int_{\mathbb{R}}\Phi\left(\frac{|f(y)|}{\lambda}\right)dy\leq C.$ (124)
Using convexity and $\Phi(0)=0$, one has $\Phi(t/C)\leq C^{-1}\Phi(t)$ for $C\geq 1$, hence (124) gives
$\|f\circ\psi\|_{L^{\Phi}}\leq C\,\|f\|_{L^{\Phi}}.$ (125)
The same change of variables yields $\|f\circ\psi\|_{L^{1}}\leq C\,\|f\|_{L^{1}}$. We now apply (125) to $\psi=\varphi_{v}$ with $v\in B$, using (123). To control higher derivatives in $W^{k,\Phi}$, we differentiate compositions directly. For $j\geq 1$, repeated differentiation of $f\circ\psi$ gives an identity of the form
$\partial_{x}^{j}(f\circ\psi)=\sum_{m=1}^{j}(f^{(m)}\circ\psi)\,Q_{j,m}\bigl(\psi^{\prime},\psi^{\prime\prime},\dots,\psi^{(j-m+1)}\bigr),$ (126)
where each $Q_{j,m}$ is a universal polynomial with no constant term and depends only on $j,m$. For $\psi=\varphi_{v}$ with $v\in B$, admissibility implies $\partial_{x}^{\ell}v\in W^{1,\Phi}$ for every $\ell\geq 0$, hence all derivatives $\partial_{x}^{\ell}v$ are bounded. Differentiating $\varphi_{v}^{\prime}=e^{v}$ shows that $\varphi_{v}^{(\ell)}$ is a finite sum of products of $e^{v}$ with bounded functions built from derivatives of $v$, hence
$\sup_{v\in B}\|\varphi_{v}^{(\ell)}\|_{L^{\infty}}<\infty,\qquad\sup_{v\in B}\|(\varphi_{v}^{-1})^{(\ell)}\|_{L^{\infty}}<\infty$ (127)
for each fixed $\ell$. Combining (126), the product estimate $\|FG\|_{L^{\Phi}}\leq\|F\|_{L^{\Phi}}\|G\|_{L^{\infty}}$, and (125) with $\psi=\varphi_{v}$, we obtain for each fixed $k$
$\sup_{v\in B}\,\|f\circ\varphi_{v}\|_{W^{k,\Phi}}\leq C_{B,k}\,\|f\|_{W^{k,\Phi}},\qquad\sup_{v\in B}\,\|f\circ\varphi_{v}\|_{L^{1}}\leq C_{B}\,\|f\|_{L^{1}}.$ (128)
Thus $f\mapsto f\circ\varphi_{v}$ acts continuously on $E_{\Phi}$, uniformly for $v\in B$.
We next prove Fréchet smoothness of $(u,v)\mapsto u\circ\varphi_{v}$. Fix $(u,v)\in E_{\Phi}\times B$. In the $u$ variable, the map is linear and its differential is
$D_{u}(u\circ\varphi_{v})\cdot\delta u=\delta u\circ\varphi_{v},\qquad\delta u\in E_{\Phi},$ (129)
which is continuous by (128). To differentiate with respect to $v$, we first record the variation of $\varphi_{v}$. From
$\varphi_{v}(x)=x+\int_{-\infty}^{x}(e^{v(t)}-1)\,dt$
we obtain for $\delta v\in E_{\Phi}$
$D_{v}\varphi_{v}\cdot\delta v\,(x)=\int_{-\infty}^{x}e^{v(t)}\,\delta v(t)\,dt,\qquad\partial_{x}\bigl(D_{v}\varphi_{v}\cdot\delta v\bigr)=e^{v}\delta v.$ (130)
Since $e^{v}$ is bounded and $\delta v\in L^{1}\cap L^{\infty}$, the function $D_{v}\varphi_{v}\cdot\delta v$ is bounded and belongs to $W^{k+1,\Phi}$ for every $k$, with seminorms controlled on $B$ by Leibniz estimates applied to $\partial_{x}^{\ell}(e^{v}\delta v)$.
For the composition, the first $v$ derivative is obtained from the one-dimensional chain rule
$D_{v}(u\circ\varphi_{v})\cdot\delta v=(u^{\prime}\circ\varphi_{v})\,(D_{v}\varphi_{v}\cdot\delta v).$ (131)
Using (128), the boundedness of $u^{\prime}$ on bounded sets of $E_{\Phi}$, and the bounds on $D_{v}\varphi_{v}\cdot\delta v$ just noted, we get continuity of the bilinear map $(u,\delta v)\mapsto D_{v}(u\circ\varphi_{v})\cdot\delta v$ as a map $E_{\Phi}\times E_{\Phi}\to E_{\Phi}$, uniformly for $v\in B$. Higher Fréchet derivatives with respect to $v$ follow by differentiating (131) repeatedly. Each differentiation either lands on a factor $u^{(m)}\circ\varphi_{v}$ or on a factor involving $D_{v}^{r}\varphi_{v}$. The resulting expressions are finite sums of terms of the form
$(u^{(m)}\circ\varphi_{v})\,\prod_{\alpha=1}^{r}\partial_{x}^{\ell_{\alpha}}\bigl(D_{v}^{s_{\alpha}}\varphi_{v}\cdot(\delta v_{\alpha,1},\dots,\delta v_{\alpha,s_{\alpha}})\bigr),$
with $m\leq r$ and $\sum s_{\alpha}=r$, multiplied by bounded coefficients depending on finitely many derivatives of $\varphi_{v}$. All such terms lie in $E_{\Phi}$ because exactly one factor is placed in $L^{\Phi}$ and the remaining factors are controlled in $L^{\infty}$, using (125), (126), (127), and admissibility. This proves that $(u,v)\mapsto u\circ\varphi_{v}$ is Fréchet smooth on $E_{\Phi}\times B$.
Finally we treat inversion. Let $v\in B$ and put $\psi_{v}=\varphi_{v}^{-1}$. The identity $\varphi_{v}\circ\psi_{v}=\mathrm{Id}$ implies, after differentiating in the $v$ direction, that for $\delta v\in E_{\Phi}$ the variation $\delta\psi=D_{v}\psi_{v}\cdot\delta v$ satisfies
$(\varphi_{v}^{\prime}\circ\psi_{v})\,\delta\psi+(D_{v}\varphi_{v}\cdot\delta v)\circ\psi_{v}=0,$ (132)
hence
$D_{v}\psi_{v}\cdot\delta v=-\,\frac{(D_{v}\varphi_{v}\cdot\delta v)\circ\psi_{v}}{\varphi_{v}^{\prime}\circ\psi_{v}}.$ (133)
The denominator is bounded above and below by (123), and the numerator is controlled in $W^{k,\Phi}$ by the already established pullback bounds applied to $\psi_{v}$ together with the bounds on $D_{v}\varphi_{v}$. Thus $v\mapsto\psi_{v}$ is $C^{1}$ as a map on $B$ into the $C^{\infty}$ topology with Orlicz control. To construct higher differentials of $v\mapsto\psi_{v}$ we differentiate the identity
$\varphi_{v}\circ\psi_{v}=\operatorname{id}$ (134)
as a relation in the parameter $v\in E_{\Phi}$. For $r\geq 1$ and directions $h_{1},\dots,h_{r}\in E_{\Phi}$ set
$\psi_{v}^{(r)}[h_{1},\dots,h_{r}]=D_{v}^{r}\psi_{v}[h_{1},\dots,h_{r}].$
The first differential is determined by (132) and (133). For $r\geq 2$, applying $D_{v}^{r}$ to (134) and isolating the unique term in which all $r$ derivatives fall on $\psi_{v}$ produces
$(\varphi_{v}^{\prime}\circ\psi_{v})\,\psi_{v}^{(r)}[h_{1},\dots,h_{r}]=-\,\mathcal{R}_{r}(v;\,h_{1},\dots,h_{r}),$ (135)
where $\mathcal{R}_{r}(v;\,h_{1},\dots,h_{r})$ is a finite sum of terms of the form
$\Big(\partial_{x}^{\ell}D_{v}^{m}\varphi_{v}[h_{i_{1}},\dots,h_{i_{m}}]\circ\psi_{v}\Big)\,\prod_{j=1}^{\ell}\psi_{v}^{(r_{j})}[h_{J_{j}}],$ (136)
with $m\geq 0$, $\ell\geq 0$, $r_{j}\geq 1$, and
$m+r_{1}+\cdots+r_{\ell}=r,\qquad\{i_{1},\dots,i_{m}\}\sqcup J_{1}\sqcup\cdots\sqcup J_{\ell}=\{1,\dots,r\}.$
In particular, $\mathcal{R}_{r}$ depends only on $D_{v}^{s}\varphi_{v}$ with $s\leq r$ and on $\psi_{v}^{(s)}$ with $s<r$.
For later reference we record the explicit formula at order $r=2$, obtained by differentiating (132) once more:
$\psi_{v}^{(2)}[h_{1},h_{2}]$ $=-(\varphi_{v}^{\prime}\circ\psi_{v})^{-1}\Big((D_{v}^{2}\varphi_{v}[h_{1},h_{2}])\circ\psi_{v}+(D_{v}\varphi_{v}[h_{1}])^{\prime}\circ\psi_{v}\ \psi_{v}^{(1)}[h_{2}]$
$+(D_{v}\varphi_{v}[h_{2}])^{\prime}\circ\psi_{v}\ \psi_{v}^{(1)}[h_{1}]+(\varphi_{v}^{\prime\prime}\circ\psi_{v})\ \psi_{v}^{(1)}[h_{1}]\psi_{v}^{(1)}[h_{2}]\Big).$ (137)
We now estimate the right hand side of (135) on a fixed bounded set $B\subset E_{\Phi}$. First, by admissibility and the bounds (123), the multiplier $(\varphi_{v}^{\prime}\circ\psi_{v})^{-1}$ belongs to $L^{\infty}(\mathbb{R})$ and satisfies
$\sup_{v\in B}\big\|(\varphi_{v}^{\prime}\circ\psi_{v})^{-1}\big\|_{L^{\infty}}<\infty.$ (138)
Second, the $v$-derivatives of $\varphi_{v}$ are explicit: for $m\geq 1$,
$D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}](x)=\int_{-\infty}^{x}\mathrm{e}^{v(t)}\,h_{1}(t)\cdots h_{m}(t)\,\mathrm{d}t,\qquad\partial_{x}D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}]=\mathrm{e}^{v}\,h_{1}\cdots h_{m}.$ (139)
Since $B$ is bounded in $E_{\Phi}$, we have $\mathrm{e}^{v}\in L^{\infty}$ uniformly on $B$ and $h_{j}\in L^{\infty}$ for each $h_{j}\in E_{\Phi}$; therefore
$\partial_{x}D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}]\in L^{\Phi}(\mathbb{R})\cap L^{1}(\mathbb{R}),\qquad D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}]\in L^{\infty}(\mathbb{R}),$
and the corresponding seminorms are controlled by the product estimate $\|fg\|_{L^{\Phi}}\leq\|f\|_{L^{\Phi}}\|g\|_{L^{\infty}}$ and the definition of $E_{\Phi}$. Moreover, the pullback bounds (125) apply to $\psi_{v}$ uniformly for $v\in B$ by (123), hence
$\sup_{v\in B}\ \big\|\partial_{x}^{\ell}D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}]\circ\psi_{v}\big\|_{L^{\Phi}}\leq C_{B,\ell}\ \big\|\partial_{x}^{\ell}D_{v}^{m}\varphi_{v}[h_{1},\dots,h_{m}]\big\|_{L^{\Phi}}.$ (140)
Assume inductively that for some $r\geq 2$ and each $s<r$ the multilinear maps $(h_{1},\dots,h_{s})\mapsto\psi_{v}^{(s)}[h_{1},\dots,h_{s}]$ exist and are continuous $E_{\Phi}^{s}\to L^{\infty}(\mathbb{R})$, uniformly for $v\in B$. Then every term in (136) lies in $L^{\Phi}(\mathbb{R})$ after applying one $x$-derivative, because one places the factor involving $\partial_{x}^{\ell}D_{v}^{m}\varphi_{v}\circ\psi_{v}$ in $L^{\Phi}$ and all remaining factors are controlled in $L^{\infty}$, using (140) and the induction hypothesis. Consequently $\mathcal{R}_{r}(v;\,h_{1},\dots,h_{r})$ defines a continuous $r$-linear map $E_{\Phi}^{r}\to L^{\Phi}(\mathbb{R})$, uniformly for $v\in B$.
Solving (135) by multiplication with $(\varphi_{v}^{\prime}\circ\psi_{v})^{-1}$ and using (138) yields existence and continuity of $\psi_{v}^{(r)}$ as an $r$-linear map $E_{\Phi}^{r}\to L^{\infty}(\mathbb{R})$, uniformly for $v\in B$. This completes the induction and proves that $v\mapsto\psi_{v}=\varphi_{v}^{-1}$ is Fréchet smooth on bounded sets. With (121), the smoothness of $u\odot v$ and $u^{-1}$ follows.
8.4. Orlicz fiber bundles
We now introduce the Orlicz density manifold and its canonical statistical bundle. In the present one-dimensional scalar setting, the crucial point is that the normalization (118) makes the Jacobian map globally invertible on the Orlicz scale, which yields a canonical global trivialization of the bundle by composition with $\varphi_{u}^{-1}$.
Definition 8.7 (Orlicz density manifold).
Define the Jacobian (pullback) map
$J\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})\to\{\text{positive Borel measures on }\mathbb{R}\},\qquad J(\varphi)=\varphi^{*}dx=\varphi^{\prime}(x)\,dx.$
The Orlicz density manifold is
$|\Omega|_{-\Phi}(\mathbb{R}):=J(\operatorname{Diff}^{-\Phi}(\mathbb{R})).$
For $u\in E_{\Phi}$ we write $\mu_{u}:=J(\varphi_{u})=e^{u(x)}\,dx$.
Proposition 8.8 (Global chart for $|\Omega|_{-\Phi}(\mathbb{R})$).
The map
$\mathfrak{J}_{\infty}\colon E_{\Phi}\to|\Omega|_{-\Phi}(\mathbb{R}),\qquad\mathfrak{J}_{\infty}(u)=\mu_{u}=e^{u}\,dx,$
is a bijection with inverse $\mu\mapsto\log(d\mu/dx)$. Transporting the Fréchet structure of $E_{\Phi}$ by $\mathfrak{J}_{\infty}$ makes $|\Omega|_{-\Phi}(\mathbb{R})$ into a Fréchet manifold, and $\mathfrak{J}_{\infty}$ is a global chart.
Proof.
If $\mu\in|\Omega|_{-\Phi}$ then $\mu=\varphi^{\prime}dx$ for some $\varphi\in\operatorname{Diff}^{-\Phi}$. Setting $u=\log(\varphi^{\prime})$ gives $u\in E_{\Phi}$ and $\mu=e^{u}dx$. Conversely, for any $u\in E_{\Phi}$, Proposition 8.6(i) produces $\varphi_{u}\in\operatorname{Diff}^{-\Phi}$ and hence $\mu_{u}=J(\varphi_{u})\in|\Omega|_{-\Phi}$. Uniqueness follows because $\varphi^{\prime}\equiv e^{u}$ determines $\varphi$ up to an additive constant, and (118) fixes that constant. The asserted manifold structure is obtained by transport of structure. ∎
Definition 8.9 (Orlicz statistical bundle).
For each $k\in\mathbb{N}_{0}\cup\{\infty\}$ define the Orlicz-Sobolev statistical bundle of order $k$ by
$\pi_{\Phi}^{(k)}\colon\mathcal{S}_{\Phi}^{(k)}\to|\Omega|_{-\Phi}(\mathbb{R}),\qquad\mathcal{S}_{\Phi}^{(k)}:=\bigsqcup_{\mu\in|\Omega|_{-\Phi}(\mathbb{R})}W^{k,\Phi}(\mu),\qquad\pi_{\Phi}^{(k)}(\mu,f)=\mu,$
where $W^{k,\Phi}(\mu)$ is the Orlicz-Sobolev space built from Luxemburg norms with respect to $\mu$. For $k=0$ we simply write $\mathcal{S}_{\Phi}:=\mathcal{S}_{\Phi}^{(0)}$.
Proposition 8.10 (Canonical trivialization).
Let $\mu=\mu_{u}\in|\Omega|_{-\Phi}(\mathbb{R})$ with $u\in E_{\Phi}$. Define the linear map
$\mathcal{T}_{u}\colon L^{\Phi}(\mu_{u})\to L^{\Phi}(dx),\qquad(\mathcal{T}_{u}f)(y)=f(\varphi_{u}^{-1}(y)).$
Then $\mathcal{T}_{u}$ is an isometric isomorphism of Banach spaces
$\|f\|_{L^{\Phi}(\mu_{u})}=\|\mathcal{T}_{u}f\|_{L^{\Phi}(dx)}.$
Consequently $\mathcal{S}_{\Phi}$ is canonically a trivial Banach bundle over $|\Omega|_{-\Phi}(\mathbb{R})$ with typical fiber $L^{\Phi}(dx)$.
Moreover, for each $k\geq 1$ the same formula defines a continuous linear isomorphism
$\mathcal{T}_{u}\colon W^{k,\Phi}(\mu_{u})\to W^{k,\Phi}(dx),$
and on bounded subsets of $E_{\Phi}$ the operator norms of $\mathcal{T}_{u}$ and $\mathcal{T}_{u}^{-1}$ on $W^{k,\Phi}$ are uniformly controlled.
Proof.
For $k=0$, by definition of $\mu_{u}=\varphi_{u}^{\prime}dx$ and the change of variables $y=\varphi_{u}(x)$,
$\int_{\mathbb{R}}\Phi\!\left(\frac{|f(x)|}{\lambda}\right)\,d\mu_{u}(x)=\int_{\mathbb{R}}\Phi\!\left(\frac{|f(\varphi_{u}^{-1}(y))|}{\lambda}\right)\,dy=\int_{\mathbb{R}}\Phi\!\left(\frac{|\mathcal{T}_{u}f(y)|}{\lambda}\right)\,dy.$
Taking the infimum over $\lambda>0$ yields $\|f\|_{L^{\Phi}(\mu_{u})}=\|\mathcal{T}_{u}f\|_{L^{\Phi}(dx)}$. Thus $\mathcal{T}_{u}$ is an isometric isomorphism with inverse $g\mapsto g\circ\varphi_{u}$.
Assume now $k\geq 1$ and write $\psi=\varphi_{u}^{-1}$. Introduce the first-order differential operator on smooth functions of $x$,
$L_{u}:=e^{-u(x)}\partial_{x}.$
We claim that for every $C^{1}$ function $h$,
$\partial_{y}(h\circ\psi)=(L_{u}h)\circ\psi.$ (141)
Indeed, writing $x=\psi(y)$, one has $\psi^{\prime}(y)=1/\varphi_{u}^{\prime}(x)=e^{-u(x)}$, hence
$\partial_{y}(h(\psi(y)))=h^{\prime}(x)\,\psi^{\prime}(y)=h^{\prime}(x)e^{-u(x)}=(L_{u}h)(x),$
which is (141). Iterating (141) yields by induction
$\partial_{y}^{j}(h\circ\psi)=(L_{u}^{j}h)\circ\psi,\qquad j\geq 0.$ (142)
We next expand $L_{u}^{j}$ in the standard derivative basis. For each $j\geq 1$ there exist smooth coefficient functions $a_{j,m}(u)$, $1\leq m\leq j$, such that
$L_{u}^{j}h=\sum_{m=1}^{j}a_{j,m}(u)\,\partial_{x}^{m}h,$ (143)
and each $a_{j,m}(u)$ is a universal polynomial expression in $e^{-u}$ and derivatives $u^{\prime},\dots,u^{(j-m)}$. This is proved by induction on $j$. For $j=1$, $L_{u}h=e^{-u}h^{\prime}$ so $a_{1,1}=e^{-u}$. Assume (143) holds at order $j$. Then
$L_{u}^{j+1}h=e^{-u}\partial_{x}\!\left(\sum_{m=1}^{j}a_{j,m}\,\partial_{x}^{m}h\right)=\sum_{m=1}^{j}e^{-u}a_{j,m}\,\partial_{x}^{m+1}h+\sum_{m=1}^{j}e^{-u}(\partial_{x}a_{j,m})\,\partial_{x}^{m}h,$
which has the form (143) at order $j+1$ with the recursion
$a_{j+1,m}=e^{-u}a_{j,m-1}+e^{-u}(\partial_{x}a_{j,m}),\qquad a_{j,0}=0=a_{j,j+1}.$
This proves (143) and the stated dependence of the coefficients on $u$ and its derivatives.
Because $u\in E_{\Phi}\subset W^{\infty,\Phi}(\mathbb{R})$ and $\Phi$ is admissible, each derivative $u^{(\ell)}$ lies in $W^{1,\Phi}(\mathbb{R})$ and is therefore bounded. Consequently, for each fixed $j$ the coefficient functions $a_{j,m}(u)$ are bounded:
$\|a_{j,m}(u)\|_{L^{\infty}}<\infty,\qquad 1\leq m\leq j.$ (144)
Now fix $f\in W^{k,\Phi}(\mu_{u})$ and set $F=\mathcal{T}_{u}f=f\circ\psi$. Combining (142) with (143) gives, for $1\leq j\leq k$,
$\partial_{y}^{j}F=\bigl(L_{u}^{j}f\bigr)\circ\psi=\left(\sum_{m=1}^{j}a_{j,m}(u)\,f^{(m)}\right)\circ\psi.$
Use the exact Luxemburg pullback identity (already proved at $k=0$) in the form
[…]
and the product estimate $\|gh\|_{L^{\Phi}(\mu_{u})}\leq\|g\|_{L^{\Phi}(\mu_{u})}\|h\|_{L^{\infty}}$. Then
$\|\partial_{y}^{j}F\|_{L^{\Phi}(dx)}=\left\|\sum_{m=1}^{j}a_{j,m}(u)\,f^{(m)}\right\|_{L^{\Phi}(\mu_{u})}\leq\sum_{m=1}^{j}\|a_{j,m}(u)\|_{L^{\infty}}\,\|f^{(m)}\|_{L^{\Phi}(\mu_{u})}.$
Summing over $0\leq j\leq k$ yields
$\|F\|_{W^{k,\Phi}(dx)}\leq C_{u,k}\,\|f\|_{W^{k,\Phi}(\mu_{u})},$
with $C_{u,k}$ depending only on finitely many $L^{\infty}$ bounds of $u$ and its derivatives up to order k1k-1. This proves continuity of $\mathcal{T}_{u}\colon W^{k,\Phi}(\mu_{u})\to W^{k,\Phi}(dx)$.
For the inverse map $g\mapsto g\circ\varphi_{u}$, apply the same argument with roles of $x$ and $y$ reversed. Equivalently, define $\widetilde{L}_{u}:=e^{u(\varphi_{u}^{-1}(y))}\partial_{y}$ on functions of $y$; then $\partial_{x}(g\circ\varphi_{u})=(\widetilde{L}_{u}g)\circ\varphi_{u}$ and hence $\partial_{x}^{j}(g\circ\varphi_{u})=(\widetilde{L}_{u}^{j}g)\circ\varphi_{u}$. Expanding $\widetilde{L}_{u}^{j}$ yields bounded coefficients depending on finitely many derivatives of $u$, and the same pullback identity (145) (with $\psi$ replaced by $\varphi_{u}$) gives a bound
$\|g\circ\varphi_{u}\|_{W^{k,\Phi}(\mu_{u})}\leq C^{\prime}_{u,k}\,\|g\|_{W^{k,\Phi}(dx)}.$
Thus $\mathcal{T}_{u}$ is a continuous linear isomorphism with continuous inverse on $W^{k,\Phi}$.
Finally, if $u$ ranges in a bounded subset of $E_{\Phi}$, then $\|u^{(\ell)}\|_{L^{\infty}}$ is uniformly bounded for $\ell\leq k-1$, so the constants in (144) and hence the operator norms of $\mathcal{T}_{u}$ and $\mathcal{T}_{u}^{-1}$ on $W^{k,\Phi}$ are uniformly controlled on that bounded set.
Theorem 8.11 (Luxemburg-Finsler structure on the Orlicz density manifold).
Let $\Phi$ be admissible. For $\mu\in|\Omega|_{-\Phi}(\mathbb{R})$ define the tangent model at $\mu$ by
$T_{\mu}|\Omega|_{-\Phi}(\mathbb{R})\cong E_{\Phi},\qquad\delta\mu\leftrightarrow\delta u\quad\text{via}\quad\mu=\mu_{u}=e^{u}dx,\ \ \delta\mu=(\delta u)\,\mu.$
Then the assignment
$\mathcal{F}_{\Phi}(\mu,\delta u):=\|\delta u\|_{L^{\Phi}(\mu)}$
defines a (weak) Finsler structure on $|\Omega|_{-\Phi}(\mathbb{R})$. The induced path-length pseudo-distance is
$d_{\Phi}(\mu_{0},\mu_{1})=\inf_{\gamma}\int_{0}^{1}\big\|\dot{u}(t)\big\|_{L^{\Phi}(\mu_{u(t)})}\,dt,$
where the infimum is taken over all piecewise $C^{1}$ curves $t\mapsto\mu_{u(t)}$ joining $\mu_{0}$ to $\mu_{1}$.
If $\Phi(t)=t^{p}/p$ then $\mathcal{F}_{\Phi}(\mu,\delta u)=p^{-1/p}\|\delta u\|_{L^{p}(\mu)}$, so $d_{\Phi}$ agrees with the $L^{p}$ Fisher-Rao length structure up to the fixed factor $p^{-1/p}$.
Proof.
By Proposition 8.8, the map
$\mathfrak{J}_{\infty}\colon E_{\Phi}\to|\Omega|_{-\Phi}(\mathbb{R}),\qquad u\mapsto\mu_{u}:=e^{u}\,dx$
is a global chart. We use this chart to identify tangent vectors and to define a fiberwise norm.
Let $u\in E_{\Phi}$ and let $u(\cdot)$ be a $C^{1}$ curve in $E_{\Phi}$ with $u(0)=u$. Set $\mu(t):=\mu_{u(t)}=e^{u(t)}dx$. Then for every test function $\phi\in C_{c}^{\infty}(\mathbb{R})$,
$\frac{d}{dt}\Big|_{t=0}\int_{\mathbb{R}}\phi\,d\mu(t)=\int_{\mathbb{R}}\phi(x)\,\dot{u}(0,x)\,e^{u(x)}\,dx=\int_{\mathbb{R}}\phi\,(\dot{u}(0))\,d\mu_{u}.$
Thus, in the chart $\mathfrak{J}_{\infty}$, the velocity of the curve $\mu(t)$ at $t=0$ is the signed measure
$\dot{\mu}(0)=(\dot{u}(0))\,\mu_{u},$
and the map $\dot{u}(0)\mapsto(\dot{u}(0))\,\mu_{u}$ identifies the model space $E_{\Phi}$ with the tangent space at $\mu_{u}$. In particular, the identification
$T_{\mu_{u}}|\Omega|_{-\Phi}(\mathbb{R})\cong E_{\Phi},\qquad\delta u\longleftrightarrow\delta\mu:=(\delta u)\,\mu_{u},$
is exactly the differential of the chart $\mathfrak{J}_{\infty}$ at $u$.
We now define the fiberwise norm. For $\delta\mu=(\delta u)\,\mu_{u}$ we set
$\mathcal{F}_{\Phi}(\mu_{u},\delta u):=\|\delta u\|_{L^{\Phi}(\mu_{u})}.$
This is well-defined because $\mu_{u}$ is a positive $\sigma$-finite measure and $L^{\Phi}(\mu_{u})$ is a Banach space with Luxemburg norm. Moreover, for each fixed base point $\mu_{u}$, the map $\delta u\mapsto\|\delta u\|_{L^{\Phi}(\mu_{u})}$ is a norm on the tangent model: it is nonnegative, vanishes only when $\delta u=0$ $\mu_{u}$-a.e., it is positively homogeneous, and it satisfies the triangle inequality. These properties are immediate from the Luxemburg definition.
Let $\mu(t)=\mu_{u(t)}$ be a piecewise $C^{1}$ curve in $|\Omega|_{-\Phi}(\mathbb{R})$. Because $\mathfrak{J}_{\infty}$ is a global chart, we represent $\mu(t)$ by a piecewise $C^{1}$ curve $u(t)\in E_{\Phi}$. The computation above shows that $\dot{\mu}(t)=(\dot{u}(t))\,\mu_{u(t)}$, hence we define the length of $\mu(\cdot)$ by
$\mathrm{Len}_{\Phi}(\mu)=\int_{0}^{1}\|\dot{u}(t)\|_{L^{\Phi}(\mu_{u(t)})}\,dt.$
This length does not depend on the choice of parametrization: if $\tau\colon[0,1]\to[0,1]$ is a $C^{1}$ increasing bijection and $\widetilde{u}(s)=u(\tau(s))$, then $\dot{\widetilde{u}}(s)=\dot{u}(\tau(s))\,\tau^{\prime}(s)$, so by homogeneity of the Luxemburg norm,
$\|\dot{\widetilde{u}}(s)\|_{L^{\Phi}(\mu_{\widetilde{u}(s)})}=\tau^{\prime}(s)\,\|\dot{u}(\tau(s))\|_{L^{\Phi}(\mu_{u(\tau(s))})}.$
Therefore
$\int_{0}^{1}\|\dot{\widetilde{u}}(s)\|_{L^{\Phi}(\mu_{\widetilde{u}(s)})}\,ds=\int_{0}^{1}\|\dot{u}(t)\|_{L^{\Phi}(\mu_{u(t)})}\,dt,$
which shows that $\mathrm{Len}_{\Phi}$ depends only on the geometric curve.
Define
$d_{\Phi}(\mu_{0},\mu_{1}):=\inf\Bigl\{\mathrm{Len}_{\Phi}(\mu)\,\Big|\,\mu\text{ piecewise }C^{1},\ \mu(0)=\mu_{0},\ \mu(1)=\mu_{1}\Bigr\}.$
Then $d_{\Phi}\geq 0$ and $d_{\Phi}(\mu,\mu)=0$ since constant curves have zero length. If $\mu$ is any curve from $\mu_{0}$ to $\mu_{1}$, its reversal is a curve from $\mu_{1}$ to $\mu_{0}$ with the same length, so $d_{\Phi}$ is symmetric. Finally, if $\mu_{01}$ joins $\mu_{0}$ to $\mu_{1}$ and $\mu_{12}$ joins $\mu_{1}$ to $\mu_{2}$, the concatenation $\mu_{02}$ joins $\mu_{0}$ to $\mu_{2}$ and satisfies $\mathrm{Len}_{\Phi}(\mu_{02})=\mathrm{Len}_{\Phi}(\mu_{01})+\mathrm{Len}_{\Phi}(\mu_{12})$, hence taking infima gives the triangle inequality
$d_{\Phi}(\mu_{0},\mu_{2})\leq d_{\Phi}(\mu_{0},\mu_{1})+d_{\Phi}(\mu_{1},\mu_{2}).$
Thus $d_{\Phi}$ is a pseudo-distance and the family of norms $\mathcal{F}_{\Phi}(\mu,\cdot)$ defines a (weak) Finsler structure on $|\Omega|_{-\Phi}(\mathbb{R})$ in the usual sense of a fiberwise norm inducing a reparametrization-invariant length functional.
For the final statement, take $\Phi(t)=t^{p}/p$. Then the Luxemburg norm on $(\mathbb{R},\mu)$ is characterized by
$\|f\|_{L^{\Phi}(\mu)}=\inf\left\{\lambda>0\ \Big|\ \frac{1}{p}\int_{\mathbb{R}}\left(\frac{|f|}{\lambda}\right)^{p}\,d\mu\leq 1\right\}=p^{-1/p}\|f\|_{L^{p}(\mu)}.$
Therefore $\mathcal{F}_{\Phi}(\mu,\delta u)=p^{-1/p}\|\delta u\|_{L^{p}(\mu)}$, and the induced length functional, hence the induced pseudo-distance, agrees with the $L^{p}$ Fisher-Rao length structure up to the fixed factor $p^{-1/p}$. ∎
8.5. Orlicz-Young duality on fibers
The bundle $\mathcal{S}_{\Phi}$ carries a canonical cotangent bundle given by complementary Orlicz spaces. We also record a fiberwise Legendre map that will serve as the intrinsic duality map for the Luxemburg norm.
Definition 8.12 (Complementary Young function).
The complementary Young function to $\Phi$ is
$\Psi(s)=\sup_{t\geq 0}\,(st-\Phi(t)).$
The pair $(\Phi,\Psi)$ satisfies Young’s inequality
$st\leq\Phi(t)+\Psi(s),\qquad s,t\geq 0.$ (146)
Proposition 8.13 (Orlicz-Hölder inequality and duality).
Let $\mu$ be $\sigma$-finite and let $(\Phi,\Psi)$ be complementary Young functions. Then for all $f\in L^{\Phi}(\mu)$ and $g\in L^{\Psi}(\mu)$ one has $fg\in L^{1}(\mu)$ and
$\left|\int_{\mathbb{R}}fg\,d\mu\right|\leq 2\,\|f\|_{L^{\Phi}(\mu)}\,\|g\|_{L^{\Psi}(\mu)}.$
If moreover $\Phi$ satisfies the $\Delta_{2}$ condition, then the Banach dual of $L^{\Phi}(\mu)$ is canonically identified with $L^{\Psi}(\mu)$ by the pairing
$\langle f,g\rangle_{\mu}=\int_{\mathbb{R}}fg\,d\mu.$
Both statements are standard in Orlicz space theory; see [27, 20].
Theorem 8.14 (Legendre map for the Luxemburg norm).
Assume in addition that $\Phi\in C^{1}((0,\infty))$ is strictly convex. Fix $\mu\in|\Omega|_{-\Phi}(\mathbb{R})$ and set $\rho_{\Phi}(f)=\int\Phi(|f|)\,d\mu$. For each nonzero $f\in L^{\Phi}(\mu)$ let $\lambda=\|f\|_{L^{\Phi}(\mu)}$, so that $\rho_{\Phi}(f/\lambda)=1$. Define
$\mathcal{J}_{\Phi}(f):=\frac{\lambda\,\Phi^{\prime}(|f|/\lambda)}{\int_{\mathbb{R}}\Phi^{\prime}(|f|/\lambda)\,|f|\,d\mu}\,\mathrm{sgn}(f).$
Then $\mathcal{J}_{\Phi}(f)\in L^{\Psi}(\mu)$ and the Luxemburg norm is Gâteaux differentiable at $f$ with derivative
$D\|f\|_{L^{\Phi}(\mu)}\cdot h=\int_{\mathbb{R}}\mathcal{J}_{\Phi}(f)\,h\,d\mu,\qquad h\in L^{\Phi}(\mu).$
Moreover $\|\mathcal{J}_{\Phi}(f)\|_{(L^{\Phi}(\mu))^{*}}=1$ and $\langle f,\mathcal{J}_{\Phi}(f)\rangle_{\mu}=\|f\|_{L^{\Phi}(\mu)}$.
Proof.
Since $\Phi$ is strictly convex, the equality $\rho_{\Phi}(f/\lambda)=1$ characterizes $\lambda=\|f\|_{L^{\Phi}(\mu)}$ uniquely. Consider the function
$F(\varepsilon,\eta)=\int_{\mathbb{R}}\Phi\!\left(\frac{|f+\varepsilon h|}{\lambda+\eta}\right)d\mu,$
defined near $(\varepsilon,\eta)=(0,0)$. By the chain rule and $\Phi\in C^{1}((0,\infty))$, $F$ is $C^{1}$ and $F(0,0)=1$. Differentiate the identity $F(\varepsilon,\eta(\varepsilon))\equiv 1$ implicitly, where $\eta(\varepsilon)$ is chosen so that $\lambda+\eta(\varepsilon)=\|f+\varepsilon h\|_{L^{\Phi}(\mu)}$. At $\varepsilon=0$ this yields
$0=\partial_{\varepsilon}F(0,0)+\eta^{\prime}(0)\,\partial_{\eta}F(0,0).$
A direct differentiation gives
$\partial_{\varepsilon}F(0,0)=\frac{1}{\lambda}\int_{\mathbb{R}}\Phi^{\prime}(|f|/\lambda)\,\mathrm{sgn}(f)\,h\,d\mu,\qquad\partial_{\eta}F(0,0)=-\frac{1}{\lambda^{2}}\int_{\mathbb{R}}\Phi^{\prime}(|f|/\lambda)\,|f|\,d\mu.$
Solving for $\eta^{\prime}(0)$ yields the stated formula for $D\|f\|\cdot h$ with the displayed $\mathcal{J}_{\Phi}(f)$. Finally, Orlicz-Hölder shows $\mathcal{J}_{\Phi}(f)\in L^{\Psi}(\mu)$ and provides $\|\mathcal{J}_{\Phi}(f)\|_{(L^{\Phi})^{*}}\leq 1$, while evaluating $D\|f\|\cdot f$ gives $\langle f,\mathcal{J}_{\Phi}(f)\rangle_{\mu}=\|f\|$, forcing $\|\mathcal{J}_{\Phi}(f)\|_{(L^{\Phi})^{*}}=1$. ∎
8.6. Koopman operators on Orlicz fibers
We record two complementary Koopman structures: (i) a fiberwise Koopman pullback giving an isometric bundle action covering the pullback action on densities, and (ii) a fixed-fiber Koopman representation on $L^{\Phi}(dx)$ (and $W^{k,\Phi}(dx)$) with explicit operator norm bounds and a clean spectral annulus estimate.
Definition 8.15 (Fiberwise Koopman pullback).
Let $\psi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$. Define the pullback action on densities by
$\psi^{*}\mu(A)=\mu(\psi(A)).$
For each $\mu\in|\Omega|_{-\Phi}(\mathbb{R})$ define the Koopman pullback map
$\mathcal{K}_{\psi}^{\mu}\colon L^{\Phi}(\mu)\to L^{\Phi}(\psi^{*}\mu),\qquad(\mathcal{K}_{\psi}^{\mu}f)(x)=f(\psi(x)).$
Theorem 8.16 (Koopman bundle isometries).
For every $\psi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$ and every $\mu\in|\Omega|_{-\Phi}(\mathbb{R})$, the map $\mathcal{K}_{\psi}^{\mu}\colon L^{\Phi}(\mu)\to L^{\Phi}(\psi^{*}\mu)$ is a well-defined linear isometry:
$\|f\circ\psi\|_{L^{\Phi}(\psi^{*}\mu)}=\|f\|_{L^{\Phi}(\mu)}.$
The assignment
$\mathcal{K}_{\psi}\colon\mathcal{S}_{\Phi}\to\mathcal{S}_{\Phi},\qquad(\mu,f)\longmapsto(\psi^{*}\mu,f\circ\psi),$
defines a bundle automorphism covering $\mu\mapsto\psi^{*}\mu$. Moreover, $\mathcal{K}_{\psi\circ\chi}=\mathcal{K}_{\chi}\circ\mathcal{K}_{\psi}$ and $\mathcal{K}_{\psi^{-1}}=\mathcal{K}_{\psi}^{-1}$.
Proof.
By definition of pullback measure,
$\int_{\mathbb{R}}\Phi\!\left(\frac{|f\circ\psi(x)|}{\lambda}\right)\,d(\psi^{*}\mu)(x)=\int_{\mathbb{R}}\Phi\!\left(\frac{|f(y)|}{\lambda}\right)\,d\mu(y).$
Taking the Luxemburg infimum over $\lambda$ gives the norm identity. Functoriality of pullback and composition yields the cocycle identities. ∎
The $W^{k,\Phi}$ mapping properties of composition operators are proved here in the same analytic framework as the $W^{\infty,1}$ case treated earlier, but we record the Orlicz-Sobolev estimates explicitly since they are needed for the bundle-level constructions below.
Definition 8.17 (Fixed-fiber Koopman representation).
Define the (right) composition operator on the reference Orlicz space by
$U_{\psi}\colon L^{\Phi}(dx)\to L^{\Phi}(dx),\qquad(U_{\psi}f)(x)=f(\psi^{-1}(x)).$
For $k\geq 1$ define $U_{\psi}$ on $W^{k,\Phi}(dx)$ by the same formula.
Proposition 8.18 (Operator norm bound).
Let $\psi$ be an orientation-preserving $C^{1}$ diffeomorphism with $1\leq\|\psi^{\prime}\|_{L^{\infty}}<\infty$. Then $U_{\psi}$ is bounded on $L^{\Phi}(dx)$ and
$\|U_{\psi}\|_{\mathcal{L}(L^{\Phi}\to L^{\Phi})}\leq\|\psi^{\prime}\|_{L^{\infty}}.$
If in addition $\|(\psi^{-1})^{\prime}\|_{L^{\infty}}<\infty$, then $U_{\psi}$ is invertible with inverse $U_{\psi^{-1}}$ and
$\|U_{\psi}^{-1}\|\leq\|(\psi^{-1})^{\prime}\|_{L^{\infty}}.$
For $\psi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$, one has $\psi^{\prime}(x)\rightarrow 1$ as $|x|\rightarrow\infty$, hence $\|\psi^{\prime}\|_{L^{\infty}}\geq 1$ and $\|(\psi^{-1})^{\prime}\|_{L^{\infty}}\geq 1$ so both bounds apply. They are moreover uniformly on bounded subsets of $E_{\Phi}$ via (123).
Proof.
Let $\psi$ be an orientation-preserving $C^{1}$ diffeomorphism with $\|\psi^{\prime}\|_{L^{\infty}}<\infty$ and define $U_{\psi}f=f\circ\psi^{-1}$. Fix $f\in L^{\Phi}(dx)$ and $\lambda>0$. With the change of variables $x=\psi(y)$, so that $dx=\psi^{\prime}(y)\,dy$, one obtains
$\int_{\mathbb{R}}\Phi\!\left(\frac{|U_{\psi}f(x)|}{\lambda}\right)\,dx=\int_{\mathbb{R}}\Phi\!\left(\frac{|f(y)|}{\lambda}\right)\,\psi^{\prime}(y)\,dy\leq\|\psi^{\prime}\|_{L^{\infty}}\int_{\mathbb{R}}\Phi\!\left(\frac{|f(y)|}{\lambda}\right)\,dy.$
If $\int_{\mathbb{R}}\Phi(|f|/\alpha)\,dy\leq 1$, then convexity of $\Phi$ and $\Phi(0)=0$ imply $\Phi(t/\|\psi^{\prime}\|_{L^{\infty}})\leq\|\psi^{\prime}\|_{L^{\infty}}^{-1}\Phi(t)$ for $\|\psi^{\prime}\|_{L^{\infty}}\geq 1$, hence
$\int_{\mathbb{R}}\Phi\!\left(\frac{|U_{\psi}f(x)|}{\alpha\|\psi^{\prime}\|_{L^{\infty}}}\right)\,dx\leq 1.$
Taking the Luxemburg infimum over $\alpha$ yields
$\|U_{\psi}f\|_{L^{\Phi}(dx)}\leq\|\psi^{\prime}\|_{L^{\infty}}\,\|f\|_{L^{\Phi}(dx)},$
so $\|U_{\psi}\|\leq\|\psi^{\prime}\|_{L^{\infty}}$.
If in addition $\|(\psi^{-1})^{\prime}\|_{L^{\infty}}<\infty$, apply the same argument to $\psi^{-1}$ to obtain $\|U_{\psi^{-1}}\|\leq\|(\psi^{-1})^{\prime}\|_{L^{\infty}}$. Since $U_{\psi^{-1}}=U_{\psi}^{-1}$, this proves invertibility and the bound on $\|U_{\psi}^{-1}\|$. ∎
Proposition 8.19 (Boundedness on Orlicz-Sobolev spaces).
Fix $k\geq 1$. On bounded subsets of $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ the map $\psi\mapsto U_{\psi}$ defines a uniformly bounded family in $\mathcal{L}(W^{k,\Phi}\to W^{k,\Phi})$. Moreover, for each fixed $k$ the map
$(\psi,f)\longmapsto U_{\psi}f$
is Fréchet smooth on bounded subsets of $\operatorname{Diff}^{-\Phi}(\mathbb{R})\times W^{k,\Phi}(\mathbb{R})$.
Proof.
Fix $k\geq 1$. Let $B\subset\operatorname{Diff}^{-\Phi}(\mathbb{R})$ be bounded in the global logarithmic chart $\Phi_{\infty}$, so that $\{\log(\psi^{\prime}):\psi\in B\}$ is bounded in $E_{\Phi}$. By (123), there exist constants $0<m_{B}\leq M_{B}<\infty$ such that for all $\psi\in B$,
$0<m_{B}\leq\psi^{\prime}(x)\leq M_{B},\qquad\|(\psi^{-1})^{\prime}\|_{L^{\infty}}\leq m_{B}^{-1}.$
Write $\eta=\psi^{-1}$ and $U_{\psi}f=f\circ\eta$.
For $j\geq 1$,
$\partial_{x}^{j}(f\circ\eta)=\sum_{m=1}^{j}(f^{(m)}\circ\eta)\,Q_{j,m}\bigl(\eta^{\prime},\eta^{\prime\prime},\dots,\eta^{(j-m+1)}\bigr),$
where $Q_{j,m}$ is a universal polynomial with no constant term depending only on $j,m$. Boundedness of $\{\log(\psi^{\prime}):\psi\in B\}$ in $E_{\Phi}$ implies uniform bounds on $\psi^{(\ell)}$ in $L^{\infty}$ for each fixed $\ell$. Implicit differentiation of $\psi\circ\eta=\operatorname{id}$ then yields uniform bounds on $\eta^{(\ell)}$ in $L^{\infty}$ for each fixed $\ell\leq k$. Hence for each fixed $j\leq k$ the coefficient $Q_{j,m}(\eta^{\prime},\dots,\eta^{(j-m+1)})$ is bounded in $L^{\infty}$ uniformly for $\psi\in B$.
The Orlicz pullback estimate (125) applied to $\eta$ gives
$\|f^{(m)}\circ\eta\|_{L^{\Phi}(dx)}\leq C_{B}\,\|f^{(m)}\|_{L^{\Phi}(dx)}.$
Using $\|FG\|_{L^{\Phi}}\leq\|F\|_{L^{\Phi}}\|G\|_{L^{\infty}}$ termwise yields
$\|\partial_{x}^{j}(f\circ\eta)\|_{L^{\Phi}}\leq C_{B,j}\sum_{m=1}^{j}\|f^{(m)}\|_{L^{\Phi}}.$
Summing over $0\leq j\leq k$ gives
$\sup_{\psi\in B}\|U_{\psi}f\|_{W^{k,\Phi}}\leq C_{B,k}\,\|f\|_{W^{k,\Phi}},$
so $\psi\mapsto U_{\psi}$ is uniformly bounded on $B$ as an operator on $W^{k,\Phi}$.
For Fréchet smoothness on bounded sets, work in the logarithmic coordinate $v=\Phi_{\infty}(\psi)\in E_{\Phi}$ so that $\psi=\varphi_{v}$ and $U_{\psi}f=f\circ\varphi_{v}^{-1}$. By Proposition 8.6(iii), the map $v\mapsto\varphi_{v}^{-1}$ is Fréchet smooth on bounded subsets of $E_{\Phi}$ in the $C^{\infty}$ topology with Orlicz control. For fixed $k$, the map $(\eta,f)\mapsto f\circ\eta$ from bounded sets with uniform $L^{\infty}$ bounds on $\eta^{\prime},\dots,\eta^{(k)}$ and $f\in W^{k,\Phi}$ into $W^{k,\Phi}$ is given by the displayed chain-rule expansion and involves only multiplication by bounded coefficients and pullback by $\eta$. Differentiating these expressions in $\eta$ produces finite sums of the same type, and the pullback and product estimates control each term. Composing with the smooth map $v\mapsto\eta=\varphi_{v}^{-1}$ yields Fréchet smoothness of $(v,f)\mapsto f\circ\varphi_{v}^{-1}$ on bounded sets, hence of $(\psi,f)\mapsto U_{\psi}f$. ∎
Corollary 8.20 (Spectral annulus bound).
Let $\psi$ be a bi-Lipschitz $C^{1}$ diffeomorphism with $\|\psi^{\prime}\|_{L^{\infty}}<\infty$ and $\|(\psi^{-1})^{\prime}\|_{L^{\infty}}<\infty$. Then $U_{\psi}$ is bounded and invertible on $L^{\Phi}(dx)$ and
$\sigma(U_{\psi})\subset\left\{\lambda\in\mathbb{C}\;\Big|\;\|(\psi^{-1})^{\prime}\|_{L^{\infty}}^{-1}\leq|\lambda|\leq\|\psi^{\prime}\|_{L^{\infty}}\right\}.$
The same inclusion holds for the spectrum of $U_{\psi}$ on $W^{k,\Phi}(\mathbb{R})$ for each fixed $k\geq 1$.
Proof.
For any bounded invertible operator $T$ on a Banach space, $\sigma(T)\subset\{\lambda:|\lambda|\leq\|T\|\}$ and $\sigma(T)\subset\{\lambda:|\lambda|\geq\|T^{-1}\|^{-1}\}$. Apply this to $T=U_{\psi}$ and use Proposition 8.18 (and Proposition 8.19 in the Sobolev case). ∎
8.7. Orlicz Bers embedding
Theorem 8.21.
Let $\Phi$ be admissible. The map
$\beta_{\Phi}\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})\longrightarrow W^{\infty,\Phi}(\mathbb{R}),\qquad\beta_{\Phi}(\varphi)=\frac{1}{2}S(\varphi),$
is Fréchet smooth. Since $S(A\circ\varphi)=S(\varphi)$ for every $A\in\operatorname{Aff}(\mathbb{R})$, the map factors through the affine quotient and induces a Fréchet smooth map
$\beta_{\Phi}\colon\operatorname{Diff}^{-\Phi}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\longrightarrow W^{\infty,\Phi}(\mathbb{R}).$
The induced map is injective and an immersion.
Proof.
Let $\varphi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$ and write $u=\log(\varphi^{\prime})\in E_{\Phi}$. As recalled in Section 6,
$S(\varphi)=u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}.$
By admissibility (ii) and Lemma 8.2 (applied on each $W^{k,\Phi}$, hence on $W^{\infty,\Phi}$), $u^{\prime}$ is bounded and $(u^{\prime})^{2}\in W^{\infty,\Phi}(\mathbb{R})$, while $u^{\prime\prime}\in W^{\infty,\Phi}(\mathbb{R})$ by definition of $E_{\Phi}$. Hence $S(\varphi)\in W^{\infty,\Phi}(\mathbb{R})$ and the map $u\mapsto u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}$ is Fréchet smooth $E_{\Phi}\to W^{\infty,\Phi}$. By Proposition 8.6(ii) this implies $\beta_{\Phi}$ is Fréchet smooth on $\operatorname{Diff}^{-\Phi}(\mathbb{R})$.
If $A\in\operatorname{Aff}(\mathbb{R})$ then $S(A)=0$ and the cocycle identity for the Schwarzian gives $S(A\circ\varphi)=S(\varphi)$, so $\beta_{\Phi}$ descends to the affine quotient.
For injectivity on the quotient, suppose $S(\varphi)=S(\psi)$. Then $\chi=\varphi\circ\psi^{-1}$ satisfies $S(\chi)=0$ by the cocycle identity, hence $\chi$ is affine. Therefore $\varphi$ and $\psi$ represent the same affine class.
For the immersion property, compute in logarithmic coordinates:
$\mathcal{S}(u)=u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2},\qquad D\mathcal{S}(u)\cdot\delta u=\delta u^{\prime\prime}-u^{\prime}\delta u^{\prime}.$
Assume $D\mathcal{S}(u)\cdot\delta u=0$ and set $w=\delta u^{\prime}$. Then $w^{\prime}-u^{\prime}w=0$, so $(e^{-u}w)^{\prime}=0$ and $w=c\,e^{u}$ for a constant $c$. Since $u\in L^{\infty}(\mathbb{R})$ one has $e^{u}\geq e^{-\|u\|_{L^{\infty}}}>0$ on $\mathbb{R}$. Because $dx(\mathbb{R})=\infty$ and $\Phi(t)>0$ for $t>0$, this implies $e^{u}\notin L^{\Phi}(\mathbb{R})$ unless $c=0$. But $\delta u\in E_{\Phi}$ implies $w=\delta u^{\prime}\in L^{\Phi}(\mathbb{R})$, hence $c=0$ and $w=0$. Therefore $\delta u$ is constant, and since $\delta u\in L^{\Phi}(\mathbb{R})$ the constant must be 0. Thus $D\mathcal{S}(u)$ is injective and $\beta_{\Phi}$ is an immersion. ∎
8.8. Orlicz stability of $S_{p}\to S$
Proposition 8.22.
Let $\Phi$ be admissible, let $k\geq 0$, and let $\varphi\in\operatorname{Diff}^{-\Phi}(\mathbb{R})$ with $u=\log(\varphi^{\prime})$ and $S=S(\varphi)$. Then
$S_{p}(\varphi)\longrightarrow S(\varphi)\qquad\text{in }W^{k,\Phi}(\mathbb{R})\text{ as }p\to\infty.$
Moreover, as $p\to\infty$ one has the asymptotic expansion
$S_{p}(\varphi)=S+\frac{1}{p}\left(uS+\frac{3}{2}(u^{\prime})^{2}\right)+\frac{1}{p^{2}}\left(\frac{u^{2}}{2}S+\frac{3}{2}u(u^{\prime})^{2}\right)+O_{W^{k,\Phi}}(p^{-3}),$ (147)
where the remainder estimate is uniform on subsets of $\operatorname{Diff}^{-\Phi}(\mathbb{R})$ on which $\|u\|_{L^{\infty}}$ and finitely many $W^{k+2,\Phi}$ seminorms of $u$ are bounded.
Proof.
Recall that
$S_{p}(\varphi)=\left(S(\varphi)+\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}\right)\varphi^{\prime 1/p}.$
Writing $\varphi^{\prime}=\mathrm{e}^{u}$ and $\varphi^{\prime\prime}/\varphi^{\prime}=u^{\prime}$ gives
$S_{p}(\varphi)=\left(S+\frac{3}{2p}(u^{\prime})^{2}\right)\mathrm{e}^{u/p},$ (148)
where $S=u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}$.
Fix $k\geq 0$. Since $\Phi$ is admissible, $u,u^{\prime}\in L^{\infty}(\mathbb{R})$ and multiplication is continuous on $W^{k,\Phi}(\mathbb{R})$. In particular,
$S\in W^{k,\Phi}(\mathbb{R}),\qquad(u^{\prime})^{2}\in W^{k,\Phi}(\mathbb{R}),$
and all products below are taken in $W^{k,\Phi}(\mathbb{R})$.
Let $z_{p}=u/p$. Then $z_{p}\in W^{k+2,\Phi}(\mathbb{R})$ and $\|z_{p}\|_{W^{k+2,\Phi}}=p^{-1}\|u\|_{W^{k+2,\Phi}}$, so $z_{p}\to 0$ in $W^{k+2,\Phi}(\mathbb{R})$ and in $L^{\infty}(\mathbb{R})$. Consider the Nemytskii map
$E\colon W^{k+2,\Phi}(\mathbb{R})\to W^{k+2,\Phi}(\mathbb{R}),\qquad E(z)=\mathrm{e}^{z}.$
By admissibility (iii), $E$ is Fréchet smooth on bounded sets, and its derivatives are given by
$DE(z)\cdot h=\mathrm{e}^{z}h,\qquad D^{2}E(z)\cdot(h_{1},h_{2})=\mathrm{e}^{z}h_{1}h_{2},\qquad D^{3}E(z)\cdot(h_{1},h_{2},h_{3})=\mathrm{e}^{z}h_{1}h_{2}h_{3}.$
Let $B\subset W^{k+2,\Phi}(\mathbb{R})$ be a bounded set on which $\|u\|_{L^{\infty}}$ and $\|u\|_{W^{k+2,\Phi}}$ are bounded; then $\{tz_{p}:t\in[0,1],\,p\geq 1\}\subset B$. Since multiplication is continuous on $W^{k+2,\Phi}(\mathbb{R})$, there is a constant $C_{B,k}$ such that for all $z\in B$ and all $h\in W^{k+2,\Phi}(\mathbb{R})$,
$\bigl\|D^{3}E(z)\cdot(h,h,h)\bigr\|_{W^{k+2,\Phi}}=\|\mathrm{e}^{z}h^{3}\|_{W^{k+2,\Phi}}\leq C_{B,k}\,\|h\|_{W^{k+2,\Phi}}^{3}.$ (149)
The Banach-space Taylor formula with integral remainder gives
$E(z_{p})=E(0)+DE(0)\cdot z_{p}+\frac{1}{2}D^{2}E(0)\cdot(z_{p},z_{p})+\frac{1}{2}\int_{0}^{1}(1-t)^{2}\,D^{3}E(tz_{p})\cdot(z_{p},z_{p},z_{p})\,\mathrm{d}t.$ (150)
Since $E(0)=1$, $DE(0)\cdot z_{p}=z_{p}$, and $D^{2}E(0)\cdot(z_{p},z_{p})=z_{p}^{2}$, we obtain
$\mathrm{e}^{u/p}=1+\frac{u}{p}+\frac{u^{2}}{2p^{2}}+R_{p},$ (151)
where
$R_{p}=\frac{1}{2}\int_{0}^{1}(1-t)^{2}\,D^{3}E(tz_{p})\cdot(z_{p},z_{p},z_{p})\,\mathrm{d}t.$
Using (149) in (150) yields
$\|R_{p}\|_{W^{k+2,\Phi}}\leq\frac{1}{2}\int_{0}^{1}(1-t)^{2}\,C_{B,k}\,\|z_{p}\|_{W^{k+2,\Phi}}^{3}\,\mathrm{d}t\leq C_{B,k}\,\|z_{p}\|_{W^{k+2,\Phi}}^{3}=O(p^{-3}),$ (152)
uniformly on $B$. In particular, $\|R_{p}\|_{W^{k,\Phi}}=O(p^{-3})$ as well.
Substituting (151) into (148) gives
$S_{p}(\varphi)=\left(S+\frac{3}{2p}(u^{\prime})^{2}\right)\left(1+\frac{u}{p}+\frac{u^{2}}{2p^{2}}+R_{p}\right).$
Expanding and grouping terms up to order $p^{-2}$ yields (147), and the remainder is $O_{W^{k,\Phi}}(p^{-3})$ by (152) and continuity of multiplication in $W^{k,\Phi}(\mathbb{R})$ on bounded sets. The convergence $S_{p}(\varphi)\to S(\varphi)$ in $W^{k,\Phi}(\mathbb{R})$ is the leading term of the expansion. ∎
9. Projective curvature of densities and forced sign-change theorems
9.1. Asymptotically Lebesgue densities
Fix $dx$ as the reference density on $\mathbb{R}$.
Definition 9.1 (Asymptotically Lebesgue densities).
The manifold of asymptotically Lebesgue densities is
$|\Omega|_{\infty,1}(\mathbb{R})=\{\mu=g\,dx\mid g>0,\;g-1\in W^{\infty,1}(\mathbb{R})\}.$
The condition $g-1\in W^{1,1}(\mathbb{R})$ forces $g(x)\to 1$ as $x\to\pm\infty$.
Lemma 9.2 (Regularity, boundedness, and asymptotics).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$. Then $g$ admits a $C^{\infty}$ representative,
$g^{(k)}\in L^{1}(\mathbb{R})\quad\forall k\geq 1,\qquad\lim_{x\to\pm\infty}g^{(k)}(x)=0\quad\forall k\geq 1,\qquad\lim_{x\to\pm\infty}g(x)=1.$
In particular there exist $0<m\leq M<\infty$ with $m\leq g\leq M$ on $\mathbb{R}$, and every derivative $g^{(k)}$ is bounded. Moreover, if $F\in C^{\infty}((0,\infty))$ has bounded derivatives on $[m,M]$, then $F(g)-F(1)\in W^{\infty,1}(\mathbb{R})$.
Proof.
Since $g-1\in W^{k,1}(\mathbb{R})$ for every $k$, the one-dimensional Sobolev embedding yields a representative in $C^{k-1}(\mathbb{R})$ with (k1)(k-1)st derivative absolutely continuous. Varying $k$ gives a $C^{\infty}$ representative and $g^{(k)}\in L^{1}(\mathbb{R})$ for all $k\geq 1$. For $k\geq 1$, we have $g^{(k)}\in W^{1,1}(\mathbb{R})$, hence $g^{(k)}$ has finite limits at $\pm\infty$. Since also $g^{(k)}\in L^{1}(\mathbb{R})$, these limits must be 0. The limit $g(\pm\infty)=1$ follows from $g-1\in W^{1,1}(\mathbb{R})$. Continuity and the asymptotic limit give global upper and lower bounds on $g$. Finally, the boundedness of $g^{(k)}$ follows from $g^{(k)}\in W^{1,1}(\mathbb{R})$, and the last statement follows by iterated chain rule together with the algebra property of $W^{k,1}(\mathbb{R})$ and boundedness of $F^{(j)}$ on $[m,M]$. ∎
9.2. The Jacobian isomorphism
Definition 9.3 (Jacobian map).
The Jacobian map is
$J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R}),\qquad J(\varphi)=\varphi^{\prime}\,dx.$
Theorem 9.4 (Jacobian correspondence).
$J$ is a Fréchet diffeomorphism with inverse
$J^{-1}(g\,dx)(x)=x+\int_{-\infty}^{x}(g(t)-1)\,dt.$
Moreover:
  1. (i)
    (Linearization intertwining) The diagram
    $\operatorname{Diff}^{-\infty}(\mathbb{R})$${\lvert\Omega\rvert}_{\infty,1}(\mathbb{R})$$W^{\infty,1}(\mathbb{R})$$W^{\infty,1}(\mathbb{R})$$\scriptstyle{J}$$\scriptstyle{\Phi_{p}}$$\scriptstyle{\Psi_{p}}$
    commutes, where $\Psi_{p}(g\,dx)=p(g^{1/p}-1)$ for $p<\infty$ and $\Psi_{\infty}(g\,dx)=\log g$.
  2. (ii)
    (Finsler isometry) $J$ is an isometry from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ to $(|\Omega|_{\infty,1}(\mathbb{R}),F_{p}^{\mathrm{FR}})$.
  3. (iii)
    (Equivariance) $J(\varphi\circ\psi)=\psi^{*}J(\varphi)$.
  4. (iv)
    (Schwarzian compatibility) Setting $S(\mu)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ for $f=\log(\mu/dx)$, we have $S(J(\varphi))=S(\varphi)$.
Proof.
Bijectivity follows from the fundamental theorem of calculus: if $\varphi^{\prime}\,dx=g\,dx$ with $g-1\in W^{\infty,1}$ then $\varphi(x)=x+\int_{-\infty}^{x}(g(t)-1)\,dt$, which lies in $\operatorname{Diff}^{-\infty}(\mathbb{R})$, and conversely $g=\varphi^{\prime}$ satisfies $g-1\in W^{\infty,1}$. Smoothness of $J$ follows from smoothness of $\varphi\mapsto\varphi^{\prime}$ and continuity of induced pullback maps in the Fréchet topologies; smoothness of $J^{-1}$ is immediate from the explicit formula since $g\mapsto\int_{-\infty}^{\cdot}(g-1)$ is continuous linear on $W^{\infty,1}$.
For the Finsler isometry, let $\delta\varphi=u\circ\varphi$. Then $D_{\varphi}J(\delta\varphi)=(u^{\prime}\circ\varphi)\varphi^{\prime}\,dx$ and
$F_{p}^{\mathrm{FR}}(J(\varphi),D_{\varphi}J(\delta\varphi))^{p}=\int_{\mathbb{R}}|u^{\prime}\circ\varphi|^{p}\,\varphi^{\prime}\,dx=\int_{\mathbb{R}}|u^{\prime}|^{p}\,dy=F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)^{p}.$
The remaining claims are direct computations. ∎
Corollary 9.5 (Transferred Fisher-Rao geodesics).
All geometric structures on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ transfer to $|\Omega|_{\infty,1}(\mathbb{R})$ via $J$. For $p<\infty$, the $L^{p}$ Fisher-Rao geodesic from $\mu_{0}=g_{0}\,dx$ to $\mu_{1}=g_{1}\,dx$ is
$\mu(t)=\Bigl((1-t)g_{0}^{1/p}+tg_{1}^{1/p}\Bigr)^{p}\,dx.$
9.3. Score functions, projective curvature, and Fisher information
Definition 9.6 (Score).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$, the score function is
$s_{\mu}=\frac{g^{\prime}}{g}=(\log g)^{\prime}.$
Proposition 9.7 (Schwarzian as score curvature).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and $s=s_{\mu}$. Then
$S(\mu)=s^{\prime}-\frac{1}{2}s^{2}.$
Moreover, $S(\mu)=0$ in $|\Omega|_{\infty,1}(\mathbb{R})$ if and only if $\mu=dx$.
Proof.
The identity is the definition of $S(\mu)$ with $f=\log g$ and $s=f^{\prime}$. If $S(\mu)=0$, then $s^{\prime}=\frac{1}{2}s^{2}$. By Lemma 9.2, $s\in C^{0}(\mathbb{R})$ with $s(\pm\infty)=0$. The ODE implies $s$ is nondecreasing. Since the limits at both ends coincide, $s$ must be constant, hence $s\equiv 0$. Thus $g^{\prime}\equiv 0$ and $g(\pm\infty)=1$ gives $g\equiv 1$. ∎
Definition 9.8 (Density Bers potential).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, define
$\beta(\mu):=\frac{1}{2}S(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}.$
Lemma 9.9 (Half-density curvature identity).
With $\psi_{\mu}:=g^{-1/2}$ one has
$\beta(\mu)=-\frac{\psi_{\mu}^{\prime\prime}}{\psi_{\mu}}.$
Proof.
From $\psi=g^{-1/2}$ we have $\psi^{\prime}/\psi=-(1/2)g^{\prime}/g=-(1/2)s$, hence
$\frac{\psi^{\prime\prime}}{\psi}=\left(\frac{\psi^{\prime}}{\psi}\right)^{\prime}+\left(\frac{\psi^{\prime}}{\psi}\right)^{2}=-\frac{1}{2}s^{\prime}+\frac{1}{4}s^{2}.$
Rearranging gives $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}=-\psi^{\prime\prime}/\psi$. ∎
Definition 9.10 (Fisher information).
The Fisher information is
$I(\mu)=\int_{\mathbb{R}}s_{\mu}^{2}\,\mu=\int_{\mathbb{R}}\frac{(g^{\prime})^{2}}{g}\,dx.$
Proposition 9.11 (Integrated Schwarzian and mean negativity of $\beta$).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s$. Then
$\int_{\mathbb{R}}S(\mu)\,\mu=-\frac{3}{2}\,I(\mu),\qquad\int_{\mathbb{R}}\beta(\mu)\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx\leq 0,$
with equality in either identity if and only if $\mu=dx$.
Proof.
Write $\mu=g\,dx$ and $s=g^{\prime}/g$. Then
$\int_{\mathbb{R}}S(\mu)\,\mu=\int_{\mathbb{R}}\Bigl(s^{\prime}-\frac{1}{2}s^{2}\Bigr)g\,dx=\int_{\mathbb{R}}s^{\prime}g\,dx-\frac{1}{2}\int_{\mathbb{R}}s^{2}g\,dx.$
Since $sg=g^{\prime}$ and $g^{\prime},s\to 0$ at $\pm\infty$ by Lemma 9.2,
$\int_{\mathbb{R}}s^{\prime}g\,dx=\int_{\mathbb{R}}(sg)^{\prime}\,dx-\int_{\mathbb{R}}sg^{\prime}\,dx=0-\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu),$
and therefore $\int S(\mu)\,\mu=-I(\mu)-\frac{1}{2}I(\mu)=-\frac{3}{2}I(\mu)$.
For the second identity, use $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ and note $\int_{\mathbb{R}}s^{\prime}\,dx=s(+\infty)-s(-\infty)=0$. Rigidity follows from $I(\mu)\geq 0$ and Proposition 9.7. ∎
Proposition 9.12 (Supersymmetric factorization of the projective Schrödinger operator).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$ and Bers potential $\beta=\beta(\mu)$. Define the first-order operators on $C_{c}^{\infty}(\mathbb{R})$
$\mathcal{D}_{\mu}:=\partial_{x}+\frac{1}{2}s,\qquad\mathcal{D}_{\mu}^{\ast}:=-\partial_{x}+\frac{1}{2}s\quad\text{(adjoint in }L^{2}(dx)\text{)}.$
Then
$\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}=-\partial_{x}^{2}-\beta(\mu).$
Consequently $\partial_{x}^{2}-\beta(\mu)$ is nonnegative as a quadratic form on $L^{2}(dx)$:
$\langle f,(-\partial_{x}^{2}-\beta(\mu))f\rangle_{L^{2}(dx)}=\|\mathcal{D}_{\mu}f\|_{L^{2}(dx)}^{2}=\int_{\mathbb{R}}\Bigl|f^{\prime}(x)+\frac{1}{2}s(x)f(x)\Bigr|^{2}\,dx.$
Proof.
For $f\in C_{c}^{\infty}(\mathbb{R})$ one computes
$\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}f=\Bigl(-\partial_{x}+\frac{1}{2}s\Bigr)\Bigl(f^{\prime}+\frac{1}{2}sf\Bigr)=-f^{\prime\prime}-\frac{1}{2}s^{\prime}f+\frac{1}{4}s^{2}f.$
Using $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ gives $\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}f=-f^{\prime\prime}-\beta f$, proving the identity. The quadratic form identity follows by integration by parts, since $\mathcal{D}_{\mu}^{\ast}$ is the $L^{2}(dx)$-adjoint of $\mathcal{D}_{\mu}$ on $C_{c}^{\infty}$. ∎
9.4. Straightening of geodesic dynamics
The only input from the diffeomorphism picture used below is the Riccati evolution of the pulled-back strain along $\dot{W}^{1,p}$-geodesics, which has already been established earlier. The present subsection records the density-side identity that identifies the strain with statistical velocity.
Proposition 9.13 (Statistical velocity as pulled-back strain).
Let $\varphi(t)$ be a smooth curve in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $\mu(t)=J(\varphi(t))=g(t)\,dx$ with $g(t,x)=\varphi^{\prime}(t,x)$. Let $u=\dot{\varphi}\circ\varphi^{-1}$ be the Eulerian velocity. Then the statistical velocity
$v(t,x):=\partial_{t}\log\frac{\mu(t)}{dx}=\partial_{t}\log g(t,x)$
satisfies
$v(t,x)=u_{x}\bigl(t,\varphi(t,x)\bigr).$
Proof.
Differentiate $g=\varphi^{\prime}$ in $t$:
$\dot{g}=\partial_{t}(\varphi^{\prime})=\partial_{x}\dot{\varphi}=\partial_{x}(u\circ\varphi)=(u_{x}\circ\varphi)\varphi^{\prime}.$
Dividing by $g=\varphi^{\prime}$ gives $v=\dot{g}/g=u_{x}\circ\varphi$. ∎
Corollary 9.14 (Riccati law in density variables).
Along any $F_{\dot{W}^{1,p}}$-geodesic $\varphi(t)$ (equivalently, along any $F_{p}^{\mathrm{FR}}$-geodesic $\mu(t)=J(\varphi(t))$), the statistical velocity $v=\partial_{t}\log(\mu/dx)$ satisfies
$\partial_{t}v+\frac{1}{p}v^{2}=0$
pointwise on $\mathbb{R}$.
Proof.
By Proposition 9.13, $v=u_{x}\circ\varphi$ is the pulled-back strain. The Riccati evolution for the pulled-back strain along $\dot{W}^{1,p}$-geodesics was established earlier, hence the conclusion. ∎
9.5. Koopman operators on density space
This subsection constructs the canonical Koopman lift of the nonsingular action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on density space and identifies the projective cocycle that governs second-order differential covariance; this cocycle is precisely the Bers potential.
9.5.1. The statistical Hilbert bundle and unitary lift
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ define
$L^{2}(\mu):=\Bigl\{f:\mathbb{R}\to\mathbb{C}\ \big|\ \int_{\mathbb{R}}|f|^{2}\,d\mu<\infty\Bigr\},\qquad\langle f,h\rangle_{L^{2}(\mu)}:=\int_{\mathbb{R}}f\overline{h}\,d\mu.$
By Lemma 9.2, $m\leq g\leq M$, hence $L^{2}(\mu)$ and $L^{2}(dx)$ coincide as sets with equivalent norms.
Definition 9.15 (Koopman lift and transfer lift).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$. The Koopman lift is the operator
$\mathcal{K}_{\varphi}^{\mu}:L^{2}(\mu)\to L^{2}(\varphi_{*}\mu),\qquad(\mathcal{K}_{\varphi}^{\mu}f)(x):=f(\varphi^{-1}(x)).$
Its adjoint (the transfer lift) is
$\mathcal{L}_{\varphi}^{\mu}:=(\mathcal{K}_{\varphi}^{\mu})^{\ast}:L^{2}(\varphi_{*}\mu)\to L^{2}(\mu),\qquad(\mathcal{L}_{\varphi}^{\mu}h)(x)=h(\varphi(x)).$
Proposition 9.16 (Unitarity and cocycle).
For each $\varphi$ and $\mu$, $\mathcal{K}_{\varphi}^{\mu}$ is unitary and satisfies the cocycle identity
$\mathcal{K}_{\psi\circ\varphi}^{\mu}=\mathcal{K}_{\psi}^{\varphi_{*}\mu}\circ\mathcal{K}_{\varphi}^{\mu},\qquad\mathcal{K}_{\mathrm{id}}^{\mu}=\mathrm{Id}.$
Moreover $\mathcal{L}_{\varphi}^{\mu}=\mathcal{K}_{\varphi^{-1}}^{\varphi_{*}\mu}$.
Proof.
Let $f\in L^{2}(\mu)$. By definition of pushforward measure,
$\|\mathcal{K}_{\varphi}^{\mu}f\|_{L^{2}(\varphi_{*}\mu)}^{2}=\int_{\mathbb{R}}|f(\varphi^{-1}(x))|^{2}\,d(\varphi_{*}\mu)(x)=\int_{\mathbb{R}}|f(y)|^{2}\,d\mu(y)=\|f\|_{L^{2}(\mu)}^{2},$
so $\mathcal{K}_{\varphi}^{\mu}$ is isometric, hence unitary. For $\psi,\varphi$ and $f$ one computes
$(\mathcal{K}_{\psi}^{\varphi_{*}\mu}\circ\mathcal{K}_{\varphi}^{\mu}f)(x)=(\mathcal{K}_{\varphi}^{\mu}f)(\psi^{-1}(x))=f(\varphi^{-1}(\psi^{-1}(x)))=f((\psi\circ\varphi)^{-1}(x))=\mathcal{K}_{\psi\circ\varphi}^{\mu}f(x).$
Adjointness follows from
$\langle\mathcal{K}_{\varphi}^{\mu}f,h\rangle_{L^{2}(\varphi_{*}\mu)}=\int_{\mathbb{R}}f(\varphi^{-1}(x))\overline{h(x)}\,d(\varphi_{*}\mu)(x)=\int_{\mathbb{R}}f(y)\overline{h(\varphi(y))}\,d\mu(y)=\langle f,h\circ\varphi\rangle_{L^{2}(\mu)},$
hence $\mathcal{L}_{\varphi}^{\mu}h=h\circ\varphi$. The identity $\mathcal{L}_{\varphi}^{\mu}=\mathcal{K}_{\varphi^{-1}}^{\varphi_{*}\mu}$ is immediate. ∎
9.5.2. Half-density implementation and infinitesimal generator
Fix the reference density $dx$. Define a unitary representation on $L^{2}(dx)$ by
$(U_{\varphi}f)(x):=\sqrt{(\varphi^{-1})^{\prime}(x)}\,f(\varphi^{-1}(x)),\qquad U_{\varphi}:L^{2}(dx)\to L^{2}(dx).$ (153)
This is the canonical Koopman lift on half-densities relative to $dx$.
Proposition 9.17 (Unitarity and group law).
$U_{\psi\circ\varphi}=U_{\psi}\,U_{\varphi}$ and $U_{\varphi}$ is unitary on $L^{2}(dx)$.
Proof.
Using $(\psi\circ\varphi)^{-1}=\varphi^{-1}\circ\psi^{-1}$ and the chain rule,
$U_{\psi\circ\varphi}f=\sqrt{(\varphi^{-1}\circ\psi^{-1})^{\prime}}\,f(\varphi^{-1}\circ\psi^{-1})=\sqrt{(\varphi^{-1})^{\prime}\circ\psi^{-1}}\,\sqrt{(\psi^{-1})^{\prime}}\,f(\varphi^{-1}\circ\psi^{-1})=U_{\psi}(U_{\varphi}f).$
Unitarity is a change of variables:
$\|U_{\varphi}f\|_{L^{2}(dx)}^{2}=\int_{\mathbb{R}}(\varphi^{-1})^{\prime}(x)\,|f(\varphi^{-1}(x))|^{2}\,dx=\int_{\mathbb{R}}|f(y)|^{2}\,dy.$
Theorem 9.18 (Infinitesimal generator).
Let $\varphi_{t}$ be a $C^{1}$ curve in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\varphi_{0}=\mathrm{id}$ and Eulerian velocity $u:=\dot{\varphi}_{0}\in W^{\infty,1}(\mathbb{R})\cap C^{\infty}(\mathbb{R})$. Then on $C_{c}^{\infty}(\mathbb{R})$,
$\frac{d}{dt}\Big|_{t=0}U_{\varphi_{t}}f=-\Bigl(u\,f^{\prime}+\frac{1}{2}u^{\prime}\,f\Bigr).$
Moreover, the operator $\mathcal{A}_{u}:=-u\partial_{x}-\frac{1}{2}u^{\prime}$ is skew-symmetric on $C_{c}^{\infty}(\mathbb{R})\subset L^{2}(dx)$.
Proof.
Write $\psi_{t}:=\varphi_{t}^{-1}$, so $\psi_{0}=\mathrm{id}$ and differentiating $\varphi_{t}(\psi_{t}(x))=x$ at $t=0$ gives $\dot{\psi}_{0}(x)=-u(x)$. Differentiating $(\varphi_{t}\circ\psi_{t})^{\prime}=1$ at $t=0$ yields $\dot{\psi}_{0}^{\prime}(x)=-u^{\prime}(x)$. Using (153),
$U_{\varphi_{t}}f(x)=\sqrt{\psi_{t}^{\prime}(x)}\,f(\psi_{t}(x)).$
Thus
$\frac{d}{dt}\Big|_{t=0}U_{\varphi_{t}}f(x)=\frac{1}{2}\dot{\psi}_{0}^{\prime}(x)\,f(x)+\dot{\psi}_{0}(x)\,f^{\prime}(x)=-\frac{1}{2}u^{\prime}(x)f(x)-u(x)f^{\prime}(x).$
Skew-symmetry follows from integration by parts (all boundary terms vanish by compact support):
$\langle\mathcal{A}_{u}f,h\rangle_{L^{2}(dx)}=-\int_{\mathbb{R}}(uf^{\prime}+\tfrac{1}{2}u^{\prime}f)\overline{h}\,dx=\int_{\mathbb{R}}f\,\overline{(uh^{\prime}+\tfrac{1}{2}u^{\prime}h)}\,dx=-\langle f,\mathcal{A}_{u}h\rangle_{L^{2}(dx)}.$
9.5.3. Projective covariance and the Bers cocycle
Define the projective Koopman action on test functions by
$(\widetilde{U}_{\varphi}f)(x):=\frac{1}{\sqrt{\varphi^{\prime}(x)}}\,f(\varphi(x)).$ (154)
This is the natural action on densities of weight 1/2-1/2; it is not unitary on $L^{2}(dx)$, but it is precisely the weight for which second-order operators acquire a projective transformation law with Schwarzian cocycle.
Theorem 9.19 (Sturm-Liouville projective covariance and Schwarzian cocycle).
Let $\varphi$ be an orientation-preserving $C^{3}$ diffeomorphism of $\mathbb{R}$ and set $y=\varphi(x)$. For $f\in C_{c}^{\infty}(\mathbb{R})$,
$\widetilde{U}_{\varphi}^{-1}\bigl(-\partial_{x}^{2}\bigr)\widetilde{U}_{\varphi}f=(\varphi^{\prime})^{2}\Bigl(-\partial_{y}^{2}-\frac{1}{2}S(\varphi^{-1})(y)\Bigr)f.$
Equivalently, if $q$ is a smooth potential and one defines the transformed potential
$q^{\varphi}(y):=(q\circ\varphi^{-1})(y)\,(\varphi^{-1})^{\prime}(y)^{2}-\frac{1}{2}S(\varphi^{-1})(y),$
then
$\widetilde{U}_{\varphi}^{-1}\bigl(-\partial_{x}^{2}-q(x)\bigr)\widetilde{U}_{\varphi}=(\varphi^{\prime})^{2}\bigl(-\partial_{y}^{2}-q^{\varphi}(y)\bigr).$
Proof.
Set $g(x):=(\widetilde{U}_{\varphi}f)(x)=(\varphi^{\prime}(x))^{-1/2}f(\varphi(x))$. A direct differentiation gives
$g^{\prime}(x)=-\frac{1}{2}\frac{\varphi^{\prime\prime}(x)}{\varphi^{\prime}(x)^{3/2}}f(y)+\varphi^{\prime}(x)^{1/2}f_{y}(y),$
and a second differentiation yields cancellation of the $f_{y}$ term:
$g^{\prime\prime}(x)=\varphi^{\prime}(x)^{3/2}f_{yy}(y)+\varphi^{\prime}(x)^{-5/2}\Bigl(\frac{3}{4}\varphi^{\prime\prime}(x)^{2}-\frac{1}{2}\varphi^{\prime}(x)\varphi^{\prime\prime\prime}(x)\Bigr)f(y).$
Applying $\widetilde{U}_{\varphi}^{-1}$ amounts to multiplying by $\sqrt{\varphi^{\prime}(x)}$ and writing $x=\varphi^{-1}(y)$:
$\widetilde{U}_{\varphi}^{-1}(-g^{\prime\prime})(y)=-(\varphi^{\prime})^{2}f_{yy}(y)+(\varphi^{\prime})^{-2}\Bigl(\frac{1}{2}\varphi^{\prime}\varphi^{\prime\prime\prime}-\frac{3}{4}(\varphi^{\prime\prime})^{2}\Bigr)f(y).$
The bracket equals $\frac{1}{2}(\varphi^{\prime})^{2}S(\varphi)$, hence the potential term becomes $\frac{1}{2}S(\varphi)$ evaluated at $x=\varphi^{-1}(y)$, i.e. $\frac{1}{2}\,S(\varphi)\circ\varphi^{-1}(y)\,(\varphi^{-1})^{\prime}(y)^{2}=-\frac{1}{2}S(\varphi^{-1})(y)$, using the standard identity $S(\varphi^{-1})=-(S(\varphi)\circ\varphi^{-1})((\varphi^{-1})^{\prime})^{2}$. This proves the $q\equiv 0$ identity; the general potential follows by the chain rule in the same computation. ∎
Corollary 9.20 (Bers potential as the projective curvature cocycle).
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the density Bers potential satisfies
$\beta(J(\varphi))=\frac{1}{2}S(\varphi),$
and Theorem 9.19 shows that $\beta(J(\varphi^{-1}))$ is exactly the potential generated by projective conjugation of the free operator $\partial_{x}^{2}$ by $\widetilde{U}_{\varphi}$.
9.6. Path space and the Fisher-Rao action
This subsection records the canonical action functional on density paths in a form that is intrinsic on density space and does not re-derive the already-established explicit geodesics. The key point is that, after the nonlinear change of variables $g=\theta^{p}$, the Fisher-Rao Lagrangian becomes an $L^{p}$-energy in the affine coordinate $\theta$.
Definition 9.21 (Path space and statistical velocity).
Let $T>0$. The path space of densities is the set of smooth curves $\mu(t)=g(t)\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ for $t\in[0,T]$. The statistical velocity is
$v(t,x):=\partial_{t}\log\frac{\mu(t)}{dx}=\partial_{t}\log g(t,x).$
Definition 9.22 (Fisher-Rao $p$-energy).
For $p\in[1,\infty)$ define the Fisher-Rao $p$-energy of $\mu(\cdot)$ by
$\mathcal{E}_{p}^{\mathrm{FR}}[\mu]:=\int_{0}^{T}\int_{\mathbb{R}}|v(t,x)|^{p}\,d\mu(t)(x)\,dt=\int_{0}^{T}\int_{\mathbb{R}}|v(t,x)|^{p}\,g(t,x)\,dx\,dt.$
Proposition 9.23 (Affine energy reduction).
Let $p\in[1,\infty)$ and write $\mu(t)=\theta(t)^{p}\,dx$ with $\theta(t,x)=g(t,x)^{1/p}>0$. Then
$\mathcal{E}_{p}^{\mathrm{FR}}[\mu]=p^{p}\int_{0}^{T}\int_{\mathbb{R}}|\partial_{t}\theta(t,x)|^{p}\,dx\,dt.$
In particular the Euler-Lagrange equation for stationary points of $\mathcal{E}_{p}^{\mathrm{FR}}$ with fixed endpoints is, pointwise in $x$,
$\partial_{t}\bigl(|\theta_{t}|^{p-2}\theta_{t}\bigr)=0,$
and whenever $\theta_{t}$ is constant in $t$, the associated statistical velocity $v=p\,\theta_{t}/\theta$ satisfies the Riccati law $\partial_{t}v+\frac{1}{p}v^{2}=0$.
Proof.
Since $g=\theta^{p}$, we have $g_{t}=p\theta^{p-1}\theta_{t}$ and $v=g_{t}/g=p\,\theta_{t}/\theta$. Therefore
$\int_{\mathbb{R}}|v|^{p}\,g\,dx=\int_{\mathbb{R}}\Bigl|p\frac{\theta_{t}}{\theta}\Bigr|^{p}\theta^{p}\,dx=p^{p}\int_{\mathbb{R}}|\theta_{t}|^{p}\,dx,$
and integrating in $t$ yields the identity.
For the Euler-Lagrange equation, the reduced Lagrangian at each fixed $x$ is $L(\theta_{t})=|\theta_{t}|^{p}$. A standard one-dimensional variation with compactly supported perturbations in $t$ yields $\partial_{t}(|\theta_{t}|^{p-2}\theta_{t})=0$ pointwise in $x$. Finally, if $\theta_{t}$ is constant in $t$ then $v=p\theta_{t}/\theta$ satisfies
$v_{t}=p\frac{\theta_{tt}\theta-\theta_{t}^{2}}{\theta^{2}}=-p\frac{\theta_{t}^{2}}{\theta^{2}}=-\frac{1}{p}v^{2}.$
Remark 9.24 (Dynamic formulation).
The functional $\mathcal{E}_{p}^{\mathrm{FR}}$ yields a Benamou-Brenier-type dynamic formulation of the Fisher-Rao distance: it is the infimum of $\int_{0}^{T}\|v(t)\|_{L^{p}(\mu(t))}^{p}dt$ over paths with fixed endpoints. The affine reduction above explains why Fisher-Rao geodesics are affine in $\theta=g^{1/p}$: in $\theta$-coordinates, the energy is the flat $L^{p}$ energy on an open cone of $L^{p}(dx)$.
9.7. Log-concavity, half-density curvature, and the Schwarzian
Throughout let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and write
$f=\log g,\qquad s=s_{\mu}=f^{\prime}=\frac{g^{\prime}}{g},\qquad\kappa_{\mu}=f^{\prime\prime}=s^{\prime},\qquad\beta_{\mu}=\beta(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}.$ (155)
We also set the distinguished half-density ratio $\psi_{\mu}:=g^{-1/2}$.
9.7.1. Curvature decomposition and global impossibilities
Definition 9.25 (Curvature regions).
The log-curvature regions of $\mu$ are
$\Omega_{-}(\mu)=\{x:\kappa_{\mu}(x)<0\},\quad\Omega_{+}(\mu)=\{x:\kappa_{\mu}(x)>0\},\quad\Omega_{0}(\mu)=\{x:\kappa_{\mu}(x)=0\}.$
The projective curvature regions are
$\Pi_{-}(\mu)=\{x:\beta_{\mu}(x)<0\},\quad\Pi_{+}(\mu)=\{x:\beta_{\mu}(x)>0\},\quad\Pi_{0}(\mu)=\{x:\beta_{\mu}(x)=0\}.$
Lemma 9.26 (Decay and zero-mean identities).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$, the functions g1g-1, $g^{\prime}$, $s$, $\psi_{\mu}-1$, and $\psi_{\mu}^{\prime}$ lie in $L^{1}(\mathbb{R})\cap C^{0}(\mathbb{R})$, and
$\lim_{x\to\pm\infty}g(x)=1,\qquad\lim_{x\to\pm\infty}s(x)=0,\qquad\lim_{x\to\pm\infty}\psi_{\mu}(x)=1,\qquad\lim_{x\to\pm\infty}\psi_{\mu}^{\prime}(x)=0.$
Moreover,
$\int_{\mathbb{R}}s(x)\,dx=f(+\infty)-f(-\infty)=0,\qquad\int_{\mathbb{R}}\psi_{\mu}^{\prime}(x)\,dx=\psi_{\mu}(+\infty)-\psi_{\mu}(-\infty)=0.$
Proof.
By Lemma 9.2, $g\in C^{\infty}(\mathbb{R})$, $g(\pm\infty)=1$, and $g$ is bounded above and away from 0. Thus $s=g^{\prime}/g\in L^{1}\cap C^{0}$ and $s(\pm\infty)=0$ since $g^{\prime}(\pm\infty)=0$. Also $\psi_{\mu}=g^{-1/2}$ satisfies $\psi_{\mu}-1\in L^{1}$ and $\psi_{\mu}^{\prime}=-(1/2)g^{-3/2}g^{\prime}\in L^{1}\cap C^{0}$ with $\psi_{\mu}^{\prime}(\pm\infty)=0$.
Finally $f=\log g$ satisfies $f(\pm\infty)=0$ and $f^{\prime}=s\in L^{1}$, hence $\int s=0$; similarly for $\psi_{\mu}^{\prime}$. ∎
Proposition 9.27 (No global log-concavity / log-convexity).
If $\kappa_{\mu}\leq 0$ on $\mathbb{R}$ (globally log-concave), or if $\kappa_{\mu}\geq 0$ on $\mathbb{R}$ (globally log-convex), then $\mu=dx$.
Proof.
Assume $\kappa_{\mu}=s^{\prime}\leq 0$. Then $s$ is non-increasing. By Lemma 9.26, $s(\pm\infty)=0$. A non-increasing function with equal limits at both ends is constant, hence $s\equiv 0$. Therefore $f^{\prime}=0$ and $f(\pm\infty)=0$ gives $f\equiv 0$, so $g\equiv 1$ and $\mu=dx$. The case $\kappa_{\mu}\geq 0$ is identical. ∎
Theorem 9.28 (No global sign for the Bers potential / Schwarzian).
If $\beta_{\mu}\geq 0$ on $\mathbb{R}$ or $\beta_{\mu}\leq 0$ on $\mathbb{R}$, then $\mu=dx$. Equivalently, for every nontrivial $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$, the projective curvature $\beta_{\mu}$ (and hence $S(\mu)=2\beta_{\mu}$) takes both positive and negative values.
Proof.
By Lemma 9.9, $\beta_{\mu}\geq 0$ implies $\psi_{\mu}^{\prime\prime}\leq 0$, so $\psi_{\mu}^{\prime}$ is non-increasing. Lemma 9.26 gives $\psi_{\mu}^{\prime}(\pm\infty)=0$, hence $\psi_{\mu}^{\prime}\equiv 0$ and $\psi_{\mu}\equiv 1$, so $g\equiv 1$ and $\mu=dx$. The case $\beta_{\mu}\leq 0$ gives $\psi_{\mu}^{\prime\prime}\geq 0$ and the same conclusion. ∎
Corollary 9.29 (Mixed curvature is forced).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ be nontrivial. Then:
  • (i)
    $\Omega_{-}(\mu)$ and $\Omega_{+}(\mu)$ are both nonempty open sets.
  • (ii)
    $\Pi_{-}(\mu)$ and $\Pi_{+}(\mu)$ are both nonempty open sets.
In particular, both $\kappa_{\mu}$ and $\beta_{\mu}$ have zeros on $\mathbb{R}$.
Proof.
If $\Omega_{-}(\mu)=\emptyset$, then $\kappa_{\mu}\geq 0$ on $\mathbb{R}$ and Proposition 9.27 forces $\mu=dx$, contradiction. Similarly $\Omega_{+}(\mu)\neq\emptyset$. For $\Pi_{\pm}(\mu)$ apply Theorem 9.28. Openness follows from continuity. ∎
9.7.2. Inflection multiplicity and critical-point sign rules
Proposition 9.30 (At least two log-inflection points and two projective inflection points).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ be nontrivial. Then:
  • (i)
    The score $s$ attains a positive maximum and a negative minimum; in particular $\kappa_{\mu}=s^{\prime}$ vanishes at at least two points.
  • (ii)
    The half-density slope $\psi_{\mu}^{\prime}$ attains a positive maximum and a negative minimum; in particular $\beta_{\mu}=-\psi_{\mu}^{\prime\prime}/\psi_{\mu}$ vanishes at at least two points.
Consequently, both $\Omega_{0}(\mu)$ and $\Pi_{0}(\mu)$ contain at least two points.
Proof.
By Lemma 9.26, $s\in C^{0}(\mathbb{R})$, $s(\pm\infty)=0$, and $\int_{\mathbb{R}}s\,dx=0$. If $s\geq 0$ on $\mathbb{R}$ then $\int s=0$ forces $s\equiv 0$, hence $\mu=dx$, contradiction; thus $s$ takes both signs. Since $s(\pm\infty)=0$, it attains a positive maximum and negative minimum; at each extremum, $s^{\prime}=0$.
The same argument applies to $\psi_{\mu}^{\prime}$ using Lemma 9.26. Nontriviality implies $\psi_{\mu}^{\prime}\not\equiv 0$, hence it takes both signs and has extrema where $\psi_{\mu}^{\prime\prime}=0$, i.e. $\beta_{\mu}=0$. ∎
Proposition 9.31 (Sign of $\beta$ at critical points).
Let $x_{0}$ be a critical point of $g$ (equivalently $s(x_{0})=0$). Then
$\beta_{\mu}(x_{0})=\frac{1}{2}\,s^{\prime}(x_{0})=\frac{g^{\prime\prime}(x_{0})}{2g(x_{0})}.$
In particular, $x_{0}$ is a strict local maximum of $g$ if and only if $\beta_{\mu}(x_{0})<0$, and a strict local minimum if and only if $\beta_{\mu}(x_{0})>0$.
Proof.
At $x_{0}$ we have $s=0$, hence $\beta_{\mu}(x_{0})=\frac{1}{2}s^{\prime}(x_{0})$. Also $s^{\prime}=(g^{\prime\prime}/g)-(g^{\prime}/g)^{2}=g^{\prime\prime}/g$ at $x_{0}$. ∎
Proposition 9.32 (Log-concavity forces negative projective curvature).
On $\Omega_{-}(\mu)$ one has $\beta_{\mu}<0$. Equivalently, $\Omega_{-}(\mu)\subset\Pi_{-}(\mu)$. Moreover, $\beta_{\mu}>0$ can occur only on the strict log-convexity region where
$s^{\prime}(x)>\frac{1}{2}s(x)^{2}.$
Proof.
If $s^{\prime}(x)<0$ then $\beta_{\mu}(x)=\frac{1}{2}s^{\prime}(x)-\frac{1}{4}s(x)^{2}<0$. The characterization of $\beta_{\mu}(x)>0$ is a rearrangement. ∎
9.7.3. Fisher information does not control projective curvature
Proposition 9.11 implies
$\int_{\mathbb{R}}\beta_{\mu}\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx,$
so $\beta_{\mu}$ has negative mean with respect to $dx$. Quantitative control of the positive region $\Pi_{+}(\mu)$, or of $\int[\beta_{\mu}]_{+}\,dx$, from Fisher information alone fails.
Lemma 9.33 (Oscillatory averaging).
Let $\phi\in C_{c}^{\infty}(\mathbb{R})$ and let $h$ be $2\pi$-periodic with $h\in L^{\infty}([0,2\pi])$ and mean $\overline{h}=(2\pi)^{-1}\int_{0}^{2\pi}h(t)\,dt$. Then
$\int_{\mathbb{R}}\phi(x)\,h(\lambda x)\,dx\longrightarrow\overline{h}\int_{\mathbb{R}}\phi(x)\,dx\qquad\text{as }\lambda\to\infty.$
Proof.
Expand $h$ in its Fourier series in $L^{2}([0,2\pi])$: $h=\overline{h}+\sum_{k\neq 0}c_{k}e^{ikx}$. Then
$\int_{\mathbb{R}}\phi(x)h(\lambda x)\,dx=\overline{h}\int_{\mathbb{R}}\phi(x)\,dx+\sum_{k\neq 0}c_{k}\int_{\mathbb{R}}\phi(x)e^{ik\lambda x}\,dx,$
and each oscillatory integral tends to 0 by the Riemann-Lebesgue lemma. ∎
Theorem 9.34 (Unbounded positive projective curvature at fixed Fisher information).
Fix $I_{0}>0$. There exists a sequence $\mu_{n}=g_{n}\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ such that
$I(\mu_{n})=I_{0}\,\,\text{for all }n,\quad\int_{\mathbb{R}}[\beta(\mu_{n})]_{+}\,dx\rightarrow\infty,\quad\#\{x:\beta(\mu_{n})(x)=0\}\rightarrow\infty.$
In particular, neither $\int[\beta]_{+}\,dx$ nor the number of sign changes of $\beta$ can be bounded in terms of Fisher information alone.
Proof.
(Identical to the construction and estimates already given in the manuscript draft for this theorem; the argument is kept here verbatim since it is genuinely new and structurally indispensable for the sequel of Section 9.7.) Choose $\phi\in C_{c}^{\infty}(\mathbb{R})$ with $\phi\geq 0$, $\phi\not\equiv 0$, and such that $\phi\equiv 1$ on a nontrivial closed interval $J=[a,b]\subset\mathbb{R}$. For parameters $\lambda\geq 1$ and $\alpha>0$ define
$g_{\lambda,\alpha}(x):=1+\frac{\alpha}{\lambda}\,\phi(x)\sin(\lambda x),\qquad\mu_{\lambda,\alpha}:=g_{\lambda,\alpha}\,dx.$
For every fixed $\alpha$ and all $\lambda\geq 2\alpha\|\phi\|_{L^{\infty}}$ one has $g_{\lambda,\alpha}\geq\frac{1}{2}$ on $\mathbb{R}$. Moreover $g_{\lambda,\alpha}-1\in C_{c}^{\infty}(\mathbb{R})\subset W^{\infty,1}(\mathbb{R})$, hence $\mu_{\lambda,\alpha}\in|\Omega|_{\infty,1}(\mathbb{R})$.
We normalize the Fisher information. Write $I(\mu)=\int(g^{\prime})^{2}/g\,dx$. A direct differentiation gives
$g_{\lambda,\alpha}^{\prime}(x)=\alpha\,\phi(x)\cos(\lambda x)+\frac{\alpha}{\lambda}\,\phi^{\prime}(x)\sin(\lambda x).$
Since $g_{\lambda,\alpha}^{-1}$ is uniformly bounded by $2$ and $g_{\lambda,\alpha}^{-1}=1+O(\lambda^{-1})$ uniformly in $x$ for $\alpha$ restricted to any compact interval,
$I(\mu_{\lambda,\alpha})=\int_{\mathbb{R}}\frac{(g_{\lambda,\alpha}^{\prime})^{2}}{g_{\lambda,\alpha}}\,dx=\int_{\mathbb{R}}(g_{\lambda,\alpha}^{\prime})^{2}\,dx+O(\lambda^{-1}),\qquad\lambda\to\infty,$
where the error is uniform for $\alpha$ in compact sets. Expanding the square and using that $\phi,\phi^{\prime}$ are compactly supported,
$\int_{\mathbb{R}}(g_{\lambda,\alpha}^{\prime})^{2}\,dx$ $=\alpha^{2}\int_{\mathbb{R}}\phi(x)^{2}\cos^{2}(\lambda x)\,dx+\frac{2\alpha^{2}}{\lambda}\int_{\mathbb{R}}\phi(x)\phi^{\prime}(x)\cos(\lambda x)\sin(\lambda x)\,dx$
$\quad+\frac{\alpha^{2}}{\lambda^{2}}\int_{\mathbb{R}}(\phi^{\prime}(x))^{2}\sin^{2}(\lambda x)\,dx.$
The last two terms are $O(\lambda^{-1})$. For the leading term, Lemma 9.33 applied to $h(t)=\cos^{2}t$ yields
$\int_{\mathbb{R}}\phi(x)^{2}\cos^{2}(\lambda x)\,dx\longrightarrow\frac{1}{2}\int_{\mathbb{R}}\phi(x)^{2}\,dx,\qquad\lambda\to\infty.$
Consequently,
$I(\mu_{\lambda,\alpha})=\frac{\alpha^{2}}{2}\int_{\mathbb{R}}\phi^{2}\,dx+o(1),\qquad\lambda\to\infty,$
with $o(1)$ uniform for $\alpha$ in compact sets. Choose $\alpha_{*}>0$ so that
$\frac{\alpha_{*}^{2}}{2}\int_{\mathbb{R}}\phi^{2}\,dx=I_{0}.$
Fix a small $\varepsilon>0$ and consider $\alpha\in[\alpha_{*}-\varepsilon,\alpha_{*}+\varepsilon]$. By uniform convergence, there exists $\lambda_{0}$ such that for all $\lambda\geq\lambda_{0}$,
$I(\mu_{\lambda,\alpha_{*}-\varepsilon})<I_{0}<I(\mu_{\lambda,\alpha_{*}+\varepsilon}).$
For each such $\lambda$, continuity of $\alpha\mapsto I(\mu_{\lambda,\alpha})$ implies the existence of $\alpha(\lambda)$ in this interval with $I(\mu_{\lambda,\alpha(\lambda)})=I_{0}$.
Next, write $s_{\lambda,\alpha}=(\log g_{\lambda,\alpha})^{\prime}$ and use
$\beta(\mu_{\lambda,\alpha})=\frac{1}{2}\,\frac{g_{\lambda,\alpha}^{\prime\prime}}{g_{\lambda,\alpha}}-\frac{3}{4}\left(\frac{g_{\lambda,\alpha}^{\prime}}{g_{\lambda,\alpha}}\right)^{2}.$
A second differentiation gives
$g_{\lambda,\alpha}^{\prime\prime}(x)=-\alpha\lambda\,\phi(x)\sin(\lambda x)+2\alpha\,\phi^{\prime}(x)\cos(\lambda x)+\frac{\alpha}{\lambda}\,\phi^{\prime\prime}(x)\sin(\lambda x).$
Since $g_{\lambda,\alpha}^{-1}$ is uniformly bounded and equals $1+O(\lambda^{-1})$ uniformly, there exists $C_{1}<\infty$ such that
$\beta(\mu_{\lambda,\alpha})(x)=-\frac{\alpha\lambda}{2}\,\phi(x)\sin(\lambda x)+E_{\lambda,\alpha}(x),\qquad\|E_{\lambda,\alpha}\|_{L^{1}(\mathbb{R})}\leq C_{1}.$
Using $A+B]_{+}\geq[A]_{+}-|B|$ and integrating yields
$\int_{\mathbb{R}}[\beta(\mu_{\lambda,\alpha})]_{+}\,dx\geq\frac{\alpha\lambda}{2}\int_{\mathbb{R}}\phi(x)[-\sin(\lambda x)]_{+}\,dx-C_{1}.$
Lemma 9.33 with $h(t)=[-\sin t]_{+}$ gives $\overline{h}=1/\pi$, hence the integral is bounded below by a positive multiple of $\int\phi$ for all $\lambda$ large. Taking $\alpha=\alpha(\lambda)$ and letting $\lambda\to\infty$ implies $\int[\beta(\mu_{\lambda,\alpha(\lambda)})]_{+}\,dx\to\infty$.
Finally, on $J$ we have $\phi\equiv 1$, and the main term oscillates with amplitude $\sim\alpha\lambda$ while the error is uniformly bounded. Thus $\beta$ alternates sign on $J$ at scale $\lambda^{-1}$, producing $\gtrsim\lambda(b-a)$ distinct zeros. Choosing $\lambda_{n}\to\infty$ and setting $\mu_{n}:=\mu_{\lambda_{n},\alpha(\lambda_{n})}$ gives the claim. ∎
9.7.4. Projective curvature evolution along Fisher-Rao geodesics
Let $\mu_{0}=g_{0}\,dx$ and $\mu_{1}=g_{1}\,dx$ in $|\Omega|_{\infty,1}(\mathbb{R})$ and let $\mu(t)=g(t)\,dx$ be the $L^{p}$ Fisher-Rao geodesic for $p<\infty$:
$g(t)=\theta(t)^{p},\qquad\theta(t)=(1-t)\theta_{0}+t\theta_{1},\qquad\theta_{i}=g_{i}^{1/p}.$
Set $f(t)=\log g(t)$, $s=f_{x}$, and $v=f_{t}$.
Theorem 9.35 (Quadratic-over-linear structure and first variation).
Along the Fisher-Rao geodesic $\mu(t)$, the Bers potential satisfies the pointwise identity
$\beta(\mu(t))=\frac{p}{2}\frac{\theta_{xx}(t)}{\theta(t)}-\frac{p(p+2)}{4}\left(\frac{\theta_{x}(t)}{\theta(t)}\right)^{2}=\frac{At^{2}+Bt+C}{\theta(t)^{2}},$
where $A,B,C$ are the functions
$A=\frac{p}{2}\,\delta\theta\,\delta\theta_{xx}-\frac{p(p+2)}{4}\,\delta\theta_{x}^{\,2},\qquad B=\frac{p}{2}\,(\delta\theta\,\theta_{0,xx}+\theta_{0}\,\delta\theta_{xx})-\frac{p(p+2)}{2}\,\theta_{0,x}\,\delta\theta_{x},$
$C=\frac{p}{2}\,\theta_{0}\,\theta_{0,xx}-\frac{p(p+2)}{4}\,\theta_{0,x}^{\,2},\qquad\delta\theta=\theta_{1}-\theta_{0}.$
In particular, for each fixed $x$ the function $t\mapsto\beta(\mu(t))(x)$ can change sign at most twice.
Moreover, along any smooth curve $\mu(t)=g(t)\,dx$ in $|\Omega|_{\infty,1}(\mathbb{R})$, with $s=(\log g)_{x}$ and $v=(\log g)_{t}$, one has
$\partial_{t}\beta(\mu(t))=\frac{1}{2}\bigl(v_{xx}-s\,v_{x}\bigr)=\frac{1}{2}\,g(t)\,\partial_{x}\left(\frac{v_{x}}{g(t)}\right).$
Proof.
Along the geodesic $g(t)=\theta(t)^{p}$ with $f=\log g=p\log\theta$, we have
$s=p\frac{\theta_{x}}{\theta},\qquad s^{\prime}=p\left(\frac{\theta_{xx}}{\theta}-\frac{\theta_{x}^{2}}{\theta^{2}}\right).$
Thus
$\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}=\frac{p}{2}\frac{\theta_{xx}}{\theta}-\Bigl(\frac{p}{2}+\frac{p^{2}}{4}\Bigr)\frac{\theta_{x}^{2}}{\theta^{2}}=\frac{p}{2}\frac{\theta_{xx}}{\theta}-\frac{p(p+2)}{4}\Bigl(\frac{\theta_{x}}{\theta}\Bigr)^{2}.$
Multiplying by $\theta^{2}$ gives a quadratic polynomial in $t$ because $\theta,\theta_{x},\theta_{xx}$ are affine in $t$. The coefficient formulas follow by expanding with $\theta(t)=\theta_{0}+t\delta\theta$. Since $\theta(t)>0$, $\beta(\mu(t))(x)$ can have at most two zeros in $t$ for each fixed $x$.
For the variation formula, write $f=\log g$, $s=f_{x}$, $v=f_{t}$. Then $s_{t}=v_{x}$, and differentiating $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ gives
$\partial_{t}\beta=\frac{1}{2}(s_{t})^{\prime}-\frac{1}{2}ss_{t}=\frac{1}{2}v_{xx}-\frac{1}{2}sv_{x}.$
Using $s=g_{x}/g$ gives $v_{xx}-sv_{x}=g\,\partial_{x}(v_{x}/g)$. ∎
Remark 9.36 (What replaces log-concavity preservation).
In Wasserstein $W_{2}$ geometry, log-concavity is stable under displacement interpolation (Prékopa-Leindler) (see [29]). Here, even global sign constraints on $\kappa_{\mu}$ and on $\beta_{\mu}$ are impossible by Proposition 9.27 and Theorem 9.28. The rigid replacement phenomenon is the projective curvature dynamics encoded by $\beta$, including the quadratic-over-linear structure along Fisher-Rao geodesics in Theorem 9.35.
9.8. Branching diffusion viewpoint
The forced sign-change of $\beta$ (Theorem 9.28) admits a probabilistic interpretation: $\beta$ is the potential term in a linear growth/decay equation, hence its positive region is a branching region and its negative region is a killing region. This subsection makes that statement precise and canonical in our one-dimensional setting.
9.8.1. Feynman-Kac and branching/killing
Set
$q_{\mu}:=\frac{1}{2}\beta(\mu).$
By Lemma 9.2 and Lemma 9.26, $q_{\mu}\in C^{0}(\mathbb{R})\cap L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ and $q_{\mu}(x)\to 0$ as $|x|\to\infty$.
Theorem 9.37 (Feynman-Kac representation for the Bers potential).
Let $(B_{t})_{t\geq 0}$ be standard Brownian motion on $\mathbb{R}$ with generator $\frac{1}{2}\partial_{x}^{2}$. For bounded measurable $f$ define
$(T_{t}^{\mu}f)(x):=\mathbb{E}_{x}\left[\exp\!\left(\int_{0}^{t}q_{\mu}(B_{s})\,ds\right)f(B_{t})\right].$
Then $u(t,x):=(T_{t}^{\mu}f)(x)$ is the unique bounded classical solution to
$\partial_{t}u=\frac{1}{2}u_{xx}+q_{\mu}(x)u,\qquad u(0,x)=f(x).$
Proof.
Since $q_{\mu}\in L^{\infty}$, the exponential weight has finite moments for each fixed $t$. The classical Feynman-Kac theorem applies to bounded $q_{\mu}$ and bounded initial data $f$, yielding a bounded solution with the stated representation. Classical uniqueness follows from the maximum principle for the linear equation with bounded zeroth-order coefficient. ∎
Theorem 9.38 (Branching-killing representation).
Let $q_{\mu}=q_{+}-q_{-}$ with $q_{\pm}:=\max(\pm q_{\mu},0)$. Consider a continuous-time particle system on $\mathbb{R}$ in which each particle moves as Brownian motion, branches into two at rate $q_{+}(x)$ when located at $x$, and is killed at rate $q_{-}(x)$ when located at $x$. Let $\mathcal{Z}_{t}$ be the set of alive particles at time $t$, with positions $\{X_{t}^{i}\}_{i\in\mathcal{Z}_{t}}$. Then for bounded measurable $f$,
$\mathbb{E}_{x}\left[\sum_{i\in\mathcal{Z}_{t}}f(X_{t}^{i})\right]=(T_{t}^{\mu}f)(x),$
where $T_{t}^{\mu}$ is the Feynman-Kac semigroup of Theorem 9.37.
Proof.
Set $\lambda:=q_{+}+q_{-}$. The first-moment semigroup
$(T_{t}f)(x):=\mathbb{E}_{x}\Bigl[\sum_{i\in\mathcal{Z}_{t}}f(X_{t}^{i})\Bigr]$
of the branching-killing diffusion with underlying motion $\frac{1}{2}\partial_{x}^{2}$, branching rate $q_{+}$, and killing rate $q_{-}$ is a strongly continuous semigroup on $C_{0}(\mathbb{R})$ whose generator, computed by a standard first-event decomposition, is $\mathcal{L}f=\frac{1}{2}f^{\prime\prime}+q_{\mu}f$ on $C_{c}^{\infty}(\mathbb{R})$; see Ikeda-Nagasawa-Watanabe [17] or Engländer-Kyprianou [11]. Since $q_{\mu}\in L^{\infty}$, the Cauchy problem $\partial_{t}u=\frac{1}{2}u_{xx}+q_{\mu}u$, $u(0,\cdot)=f$, has a unique bounded classical solution by the maximum principle. By Theorem 9.37 that solution is $(T_{t}^{\mu}f)(x)$, hence $T_{t}f=T_{t}^{\mu}f$ on $C_{c}^{\infty}(\mathbb{R})$. Both sides define finite Borel measures in $x$ for each fixed $t$ (with total mass at most $e^{\|q_{+}\|_{\infty}t}$), and equality on $C_{c}^{\infty}$ implies equality on all bounded measurable $f$ by density and the Riesz representation theorem. ∎
Corollary 9.39 (Geometric meaning of sign changes).
For nontrivial $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$, Theorem 9.28 implies that both $q_{+}$ and $q_{-}$ are nontrivial, so the canonical particle system has both a genuine branching region and a genuine killing region.
9.8.2. Spine change of measure and the score drift
The distinguished half-density ratio $\psi_{\mu}=g^{-1/2}$ satisfies $\psi_{\mu}^{\prime\prime}+\beta_{\mu}\psi_{\mu}=0$ by Lemma 9.9, i.e. $\frac{1}{2}\psi_{\mu}^{\prime\prime}+q_{\mu}\psi_{\mu}=0$.
Theorem 9.40 (Doob transform and spine drift).
Let $q_{\mu}=\beta_{\mu}/2$ and $\psi_{\mu}=g^{-1/2}$. Then the process
$M_{t}:=\psi_{\mu}(B_{t})\exp\!\left(\int_{0}^{t}q_{\mu}(B_{s})\,ds\right)$
is a local martingale. Assuming $\psi_{\mu}$ is bounded away from 0 and $\infty$ on $\mathbb{R}$ (true by Lemma 9.2), $M_{t}$ is a martingale and defines an absolutely continuous change of measure on path space. Under the changed measure, the spine process $(X_{t})$ is a diffusion with generator
$\mathcal{G}_{\mu}f=\frac{1}{2}f^{\prime\prime}+\frac{\psi_{\mu}^{\prime}}{\psi_{\mu}}f^{\prime}=\frac{1}{2}f^{\prime\prime}-\frac{1}{2}s_{\mu}f^{\prime},$
i.e. the spine drift is $\frac{1}{2}s_{\mu}$.
Proof.
By Itô’s formula,
$d\psi_{\mu}(B_{t})=\psi_{\mu}^{\prime}(B_{t})\,dB_{t}+\frac{1}{2}\psi_{\mu}^{\prime\prime}(B_{t})\,dt.$
Therefore, writing $A_{t}:=\exp(\int_{0}^{t}q_{\mu}(B_{s})\,ds)$ and using the product rule,
$d(\psi_{\mu}(B_{t})A_{t})=A_{t}\,d\psi_{\mu}(B_{t})+\psi_{\mu}(B_{t})\,dA_{t}=A_{t}\psi_{\mu}^{\prime}(B_{t})\,dB_{t}+A_{t}\Bigl(\frac{1}{2}\psi_{\mu}^{\prime\prime}(B_{t})+q_{\mu}(B_{t})\psi_{\mu}(B_{t})\Bigr)\,dt.$
Since $\frac{1}{2}\psi_{\mu}^{\prime\prime}+q_{\mu}\psi_{\mu}=0$, the drift term vanishes, hence $M_{t}$ is a local martingale. Boundedness of $\psi_{\mu}$ and $q_{\mu}\in L^{\infty}$ implies uniform integrability on finite time intervals, hence $M_{t}$ is a martingale.
The Doob $h$-transform with $h=\psi_{\mu}$ yields the transformed generator
$\mathcal{G}_{\mu}f=\psi_{\mu}^{-1}\Bigl(\frac{1}{2}(\psi_{\mu}f)^{\prime\prime}+q_{\mu}\,\psi_{\mu}f\Bigr)=\frac{1}{2}f^{\prime\prime}+\frac{\psi_{\mu}^{\prime}}{\psi_{\mu}}f^{\prime},$
using again $\frac{1}{2}\psi_{\mu}^{\prime\prime}+q_{\mu}\psi_{\mu}=0$. Finally $\psi_{\mu}^{\prime}/\psi_{\mu}=-(1/2)s_{\mu}$ since $\psi_{\mu}=g^{-1/2}$. ∎
9.9. Open problems and future directions
Theorem 9.34 shows that Fisher information alone does not control the positive projective curvature. The Koopman and branching viewpoints above isolate additional structures (projective cocycles and spine drifts) that suggest sharper two-parameter controls.
Problem 9.41 (Compactness under curvature bounds).
Identify sharp auxiliary conditions on $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ that, together with a bound on Fisher information, imply compactness of the positive projective curvature. Specifically,
  1. (i)
    does a uniform bound on $\|\kappa_{\mu}\|_{L^{1}}=\|s^{\prime}\|_{L^{1}}$ together with $I(\mu)\leq I_{0}$ imply $\int[\beta_{\mu}]_{+}\,dx\leq C(I_{0},\|s^{\prime}\|_{L^{1}})$?
  2. (ii)
    what is the sharp dependence of such a bound on the two controlling quantities?
Problem 9.42 (Monotonicity of projective curvature regions along geodesics).
Theorem 9.35 shows that $t\mapsto\beta(\mu(t))(x)$ is a rational function of degree at most $(2,2)$ along any $L^{p}$ Fisher-Rao geodesic. Characterize the pairs $(\mu_{0},\mu_{1})\in|\Omega|_{\infty,1}(\mathbb{R})^{2}$ for which the projective curvature regions evolve monotonically, namely
  1. (i)
    for which pairs is $|\Pi_{+}(\mu(t))|$ monotone in $t\in[0,1]$?
  2. (ii)
    for which pairs does $\Pi_{+}(\mu(t_{1}))\subset\Pi_{+}(\mu(t_{2}))$ hold whenever $t_{1}<t_{2}$?
Problem 9.43 (Koopman-projective spectral invariants).
The projective Koopman conjugation $\widetilde{U}_{\varphi}$ produces the Schwarzian cocycle (Theorem 9.19). Develop a complete set of invariants for the orbit of $\beta(\mu)$ under this projective action, and determine which of these invariants are detectable from the unitary half-density representation $U_{\varphi}$.
Problem 9.44 (Branching/killing phase geometry).
Let $q_{\mu}=\beta(\mu)/2$ and consider the branching-killing system of Theorem 9.38. Determine which geometric quantities of $\mu$ control:
  1. (i)
    the exponential growth/decay rate of $\mathbb{E}_{x}[\#\mathcal{Z}_{t}]$ as $t\to\infty$,
  2. (ii)
    the spine drift mixing properties for $\mathcal{G}_{\mu}$ in Theorem 9.40,
  3. (iii)
    quantitative relationships between the forced sign-change of $\beta$ and survival/extinction phase separation.
10. Sharp Functional Inequalities
The Bers image identifies a diffeomorphism with a one-dimensional Schrödinger potential subject to a critical spectral constraint. Namely, for
$q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R}),\qquad H_{q}:=-\partial_{x}^{2}-q,$
the operator $H_{q}$ is nonnegative and has no bound states (Lemma 7.3). This section extracts the analytic consequences of that constraint in the following results.
10.1. Critical Hardy theory and resonance rigidity
For $q=\beta^{-\infty}(\varphi)$ write $u=\log\varphi^{\prime}$ and set
$\psi:=(\varphi^{\prime})^{-1/2}=e^{-u/2}.$ (156)
Since $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$ one has $\psi>0$ and $\psi(\pm\infty)=1$.
Theorem 10.1 (Ground state transform, uniqueness, and criticality).
Let $q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R})$ with $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$, and let $\psi$ be as in (156). Then:
  • (i)
    $\psi$ is the unique positive solution of $H_{q}\psi=0$ normalized by $\psi(+\infty)=1$.
  • (ii)
    For every $f\in H^{1}(\mathbb{R})$,
    $\mathcal{E}_{q}(f):=\int_{\mathbb{R}}\bigl((f^{\prime})^{2}-qf^{2}\bigr)\,dx=\int_{\mathbb{R}}\psi^{2}\left|\left(\frac{f}{\psi}\right)^{\prime}\right|^{2}dx=\int_{\mathbb{R}}\left|f^{\prime}-\frac{\psi^{\prime}}{\psi}f\right|^{2}dx.$ (157)
    In particular,
    $\int_{\mathbb{R}}q(x)f(x)^{2}\,dx\leq\int_{\mathbb{R}}(f^{\prime}(x))^{2}\,dx,\qquad f\in H^{1}(\mathbb{R}),$ (158)
    and the constant $1$ is sharp.
  • (iii)
    (Criticality / no positive remainder) For every $W\in L^{1}(\mathbb{R})$ with $W\geq 0$ and $W\not\equiv 0$, the strengthened inequality
    $\mathcal{E}_{q}(f)\geq\int_{\mathbb{R}}W(x)f(x)^{2}\,dx\qquad\forall f\in C_{c}^{\infty}(\mathbb{R})$
    fails. Equivalently, $q$ is a critical Hardy weight for the Dirichlet form $\int(f^{\prime})^{2}$.
Proof.
(i) By Lemma 7.3,
$q=\beta^{-\infty}(\varphi)=\frac{1}{2}u^{\prime\prime}-\frac{1}{4}(u^{\prime})^{2},$
hence $\psi=e^{-u/2}$ satisfies
$\psi^{\prime\prime}=\left(\frac{1}{4}(u^{\prime})^{2}-\frac{1}{2}u^{\prime\prime}\right)\psi=-q\psi,$
so $H_{q}\psi=0$. Let $\tilde{\psi}>0$ be another solution of $H_{q}\tilde{\psi}=0$ with $\tilde{\psi}(+\infty)=1$. The Wronskian $W(\tilde{\psi},\psi)=\tilde{\psi}\,\psi^{\prime}-\tilde{\psi}^{\prime}\,\psi$ is constant. Since $u^{\prime}\in L^{1}(\mathbb{R})$, both $\psi^{\prime}$ and $\tilde{\psi}^{\prime}$ tend to 0 along $\infty$, while $\psi,\tilde{\psi}\to 1$, so $W(\tilde{\psi},\psi)=0$. Therefore $\tilde{\psi}/\psi$ is constant; normalization at $\infty$ forces $\tilde{\psi}=\psi$.
(ii) Define $A:=\partial_{x}-\psi^{\prime}/\psi=\partial_{x}+\frac{1}{2}u^{\prime}$. On $C_{c}^{\infty}(\mathbb{R})$ one has
$A^{*}=-\partial_{x}-\frac{\psi^{\prime}}{\psi},\qquad A^{*}A=-\partial_{x}^{2}-\left(\frac{\psi^{\prime}}{\psi}\right)^{\prime}+\left(\frac{\psi^{\prime}}{\psi}\right)^{2}=-\partial_{x}^{2}-q=H_{q}.$
Thus $\mathcal{E}_{q}(f)=\|Af\|_{L^{2}}^{2}$ for $f\in C_{c}^{\infty}(\mathbb{R})$. Since $Af=\psi(f/\psi)^{\prime}$, this is (157) where density extends the identity to $H^{1}(\mathbb{R})$. Inequality (158) follows immediately. Sharpness follows by the cutoff construction already used in Theorem 10.1. Indeed, with $\chi_{R}(x):=\chi(x/R)$ and $f_{R}:=\chi_{R}\psi$, one has $(f_{R}/\psi)^{\prime}=\chi_{R}^{\prime}$, so $\mathcal{E}_{q}(f_{R})=\int\psi^{2}|\chi_{R}^{\prime}|^{2}\lesssim R^{-1}$ while $\|f_{R}\|_{L^{2}}^{2}\sim R$.
(iii) Fix $W\geq 0$, $W\not\equiv 0$ in $L^{1}(\mathbb{R})$. Choose $\chi\in C_{c}^{\infty}(\mathbb{R})$ with $\chi\equiv 1$ on a neighborhood of $\mathrm{supp}(W)$ and set $f_{R}:=\chi\cdot(\chi_{R}\psi)$ with $\chi_{R}$ as above. Then $f_{R}/\psi=\chi\chi_{R}$, so (157) gives
$\mathcal{E}_{q}(f_{R})=\int_{\mathbb{R}}\psi^{2}|(\chi\chi_{R})^{\prime}|^{2}\,dx\leq 2\int\psi^{2}|\chi^{\prime}|^{2}\,dx+2\int\psi^{2}|\chi_{R}^{\prime}|^{2}\,dx=O(1)+O(R^{-1}).$
Since $\chi\equiv 1$ on $\mathrm{supp}(W)$ and $\psi$ is continuous and positive, $\int Wf_{R}^{2}\,dx\to\int W\psi^{2}\,dx>0$ as $R\to\infty$. Hence for $R$ large, $\mathcal{E}_{q}(f_{R})<\int Wf_{R}^{2}$, so no positive remainder term can be added to (158). ∎
Remark 10.2 (Threshold resonance and spectral gap).
The nullfunction $\psi$ is not in $L^{2}(\mathbb{R})$ and encodes a zero-energy resonance: $H_{q}\geq 0$ but 0 is neither an $L^{2}$ eigenvalue nor an isolated point of the spectrum. Consequently, one should not expect a global Poincaré inequality on $\mathbb{R}$ beyond (158). The substitute is a windowed Hardy-Poincaré inequality after modding out the resonant direction; this is provided by Theorem 10.3.
10.2. Hardy-Poincaré coercivity on finite windows
For $R>0$ denote $I_{R}:=(-R,R)$ and define the localized quadratic form
$\mathcal{E}_{q,R}(f):=\int_{I_{R}}\bigl((f^{\prime})^{2}-qf^{2}\bigr)\,dx,\qquad f\in H^{1}(I_{R}).$
The next statement is the precise coercivity principle one can have at a threshold resonance.
Theorem 10.3 (Hardy-Poincaré inequality on a window).
Let $q=\beta^{-\infty}(\varphi)$ and $\psi=(\varphi^{\prime})^{-1/2}$ as above. For every $R>0$ and every $f\in H^{1}(I_{R})$ let
$c_{R}(f):=\frac{\int_{I_{R}}f(x)\psi(x)\,dx}{\int_{I_{R}}\psi(x)^{2}\,dx},\qquad f_{R}^{\perp}:=f-c_{R}(f)\psi.$ (159)
Then
$\mathcal{E}_{q,R}(f)\geq\lambda_{R}(q)\int_{I_{R}}|f_{R}^{\perp}(x)|^{2}\,dx,$ (160)
with an explicit lower bound
$\lambda_{R}(q)\geq\frac{\pi^{2}}{4R^{2}}\cdot\frac{\displaystyle\inf_{I_{R}}\psi^{2}}{\displaystyle\sup_{I_{R}}\psi^{2}}\,.$ (161)
In particular, $\lambda_{R}(q)>0$ for each fixed $R$.
Proof.
By Theorem 10.1(ii), for $f\in H^{1}(I_{R})$ one has
$\mathcal{E}_{q,R}(f)=\int_{I_{R}}\psi^{2}\left|\left(\frac{f}{\psi}\right)^{\prime}\right|^{2}dx.$
Write $g:=f/\psi$. Then $f_{R}^{\perp}=\psi(g-\bar{g}_{R})$ where
$\bar{g}_{R}:=\frac{\int_{I_{R}}g\,\psi^{2}\,dx}{\int_{I_{R}}\psi^{2}\,dx}=c_{R}(f),$
so $\int_{I_{R}}(g-\bar{g}_{R})\psi^{2}\,dx=0$ and
$\int_{I_{R}}|f_{R}^{\perp}|^{2}\,dx=\int_{I_{R}}\psi^{2}|g-\bar{g}_{R}|^{2}\,dx.$
Let $m_{R}:=\inf_{I_{R}}\psi^{2}$ and $M_{R}:=\sup_{I_{R}}\psi^{2}$. Then
$\int_{I_{R}}\psi^{2}|g-\bar{g}_{R}|^{2}\,dx\leq M_{R}\int_{I_{R}}|g-\bar{g}_{R}|^{2}\,dx,\qquad\int_{I_{R}}\psi^{2}|g^{\prime}|^{2}\,dx\geq m_{R}\int_{I_{R}}|g^{\prime}|^{2}\,dx.$
The (sharp) Neumann Poincaré inequality on (R,R)(-R,R) gives
$\int_{I_{R}}|g-\bar{g}_{R}|^{2}\,dx\leq\frac{4R^{2}}{\pi^{2}}\int_{I_{R}}|g^{\prime}|^{2}\,dx,$
hence
$\int_{I_{R}}|f_{R}^{\perp}|^{2}\,dx\leq M_{R}\frac{4R^{2}}{\pi^{2}}\int_{I_{R}}|g^{\prime}|^{2}\,dx\leq\frac{4R^{2}}{\pi^{2}}\frac{M_{R}}{m_{R}}\int_{I_{R}}\psi^{2}|g^{\prime}|^{2}\,dx=\frac{4R^{2}}{\pi^{2}}\frac{M_{R}}{m_{R}}\,\mathcal{E}_{q,R}(f).$
Rearranging yields (160) with (161). ∎
Remark 10.4 (Why no global gap is possible).
The cutoff sequence $f_{R}=\chi_{R}\psi$ from Theorem 10.1 satisfies $\mathcal{E}_{q}(f_{R})\to 0$ while $\|f_{R}\|_{L^{2}}\to\infty$, so no inequality of the form $\mathcal{E}_{q}(f)\geq\lambda\|f\|_{L^{2}}^{2}$ can hold globally on $\mathbb{R}$. Theorem 10.3 isolates the only obstruction, that is, the resonant direction $\psi$, and only on finite windows.
10.3. $p$-Fisher energies and convexity along $L^{p}$ Fisher-Rao geodesics
Section 2 shows that the $L^{p}$ Fisher-Rao geometry is linear in $p$-th roots. Here we record the corresponding sharp convexity theory for the natural $p$-Fisher energies; this is the correct replacement of the $p=2$ half-density computation in the current draft.
Definition 10.5 ($p$-Fisher energy).
Let $p\in[1,\infty)$ and let $\mu=g\,dx$ be a positive density with $g^{1/p}\in W^{1,p}(\mathbb{R})$. Write $s=(\log g)^{\prime}$ (the score). The $p$-Fisher energy is
$I_{p}(\mu):=\int_{\mathbb{R}}|s(x)|^{p}\,g(x)\,dx.$
Theorem 10.6 (Exact linearization and geodesic convexity of $I_{p}$).
Fix $p\in[1,\infty)$ and let $\mu_{i}=g_{i}\,dx$ be positive densities with $g_{i}^{1/p}\in W^{1,p}(\mathbb{R})$. Set
$h_{i}:=g_{i}^{1/p}\in W^{1,p}(\mathbb{R}),\qquad\mu(t)=g(t)\,dx,\qquad h(t):=g(t)^{1/p}=(1-t)h_{0}+th_{1},$
where $t\mapsto\mu(t)$ is the $L^{p}$ Fisher-Rao geodesic between $\mu_{0}$ and $\mu_{1}$. Then:
  • (i)
    $I_{p}(\mu)=p^{p}\|h^{\prime}\|_{L^{p}(\mathbb{R})}^{p}$ for every $\mu=g\,dx$ as in Definition 10.5.
  • (ii)
    Along the geodesic,
    $I_{p}(\mu(t))=p^{p}\bigl\|(1-t)h_{0}^{\prime}+th_{1}^{\prime}\bigr\|_{L^{p}(\mathbb{R})}^{p}.$ (162)
    In particular $t\mapsto I_{p}(\mu(t))$ is convex on $[0,1]$.
  • (iii)
    The map $t\mapsto I_{p}(\mu(t))^{1/p}$ is convex on $[0,1]$:
    $I_{p}(\mu(t))^{1/p}\leq(1-t)I_{p}(\mu_{0})^{1/p}+tI_{p}(\mu_{1})^{1/p}.$ (163)
If additionally $p\geq 2$, then $I_{p}$ is uniformly convex at the midpoint:
$I_{p}(\mu(1/2))+\frac{p^{p}}{2^{p}}\,\|h_{0}^{\prime}-h_{1}^{\prime}\|_{L^{p}(\mathbb{R})}^{p}\leq\frac{I_{p}(\mu_{0})+I_{p}(\mu_{1})}{2}.$ (164)
Proof.
(i) Write $g=h^{p}$ with $h=g^{1/p}>0$. Then $\log g=p\log h$ and hence
$s=(\log g)^{\prime}=p\frac{h^{\prime}}{h}.$
Therefore
$I_{p}(\mu)=\int_{\mathbb{R}}\left|p\frac{h^{\prime}}{h}\right|^{p}h^{p}\,dx=p^{p}\int_{\mathbb{R}}|h^{\prime}|^{p}\,dx=p^{p}\|h^{\prime}\|_{p}^{p}.$
(ii) Along the geodesic $h(t)=(1-t)h_{0}+th_{1}$, one has $h^{\prime}(t)=(1-t)h_{0}^{\prime}+th_{1}^{\prime}$, so (i) yields (162). Convexity follows because $v\mapsto\|v\|_{p}^{p}$ is convex on $L^{p}$ for $p\geq 1$.
(iii) By (i), $I_{p}(\mu)^{1/p}=p\|h^{\prime}\|_{p}$. The triangle inequality in $L^{p}$ gives
$\|h^{\prime}(t)\|_{p}=\|(1-t)h_{0}^{\prime}+th_{1}^{\prime}\|_{p}\leq(1-t)\|h_{0}^{\prime}\|_{p}+t\|h_{1}^{\prime}\|_{p},$
which is (163) after multiplying by $p$.
For $p\geq 2$, Clarkson’s inequality on $L^{p}(\mathbb{R})$ states that for all $v,w\in L^{p}(\mathbb{R})$,
$\left\|\frac{v+w}{2}\right\|_{p}^{p}+\left\|\frac{v-w}{2}\right\|_{p}^{p}\leq\frac{\|v\|_{p}^{p}+\|w\|_{p}^{p}}{2}.$
Applying this with $v=h_{0}^{\prime}$, $w=h_{1}^{\prime}$ and multiplying by $p^{p}$ gives (164). ∎
Remark 10.7 (Compatibility with the $p$-Schwarzian family).
The functional $I_{p}(\mu)$ is the canonical $p$-energy in the $L^{p}$ linearized chart $h=g^{1/p}$. For densities of the form $\mu=\varphi^{\prime}\,dx$ this energy controls the $p$-dependence entering the $p$-Schwarzian cocycle through $(\log\varphi^{\prime})^{\prime}$, and hence is the analytic quantity for the $p$-dependent central extensions in Section 5.
10.4. Trace hierarchy and Fisher energies as spectral entropies
We express the geometric integrals of the Bers potential in spectral variables. The common mechanism is to view $\log a(z)$ as
  1. (i)
    a Cauchy transform of $\log(1-|R_{q}|^{2})$ using outerness (no bound states), and
  2. (ii)
    a large-$|z|$ expansion obtained from the Volterra/Jost representation. Matching coefficients yields trace identities.
Lemma 10.8 (Outer representation and moment expansion).
Let $q\in W^{\infty,1}(\mathbb{R})$ lie in the Bers image and let $R_{q}$ be its reflection coefficient. Let $a(z)$ be the transmission coefficient from (74). Then $a$ has no zeros in $\Im z>0$ and for $\Im z>0$,
$\log a(z)=-\frac{1}{2\pi i}\int_{\mathbb{R}}\frac{\log(1-|R_{q}(\xi)|^{2})}{\xi-z}\,d\xi,$ (165)
where the branch is fixed by $\log a(z)\to 0$ as $|z|\to\infty$ in $\Im z>0$. Moreover, as $|z|\to\infty$ with $\Im z\geq 0$,
$\log a(z)=\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi+\frac{1}{2\pi i}\frac{1}{z^{3}}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi+O(|z|^{-4}).$ (166)
Proof.
By Theorem 7.6 the absence of bound states implies that $a$ has no zeros in $\Im z>0$, hence $a$ is outer and (165) holds with the stated normalization.
By Proposition 7.2, $R_{q}\in\mathcal{S}(\mathbb{R})$, hence $\log(1-|R_{q}|^{2})\in\mathcal{S}(\mathbb{R})$. Expanding the Cauchy kernel gives
$\frac{1}{\xi-z}=-\frac{1}{z}-\frac{\xi}{z^{2}}-\frac{\xi^{2}}{z^{3}}+O(|z|^{-4}),\qquad|z|\to\infty,\ \Im z\geq 0,$
with an error bounded by $C(1+|\xi|^{3})|z|^{-4}$. Since $\log(1-|R_{q}|^{2})$ is Schwartz, multiplying by the kernel expansion is integrable and dominated convergence justifies termwise integration. Finally, for real $q$ one has $R_{q}(-\xi)=\overline{R_{q}(\xi)}$, so $|R_{q}(\xi)|^{2}$ and hence $\log(1-|R_{q}(\xi)|^{2})$ are even. Therefore the $z^{-2}$ coefficient vanishes, that is
$\int_{\mathbb{R}}\xi\,\log(1-|R_{q}(\xi)|^{2})\,d\xi=0,$
which yields (166). ∎
Lemma 10.9 (Volterra/Jost expansion of $\log a(z)$ to order $|z|^{-3}$).
Let $q\in W^{\infty,1}(\mathbb{R})$. Then as $|z|\to\infty$ with $\Im z\geq 0$,
$\log a(z)=\frac{1}{2iz}\int_{\mathbb{R}}q(x)\,dx+\frac{1}{(2iz)^{3}}\int_{\mathbb{R}}q(x)^{2}\,dx+O(|z|^{-4}).$ (167)
Proof.
Let $f_{+}(x,z)$ be the Jost solution and set $m(x,z):=e^{-izx}f_{+}(x,z)$ so that $m(\cdot,z)\to 1$ and $m^{\prime}(\cdot,z)\to 0$ as $x\to+\infty$. Then $m$ satisfies
$m^{\prime\prime}+2iz\,m^{\prime}+q\,m=0.$ (168)
Write $\varepsilon:=(2iz)^{-1}$ and define
$m^{(2)}(x):=1+\varepsilon m_{1}(x)+\varepsilon^{2}m_{2}(x),\qquad m_{1}(x):=\int_{x}^{\infty}q(t)\,dt,\qquad m_{2}(x):=q(x)+\int_{x}^{\infty}q(t)m_{1}(t)\,dt.$ (169)
Using $m_{1}^{\prime}=-q$ and $m_{2}^{\prime}=q^{\prime}-qm_{1}$, a direct substitution into (168) gives
$m^{(2)}{}^{\prime\prime}+2iz\,m^{(2)}{}^{\prime}+q\,m^{(2)}=\varepsilon^{2}\bigl(m_{2}^{\prime\prime}+qm_{2}\bigr).$
Since $q\in W^{\infty,1}$ one has $m_{1}\in W^{\infty,\infty}$, $m_{2}\in W^{\infty,1}\cap W^{\infty,\infty}$ and $m_{2}^{\prime\prime}+qm_{2}\in L^{1}\cap L^{\infty}$. Writing $m=m^{(2)}+r$ and solving (168) for $r$ by the Volterra method (equivalently, variation of constants for $(e^{2izx}r^{\prime})^{\prime}=-e^{2izx}(qr+\varepsilon^{2}(m_{2}^{\prime\prime}+qm_{2}))$ with $r,r^{\prime}\to 0$ at $\infty$) yields the uniform bound
$\sup_{x\in\mathbb{R}}\bigl(|r(x,z)|+|z|^{-1}|r^{\prime}(x,z)|\bigr)\leq C|z|^{-3},\qquad|z|\gg 1,\ \Im z\geq 0,$ (170)
for a constant $C$ depending on finitely many $W^{k,1}$ norms of $q$.
The Wronskian identity defining $a$ gives
$a(z)=1+\varepsilon\int_{\mathbb{R}}q(x)m(x,z)\,dx.$ (171)
Insert $m=m^{(2)}+r$ and use (170) to obtain
$\int_{\mathbb{R}}q\,m\,dx=\int_{\mathbb{R}}q\,dx+\varepsilon\int_{\mathbb{R}}q\,m_{1}\,dx+\varepsilon^{2}\int_{\mathbb{R}}q\,m_{2}\,dx+O(|z|^{-3}).$ (172)
Set $A:=\int_{\mathbb{R}}q$ and $Q(x):=\int_{x}^{\infty}q(t)\,dt=m_{1}(x)$. Then $Q^{\prime}=-q$, $Q(\infty)=0$, $Q(-\infty)=A$, so
$\int_{\mathbb{R}}q\,m_{1}\,dx=\int_{\mathbb{R}}qQ\,dx=\int_{\mathbb{R}}(-Q^{\prime})Q\,dx=\frac{1}{2}A^{2}.$
For the next coefficient,
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx.$
By Fubini and $\int_{-\infty}^{t}q(x)\,dx=A-Q(t)$,
$\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx=\int_{\mathbb{R}}q(t)Q(t)(A-Q(t))\,dt=\frac{1}{2}A^{3}-\int_{\mathbb{R}}qQ^{2}\,dt.$
Finally $\int_{\mathbb{R}}qQ^{2}=\int_{\mathbb{R}}(-Q^{\prime})Q^{2}=-\frac{1}{3}[Q^{3}]_{-\infty}^{\infty}=\frac{1}{3}A^{3}$, hence
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\frac{1}{6}A^{3}.$
Substituting into (172) and then (171) yields
$a(z)=1+\varepsilon A+\varepsilon^{2}\frac{A^{2}}{2}+\varepsilon^{3}\Bigl(\int_{\mathbb{R}}q^{2}\,dx+\frac{A^{3}}{6}\Bigr)+O(|z|^{-4}).$
Expanding $\log(1+x)=x-\frac{1}{2}x^{2}+\frac{1}{3}x^{3}+O(x^{4})$ and using the displayed expansion shows the even powers cancel, and one obtains
$\log a(z)=\varepsilon A+\varepsilon^{3}\int_{\mathbb{R}}q^{2}\,dx+O(|z|^{-4}),$
which is (167). ∎
Theorem 10.10 (First trace identity).
Let $q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R})$ and let $R_{q}$ be its reflection coefficient. Then
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk,$ (173)
equivalently
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}\,dk.$ (174)
If $q=\beta(\mu)$ for $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, then
$\int_{\mathbb{R}}s(x)^{2}\,dx=\frac{4}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}\,dk.$ (175)
Proof.
By Lemma 10.8 and Lemma 10.9, the coefficient of $z^{-1}$ in the large-$|z|$ expansion of $\log a(z)$ is simultaneously
$\frac{1}{2iz}\int_{\mathbb{R}}q(x)\,dx\qquad\text{and}\qquad\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi$
(we provide more explicit analysis for the coefficient of $z^{-3}$ in the proof of Theorem 10.11). Equating these coefficients yields (173), and (174) is its rearrangement.
If $q=\beta(\mu)$, Proposition 9.11 gives $\int_{\mathbb{R}}q\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx$, and inserting this into (174) gives (175). ∎
Theorem 10.10 converts the linear functional $\int q$ into a spectral entropy integral. The next identity upgrades this to the quadratic energy $\int q^{2}$ and yields, on the density side, a mixed Sobolev/Orlicz functional of the score.
Theorem 10.11 (Second trace identity).
Under the hypotheses of Theorem 10.10,
$\int_{\mathbb{R}}q(x)^{2}\,dx=-\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk=\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}dk.$ (176)
If $q=\beta(\mu)$ for $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, then
$\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}(x)|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s(x)|^{4}\,dx=\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}dk.$ (177)
Proof.
By Proposition 7.2 one has $R_{q}\in\mathcal{S}(\mathbb{R})$, hence $\log(1-|R_{q}|^{2})\in\mathcal{S}(\mathbb{R})$ and all moment integrals below are absolutely convergent. Let $a(z)$ be the transmission coefficient from (74). Since $q$ lies in the Bers image, $H_{q}=-\partial_{x}^{2}-q$ has no bound states, so $a$ has no zeros in $\Im z>0$ and is an outer function (Theorem 7.6). Consequently, for $\Im z>0$,
$\log a(z)=-\frac{1}{2\pi i}\int_{\mathbb{R}}\frac{\log(1-|R_{q}(\xi)|^{2})}{\xi-z}\,d\xi,$ (178)
where the branch is fixed by $\log a(z)\to 0$ as $|z|\to\infty$ in $\Im z>0$. Expanding the Cauchy kernel gives, as $|z|\to\infty$ with $\Im z\geq 0$,
$\frac{1}{\xi-z}=-\frac{1}{z}-\frac{\xi}{z^{2}}-\frac{\xi^{2}}{z^{3}}+O(|z|^{-4}),$
with an error dominated by $C(1+|\xi|^{3})|z|^{-4}$. Since $\log(1-|R_{q}|^{2})$ is Schwartz, termwise integration is justified, and because $q$ is real one has $R_{q}(-\xi)=\overline{R_{q}(\xi)}$, hence $\log(1-|R_{q}(\xi)|^{2})$ is even and the $z^{-2}$ moment vanishes. Therefore
$\log a(z)=\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi+\frac{1}{2\pi i}\frac{1}{z^{3}}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi+O(|z|^{-4}).$ (179)
We now compute the same large-$|z|$ expansion from the Jost/Volterra side, keeping the first two nontrivial odd orders. Let $f_{+}(x,z)$ be the Jost solution and set $m(x,z):=e^{-izx}f_{+}(x,z)$, so that $m(\cdot,z)\to 1$ and $m^{\prime}(\cdot,z)\to 0$ as $x\to+\infty$. Then $m$ satisfies
$m^{\prime\prime}+2iz\,m^{\prime}+q\,m=0.$ (180)
Write $\varepsilon:=(2iz)^{-1}$. Define
$m^{(2)}(x):=1+\varepsilon m_{1}(x)+\varepsilon^{2}m_{2}(x),\qquad m_{1}(x):=\int_{x}^{\infty}q(t)\,dt,\qquad m_{2}(x):=q(x)+\int_{x}^{\infty}q(t)m_{1}(t)\,dt.$ (181)
Using $m_{1}^{\prime}=-q$ and $m_{2}^{\prime}=q^{\prime}-qm_{1}$, a direct substitution of (181) into (180) yields
$m^{(2)}{}^{\prime\prime}+2iz\,m^{(2)}{}^{\prime}+q\,m^{(2)}=\varepsilon^{2}\bigl(m_{2}^{\prime\prime}+qm_{2}\bigr).$
Because $q\in W^{\infty,1}$, one has $m_{1}\in W^{\infty,\infty}$ and $m_{2}\in W^{\infty,1}\cap W^{\infty,\infty}$, hence $m_{2}^{\prime\prime}+qm_{2}\in L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$. Writing $m=m^{(2)}+r$, the remainder $r$ solves
$r^{\prime\prime}+2iz\,r^{\prime}+q\,r=-\varepsilon^{2}(m_{2}^{\prime\prime}+qm_{2}),\qquad r,r^{\prime}\to 0\ \text{as }x\to+\infty.$
By the Volterra estimates from Proposition 7.2 (applied to the inhomogeneous equation),
$\sup_{x\in\mathbb{R}}\bigl(|r(x,z)|+|z|^{-1}|r^{\prime}(x,z)|\bigr)\leq C|z|^{-3},\qquad|z|\gg 1,\ \Im z\geq 0.$ (182)
with $C$ depending only on finitely many $W^{k,1}$ norms of $q$.
The Wronskian identity defining $a$ (the same identity used to obtain the $z^{-1}$ coefficient in the first trace formula) gives
$a(z)=1+\varepsilon\int_{\mathbb{R}}q(x)\,m(x,z)\,dx.$ (183)
Insert $m=m^{(2)}+r$ and use (182) to obtain
$\int_{\mathbb{R}}q\,m\,dx=\int_{\mathbb{R}}q\,dx+\varepsilon\int_{\mathbb{R}}q\,m_{1}\,dx+\varepsilon^{2}\int_{\mathbb{R}}q\,m_{2}\,dx+O(|z|^{-3}).$ (184)
Set $A:=\int_{\mathbb{R}}q$ and $Q(x):=\int_{x}^{\infty}q(t)\,dt=m_{1}(x)$. Then $Q^{\prime}=-q$, $Q(\infty)=0$, and $Q(-\infty)=A$. Consequently
$\int_{\mathbb{R}}q\,m_{1}\,dx=\int_{\mathbb{R}}qQ\,dx=\int_{\mathbb{R}}(-Q^{\prime})Q\,dx=\frac{1}{2}A^{2}.$
Moreover,
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx.$
By Fubini, using $\int_{-\infty}^{t}q(x)\,dx=A-Q(t)$,
$\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx=\int_{\mathbb{R}}q(t)Q(t)\bigl(A-Q(t)\bigr)\,dt=\frac{1}{2}A^{3}-\int_{\mathbb{R}}qQ^{2}\,dt.$
Finally $\int_{\mathbb{R}}qQ^{2}=\int_{\mathbb{R}}(-Q^{\prime})Q^{2}=-\frac{1}{3}[Q^{3}]_{-\infty}^{\infty}=\frac{1}{3}A^{3}$, hence
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\frac{1}{6}A^{3}.$
Substituting into (184) and then (183) yields
$a(z)=1+\varepsilon A+\varepsilon^{2}\frac{A^{2}}{2}+\varepsilon^{3}\Bigl(\int_{\mathbb{R}}q^{2}\,dx+\frac{A^{3}}{6}\Bigr)+O(|z|^{-4}).$
Expanding $\log(1+x)=x-\frac{1}{2}x^{2}+\frac{1}{3}x^{3}+O(x^{4})$ and inserting the displayed asymptotic shows that the even powers cancel and the cubic $A^{3}$ contributions cancel as well, leaving
$\log a(z)=\varepsilon A+\varepsilon^{3}\int_{\mathbb{R}}q(x)^{2}\,dx+O(|z|^{-4})=\frac{1}{2iz}\int_{\mathbb{R}}q\,dx+\frac{1}{(2iz)^{3}}\int_{\mathbb{R}}q^{2}\,dx+O(|z|^{-4}).$ (185)
Comparing the $z^{-3}$ coefficients in (179) and (185) gives
$\frac{1}{(2i)^{3}}\int_{\mathbb{R}}q^{2}\,dx=\frac{1}{2\pi i}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi,$
hence, since $(2i)^{3}=-8i$,
$\int_{\mathbb{R}}q(x)^{2}\,dx=-\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk,$
which is the first equality in (176); the second is its reformulation.
If $q=\beta(\mu)$ with $\mu=g\,dx$ and score $s=(\log g)^{\prime}$, then $q=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ and $s(\pm\infty)=0$. Expanding $q^{2}$ and integrating by parts eliminates the mixed term:
$\int_{\mathbb{R}}q^{2}\,dx=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}|^{2}\,dx-\frac{1}{4}\int_{\mathbb{R}}s^{\prime}s^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s|^{4}\,dx=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s|^{4}\,dx,$
since $\int s^{\prime}s^{2}=\frac{1}{3}\int(s^{3})^{\prime}=0$. Substituting this into (176) gives (177). ∎
Remark 10.12 (Hierarchy).
The two identities above are the first two nontrivial members of the odd-order large-$|z|$ expansion of $\log a(z)$. Namely, the $z^{-2}$ term is absent by parity, and each subsequent odd coefficient yields a conserved differential polynomial in $q$ equated to a spectral moment of $\log(1-|R_{q}|^{2})$. In this paper, we stop at $n=0,1$ ($n$ indexes the coefficient of $z^{-(2n+1)}$), which already produces the quadratic/superquadratic score functional (177) needed as an analytic input for Part II.
Remark 10.13 (Relation to mKdV and the Miura transform).
The mixed score functional in (177),
$\mathcal{H}(s):=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}(x)|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s(x)|^{4}\,dx,$
is not the usual information-geometric energy (compare the Fisher information $\int s^{2}g\,dx$): it is unweighted and mixes an $H^{1}$ term with an $L^{4}$ term. It is nevertheless familiar from one-dimensional integrable PDE. Indeed, setting $v:=\frac{s}{2}$ one has
$\mathcal{H}(s)=\int_{\mathbb{R}}\bigl(|v_{x}|^{2}+|v|^{4}\bigr)\,dx,$
which, up to conventional normalizations of the flow, is the standard Hamiltonian energy for the (defocusing) modified KdV equation [12]. Moreover, the identity
$q=\beta(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}=v_{x}-v^{2}$
is exactly the Miura transform (up to sign conventions) linking mKdV and KdV variables [23]. In particular, $\int_{\mathbb{R}}q^{2}\,dx=\mathcal{H}(s)$, so (177) can be viewed as a scattering-theoretic representation of a classical integrable Hamiltonian in terms of the reflection coefficient. From the inverse-scattering viewpoint, this is a low-order instance of the trace/sum-rule hierarchy expressing conserved differential polynomials in $q$ as spectral moments of $\log(1-|R_{q}|^{2})$ [7, 10].
11. Background on $L^{p}$ Fisher-Rao geometry
We recall the $L^{p}$ Fisher-Rao framework on compact manifolds, following Bauer-Bruveris-Harms-Michor [6].
Let $M$ be a closed oriented $n$-manifold with density bundle $|\Omega|=|\bigwedge^{n}T^{*}M|$, positive sections $|\Omega|_{+}(M)=\{\mu\in\Gamma(|\Omega|)\mid\mu>0\}$, and probability densities $\operatorname{Prob}(M)=\{\mu\in|\Omega|_{+}(M)\mid\int_{M}\mu=1\}$. Fix a reference density $\mathrm{d}x\in|\Omega|_{+}(M)$.
Definition 11.1 ([6]).
For $p\in[1,\infty)$, the $L^{p}$ Fisher-Rao Finsler structure on $|\Omega|_{+}(M)$ is
$F_{p}(\mu,a)=\left(\int_{M}\left|\frac{a}{\mu}\right|^{p}\mu\right)^{1/p},\qquad a\in T_{\mu}|\Omega|_{+}(M)=\Gamma(|\Omega|).$
For $p=2$ this squares to the classical Fisher-Rao Riemannian metric. The Finsler structure is $\operatorname{Diff}(M)$-invariant: $F_{p}(\varphi^{*}\mu,\varphi^{*}a)=F_{p}(\mu,a)$ for every $\varphi\in\operatorname{Diff}(M)$.
Theorem 11.2 ([6]).
The map $\Phi_{p}(\mu)=p(\mu/\mathrm{d}x)^{1/p}$ is an isometric embedding $(|\Omega|_{+}(M),F_{p})\hookrightarrow(L^{p}(M,\mathrm{d}x),\|\cdot\|_{L^{p}})$ onto the positive cone; on $\operatorname{Prob}(M)$ the image lies in the $L^{p}$ sphere of radius $p$. The unique $F_{p}$-geodesic from $\mu_{0}=\rho_{0}\,\mathrm{d}x$ to $\mu_{1}=\rho_{1}\,\mathrm{d}x$ in $|\Omega|_{+}(M)$ is
$\mu(t)=\bigl((1-t)\rho_{0}^{1/p}+t\rho_{1}^{1/p}\bigr)^{p}\mathrm{d}x.$
The classification theorem of Bauer-Bruveris-Modin [5] shows that in dimension $\dim M\geq 2$, every smooth $\operatorname{Diff}(M)$-invariant weak Riemannian metric on $|\Omega|_{+}(M)$ is a linear combination of the Fisher-Rao metric and the total-mass inner product. On $\operatorname{Prob}(M)$, only the Fisher-Rao metric survives. This classification is intrinsically Hilbertian in the sense that the $L^{p}$ Fisher-Rao structures for $p\neq 2$ are Finsler and fall outside its scope, which is the starting point of the present paper.
12. Remarks
This final section records several structural consequences of the results in Sections 2-10. A recurring theme is that the theory admits three complementary coordinate systems-affine ($p$-root), projective (logarithmic/Schwarzian), and spectral (scattering)-each of which linearizes a different aspect of the nonlinear geometry.
The geometric meaning of blow-up: boundary hitting in $p$-root coordinates
Fix $p\in[1,\infty)$. By Theorem 2.8, the $p$-root chart
$\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$
identifies $\operatorname{Diff}^{-\infty}(\mathbb{R})$ with the open half-space $\{f\in W^{\infty,1}(\mathbb{R})\mid f>-p\}$. Moreover, $F_{\dot{W}^{1,p}}$-geodesics are affine in these coordinates. Let $\varphi(t)$ be the geodesic with initial data $(\varphi_{0},u_{0})$ and set $f(t):=\Phi_{p}(\varphi(t))$. Then
$f(t)=f_{0}+t\dot{f}_{0},\qquad f_{0}=\Phi_{p}(\varphi_{0}),\qquad\dot{f}_{0}=\partial_{t}f(t)\big|_{t=0}.$
The positivity constraint $\varphi^{\prime}(t)>0$ is equivalent to f(t)>pf(t)>-p. It is convenient to translate the half-space to the positive cone by setting
$g(t):=f(t)+p=p\,\varphi^{\prime}(t)^{1/p}.$
Since $f(t)$ is affine and $p$ is constant, $g(t)$ is affine as well: $g(t)=g_{0}+t\dot{g}_{0}$ with $g_{0}=p\,\varphi_{0}^{\prime 1/p}>0$. The chain rule applied to $\Phi_{p}$ yields
$\dot{f}_{0}(x)=\varphi_{0}^{\prime}(x)^{1/p-1}\,\partial_{x}(u_{0}\circ\varphi_{0})(x)=(v_{0}\circ\varphi_{0})(x)\,\varphi_{0}^{\prime}(x)^{1/p},\qquad v_{0}:=\partial_{x}u_{0},$
where $v_{0}\circ\varphi_{0}$ is the Lagrangian strain and $\partial_{x}(u_{0}\circ\varphi_{0})=(v_{0}\circ\varphi_{0})\,\varphi_{0}^{\prime}$ by the chain rule. Since the additive constant cancels, $\dot{g}_{0}=\dot{f}_{0}$. Therefore the maximal existence time is the first time the affine path $g(t)$ hits the boundary $\{g=0\}$:
$T^{*}=\inf_{x:\,\dot{g}_{0}(x)<0}\frac{-g_{0}(x)}{\dot{g}_{0}(x)}=\inf_{x:\,v_{0}(\varphi_{0}(x))<0}\frac{-p}{v_{0}(\varphi_{0}(x))}=-\frac{p}{\inf_{x\in\mathbb{R}}v_{0}(x)}.$
This recovers Theorem 3.3. In particular, the blow-up mechanism is literally boundary hitting in an affine chart: the only nonlinearity is the positivity wall $\varphi^{\prime}>0$, and the gradient catastrophe occurs at the first contact point where $g(t,\cdot)$ touches 0, i.e. along characteristics issuing from minimizers of $v_{0}$.
From this viewpoint, the LDP in Section 3.6 is a small-noise theory for the exit time of an affine motion from a convex cone in a Banach space; the spine direction in Theorem 3.18 selects the most likely infinitesimal perturbation that accelerates the exit.
Information-theoretic content of the Bers potential
On the density side (Section 9), let $\mu=g\,\mathrm{d}x$ with score $s=(\log g)^{\prime}$ and density Schwarzian $S(\mu):=s^{\prime}-\frac{1}{2}s^{2}$. The Bers potential is
$\beta(\mu)=\frac{1}{2}S(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2},$ (186)
agreeing with $\beta^{-\infty}([\varphi])=\frac{1}{2}S(\varphi)$ for $\mu=\varphi^{\prime}\,\mathrm{d}x$.
Several interpretations are worth isolating.
  1. (i)
    Ground state transform / canonical nullfunction. Set $\psi:=g^{-1/2}$. A direct calculation gives
    $(\partial_{x}^{2}+\beta(\mu))\,\psi=0.$
    Thus $\beta(\mu)$ is the unique potential for which the half-density $\psi$ is a positive normalized zero-energy state. This is the analytic reason the Bers image admits a Sturm-Liouville characterization (Section 6) and why the Hardy form is critical (Theorem 10.1).
  2. (ii)
    Projective deviation from Lebesgue. In the decay class $|\Omega|_{\infty,1}(\mathbb{R})$, $\beta(\mu)\equiv 0$ forces $s\equiv 0$, hence $\mu=\mathrm{d}x$. Equivalently, on the diffeomorphism side, $S(\varphi)\equiv 0$ forces $\varphi\in\operatorname{Aff}(\mathbb{R})$.
  3. (iii)
    Curvature obstructions and forced sign change. Theorems 9.28 and Corollary 9.29 show that for every nontrivial $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ the Bers potential cannot have a global sign: both $\{x:\beta(\mu)(x)>0\}$ and $\{x:\beta(\mu)(x)<0\}$ are nonempty. This is a one-dimensional projective rigidity phenomenon: the projective curvature encoded by $\beta$ must oscillate.
  4. (iv)
    Miura structure and integrable energies. Writing $v:=\frac{s}{2}$, one has $\beta(\mu)=v_{x}-v^{2}$, i.e. the Miura transform. The induced mixed $H^{1}/L^{4}$ Hamiltonian and its scattering representation are discussed in Remark 10.13.
Finally, since $\mu=\varphi^{\prime}\,\mathrm{d}x$ is exactly the pushforward of $\mathrm{d}x$ by the monotone map $\varphi$, one may also view $\beta(\mu)$ as the Schwarzian curvature of the one-dimensional transport map from $\mathrm{d}x$ to $\mu$. This gives geometric intuition for why $\beta$ naturally couples to Fisher-type energies, even though Fisher-Rao and Wasserstein geometries encode fundamentally different geodesic interpolations.
The $p\to\infty$ limit and logarithmic coordinates
The family of charts $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ interpolates between three natural coordinate systems:
  • $p=1$ (velocity coordinate). $\Phi_{1}(\varphi)=\varphi^{\prime}-1$ measures Jacobian deviation from the identity.
  • $p=2$ (half-density coordinate). $\Phi_{2}(\varphi)=2(\sqrt{\varphi^{\prime}}-1)$ is adapted to the Hilbertian Fisher-Rao case.
  • $p=\infty$ (logarithmic coordinate). $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is additive under composition, and the Schwarzian becomes $S(\varphi)=u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}$ with $u=\Phi_{\infty}(\varphi)$.
The passage $p\to\infty$ is a classical “tropical” limit
$p\bigl(x^{1/p}-1\bigr)\longrightarrow\log x,$
which converts multiplicative Jacobian cocycles into additive logarithmic cocycles and explains why a projective invariant (the Schwarzian) emerges from an affine linearization.
On the cohomological side, the Bott-Thurston/Virasoro class is rigid across $p$: Theorem 5.7 shows that $\omega_{p}=p^{-2}\omega$, hence all $\omega_{p}$ span the same one-dimensional subspace in $H^{2}(\mathfrak{g}^{-\infty};\mathbb{R})$. This rigidity is what ultimately allows a single central extension to absorb all $p$-geometric multipliers.
The scattering transform as nonlinear Fourier analysis
For $q\in W^{\infty,1}(\mathbb{R})$ in the Bers image, the associated scattering data provide a global spectral coordinate system. In the no-bound-state regime (Lemma 7.3), the transmission coefficient $a(z)$ is an outer function in the upper half-plane (i.e. it has no zeros in $\Im z>0$ and is determined by its boundary modulus via the Poisson integral), and the reflection coefficient satisfies $R_{q}\in\mathcal{S}(\mathbb{R})$ (Section 7). Marchenko inversion reconstructs $q$ from $R_{q}$, so the composite
$\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\xrightarrow{\;\beta^{-\infty}\;}\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))\xrightarrow{\;q\mapsto R_{q}\;}\mathcal{S}(\mathbb{R})$
may be regarded as a nonlinear Fourier transform on the quotient.
Two features are particularly relevant downstream.
  1. (i)
    Spectral entropies and trace identities. The trace identities in Section 10.4 express geometric integrals of the Bers potential as spectral moments of $\log(1-|R_{q}|^{2})$. In particular, (175) and (177) identify low-order Fisher/score functionals with explicit spectral “entropies” of the reflection coefficient. These are the analytic control quantities used in the functional estimates of Section 10.
  2. (ii)
    Affine subgroup as trivial scattering. In our class, $R_{q}\equiv 0$ forces $q\equiv 0$ (since there are no bound states), hence $\beta^{-\infty}([\varphi])=0$ and $\varphi\in\operatorname{Aff}(\mathbb{R})$. Thus $R_{q}$ measures precisely the “non-affine” part of a diffeomorphism.
Central extensions and multiplier-free Koopman fields
The Bott cocycle $B_{p}$ and its Lie algebra representative $\omega_{p}$ define the Virasoro-Bott central extension of $\operatorname{Diff}^{-\infty}(\mathbb{R})$. Two points from Section 5 deserve emphasis.
  1. (i)
    Rigidity across $p$. The class $\omega_{p}]$ scales as $p^{-2}[\omega]$ (Theorem 5.7), so the extension is intrinsic to the group and not to a particular $L^{p}$ geometry. From the semiclassical viewpoint, this reflects the fact that the Schwarzian appears universally as the leading correction in symbol calculi; the central extension is the algebraic shadow of this analytic universality.
  2. (ii)
    Cumulant transgression. The cumulant functional (Definition 5.21) produces an explicit coboundary on an appropriate groupoid, and the resulting central extension absorbs the Bott multiplier in Koopman-type formulas (Theorem 5.25). This is the prototype for the multiplier-removal mechanism used in Part II for nonsingular dynamics.
A curvature balance identity
Although $\beta$ must change sign (Theorem 9.28), its mean against the density is always strictly negative unless $\mu=\mathrm{d}x$. Indeed, with $\mu=g\,\mathrm{d}x$ and $s=(\log g)^{\prime}$, integration by parts (using the decay built into $|\Omega|_{\infty,1}(\mathbb{R})$) gives
$\int_{\mathbb{R}}S(\mu)\,\mu=\int_{\mathbb{R}}\Bigl(s^{\prime}-\tfrac{1}{2}s^{2}\Bigr)g\,\mathrm{d}x=\int_{\mathbb{R}}s^{\prime}g\,\mathrm{d}x-\frac{1}{2}\int_{\mathbb{R}}s^{2}g\,\mathrm{d}x=-\frac{3}{2}\int_{\mathbb{R}}s^{2}g\,\mathrm{d}x=-\frac{3}{2}\,I(\mu),$ (187)
where $I(\mu)=\int s^{2}\,\mu$ is the Fisher information. Thus $S(\mu)$ has negative average in the intrinsic measure $\mu$, even though it must take both signs pointwise. The combination of (187) with the forced sign-change theorems provides a concise summary of the projective rigidity detected in Section 9.
Geodesic convexity and comparison with optimal transport
The affine $p$-root structure yields sharp convexity principles for Fisher-type energies. Theorem 10.6 shows that along $L^{p}$ Fisher-Rao geodesics the $p$-Fisher energy is
$I_{p}(\mu(t))=p^{p}\bigl\|(1-t)h_{0}^{\prime}+th_{1}^{\prime}\bigr\|_{L^{p}}^{p},\qquad h_{i}=g_{i}^{1/p},$
hence $t\mapsto I_{p}(\mu(t))$ is convex for every $p\geq 1$, and $t\mapsto I_{p}(\mu(t))^{1/p}$ is convex as well (by the triangle inequality in $L^{p}$). For $p\geq 2$, Clarkson’s inequality upgrades this to uniform convexity at the midpoint. This convexity is immediate from the linearization, in contrast to displacement convexity in Wasserstein geometry which relies on curvature-dimension mechanisms [3, 29]. In one dimension, the monotone transport map from $\mathrm{d}x$ to $\mu$ is precisely the cumulative distribution inverse, so the identification $\mu=\varphi^{\prime}\,\mathrm{d}x$ links the Fisher-Rao and transport pictures at the level of the transport map; understanding how these two convexity theories interact is a natural direction for further work.
Preview of Part II: Koopman operators and exponential charts
Part II [21] develops an Orlicz-Sobolev statistical bundle over a manifold of probability measures equivalent to a fixed reference $\pi$ on a general measurable space $(M,\mathcal{F})$, with nonsingular transformations $T\colon M\to M$ acting by bundle isometries (Koopman operators in Luxemburg geometry). The following table records the dictionary between the one-dimensional scalar theory of Part I and the bundle objects of Part II.
Part I: diffeomorphisms/densities on $\mathbb{R}$ Part II: Koopman-equivariant statistical bundle
$\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $|\Omega|_{\infty,1}(\mathbb{R})$ linked by the Jacobian map Manifold of measures equivalent to a reference probability $\pi$ on $(M,\mathcal{F})$
Linearizing chart $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ and $p$-root chart $g\mapsto g^{1/p}$ Exponential (centered log-density) chart $q\mapsto\log(q/\pi)-\mathbb{E}_{\pi}[\log(q/\pi)]$
Right-invariant $\dot{W}^{1,p}$ / $L^{p}$ Fisher-Rao Finsler structures Luxemburg-Finsler score geometry on Orlicz-Sobolev fibers
Nonlinearity as a positivity wall ($\varphi^{\prime}>0\Leftrightarrow\Phi_{p}(\varphi)>-p$) Absolute continuity/positivity constraint for Radon-Nikodym derivatives
Schwarzian and Bers potential ($\beta=\frac{1}{2}S$) as projective curvature Koopman cocycles (log-Jacobians) and associated curvature potentials
Real Bers embedding and Orlicz Bers embedding $\beta_{\Phi}$ Statistical projection of the bundle into a space of potentials
Scattering coordinate $q\mapsto R_{q}$, Marchenko inversion, trace hierarchy Spectral coordinates via kernel methods and Galerkin hierarchies
Hunter-Saxton catastrophe $T^{*}$ as boundary hitting; LDP and spine action Loss of absolute continuity / cocycle blow-up; large deviations and spine variational principles
Virasoro-Bott extension; cumulant transgression removes multipliers Cumulant-Koopman central extension and multiplier-free Koopman fields
The guiding principle is that Part I supplies a complete one-dimensional model in which the affine, projective, and spectral pictures can be identified explicitly. Part II lifts these mechanisms to infinite-dimensional Orlicz bundles, where the same three pictures reappear as exponential charts, cocycle curvatures, and spectral decompositions.