### v2

# The Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

Abstract.
We develop a real-analytic counterpart of the Bers embedding for the Fréchet Lie group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ of decay-controlled diffeomorphisms of the line. For $p\in[1,\infty)$, the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically linearizes the homogeneous $\dot{W}^{1,p}$ Finsler metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, yielding explicit geodesics and a canonical flat connection whose Eulerian geodesic equation is the generalized Hunter-Saxton equation; for $p=\infty$, logarithmic coordinates provide a global isometry and the Schwarzian derivative emerges as the projective curvature. We construct a real Bers map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ via this Schwarzian, prove it is a Fréchet-smooth injective immersion whose linearization admits a tame right inverse given by an explicit Volterra operator, and characterize its image through Sturm-Liouville spectral theory and one-dimensional scattering. Through the Jacobian correspondence, these structures transfer to a manifold of densities asymptotic to Lebesgue measure, where the Schwarzian becomes the score curvature governing Fisher information. We prove that for every nontrivial density, the projective curvature necessarily takes both signs (forced sign-change), while Fisher information alone does not control the positive projective curvature. Trace identities express Fisher information and its higher analogues as spectral entropies of the scattering reflection coefficient.
Key words and phrases:
Schwarzian derivative, Bers embedding, Fisher-Rao metric, diffeomorphism groups, scattering theory
2020 Mathematics Subject Classification:
58D05, 58B20, 53C21, 34B24, 46E30
1. Introduction
The Schwarzian derivative
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}$
is a classical object in complex analysis and projective geometry, encoding the failure of a diffeomorphism to be a Möbius transformation. Its appearance in diverse contexts, from univalent function theory and Teichmüller spaces to conformal field theory and one-dimensional dynamics, reflects a fundamental principle: the Schwarzian is the unique differential invariant of third order that vanishes precisely on projective transformations and satisfies the chain rule cocycle
$S(\varphi\circ\psi)=(S(\varphi)\circ\psi)\,(\psi^{\prime})^{2}+S(\psi).$
The classical Bers embedding realizes the universal Teichmüller space as a domain in the Banach space of bounded holomorphic quadratic differentials on the lower half-plane. This embedding, fundamental to Teichmüller theory, is constructed via the Schwarzian derivative of quasiconformal extensions.
In this paper we develop a real-analytic counterpart of the Bers embedding for a Fréchet Lie group of diffeomorphisms of $\mathbb{R}$ with controlled decay at infinity and construct, around it, a unified geometric theory that interweaves four subjects:
  1. (i)
    the $L^{p}$ Finsler geometry of diffeomorphism groups and its complete isometric linearization,
  2. (ii)
    the Schwarzian derivative as a projective curvature that governs the passage from affine to projective structures,
  3. (iii)
    one-dimensional scattering theory for Schrödinger operators, which furnishes nonlinear spectral coordinates on the quotient $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$, and
  4. (iv)
    the Fisher-Rao geometry of positive densities, where the Schwarzian reappears as a score curvature and Fisher information becomes a spectral entropy.
These threads are linked by a single mechanism: the $p$-root map $\varphi\mapsto p(\varphi^{\prime 1/p}-1)$ isometrically flattens the right-invariant Finsler geometry on diffeomorphisms and, in the limit $p\to\infty$, produces logarithmic coordinates in which the Schwarzian takes the potential form $S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$, $f=\log(\varphi^{\prime})$. The real Bers embedding, the Sturm-Liouville spectral characterization of its image, the scattering transform, the Virasoro-Bott central extension, the $L^{p}$-Schwarzian interpolation, the Jacobian correspondence to densities, the forced sign-change theorems for projective curvature, and the sharp functional inequalities all follow from this single organizing principle.
1.1. Main results
Let $W^{\infty,1}(\mathbb{R})=\bigcap_{k\geq 0}W^{k,1}(\mathbb{R})$ denote the Fréchet space of smooth functions whose derivatives of all orders lie in $L^{1}(\mathbb{R})$. Define the decay-controlled diffeomorphism group
$\operatorname{Diff}^{-\infty}(\mathbb{R})=\left\{\varphi=\operatorname{id}+f\mid f^{\prime}\in W^{\infty,1}(\mathbb{R}),\;f^{\prime}>-1,\;\lim_{x\to-\infty}f(x)=0\right\}.$
This is a Fréchet Lie group with Lie algebra $\mathfrak{g}^{-\infty}=\{u\mid u^{\prime}\in W^{\infty,1}(\mathbb{R}),\;\lim_{x\to-\infty}u(x)=0\}$, equipped with the right-invariant homogeneous $\dot{W}^{1,p}$ Finsler structure
$F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\left\|u^{\prime}\right\|_{L^{p}(\mathbb{R})},\qquad u=\delta\varphi\circ\varphi^{-1}.$
A.  Isometric linearization, flat connections, and geodesic dynamics
Our first main result establishes that this Finsler geometry admits a global linearization.
Theorem 1.1 (Isometric linearization, Theorem 2.8).
For $p\in[1,\infty)$, the map
$\Phi_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$
is an isometric embedding from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ into $(W^{\infty,1}(\mathbb{R}),\left\|\cdot\right\|_{L^{p}})$. Its image is the open set $\{f\in W^{\infty,1}(\mathbb{R})\mid f>-p\}$. For $p=\infty$, the map $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is an isometric diffeomorphism onto $W^{\infty,1}(\mathbb{R})$.
The linearization trivializes the geodesic problem: minimizing $F_{\dot{W}^{1,p}}$-geodesics become affine segments in the target space. We derive from this a canonical flat connection $\nabla^{(p)}$ on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, compute its parallel transport explicitly (which has trivial holonomy), and show that the geodesic equation in Eulerian form is the strain equation
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0,$
whose $x$-derivative gives the generalized Hunter-Saxton equation. For $p=2$ this is the classical Hunter-Saxton equation. As $p\to\infty$, the quadratic nonlinearity vanishes and geodesics become solutions of the pure transport equation $\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})=0$. The connections $\nabla^{(p)}$ admit a two-parameter deformation $\nabla^{(p,\alpha)}$ by a projective tensor $\mathcal{C}(u,a)=u_{xx}a-ua_{xx}$; the resulting curvature involves the Lie bracket structure of vector fields through the commutator $\partial_{x}^{2},\mathcal{L}_{[u,v]}]$.
The family $\{\Phi_{p}\}_{p\in[1,\infty]}$ admits a full asymptotic expansion in Fréchet seminorms as $p\to\infty$ (Proposition 2.16).
B.  The Schwarzian as projective curvature
The Schwarzian derivative emerges naturally in the $p\to\infty$ limit through the potential identity
$S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2},\qquad f=\Phi_{\infty}(\varphi)=\log(\varphi^{\prime}).$
We develop the Schwarzian as a score curvature on densities: if $\mu=g\,dx$ is a positive density with score $s=(\log g)^{\prime}$, then $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. The mean projective curvature satisfies the identity $\int S(\mu)\,g\,dx=-\frac{3}{2}I(\mu)$ where $I(\mu)=\int s^{2}\,g\,dx$ is the Fisher information (Proposition 3.8), and the Schwarzian admits a projective decomposition inside the second-order score composite that yields a Bhattacharyya-type variance bound for unbiased estimation (Theorem 3.10).
We introduce the $L^{p}$-Schwarzian to interpolate between the affine linearizations $\Phi_{p}$ and the projective Schwarzian.
Theorem 1.2 ($L^{p}$-Schwarzian, Propositions 3.20 and 3.24).
For $p\in[1,\infty)$, the quantity
$S_{p}(\varphi)=\left(\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}+S(\varphi)\right)\varphi^{\prime 1/p}$
satisfies the cocycle identity $S_{p}(\varphi\circ\psi)=(S_{p}(\varphi)\circ\psi)\,\psi^{\prime 1/p}+S_{p}(\psi)$ and converges to $S(\varphi)$ as $p\to\infty$ with full asymptotic expansion in $W^{k,1}(\mathbb{R})$:
$S_{p}(\varphi)=S(\varphi)+\frac{1}{p}\left(uS+\frac{3}{2}(u^{\prime})^{2}\right)+O_{W^{k,1}}\!\left(\frac{1}{p^{2}}\right),\qquad u=\log(\varphi^{\prime}).$
The Sturm-Liouville equation $y^{\prime\prime}+qy=0$ provides a canonical differential-geometric framework for the Schwarzian.
C.  Central extensions and cocycle geometry
The Schwarzian is the fundamental obstruction to lifting the composition law on diffeomorphisms to an affine structure. We develop the cohomological content of this obstruction through several constructions. The Gelfand-Fuchs $2$-cocycle $\omega(u,v)=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx$ is shown to represent a nontrivial class in $H^{2}(\mathfrak{g}^{-\infty};\mathbb{R})$ (Proposition 4.2), and it integrates to the Virasoro-Bott central extension $\widehat{\operatorname{Diff}}^{-\infty}(\mathbb{R})$ via the Bott-Thurston group cocycle $B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})$.
We prove a rigidity theorem: the $p$-root Jacobian cocycle produces no new Lie algebra cohomology classes beyond the Gelfand-Fuchs class, up to explicit rescaling (Section 4). In contrast, the $L^{p}$-Schwarzian fails to satisfy a two-term cocycle identity; the unavoidable cross-term in its composition law is a manifestation of the fact that projective curvature cannot be captured by a simple $L^{p}$ cocycle.
D.  The real Bers embedding
Since $S(\varphi)=0$ if and only if $\varphi$ is affine, the Schwarzian descends to the quotient $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
Theorem 1.3 (Real Bers embedding, Theorem 5.7).
The map
$\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\longrightarrow W^{\infty,1}(\mathbb{R}),\qquad\beta^{-\infty}([\varphi])=\frac{1}{2}S(\varphi)$
is a Fréchet smooth injective immersion. Its differential at $\varphi]$ admits a continuous tame right inverse $R_{u}\colon W^{k,1}(\mathbb{R})\to W^{k+2,1}(\mathbb{R})$ for every $k\geq 0$, given by the explicit Volterra operator
$(R_{u}v)(x)=\int_{-\infty}^{x}\mathrm{e}^{u(t)}\int_{-\infty}^{t}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y\,\mathrm{d}t,\qquad u=\log(\varphi^{\prime}).$
The Bers potential $q=\beta^{-\infty}(\varphi)$ admits a Miura-type factorization $q=\frac{1}{2}u^{\prime\prime}-\frac{1}{4}(u^{\prime})^{2}$ (Proposition 5.6), where $u=\log(\varphi^{\prime})$ plays the role of a Miura variable. This factorization guarantees that the associated Schrödinger operator $H_{q}=-\partial_{x}^{2}-q$ is nonnegative and has no bound states: $H_{q}=A^{*}A$ where $A=\partial_{x}+\frac{1}{2}u^{\prime}$. The image of $\beta^{-\infty}$ admits a spectral characterization: a potential $q\in W^{\infty,1}(\mathbb{R})$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the Sturm-Liouville equation $y^{\prime\prime}+qy=0$ admits a positive solution $y_{2}$ with $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$, and the reconstruction is $\varphi=y_{1}/y_{2}$ where $y_{1}$ is the second solution with Wronskian $W(y_{1},y_{2})\equiv 1$ (Theorem 5.8).
E.  Scattering coordinates and nonlinear spectral analysis
The Miura factorization implies that the Schrödinger operator $H_{q}$ for a Bers potential $q=\beta^{-\infty}(\varphi)$ has essential spectrum $0,\infty)$ with no negative eigenvalues. The scattering data therefore consist of the reflection coefficient alone, and this defines a nonlinear spectral coordinate.
Theorem 1.4 (Scattering Bers map, Theorem 6.6).
The map $\mathcal{S}(\varphi):=R_{\beta^{-\infty}(\varphi)}$, assigning to each diffeomorphism the reflection coefficient of its Bers potential, descends to an injective map
$\mathcal{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\hookrightarrow\mathscr{S}(\mathbb{R}).$
The reflection coefficient satisfies the reality condition $\overline{R(-k)}=R(k)$, the strict sub-unitarity $|R(k)|<1$ for $k\neq 0$, and the transmission denominator is the outer function (35) determined by $|R|$.
We develop the scattering theory in detail. Jost solutions and their Volterra integral equations are constructed for potentials in $W^{\infty,1}(\mathbb{R})$, yielding Schwartz-class scattering data (Section 6).
F.  Density geometry: projective curvature and forced sign-change
The Jacobian map $J(\varphi)=\varphi^{\prime}\,\mathrm{d}x$ establishes a Fréchet diffeomorphism and Finsler isometry between $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the manifold $|\Omega|_{\infty,1}(\mathbb{R})$ of positive densities asymptotic to Lebesgue measure (Theorem 7.4). This correspondence intertwines all linearizations: the $p$-root map $\Phi_{p}$ on diffeomorphisms corresponds to the statistical $p$-root map $\Psi_{p}(g\,dx)=p(g^{1/p}-1)$ on densities, and the flat connection, geodesics, and Fisher-Rao metric transfer canonically.
On the density side, the Schwarzian becomes the score curvature $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$, the Bers potential is $\beta(\mu)=\frac{1}{2}S(\mu)$, and the generalized Hunter-Saxton equation becomes the pointwise Riccati law $\partial_{t}v+\frac{1}{p}v^{2}=0$ for the statistical velocity $v=\partial_{t}\log(g)$.
A central qualitative result is the forced sign-change theorem: for every nontrivial density $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$, the Bers potential $\beta(\mu)$ necessarily takes both positive and negative values (Theorem 7.18). We derive a hierarchy of structural consequences: every nontrivial density has at least two projective inflection points (Proposition 7.20), the Bers potential is negative at local maxima and positive at local minima of the density (Proposition 7.21), and log-concavity regions are always contained in the negative projective curvature region (Proposition 7.22). Perhaps most strikingly, Fisher information alone does not control the positive projective curvature (Theorem 7.24).
G.  Sharp functional inequalities
The Bers image identifies a diffeomorphism with a one-dimensional Schrödinger potential subject to a critical spectral constraint: the operator $H_{q}=-\partial_{x}^{2}-q$ is nonnegative with no bound states. We extract several precise analytic consequences.
The ground state transform $\psi=(\varphi^{\prime})^{-1/2}$ gives a sharp Hardy inequality $\int qf^{2}\,dx\leq\int(f^{\prime})^{2}\,dx$ in which $q$ is a critical Hardy weight: no positive remainder can be added (Theorem 8.1). On finite windows, the correct substitute for a spectral gap is a Hardy-Poincaré inequality after projection onto the orthogonal complement of the resonant direction, with an explicit lower bound (Theorem 8.3).
The main analytic results in this section are trace identities that express the geometric integrals of the Bers potential in spectral variables. The first trace identity equates $\int q\,dx$ with a spectral entropy of the reflection coefficient (Theorem 8.7):
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log(1-|R_{q}(k)|^{2})\,dk.$
On the density side, this becomes the identity $\int s^{2}\,dx=\frac{4}{\pi}\int\log(1-|R_{q}(k)|^{2})^{-1}\,dk$, expressing Fisher information as a spectral entropy. The second trace identity upgrades this to the quadratic level, expressing $\int q^{2}\,dx$ as a frequency-weighted spectral entropy (Theorem 8.8); on the density side this yields an identity for the mixed Sobolev/Orlicz functional $\frac{1}{4}\int|s^{\prime}|^{2}\,dx+\frac{1}{16}\int|s|^{4}\,dx$ in terms of $\int k^{2}\log(1-|R_{q}|^{2})^{-1}\,dk$.
1.2. Relation to prior work
The $L^{p}$ Fisher-Rao geometry on spaces of probability densities over compact manifolds was systematically developed by Bauer, Bruveris, Harms, and Michor [4], who proved that the $p$-root map $\mu\mapsto p(\mu/\mathrm{d}x)^{1/p}$ isometrically embeds the space of positive densities into $L^{p}$. Our treatment of $L^{p}$ geometry on densities in Section 9 recalls their framework and extends it to the noncompact setting $\mathbb{R}$ with decay conditions. The classification of diffeomorphism-invariant Riemannian metrics on densities is due to Bauer, Bruveris, and Michor [3].
The Fréchet Lie group structure on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the existence of right-invariant Sobolev metrics were established by Bauer, Bruveris, and Michor [2]. The classical Bers embedding in Teichmüller theory is treated in Lehto [14]. The Schwarzian derivative and its role in projective geometry are classical; see Ovsienko and Tabachnikov [16] for a modern treatment and Fuks [10] for the cohomological aspects. The Hunter-Saxton equation as a geodesic equation on diffeomorphism groups was identified by Khesin and Misiołek [11]; the geometry of infinite-dimensional groups is treated comprehensively in Khesin and Wendt [12]. The one-dimensional inverse scattering theory on the line is due to Deift and Trubowitz [7]; see also Beals, Deift, and Tomei [5] and Faddeev and Takhtajan [8]. The Miura transformation connecting KdV and mKdV equations originates in [15]. The Bott-Thurston cocycle and its role in the theory of characteristic classes of diffeomorphism groups appears in Bott [6] and the Virasoro extension in Pressley and Segal [17]; the cohomological framework is developed in Feigin and Fuchs [9] and Fuks [10]. Orlicz space theory follows Rao and Ren [18] and the Sobolev embedding theorems are from Adams and Fournier [1].
What is new in the present work, beyond the construction of the real Bers embedding itself, is the synthesis: the simultaneous development of the $L^{p}$ linearization, the $L^{p}$-Schwarzian interpolation, the scattering coordinate, the Orlicz-Sobolev extension, the density-side projective curvature theory with forced sign-change and Fisher non-control phenomena, the central extension gerbe on the density quotient with cumulant transgression, the complete blow-up theory with large deviations and spine decomposition, and the trace identities expressing Fisher energies as spectral entropies-all derived from the single organizing principle of $p$-root coordinates and their $p\to\infty$ limit.
1.3. Organization
The paper is organized as follows. Section 2 establishes the isometric linearization of $\operatorname{Diff}^{-\infty}(\mathbb{R})$, derives the flat connection, geodesics, and the Fréchet asymptotics as $p\to\infty$. Section 3 develops the Schwarzian derivative as a projective potential, introduces the score curvature and its Fisher-Rao identities, constructs the $L^{p}$-Schwarzian family, and gives the Sturm-Liouville interpretation. Section 4 treats the Gelfand-Fuchs cocycle, the Virasoro-Bott group extension, and $L^{p}$ cocycle rigidity. Section 5 constructs the real Bers embedding, proves injectivity and the tame right inverse theorem, establishes the Miura factorization, and gives the spectral characterization of the image. Section 6 develops the scattering theory for Bers potentials: Jost solutions, the absence of bound states, and the scattering transform as a nonlinear Fourier map. Section 7 develops the density-side theory: the Jacobian isomorphism, score functions and projective curvature, the forced sign-change theorems with their inflection-point and Fisher non-control consequences. Section 8 establishes the sharp functional inequalities: critical Hardy theory, and the trace hierarchy expressing Fisher energies as spectral entropies of the reflection coefficient. An appendix recalls the $L^{p}$ Fisher-Rao framework on compact manifolds.
Broader programme. This paper is the first of a two-part work. The blow-up analysis for the generalized Hunter-Saxton equation (large deviations and spine decomposition), the extension to Orlicz-Sobolev scales, gerbe structures on the density quotient, and the Koopman-equivariant statistical bundle over general non-singular dynamical systems will appear in a companion paper and in Part II [13].
Acknowledgments
The present paper was initiated during the author’s doctoral studies at Northwestern University and continued during his appointment at Worcester Polytechnic Institute. The geometric and microlocal analysis traditions at Northwestern shaped the analytical perspective underlying this work. The author thanks the late Steve Zelditch, Jared Wunsch, and Elton P. Hsu for many inspiring conversations. Partially supported by NSF RTG grant DMS-1502632.
2. Isometric linearization and geodesic structure of the diffeomorphism group
2.1. The group $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and the $\dot{W}^{1,p}$ metric
Let $W^{k,1}(\mathbb{R})$ denote the Sobolev space of functions with $k$ derivatives in $L^{1}(\mathbb{R})$, and let $W^{\infty,1}(\mathbb{R})=\bigcap_{k\geq 0}W^{k,1}(\mathbb{R})$ with the Fréchet topology induced by the seminorms $\left\|\cdot\right\|_{W^{k,1}}$.
Definition 2.1.
The decay-controlled diffeomorphism group is
$\operatorname{Diff}^{-\infty}(\mathbb{R})=\left\{\varphi=\operatorname{id}+f\ \middle|\ f^{\prime}\in W^{\infty,1}(\mathbb{R}),\ f^{\prime}>-1,\ \lim_{x\to-\infty}f(x)=0\right\}.$
Its Lie algebra is
$\mathfrak{g}^{-\infty}=\left\{u\ \middle|\ u^{\prime}\in W^{\infty,1}(\mathbb{R}),\ \lim_{x\to-\infty}u(x)=0\right\}.$
The condition $f^{\prime}\in W^{1,1}(\mathbb{R})$ implies that $f^{\prime}$ extends continuously to $\infty$$\infty$] with $f^{\prime}(\pm\infty)=0$, hence $\varphi^{\prime}(x)\to 1$ as $x\to\pm\infty$. Together with $f^{\prime}>-1$ this gives $\varphi^{\prime}>0$ on $\mathbb{R}$, so $\varphi$ is a diffeomorphism.
Proposition 2.2 ([2]).
$\operatorname{Diff}^{-\infty}(\mathbb{R})$ is a Fréchet Lie group with Lie algebra $\mathfrak{g}^{-\infty}$.
2.2. Fiberwise trivialization and proto-bundle geometry
Proposition 2.3 (Tangent bundle trivialization).
The right-trivialization
$\Psi\colon T\operatorname{Diff}^{-\infty}(\mathbb{R})\to\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty},\qquad\Psi(\varphi,\delta\varphi)=\bigl(\varphi,\,\delta\varphi\circ\varphi^{-1}\bigr)$
is a Fréchet diffeomorphism, with inverse
$\Psi^{-1}(\varphi,u)=(\varphi,u\circ\varphi).$
Under this trivialization:
  1. (i)
    for $p\in[1,\infty]$ the right-invariant $\dot{W}^{1,p}$ Finsler structure satisfies
    $F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\|u^{\prime}\|_{L^{p}(\mathbb{R})}$
    for $(\varphi,u)=\Psi(\varphi,\delta\varphi)$, and is independent of $\varphi$;
  2. (ii)
    for $p\in[1,\infty)$ one has the fiberwise identity
    $D_{\varphi}\Phi_{p}(u\circ\varphi)=(u^{\prime}\circ\varphi)\,\varphi^{\prime 1/p}.$
Proof.
Because $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is a Fréchet Lie group (Proposition 2.2), right translation $R_{\varphi}(\psi)=\psi\circ\varphi$ and inversion $\varphi\mapsto\varphi^{-1}$ are smooth maps. The differential of $R_{\varphi}$ at the identity is $T_{\operatorname{id}}R_{\varphi}(u)=u\circ\varphi$. Consequently the map
$\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty}\to T\operatorname{Diff}^{-\infty}(\mathbb{R}),\qquad(\varphi,u)\longmapsto(\varphi,u\circ\varphi)$
is smooth. Composing with inversion gives the smooth map $\Psi$, and the displayed formula $\Psi^{-1}(\varphi,u)=(\varphi,u\circ\varphi)$ is its inverse. This proves that $\Psi$ is a Fréchet diffeomorphism.
For (i), by definition $u=\delta\varphi\circ\varphi^{-1}$ is the Eulerian velocity, hence $F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\|u^{\prime}\|_{L^{p}}$.
For (ii), differentiate $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ in the direction $\delta\varphi=u\circ\varphi$. As in the computation in Theorem 2.8, one has $D_{\varphi}\Phi_{p}(\delta\varphi)=\varphi^{\prime 1/p-1}\delta\varphi^{\prime}$, and $\delta\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, which yields the stated identity. ∎
Definition 2.4.
For $p\in[1,\infty]$, the $\dot{W}^{1,p}$ right-invariant Finsler structure on $T\operatorname{Diff}^{-\infty}(\mathbb{R})$ is
$F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)=\bigl\|(\delta\varphi\circ\varphi^{-1})^{\prime}\bigr\|_{L^{p}(\mathbb{R})}.$
Definition 2.5 (Jacobian density manifold).
Let $|\Omega|_{\infty,1}(\mathbb{R})$ be the Fréchet manifold of positive densities $\mu=\rho\,dx$ with $\rho>0$ and $\rho-1\in W^{\infty,1}(\mathbb{R})$. We identify $|\Omega|_{\infty,1}(\mathbb{R})$ with the open subset $\{\rho\in W^{\infty,1}(\mathbb{R})\mid\rho>0\}$ via $\mu\leftrightarrow\rho$.
Proposition 2.6 (Jacobian map as a global chart).
The Jacobian map
$J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R}),\qquad J(\varphi)=\varphi^{\prime}\,dx$
is a Fréchet diffeomorphism. Its inverse is
$J^{-1}(\rho\,dx)(x)=x+\int_{-\infty}^{x}(\rho(s)-1)\,ds.$
Proof.
If $\varphi=\operatorname{id}+f\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, then $\rho=\varphi^{\prime}=1+f^{\prime}$ satisfies $\rho-1=f^{\prime}\in W^{\infty,1}(\mathbb{R})$ and $\rho>0$, hence $J(\varphi)\in|\Omega|_{\infty,1}(\mathbb{R})$.
Conversely, let $\rho\in W^{\infty,1}(\mathbb{R})$ satisfy $\rho>0$ and $\rho-1\in W^{\infty,1}(\mathbb{R})$, and define
$\varphi(x)=x+\int_{-\infty}^{x}(\rho(s)-1)\,ds.$
Then $\varphi^{\prime}=\rho$, and $\varphi-\operatorname{id}$ has derivative $\rho-1\in W^{\infty,1}(\mathbb{R})$ and satisfies $\lim_{x\to-\infty}(\varphi(x)-x)=0$. Therefore $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $J(\varphi)=\rho\,dx$. This proves bijectivity with the displayed inverse.
Smoothness follows because $\varphi\mapsto\varphi^{\prime}$ is a continuous linear map $\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ in the model topology, and the inverse is the continuous linear operator $h\mapsto\int_{-\infty}^{x}h$ composed with addition of the identity. ∎
Definition 2.7 (Proto-statistical bundle).
Define the product bundle
$\pi\colon E^{-\infty}:=|\Omega|_{\infty,1}(\mathbb{R})\times W^{\infty,1}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R})$
with fiber $E_{\mu}=W^{\infty,1}(\mathbb{R})$. The derivative map on Eulerian fields induces a bundle isomorphism
$\widetilde{J}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\times\mathfrak{g}^{-\infty}\to E^{-\infty},\qquad\widetilde{J}(\varphi,u)=\bigl(J(\varphi),\,u^{\prime}\bigr),$
covering $J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R})$.
Proof.
The map $\widetilde{J}$ is smooth because $J$ is smooth and $u\mapsto u^{\prime}$ is a continuous linear map $\mathfrak{g}^{-\infty}\to W^{\infty,1}(\mathbb{R})$.
To prove bijectivity, let $(\mu,w)\in E^{-\infty}$ with $\mu=\rho\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and $w\in W^{\infty,1}(\mathbb{R})$. Define $\varphi:=J^{-1}(\mu)$ using Proposition 2.6, and define
$u(x):=\int_{-\infty}^{x}w(s)\,ds.$
Then $u\in\mathfrak{g}^{-\infty}$ and $u^{\prime}=w$. Moreover $\widetilde{J}(\varphi,u)=(\mu,w)$ by construction. This provides a two-sided inverse, and the inverse map is smooth because it is given by the smooth chart $J^{-1}$ and a continuous linear integration operator. ∎
2.3. Isometric embedding via $p$-root coordinates
Theorem 2.8.
For $p\in[1,\infty)$, the map
$\Phi_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$
is an isometric embedding from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ into $(W^{\infty,1}(\mathbb{R}),\left\|\cdot\right\|_{L^{p}})$. Its image is $U_{p}=\{f\in W^{\infty,1}(\mathbb{R})\mid f>-p\}$, and the inverse is
$\Phi_{p}^{-1}(f)(x)=x+\int_{-\infty}^{x}\left((1+f(\tilde{x})/p)^{p}-1\right)\mathrm{d}\tilde{x}$.
Proof.
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\varphi^{\prime}=1+h^{\prime}$ where $h^{\prime}\in W^{\infty,1}(\mathbb{R})$ and $h^{\prime}>-1$. Since $W^{\infty,1}(\mathbb{R})\subset C^{\infty}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$, there exist constants $0<c\leq\varphi^{\prime}\leq C<\infty$. The function $\rho\mapsto\rho^{1/p}$ is smooth on $(0,\infty)$, so $\varphi^{\prime 1/p}$ is smooth with
$\partial_{x}(\varphi^{\prime 1/p})=\frac{1}{p}\varphi^{\prime 1/p-1}\varphi^{\prime\prime}.$
Since $\varphi^{\prime\prime}\in L^{1}(\mathbb{R})$ and $\varphi^{\prime 1/p-1}$ is bounded, we have $\partial_{x}(\varphi^{\prime 1/p})\in L^{1}(\mathbb{R})$. Higher derivatives are handled similarly using Faà di Bruno’s formula. Each $\partial_{x}^{k}(\varphi^{\prime 1/p})$ is a finite sum of products of bounded functions of $\varphi^{\prime}$ with derivatives $\varphi^{(j)}$ for $j\geq 2$, all of which lie in $L^{1}(\mathbb{R})$. Thus $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)\in W^{\infty,1}(\mathbb{R})$.
The image characterization follows from $\varphi^{\prime 1/p}=1+f/p>0$ if and only if f>pf>-p. The inverse formula is verified by differentiation. If $\varphi=\Phi_{p}^{-1}(f)$, then $\varphi^{\prime}=(1+f/p)^{p}$, so $\varphi^{\prime 1/p}=1+f/p$ and $\Phi_{p}(\varphi)=f$.
For the isometry, let $\delta\varphi\in T_{\varphi}\operatorname{Diff}^{-\infty}(\mathbb{R})$ and write $\delta\varphi=u\circ\varphi$ with $u\in\mathfrak{g}^{-\infty}$. Differentiating $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ gives
$D_{\varphi}\Phi_{p}(\delta\varphi)=\varphi^{\prime 1/p-1}\delta\varphi^{\prime}.$
Since $\delta\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, we have
$D_{\varphi}\Phi_{p}(\delta\varphi)=(u^{\prime}\circ\varphi)\varphi^{\prime 1/p}.$
Computing the $L^{p}$ norm and changing variables $y=\varphi(x)$,
$\left\|D_{\varphi}\Phi_{p}(\delta\varphi)\right\|_{L^{p}}^{p}=\int_{\mathbb{R}}|u^{\prime}(\varphi(x))|^{p}\varphi^{\prime}(x)\,\mathrm{d}x=\int_{\mathbb{R}}|u^{\prime}(y)|^{p}\,\mathrm{d}y=\left\|u^{\prime}\right\|_{L^{p}}^{p}.$
Thus $\left\|D_{\varphi}\Phi_{p}(\delta\varphi)\right\|_{L^{p}}=F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)$.
Smoothness of $\Phi_{p}$ and $\Phi_{p}^{-1}$ follows from smoothness of $\varphi\mapsto\varphi^{\prime}$ and the Nemytskii operators $\rho\mapsto\rho^{1/p}$ and $f\mapsto(1+f/p)^{p}$ on the relevant open sets. ∎
Corollary 2.9.
The map $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is an isometric diffeomorphism from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),\dot{W}^{1,\infty})$ onto $(W^{\infty,1}(\mathbb{R}),L^{\infty})$ with inverse
$\Phi_{\infty}^{-1}(f)(x)=x+\int_{-\infty}^{x}(\mathrm{e}^{f(\tilde{x})}-1)\,\mathrm{d}\tilde{x}$.
Moreover, $\Phi_{p}(\varphi)\to\Phi_{\infty}(\varphi)$ pointwise as $p\to\infty$.
Remark 2.10.
The Jacobian $\varphi^{\prime}$ is the Radon-Nikodym derivative $\mathrm{d}(\varphi^{*}\mathrm{d}x)/\mathrm{d}x$. The coordinate $\Phi_{p}(\varphi)=p(\varphi^{\prime 1/p}-1)$ is the centered $p$-root of this density ratio, with centering chosen to ensure membership in $W^{\infty,1}(\mathbb{R})$ given the normalization $\varphi^{\prime}(\pm\infty)=1$. If $A=a\circ\varphi\in T_{\varphi}\operatorname{Diff}^{-\infty}(\mathbb{R})$, then
$D_{\varphi}\Phi_{p}(A)=(a_{x}\circ\varphi)\,\varphi^{\prime 1/p}.$
2.4. The multiplicative cocycle
Proposition 2.11.
Define $\Theta_{p}(\varphi)=1+\Phi_{p}(\varphi)/p=\varphi^{\prime 1/p}$ for $p<\infty$. Then
$\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\,\Theta_{p}(\psi).$
For $p=\infty$, this becomes the additive cocycle $\Phi_{\infty}(\varphi\circ\psi)=(\Phi_{\infty}(\varphi)\circ\psi)+\Phi_{\infty}(\psi)$.
Proof.
The chain rule gives $(\varphi\circ\psi)^{\prime}=(\varphi^{\prime}\circ\psi)\psi^{\prime}$. Taking $p$-th roots yields $\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\Theta_{p}(\psi)$. The $p=\infty$ case follows by taking logarithms. ∎
2.5. Distance and geodesics
Corollary 2.12.
For $p\in(1,\infty)$, the induced path metric on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is
$d_{p}(\varphi_{0},\varphi_{1})=\left\|\Phi_{p}(\varphi_{0})-\Phi_{p}(\varphi_{1})\right\|_{L^{p}}.$
The unique minimizing geodesic from $\varphi_{0}$ to $\varphi_{1}$ is $\varphi(t)=\Phi_{p}^{-1}((1-t)\Phi_{p}(\varphi_{0})+t\Phi_{p}(\varphi_{1}))$, explicitly
$\varphi(t,x)=x+\int_{-\infty}^{x}\left(\left((1-t)\varphi_{0}^{\prime 1/p}(s)+t\varphi_{1}^{\prime 1/p}(s)\right)^{p}-1\right)\mathrm{d}s.$
Proof.
The map $\Phi_{p}$ is an isometric embedding into the strictly convex Banach space $L^{p}(\mathbb{R})$ for $p\in(1,\infty)$. The induced path distance therefore equals the ambient norm distance, and the unique minimizing curve is the pullback of the affine segment between the endpoints. ∎
2.6. The canonical flat connection and a projective deformation
Proposition 2.13.
For $p\in(1,\infty]$, there exists a unique torsion-free flat connection $\nabla^{(p)}$ on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ characterized by
$D_{\varphi(t)}\Phi_{p}\left(\nabla^{(p)}_{\dot{\varphi}}A\right)=\frac{\mathrm{d}}{\mathrm{d}t}\left(D_{\varphi(t)}\Phi_{p}(A(t))\right)$
for smooth curves $\varphi(t)$ and vector fields $A(t)$ along them. In Eulerian variables $A=a\circ\varphi$, $\dot{\varphi}=u\circ\varphi$, one has
$\nabla^{(p)}_{\dot{\varphi}}A=\left(\partial_{t}a+u\,\partial_{x}a-\frac{p-1}{p}\,\partial_{x}^{-1}(a_{x}u_{x})\right)\circ\varphi,\qquad(\partial_{x}^{-1}f)(x):=\int_{-\infty}^{x}f(s)\,ds.$
The geodesic equation $\nabla^{(p)}_{\dot{\varphi}}\dot{\varphi}=0$ is equivalent to the strain equation
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0.$
Proof.
Since $\Phi_{p}$ is a global chart from $\operatorname{Diff}^{-\infty}(\mathbb{R})$ onto the open subset $U_{p}\subset W^{\infty,1}(\mathbb{R})$, define $\nabla^{(p)}$ by requiring that for every smooth curve $\varphi(t)$ and every smooth vector field $A(t)$ along $\varphi(t)$,
$D_{\varphi(t)}\Phi_{p}\!\left(\nabla^{(p)}_{\dot{\varphi}}A\right)=\frac{d}{dt}\Bigl(D_{\varphi(t)}\Phi_{p}(A(t))\Bigr).$
This determines a unique affine connection. The connection is torsion-free and flat because it is obtained from the canonical flat connection on the affine space $W^{\infty,1}(\mathbb{R})$ by transport through the chart $\Phi_{p}$.
To obtain the Eulerian formula, write $A=a\circ\varphi$ and $\dot{\varphi}=u\circ\varphi$. By Remark 2.10,
$D_{\varphi}\Phi_{p}(A)=(a_{x}\circ\varphi)\,\varphi^{\prime 1/p}.$
Differentiate in $t$:
$\frac{d}{dt}\Bigl((a_{x}\circ\varphi)\,\varphi^{\prime 1/p}\Bigr)=\Bigl((\partial_{t}a_{x}+u\,\partial_{x}a_{x})\circ\varphi\Bigr)\varphi^{\prime 1/p}+(a_{x}\circ\varphi)\,\frac{d}{dt}\bigl(\varphi^{\prime 1/p}\bigr).$
Since $\partial_{t}\varphi^{\prime}=(u^{\prime}\circ\varphi)\varphi^{\prime}$, one has
$\frac{d}{dt}\bigl(\varphi^{\prime 1/p}\bigr)=\frac{1}{p}\,\varphi^{\prime 1/p-1}\,\partial_{t}\varphi^{\prime}=\frac{1}{p}\,(u^{\prime}\circ\varphi)\,\varphi^{\prime 1/p}.$
Therefore
$\frac{d}{dt}\Bigl(D_{\varphi}\Phi_{p}(A)\Bigr)=\Bigl(\partial_{t}a_{x}+u\,\partial_{x}a_{x}+\frac{1}{p}a_{x}u_{x}\Bigr)\circ\varphi\cdot\varphi^{\prime 1/p}.$
Write $\nabla^{(p)}_{\dot{\varphi}}A=b\circ\varphi$. Then $D_{\varphi}\Phi_{p}(b\circ\varphi)=(b_{x}\circ\varphi)\varphi^{\prime 1/p}$, so the defining identity yields
$b_{x}=\partial_{t}a_{x}+u\,\partial_{x}a_{x}+\frac{1}{p}a_{x}u_{x}=\partial_{x}(\partial_{t}a+ua_{x})-\frac{p-1}{p}a_{x}u_{x}.$
Imposing the normalization $b(-\infty)=0$ and integrating from -$\infty$ gives
$b=\partial_{t}a+ua_{x}-\frac{p-1}{p}\,\partial_{x}^{-1}(a_{x}u_{x}),$
which is the stated formula.
For the geodesic equation, take $A=\dot{\varphi}$, so $a=u$. Then $\nabla^{(p)}_{\dot{\varphi}}\dot{\varphi}=0$ is equivalent to
$u_{t}+uu_{x}-\frac{p-1}{p}\,\partial_{x}^{-1}(u_{x}^{2})=0.$
Differentiating in $x$ and using $\partial_{x}\partial_{x}^{-1}=\mathrm{Id}$ gives
$\partial_{t}(u_{x})+u\,\partial_{x}(u_{x})+\frac{1}{p}(u_{x})^{2}=0.$
Corollary 2.14 (Generalized Hunter-Saxton equation, third-order form).
The Eulerian velocity $u$ of an $F_{\dot{W}^{1,p}}$-geodesic satisfies
$u_{txx}+uu_{xxx}+\left(1+\frac{2}{p}\right)u_{x}u_{xx}=0.$
For $p=2$, this is the classical Hunter-Saxton equation. For $p=\infty$, it reduces to $u_{txx}+uu_{xxx}+u_{x}u_{xx}=0$.
Proof.
Set $v:=u_{x}$. The strain equation in Proposition 2.13 is
$v_{t}+uv_{x}+\frac{1}{p}v^{2}=0.$
Differentiating $v_{t}+uv_{x}+\frac{1}{p}v^{2}=0$ with respect to $x$ and then substituting $v=u_{x}$ (so that $v_{tx}=u_{txx}$, $v_{xx}=u_{xxx}$, and $v_{x}=u_{xx}$) yields the equation.
Remark 2.15 (Projective deformation).
The flat connection $\nabla^{(p)}$ admits a one-parameter deformation $\nabla^{(p,\alpha)}$ by the projective tensor $\mathcal{C}(u,a)=u_{xx}a-ua_{xx}$. For $\alpha\neq 0$ this introduces torsion and curvature; the leading curvature term involves the Lie bracket through $\partial_{x}^{2},\mathcal{L}_{[u,v]}]$. The detailed analysis of this deformation family will appear in a companion paper.
2.7. Asymptotics as $p\to\infty$
Proposition 2.16.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $f_{\infty}=\log(\varphi^{\prime})\in W^{\infty,1}(\mathbb{R})$,
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+O_{W^{k,1}}\!\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
Set $\rho:=\varphi^{\prime}>0$ and $f_{\infty}:=\log\rho\in W^{\infty,1}(\mathbb{R})$. Then for each $p\geq 1$,
$\Phi_{p}(\varphi)=p\bigl(\rho^{1/p}-1\bigr)=p\bigl(e^{f_{\infty}/p}-1\bigr).$
Fix $k\in\mathbb{N}$. Let $\mathcal{B}\subset\mathrm{Diff}^{-\infty}(\mathbb{R})$ be a bounded set. By definition of the Fréchet structure, boundedness of $\mathcal{B}$ implies that for every $m\leq k+1$ there is a constant $M_{m}(\mathcal{B})$ such that
$\sup_{\varphi\in\mathcal{B}}\|f_{\infty}\|_{W^{m,1}(\mathbb{R})}\leq M_{m}(\mathcal{B}).$
In particular, for each $j\leq k$ the derivative $f_{\infty}^{(j)}$ belongs to $W^{1,1}(\mathbb{R})$ uniformly on $\mathcal{B}$, hence it is bounded uniformly in $L^{\infty}(\mathbb{R})$. Indeed, if $g\in W^{1,1}(\mathbb{R})$ and $g\in L^{1}(\mathbb{R})$, then $g$ has limits $g(\pm\infty)=0$ and for every $x\in\mathbb{R}$,
$|g(x)|=\left|\int_{x}^{\infty}g^{\prime}(t)\,dt\right|\leq\|g^{\prime}\|_{L^{1}(\mathbb{R})},$
so $\|g\|_{L^{\infty}}\leq\|g^{\prime}\|_{L^{1}}\leq\|g\|_{W^{1,1}}$. Applying this with $g=f_{\infty}^{(j)}$ yields uniform bounds
$\sup_{\varphi\in\mathcal{B}}\|f_{\infty}^{(j)}\|_{L^{\infty}(\mathbb{R})}\leq M_{j+1}(\mathcal{B}),\qquad 0\leq j\leq k.$
We now expand the exponential with an integral remainder. For every real $z$,
$e^{z}=1+z+\frac{z^{2}}{2}+\frac{z^{3}}{6}+\frac{z^{4}}{6}\int_{0}^{1}(1-t)^{3}e^{tz}\,dt,$
which follows by applying Taylor’s theorem with integral remainder to the $C^{\infty}$ function $e^{z}$. With $z=f_{\infty}/p$ this gives the pointwise identity
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+R_{p},\qquad R_{p}(x)=\frac{1}{6p^{3}}f_{\infty}(x)^{4}E_{p}(x),$
where
$E_{p}(x):=\int_{0}^{1}(1-t)^{3}e^{tf_{\infty}(x)/p}\,dt.$
Since $0\leq(1-t)^{3}\leq 1$ and $e^{tf_{\infty}/p}\leq e^{\|f_{\infty}\|_{L^{\infty}}/p}\leq e^{\|f_{\infty}\|_{L^{\infty}}}$, we have the uniform bound
$\|E_{p}\|_{L^{\infty}(\mathbb{R})}\leq e^{\|f_{\infty}\|_{L^{\infty}(\mathbb{R})}}.$
In particular, for $m=0$,
$\|R_{p}\|_{L^{1}(\mathbb{R})}\leq\frac{1}{6p^{3}}\|E_{p}\|_{L^{\infty}}\|f_{\infty}^{4}\|_{L^{1}}\leq\frac{C(\mathcal{B})}{p^{3}},$
because $f_{\infty}\in L^{\infty}\cap L^{1}$ uniformly on $\mathcal{B}$, hence $\|f_{\infty}^{4}\|_{L^{1}}\leq\|f_{\infty}\|_{L^{\infty}}^{3}\|f_{\infty}\|_{L^{1}}$ is uniformly bounded on $\mathcal{B}$.
To control $\partial_{x}^{m}R_{p}$ in $L^{1}$ for $1\leq m\leq k$, we estimate $\partial_{x}^{m}(f_{\infty}^{4}E_{p})$ and then divide by $p^{3}$. By Leibniz,
$\partial_{x}^{m}(f_{\infty}^{4}E_{p})=\sum_{a=0}^{m}\binom{m}{a}\partial_{x}^{a}(f_{\infty}^{4})\,\partial_{x}^{m-a}E_{p}.$
The derivatives $\partial_{x}^{a}(f_{\infty}^{4})$ are finite sums of products of $f_{\infty}^{(j)}$ with total differentiation order $a$ and with at most four factors. For each such product, one factor can be placed in $L^{1}$ (choose a factor carrying at least one derivative when $a\geq 1$; when $a=0$ use $f_{\infty}\in L^{1}$), and all remaining factors are placed in $L^{\infty}$. Using the uniform $L^{\infty}$ bounds for $f_{\infty}^{(j)}$ recorded above, it follows that for each $a\leq k$,
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{a}(f_{\infty}^{4})\|_{L^{1}(\mathbb{R})}\leq C_{a}(\mathcal{B}).$
It remains to bound $\partial_{x}^{r}E_{p}$ in $L^{\infty}$ for $0\leq r\leq k$. Write $E_{p}=F_{p}\circ f_{\infty}$, where
$F_{p}(s):=\int_{0}^{1}(1-t)^{3}e^{ts/p}\,dt.$
Then $F_{p}^{(n)}(s)=\int_{0}^{1}(1-t)^{3}(t/p)^{n}e^{ts/p}\,dt$, hence for every $n\geq 0$ and $s\in\mathbb{R}$,
$|F_{p}^{(n)}(s)|\leq\frac{1}{p^{n}}e^{|s|/p}\leq e^{|s|},$
using $p\geq 1$. Applying the Faà di Bruno formula to the composition $F_{p}\circ f_{\infty}$ yields, for each $r\geq 1$,
$\partial_{x}^{r}E_{p}(x)=\sum_{n=1}^{r}F_{p}^{(n)}(f_{\infty}(x))\sum_{\begin{subarray}{c}\alpha_{1}+\cdots+\alpha_{r}=n\\ \alpha_{1}+2\alpha_{2}+\cdots+r\alpha_{r}=r\end{subarray}}c_{\alpha}\prod_{j=1}^{r}\bigl(f_{\infty}^{(j)}(x)\bigr)^{\alpha_{j}},$
with combinatorial coefficients $c_{\alpha}$ depending only on $r$. Each product on the right involves derivatives $f_{\infty}^{(j)}$ with $1\leq j\leq r$, hence every factor is bounded in $L^{\infty}$ uniformly on $\mathcal{B}$. Since $F_{p}^{(n)}(f_{\infty})$ is also bounded in $L^{\infty}$ uniformly on $\mathcal{B}$ by the previous inequality and the uniform $L^{\infty}$ bound for $f_{\infty}$, we obtain
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{r}E_{p}\|_{L^{\infty}(\mathbb{R})}\leq C_{r}(\mathcal{B}),\qquad 0\leq r\leq k,$
where $r=0$ is the earlier bound on $\|E_{p}\|_{L^{\infty}}$.
Combining these bounds with the Leibniz expansion gives, for each $m\leq k$,
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m}(f_{\infty}^{4}E_{p})\|_{L^{1}(\mathbb{R})}\leq\sum_{a=0}^{m}\binom{m}{a}\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{a}(f_{\infty}^{4})\|_{L^{1}}\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m-a}E_{p}\|_{L^{\infty}}\leq C_{m}(\mathcal{B}).$
Therefore
$\sup_{\varphi\in\mathcal{B}}\|\partial_{x}^{m}R_{p}\|_{L^{1}(\mathbb{R})}\leq\frac{1}{6p^{3}}C_{m}(\mathcal{B}),\qquad 0\leq m\leq k,$
and hence
$\sup_{\varphi\in\mathcal{B}}\|R_{p}\|_{W^{k,1}(\mathbb{R})}\leq\frac{C_{k}(\mathcal{B})}{p^{3}}.$
Therefore
$\Phi_{p}(\varphi)=f_{\infty}+\frac{1}{2p}f_{\infty}^{2}+\frac{1}{6p^{2}}f_{\infty}^{3}+O_{W^{k,1}}\!\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly for $\varphi$ ranging in bounded subsets of $\mathrm{Diff}^{-\infty}(\mathbb{R})$.
3. The Schwarzian derivative: projective curvature and $L^{p}$ deformation
3.1. Definition and basic properties
Definition 3.1.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the Schwarzian derivative is
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}.$
Proposition 3.2.
Let $f=\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$. Then
$S(\varphi)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}.$
The map $S\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth with differential
$D_{\varphi}S(\delta\varphi)=\delta f^{\prime\prime}-f^{\prime}\delta f^{\prime},\qquad\delta f=\frac{\delta\varphi^{\prime}}{\varphi^{\prime}}.$
Proof.
From $f=\log(\varphi^{\prime})$ we have $f^{\prime}=\varphi^{\prime\prime}/\varphi^{\prime}$ and $f^{\prime\prime}=\varphi^{\prime\prime\prime}/\varphi^{\prime}-(\varphi^{\prime\prime}/\varphi^{\prime})^{2}$, which implies
$S(\varphi)=f^{\prime\prime}+(f^{\prime})^{2}-\frac{3}{2}(f^{\prime})^{2}=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}.$
Smoothness follows from $\varphi\mapsto f$ being smooth and $W^{\infty,1}$ being closed under differentiation and multiplication. The differential is computed by varying $f$. ∎
Proposition 3.3 (Cocycle identity).
For $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$.
$S(\varphi\circ\psi)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi).$
In particular, $S(\varphi)=0$ if and only if $\varphi\in\operatorname{Aff}(\mathbb{R})$.
Proof.
Direct computation using the chain rule. The kernel statement follows since $S(ax+b)=0$ for $a>0$, $b\in\mathbb{R}$, and if $S(\varphi)=0$ then $f=\log(\varphi^{\prime})$ satisfies $f^{\prime\prime}=\frac{1}{2}(f^{\prime})^{2}$. The general solution is $f^{\prime}=-2/(x-x_{0})$ or $f^{\prime}\equiv 0$. Only $f^{\prime}\equiv 0$ gives $\varphi^{\prime}-1\in W^{1,1}(\mathbb{R})$, hence $\varphi^{\prime}=c$ constant, which with $\varphi^{\prime}(\pm\infty)=1$ forces $\varphi^{\prime}=1$, so $\varphi(x)=x+b$. ∎
Remark 3.4.
The additive cocycle $\Phi_{\infty}(\varphi\circ\psi)=(\Phi_{\infty}(\varphi)\circ\psi)+\Phi_{\infty}(\psi)$ is the linear shadow of the projective cocycle for $S$. The operator $f\mapsto f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ is the one-dimensional projective curvature.
3.2. Score curvature and Fisher-Rao identities
The potential form of Proposition 3.2 extends verbatim from Jacobians $\varphi^{\prime}$ to arbitrary positive densities. This recasts the Schwarzian as an intrinsic “score curvature” and connects it to Fisher information via identities that are invisible in the purely group-theoretic formulation.
Definition 3.5 (Score, Fisher information, and density Schwarzian).
Let $\mu=g\,dx$ be a probability measure on $\mathbb{R}$ with $g>0$, $\log g\in W^{\infty,1}(\mathbb{R})$, and sufficient decay so that all integrations by parts below are legitimate. Its score is
$s_{\mu}:=(\log g)^{\prime}\in W^{\infty,1}(\mathbb{R}).$
The (location) Fisher information is
$I(\mu):=\int_{\mathbb{R}}s_{\mu}(x)^{2}\,g(x)\,dx=\|s_{\mu}\|_{L^{2}(\mu)}^{2}.$
The Schwarzian curvature of $\mu$ is the projective curvature of its score,
$S(\mu):=s_{\mu}^{\prime}-\frac{1}{2}s_{\mu}^{2}\in W^{\infty,1}(\mathbb{R}).$
Remark 3.6.
If $g=\varphi^{\prime}$ for some $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, then $S(\mu)=S(\varphi)$ by Proposition 3.2 applied to $f=\log g$. Thus $S(\mu)$ is the density-side incarnation of the group Schwarzian.
Proposition 3.7 (Square-root representation).
Let $\rho:=\sqrt{g}$. Then
$I(\mu)=4\int_{\mathbb{R}}|\rho^{\prime}(x)|^{2}\,dx.$
Proof.
Since $\rho^{\prime}/\rho=\frac{1}{2}(\log g)^{\prime}=\frac{1}{2}s_{\mu}$, we have $s_{\mu}=2\rho^{\prime}/\rho$ and hence
$I(\mu)=\int_{\mathbb{R}}s_{\mu}^{2}g\,dx=\int_{\mathbb{R}}\frac{4(\rho^{\prime})^{2}}{\rho^{2}}\,\rho^{2}\,dx=4\int_{\mathbb{R}}(\rho^{\prime})^{2}\,dx.$
Proposition 3.8 (Mean projective curvature and Fisher information).
Assume additionally that $g^{\prime}\in L^{1}(\mathbb{R})$ and $g^{\prime}(x)\to 0$ as $x\to\pm\infty$. Then
$\int_{\mathbb{R}}S(\mu)(x)\,g(x)\,dx\;=\;-\frac{3}{2}\,I(\mu).$ (1)
Proof.
Write $s=s_{\mu}$ and note $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. Using $(\log g)^{\prime\prime}=g^{\prime\prime}/g-(g^{\prime}/g)^{2}$, we obtain $s^{\prime}g=g^{\prime\prime}-(g^{\prime})^{2}/g$ and therefore
$\int_{\mathbb{R}}s^{\prime}g\,dx=\int_{\mathbb{R}}g^{\prime\prime}\,dx-\int_{\mathbb{R}}\frac{(g^{\prime})^{2}}{g}\,dx=-\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu),$
since $\int g^{\prime\prime}=g^{\prime}(+\infty)-g^{\prime}(-\infty)=0$ by hypothesis and $\frac{(g^{\prime})^{2}}{g}=s^{2}g$. Consequently,
$\int_{\mathbb{R}}S(\mu)\,g\,dx=\int_{\mathbb{R}}s^{\prime}g\,dx-\frac{1}{2}\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu)-\frac{1}{2}I(\mu)=-\frac{3}{2}I(\mu).$
Corollary 3.9.
Under the hypotheses of Proposition 3.8,
$\|S(\mu)\|_{L^{2}(\mu)}\geq\frac{3}{2}\,I(\mu).$
Proof.
By Cauchy-Schwarz and (1),
$\frac{3}{2}I(\mu)=\left|\int_{\mathbb{R}}S(\mu)\,d\mu\right|\leq\|S(\mu)\|_{L^{2}(\mu)}\cdot\|1\|_{L^{2}(\mu)}=\|S(\mu)\|_{L^{2}(\mu)}.$
Theorem 3.10 (Second-order score constraint and Bhattacharyya-type bound).
[…]
$u_{1}:=\partial_{\theta}\log g(X-\theta)\big|_{\theta=0}=-s_{\mu}(X),\qquad v_{2}:=\partial_{\theta}^{2}\log g(X-\theta)\big|_{\theta=0}+u_{1}(X)^{2}=s_{\mu}^{\prime}(X)+s_{\mu}(X)^{2}.$
Then the unbiasedness conditions imply the linear constraints in $L^{2}(\mu)$:
$\langle T,u_{1}\rangle_{\mu}=1,\qquad\langle T,v_{2}\rangle_{\mu}=0,$ (2)
and the variance satisfies
$\mathrm{Var}_{0}(\hat{\theta})=\|T\|_{L^{2}(\mu)}^{2}\;\geq\;\frac{\|v_{2}\|_{L^{2}(\mu)}^{2}}{\|u_{1}\|_{L^{2}(\mu)}^{2}\,\|v_{2}\|_{L^{2}(\mu)}^{2}-\langle u_{1},v_{2}\rangle_{\mu}^{2}}.$ (3)
Moreover,
$v_{2}\;=\;S(\mu)+\frac{3}{2}\,s_{\mu}^{2},$ (4)
so the only projectively invariant contribution to the second score composite is $S(\mu)$.
Proof.
The first constraint in (2) is the standard Cramér-Rao identity. Indeed $E_{\theta}[T]=0$, hence $0=\partial_{\theta}E_{\theta}[T]=E_{\theta}[\partial_{\theta}T]+E_{\theta}[T\,\partial_{\theta}\log p_{\theta}]$, and since $\partial_{\theta}T=-1$ and $\partial_{\theta}\log p_{\theta}=u_{1}$ this gives $E_{\theta}[Tu_{1}]=1$ at $\theta=0$.
For the second constraint, differentiate once more: $0=\partial_{\theta}^{2}E_{\theta}[T]=\partial_{\theta}E_{\theta}[Tu_{1}]$. Using $\partial_{\theta}E_{\theta}[F]=E_{\theta}[Fu_{1}]$ for $\theta$-independent $F$ and the identity $\partial_{\theta}u_{1}=\partial_{\theta}^{2}\log p_{\theta}=u_{2}$, one obtains
$0=\partial_{\theta}E_{\theta}[Tu_{1}]=E_{\theta}[Tu_{2}]+E_{\theta}[Tu_{1}^{2}]+E_{\theta}[(\partial_{\theta}T)u_{1}].$
Since $\partial_{\theta}T=-1$ and $E_{\theta}[u_{1}]=0$, the last term vanishes and we obtain $E_{0}[T(u_{2}+u_{1}^{2})]=0$, i.e. $\langle T,v_{2}\rangle_{\mu}=0$.
Now consider the Hilbert space $H=L^{2}(\mu)$ with inner product $\langle\cdot,\cdot\rangle_{\mu}$. Among all $T\in H$ satisfying (2), the minimizer of $\|T\|_{H}$ is the unique element $T_{0}\in\mathrm{span}\{u_{1},v_{2}\}$ satisfying the same constraints. Writing $T_{0}=\alpha u_{1}+\beta v_{2}$ and imposing (2) yields the $2\times 2$ linear system
$\alpha\|u_{1}\|^{2}+\beta\langle v_{2},u_{1}\rangle=1,\qquad\alpha\langle u_{1},v_{2}\rangle+\beta\|v_{2}\|^{2}=0,$
whose determinant is $\|u_{1}\|^{2}\|v_{2}\|^{2}-\langle u_{1},v_{2}\rangle^{2}>0$. A direct elimination gives
$\|T_{0}\|^{2}=\frac{\|v_{2}\|^{2}}{\|u_{1}\|^{2}\|v_{2}\|^{2}-\langle u_{1},v_{2}\rangle^{2}}.$
For any admissible $T$, the difference $T-T_{0}$ is orthogonal to $\mathrm{span}\{u_{1},v_{2}\}$ by the normal equations for constrained minimization, hence by Pythagoras $\|T\|^{2}=\|T_{0}\|^{2}+\|T-T_{0}\|^{2}\geq\|T_{0}\|^{2}$, which is exactly (3).
Finally, (4) is the algebraic identity $s^{\prime}+s^{2}=(s^{\prime}-\frac{1}{2}s^{2})+\frac{3}{2}s^{2}$ with $S(\mu)=s^{\prime}-\frac{1}{2}s^{2}$. ∎
Remark 3.11.
The numerical lower bound (3) coincides with the classical second-order Bhattacharyya bound for the location model. The point of the present formulation is the projective decomposition (4): it isolates $S(\mu)$ as the projectively invariant component inside the second-order score composite.
3.3. Projective connections and cocycle lifting
Definition 3.12 (Quadratic differential pullback).
Let $\mathcal{Q}^{\infty,1}(\mathbb{R}):=W^{\infty,1}(\mathbb{R})$, viewed as coefficients of quadratic differentials $q(x)\,dx^{2}$. For $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ define the pullback
$\psi^{*}q:=(q\circ\psi)(\psi^{\prime})^{2}.$
Then pullback is contravariant:
$(\psi_{1}\circ\psi_{2})^{*}=\psi_{2}^{*}\circ\psi_{1}^{*},$
so $q\mapsto q\cdot\psi:=\psi^{*}q$ is a right action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $\mathcal{Q}^{\infty,1}(\mathbb{R})$.
Proposition 3.13 (Affine action on projective connections).
Define, for $q\in\mathcal{Q}^{\infty,1}(\mathbb{R})$ and $\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$,
$q\star\psi:=\psi^{*}q+S(\psi).$
Then $\star$ is a right action:
$(q\star\psi_{1})\star\psi_{2}=q\star(\psi_{1}\circ\psi_{2}).$
Proof.
By contravariance of pullback and Proposition 3.3,
$(q\star\psi_{1})\star\psi_{2}$ $=\psi_{2}^{*}(\psi_{1}^{*}q+S(\psi_{1}))+S(\psi_{2})$
$=(\psi_{1}\circ\psi_{2})^{*}q+(S(\psi_{1})\circ\psi_{2})(\psi_{2}^{\prime})^{2}+S(\psi_{2})$
$=(\psi_{1}\circ\psi_{2})^{*}q+S(\psi_{1}\circ\psi_{2})=q\star(\psi_{1}\circ\psi_{2}).$
Definition 3.14 (Schwarzian bundle over the affine quotient).
Let $\operatorname{Aff}(\mathbb{R})\subset\operatorname{Diff}^{-\infty}(\mathbb{R})$ be the affine subgroup (in the decay class, only translations survive, but we keep the notation). Define the associated bundle
$\mathcal{Q}_{\mathrm{proj}}:=\operatorname{Diff}^{-\infty}(\mathbb{R})\times_{\operatorname{Aff}(\mathbb{R})}\mathcal{Q}^{\infty,1}(\mathbb{R}),$
where $(\varphi,q)\sim(\varphi\circ A,A^{*}q)$ for $A\in\operatorname{Aff}(\mathbb{R})$, with $A^{*}q=(q\circ A)(A^{\prime})^{2}$. We write elements as $\varphi,q]$ and the base point as $\varphi]\in\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
Theorem 3.15 (Cocycle lift as canonical section).
The Schwarzian derivative defines a well-defined section
$\sigma_{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\to\mathcal{Q}_{\mathrm{proj}},\qquad\sigma_{S}([\varphi])=[\varphi,S(\varphi)].$
Moreover, the left action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on the base by left composition,
$\psi\cdot[\varphi]=[\psi\circ\varphi],$
lifts to an action on $\mathcal{Q}_{\mathrm{proj}}$ given by
$\psi\cdot[\varphi,q]:=[\psi\circ\varphi,\;q+\varphi^{*}S(\psi)],$ (5)
and $\sigma_{S}$ is equivariant:
$\psi\cdot\sigma_{S}([\varphi])=\sigma_{S}(\psi\cdot[\varphi]).$
Proof.
Well-definedness of $\sigma_{S}$. If $\varphi$ is replaced by $\varphi\circ A$ with $A\in\operatorname{Aff}(\mathbb{R})$, then by Proposition 3.3 and $S(A)=0$,
$S(\varphi\circ A)=(S(\varphi)\circ A)(A^{\prime})^{2}=A^{*}S(\varphi),$
so $\varphi\circ A,S(\varphi\circ A)]=[\varphi,A^{*}S(\varphi)]=[\varphi,S(\varphi)]$ in $\mathcal{Q}_{\mathrm{proj}}$.
Well-definedness of the lifted action. If $\varphi,q]=[\varphi\circ A,A^{*}q]$, then using contravariance of pullback,
$(\varphi\circ A)^{*}S(\psi)=A^{*}(\varphi^{*}S(\psi)),$
hence
$\psi\cdot[\varphi\circ A,A^{*}q]=[\psi\circ\varphi\circ A,\;A^{*}q+(\varphi\circ A)^{*}S(\psi)]=[\psi\circ\varphi\circ A,\;A^{*}(q+\varphi^{*}S(\psi))]=\psi\cdot[\varphi,q],$
so (5) descends to the quotient.
Equivariance. By Proposition 3.3,
$S(\psi\circ\varphi)=\varphi^{*}S(\psi)+S(\varphi),$
therefore
$\psi\cdot\sigma_{S}([\varphi])=\psi\cdot[\varphi,S(\varphi)]=[\psi\circ\varphi,\;S(\varphi)+\varphi^{*}S(\psi)]=[\psi\circ\varphi,\;S(\psi\circ\varphi)]=\sigma_{S}([\psi\circ\varphi]).$
Proposition 3.16 (Infinitesimal Schwarzian cocycle).
Let $u$ be a vector field on $\mathbb{R}$ such that $u^{\prime},u^{\prime\prime},u^{\prime\prime\prime}\in L^{1}(\mathbb{R})$ and $u^{\prime},u^{\prime\prime}\to 0$ at $\pm\infty$, and let $\varphi_{t}$ be the corresponding local flow with $\varphi_{0}=\operatorname{id}$. Then
$S(\varphi_{t})=t\,u^{\prime\prime\prime}+O_{W^{k,1}}(t^{2})\qquad(t\to 0),$
for each $k\geq 0$. In particular, $D_{\operatorname{id}}S(u)=u^{\prime\prime\prime}$.
Proof.
Write $\varphi_{t}(x)=x+tu(x)+O(t^{2})$ in $W^{k,1}$ for every $k$. Then $\varphi_{t}^{\prime}=1+tu^{\prime}+O(t^{2})$, $\varphi_{t}^{\prime\prime}=tu^{\prime\prime}+O(t^{2})$, $\varphi_{t}^{\prime\prime\prime}=tu^{\prime\prime\prime}+O(t^{2})$. Substituting into Definition 3.1 and expanding to first order yields $S(\varphi_{t})=tu^{\prime\prime\prime}+O(t^{2})$ in $W^{k,1}$. ∎
Definition 3.17 (Gelfand-Fuchs cocycle).
On the Lie algebra $\mathfrak{X}^{-\infty}(\mathbb{R})$ of such vector fields, define
$\omega_{\mathrm{GF}}(u,v):=\frac{1}{2}\int_{\mathbb{R}}\bigl(u^{\prime}v^{\prime\prime}-u^{\prime\prime}v^{\prime}\bigr)\,dx.$
Proposition 3.18 (Normalization and cocycle property).
For $u,v\in\mathfrak{X}^{-\infty}(\mathbb{R})$,
$\omega_{\mathrm{GF}}(u,v)=\int_{\mathbb{R}}u^{\prime\prime\prime}(x)\,v(x)\,dx=-\int_{\mathbb{R}}u(x)\,v^{\prime\prime\prime}(x)\,dx.$ (6)
In particular,
$\omega_{\mathrm{GF}}(u,v)=\int_{\mathbb{R}}\bigl(D_{\operatorname{id}}S(u)\bigr)(x)\,v(x)\,dx.$ (7)
Moreover, $\omega_{\mathrm{GF}}$ is a Lie algebra $2$-cocycle:
$\omega_{\mathrm{GF}}([u,v],w)+\omega_{\mathrm{GF}}([v,w],u)+\omega_{\mathrm{GF}}([w,u],v)=0.$
Proof.
Since $u^{\prime},v^{\prime}\to 0$ at $\pm\infty$ and $u^{\prime\prime},v^{\prime\prime}\in L^{1}$, integration by parts gives
$\int_{\mathbb{R}}u^{\prime\prime}v^{\prime}\,dx=-\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx,$
so
$\omega_{\mathrm{GF}}(u,v)=\frac{1}{2}\int_{\mathbb{R}}(u^{\prime}v^{\prime\prime}-u^{\prime\prime}v^{\prime})\,dx=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx.$
A second integration by parts yields $\int u^{\prime}v^{\prime\prime}=\int u^{\prime\prime\prime}v=-\int uv^{\prime\prime\prime}$, proving (6). Then (7) follows from Proposition 3.16.
For the cocycle identity, write $\omega_{\mathrm{GF}}(a,b)=\int a\,b^{\prime\prime\prime}$ using (6) and the bracket $u,v]=uv^{\prime}-vu^{\prime}$. The cyclic sum of the integrands is
$(uv^{\prime}-vu^{\prime})w^{\prime\prime\prime}+(vw^{\prime}-wv^{\prime})u^{\prime\prime\prime}+(wu^{\prime}-uw^{\prime})v^{\prime\prime\prime}=\partial_{x}\Bigl(uv^{\prime}w^{\prime\prime}-uv^{\prime\prime}w^{\prime}+vw^{\prime}u^{\prime\prime}-vw^{\prime\prime}u^{\prime}+wu^{\prime}v^{\prime\prime}-wu^{\prime\prime}v^{\prime}\Bigr),$
an exact derivative. The boundary term vanishes by the decay hypotheses, hence the cyclic integral is zero. ∎
3.4. The $L^{p}$-Schwarzian
We introduce a family interpolating between the affine cocycle structure of $\Phi_{p}$ and the projective cocycle of $S$.
Definition 3.19.
For $p\in[1,\infty)$ and $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, the $L^{p}$-Schwarzian is
$S_{p}(\varphi)=\left(\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}+S(\varphi)\right)\varphi^{\prime 1/p}.$
Proposition 3.20.
The $L^{p}$-Schwarzian satisfies
  • (i)
    $S_{p}(A)=0$ for affine $A$.
  • (ii)
    $S_{p}(\varphi\circ\psi)=(S_{p}(\varphi)\circ\psi)(\psi^{\prime})^{2+1/p}+S_{p}(\psi)(\varphi^{\prime}\circ\psi)^{1/p}+\frac{3}{p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}\,((\varphi\circ\psi)^{\prime})^{1/p}$.
  • (iii)
    $S_{p}(\varphi)(x)\to S(\varphi)(x)$ pointwise as $p\to\infty$.
Proof.
We use the pointwise identity
$S_{p}(\theta)=\left(S(\theta)+\frac{3}{2p}\left(\frac{\theta^{\prime\prime}}{\theta^{\prime}}\right)^{2}\right)(\theta^{\prime})^{1/p},$ (8)
valid for any orientation-preserving $C^{3}$ diffeomorphism $\theta$. For (i), if $A(x)=ax+b$ with $a>0$, then $S(A)=0$ and $A^{\prime\prime}=0$, hence (8) gives $S_{p}(A)=0$. For (ii), set $\Theta=\varphi\circ\psi$. Then $\Theta^{\prime}=(\varphi^{\prime}\circ\psi)\psi^{\prime}$ and
$\frac{\Theta^{\prime\prime}}{\Theta^{\prime}}=\partial_{x}\log\Theta^{\prime}=\partial_{x}\bigl(\log(\varphi^{\prime}\circ\psi)+\log\psi^{\prime}\bigr)=\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime}+\frac{\psi^{\prime\prime}}{\psi^{\prime}}.$
Squaring yields
$\left(\frac{\Theta^{\prime\prime}}{\Theta^{\prime}}\right)^{2}=\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)^{2}(\psi^{\prime})^{2}+2\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}+\left(\frac{\psi^{\prime\prime}}{\psi^{\prime}}\right)^{2}.$
The classical Schwarzian satisfies $S(\Theta)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi)$. Inserting these identities into (8) for $\theta=\Theta$ and using $\Theta^{\prime 1/p}=(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}$, we obtain
$S_{p}(\Theta)$ $=\left(\left(S(\varphi)+\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}\right)\circ\psi\right)(\psi^{\prime})^{2+1/p}(\varphi^{\prime}\circ\psi)^{1/p}$
$\quad+\left(S(\psi)+\frac{3}{2p}\left(\frac{\psi^{\prime\prime}}{\psi^{\prime}}\right)^{2}\right)(\psi^{\prime})^{1/p}(\varphi^{\prime}\circ\psi)^{1/p}$
$\quad+\frac{3}{p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\circ\psi\right)\psi^{\prime\prime}\,(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}.$
The first two terms are exactly $(S_{p}(\varphi)\circ\psi)(\psi^{\prime})^{2+1/p}$ and $S_{p}(\psi)(\varphi^{\prime}\circ\psi)^{1/p}$ by (8), and the last term equals the stated formula since $(\varphi^{\prime}\circ\psi)^{1/p}(\psi^{\prime})^{1/p}=\Theta^{\prime 1/p}$. For (iii), (8) gives
$S_{p}(\varphi)=S(\varphi)\,(\varphi^{\prime})^{1/p}+\frac{3}{2p}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}(\varphi^{\prime})^{1/p},$
and pointwise $(\varphi^{\prime})^{1/p}\to 1$ while $\frac{1}{p}(\varphi^{\prime\prime}/\varphi^{\prime})^{2}(\varphi^{\prime})^{1/p}\to 0$, hence $S_{p}(\varphi)\to S(\varphi)$ pointwise. ∎
Remark 3.21 (Cocycle defect at finite $p$).
The third term in Proposition 3.20(ii) shows that $S_{p}$ is not a $1$-cocycle for finite $p$. It is precisely the obstruction to functoriality of the nonlinear curvature map $\varphi\mapsto S_{p}(\varphi)$ under composition. Two features justify $S_{p}$ as the correct interpolation between affine and projective geometry: (i) the defect term vanishes in the classical limit $p\to\infty$, recovering the Schwarzian cocycle; (ii) in $p$-root coordinates (Proposition 3.22) the expression for $S_{p}$ becomes a clean second-order curvature functional.
Proposition 3.22.
In $p$-root coordinates $\theta=\varphi^{\prime 1/p}$.
$S_{p}(\varphi)=p\theta^{\prime\prime}-\frac{p(p-1)}{2}\frac{(\theta^{\prime})^{2}}{\theta}.$
Equivalently, with $f=\Phi_{p}(\varphi)=p(\theta-1)$.
$S_{p}(\varphi)=f^{\prime\prime}-\frac{p-1}{2p}\frac{(f^{\prime})^{2}}{1+f/p}.$
Proof.
Let $u=\log(\varphi^{\prime})=p\log\theta$. Then $u^{\prime}=p\theta^{\prime}/\theta$ and $u^{\prime\prime}=p\theta^{\prime\prime}/\theta-p(\theta^{\prime})^{2}/\theta^{2}$. Substituting into the definition and simplifying yields both formulas. ∎
Corollary 3.23.
The map $S_{p}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth. On bounded subsets, it is locally Lipschitz as a map into $W^{k,1}(\mathbb{R})$.
3.5. Asymptotics of $S_{p}$ as $p\to\infty$
Proposition 3.24.
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $u=\log(\varphi^{\prime})$ and $S=S(\varphi)$.
$S_{p}(\varphi)=S+\frac{1}{p}\left(uS+\frac{3}{2}(u^{\prime})^{2}\right)+\frac{1}{p^{2}}\left(\frac{u^{2}S}{2}+\frac{3u(u^{\prime})^{2}}{2}\right)+O_{W^{k,1}}\left(\frac{1}{p^{3}}\right)$
as $p\to\infty$, uniformly on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
From Definition 3.19, $S_{p}(\varphi)=(S+\frac{3}{2p}(u^{\prime})^{2})\mathrm{e}^{u/p}$. Expanding $\mathrm{e}^{u/p}=1+u/p+u^{2}/(2p^{2})+O(p^{-3})$ and collecting terms yields the expansion. The $W^{k,1}$ remainder estimate uses Faà di Bruno and Leibniz as in Proposition 2.16. ∎
3.6. Sturm-Liouville interpretation
The Schwarzian governs a natural Sturm-Liouville problem.
Proposition 3.25.
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $S=S(\varphi)$. Define $y_{2}=(\varphi^{\prime})^{-1/2}$ and $y_{1}=\varphi(\varphi^{\prime})^{-1/2}$. Then $y_{1},y_{2}$ are linearly independent solutions of
$y^{\prime\prime}+\frac{1}{2}Sy=0$
with Wronskian $W(y_{1},y_{2})\equiv 1$. Conversely, if $y_{1},y_{2}$ solve $y^{\prime\prime}+qy=0$ with $y_{2}>0$ and $W(y_{1},y_{2})\equiv 1$, then $\varphi=y_{1}/y_{2}$ satisfies $\varphi^{\prime}>0$ and $S(\varphi)=2q$.
Proof.
Let $u=\log(\varphi^{\prime})$. For $y_{2}=\mathrm{e}^{-u/2}$. $y_{2}^{\prime}=-\frac{1}{2}u^{\prime}y_{2}$ and $y_{2}^{\prime\prime}=(\frac{1}{4}(u^{\prime})^{2}-\frac{1}{2}u^{\prime\prime})y_{2}$. Then
$y_{2}^{\prime\prime}+\frac{S}{2}y_{2}=\left(\frac{(u^{\prime})^{2}}{4}-\frac{u^{\prime\prime}}{2}+\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}\right)y_{2}=0.$
For $y_{1}=\varphi y_{2}$. $y_{1}^{\prime\prime}=\varphi y_{2}^{\prime\prime}$ after simplification, so $y_{1}^{\prime\prime}+\frac{S}{2}y_{1}=0$. The Wronskian is $W(y_{1},y_{2})=y_{1}^{\prime}y_{2}-y_{1}y_{2}^{\prime}=\varphi^{\prime}y_{2}^{2}=\varphi^{\prime}(\varphi^{\prime})^{-1}=1$. Conversely, $\varphi^{\prime}=W(y_{1},y_{2})/y_{2}^{2}=1/y_{2}^{2}>0$. Computing $S(\varphi)$ from $\log(\varphi^{\prime})=-2\log y_{2}$ gives $S(\varphi)=2q$. ∎
Theorem 3.26 (Liouville conjugation).
Define $(\mathcal{L}_{\varphi}u)(x)=(\varphi^{\prime}(x))^{-1/2}u(\varphi(x))$. For any $q\in C^{\infty}(\mathbb{R})$.
$\mathcal{L}_{\varphi}\circ(\partial_{y}^{2}+q(y))\circ\mathcal{L}_{\varphi}^{-1}=\partial_{x}^{2}+q(\varphi(x))\varphi^{\prime}(x)^{2}+\frac{1}{2}S(\varphi)(x).$
Proof.
Set $v=\mathcal{L}_{\varphi}^{-1}u$, so $v(\varphi(x))=\varphi^{\prime}(x)^{1/2}u(x)$. Differentiate twice and simplify using $\varphi^{\prime\prime}=\varphi^{\prime}(\log\varphi^{\prime})^{\prime}$. ∎
Remark 3.27.
The Schwarzian is the unique scalar correction ensuring functoriality of Liouville conjugation for second-order operators. In semiclassical analysis, it is the subprincipal correction when conjugating $h^{2}\partial_{x}^{2}$ by the Fourier integral operator implementing $\varphi$.
4. Central Extensions and Cocycle Geometry
We now develop the cohomological content of the Schwarzian obstruction through Lie algebra cohomology and central extensions.
The main results of this section are the nontriviality of the Gelfand-Fuchs class (Proposition 4.2), its integration to the Virasoro-Bott group extension via the Bott cocycle (Proposition 4.5), and the rigidity of $L^{p}$ cocycles (Theorem 4.7).
4.1. The Gelfand-Fuchs cocycle
Definition 4.1 (Gelfand-Fuchs $2$-cocycle).
On the Lie algebra $\mathfrak{g}^{-\infty}$, define the bilinear form
$\omega(u,v):=\int_{\mathbb{R}}u^{\prime}(x)\,v^{\prime\prime}(x)\,\mathrm{d}x.$
Proposition 4.2 (Nontriviality of the Gelfand-Fuchs class).
The form $\omega$ is an alternating Lie algebra $2$-cocycle on $\mathfrak{g}^{-\infty}$; we refer to the verification of antisymmetry and the cocycle identity in §4 (projective cocycle lifting). Moreover, $\omega$ is not a coboundary: there is no continuous linear functional $\Lambda\colon\mathfrak{g}^{-\infty}\to\mathbb{R}$ such that $\omega(u,v)=\Lambda([u,v])$ for all $u,v$. Hence $\omega]\neq 0$ in $H^{2}(\mathfrak{g}^{-\infty};\mathbb{R})$.
Proof.
Assume for contradiction that $\omega$ is a coboundary, that is, there exists a continuous linear functional $\Lambda:\mathfrak{g}^{-\infty}\to\mathbb{R}$ such that
$\omega(u,v)=\Lambda([u,v])\qquad\text{for all }u,v\in\mathfrak{g}^{-\infty}.$
Restricting to $C_{c}^{\infty}(\mathbb{R})\subset\mathfrak{g}^{-\infty}$, fix arbitrary $v\in C_{c}^{\infty}(\mathbb{R})$ and choose $u\in C_{c}^{\infty}(\mathbb{R})$ with $u\equiv 1$ on a neighborhood of $\mathrm{supp}(v)$. Then $u^{\prime}=0$ on $\mathrm{spt}(v)$, hence $u^{\prime}v^{\prime\prime}\equiv 0$ and $\omega(u,v)=0$. On the other hand, on $\mathrm{supp}(v)$ one has $u,v]=uv^{\prime}-u^{\prime}v=v^{\prime}$, so
$0=\omega(u,v)=\Lambda([u,v])=\Lambda(v^{\prime}).$
Therefore,
$\Lambda(\partial_{x}f)=0\qquad\text{for all }f\in C_{c}^{\infty}(\mathbb{R}).$ (9)
If $g\in C_{c}^{\infty}(\mathbb{R})$ satisfies $\int_{\mathbb{R}}g\,dx=0$, then $g=f^{\prime}$ for $f(x):=\int_{-\infty}^{x}g(t)\,dt\in C_{c}^{\infty}(\mathbb{R})$, so (9) gives $\Lambda(g)=0$. Hence $\Lambda$ vanishes on the codimension-one subspace $\{g\in C_{c}^{\infty}(\mathbb{R}):\int g=0\}$, and there exists $c\in\mathbb{R}$ with
$\Lambda(g)=c\int_{\mathbb{R}}g(x)\,dx\qquad\text{for all }g\in C_{c}^{\infty}(\mathbb{R}).$ (10)
Consequently, for $u,v\in C_{c}^{\infty}(\mathbb{R})$,
$\omega(u,v)=\Lambda([u,v])=c\int_{\mathbb{R}}(uv^{\prime}-u^{\prime}v)\,dx=-2c\int_{\mathbb{R}}u^{\prime}v\,dx,$
where $\int uv^{\prime}=-\int u^{\prime}v$ by compact support.
On the other hand, integrating by parts twice gives, for $u,v\in C_{c}^{\infty}(\mathbb{R})$,
$\omega(u,v)=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx=\int_{\mathbb{R}}u^{\prime\prime\prime}v\,dx.$
Combining the two identities yields
$\int_{\mathbb{R}}(u^{\prime\prime\prime}+2c\,u^{\prime})\,v\,dx=0\qquad\text{for all }u,v\in C_{c}^{\infty}(\mathbb{R}).$
Fixing $u$ and using the arbitrariness of $v$ implies $u^{\prime\prime\prime}+2c\,u^{\prime}=0$ as a distribution, hence as a smooth identity. Writing $w:=u^{\prime}$, we obtain
$w^{\prime\prime}+2c\,w=0\quad\text{on }\mathbb{R}.$
Every $C^{\infty}$ solution of this ODE is real-analytic, so if $w$ has compact support then $w\equiv 0$. Thus $u^{\prime}\equiv 0$, so $u$ is constant; since $u\in C_{c}^{\infty}(\mathbb{R})$, it follows that $u\equiv 0$. This contradiction shows that no such $\Lambda$ exists. ∎
Remark 4.3.
We will use the already established linearization $D_{\operatorname{id}}S(u)=u^{\prime\prime\prime}$ and the identity $\omega(u,v)=\int_{\mathbb{R}}u^{\prime\prime\prime}(x)v(x)\,dx$ (proved in §4) to identify $\omega$ as the transgression of the infinitesimal Schwarzian.
4.2. The Virasoro-Bott extension
Definition 4.4 (Central extension).
The Virasoro-Bott central extension of $\mathfrak{g}^{-\infty}$ is the Lie algebra
$\widehat{\mathfrak{g}}^{-\infty}=\mathfrak{g}^{-\infty}\oplus\mathbb{R}c$
with bracket
$(u,\alpha),(v,\beta)]=\left([u,v],\omega(u,v)\right).$
The element $c=(0,1)$ is central. $c,\cdot]=0$.
Proposition 4.5 (Integrability).
The central extension $\widehat{\mathfrak{g}}^{-\infty}$ integrates to a central extension of Fréchet Lie groups
$1\to\mathbb{R}\to\widehat{\operatorname{Diff}}^{-\infty}(\mathbb{R})\to\operatorname{Diff}^{-\infty}(\mathbb{R})\to 1$
where the group $2$-cocycle is the Bott cocycle
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,\mathrm{d}\log(\psi^{\prime}).$
Proof.
We first verify that the Bott functional
$B(\varphi,\psi)=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)(x)\,\frac{\psi^{\prime\prime}(x)}{\psi^{\prime}(x)}\,dx$
is well-defined on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ under the standing decay assumptions: by definition $\psi^{\prime}-1$ is integrable together with all derivatives, hence $\log(\psi^{\prime})$ and $(\log\psi^{\prime})^{\prime}=\psi^{\prime\prime}/\psi^{\prime}$ are integrable; similarly $\log(\varphi^{\prime}\circ\psi)$ is bounded with integrable derivatives, so the product is integrable and the integral converges. Normalization is immediate: $B(\operatorname{id},\psi)=0$ and $B(\varphi,\operatorname{id})=0$ since $\log(\operatorname{id}^{\prime})=0$ and $d\log(\operatorname{id}^{\prime})=0$.
Group cocycle identity. For $\varphi,\psi,\chi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$, write the group coboundary
$(\delta B)(\varphi,\psi,\chi):=B(\psi,\chi)-B(\varphi\circ\psi,\chi)+B(\varphi,\psi\circ\chi)-B(\varphi,\psi).$
Using the chain rule and logarithms,
$\log\bigl((\psi\circ\chi)^{\prime}\bigr)=\log(\psi^{\prime}\circ\chi)+\log(\chi^{\prime}),\qquad d\log\bigl((\psi\circ\chi)^{\prime}\bigr)=d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)+d\log(\chi^{\prime}),$
we expand
$B(\varphi,\psi\circ\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log\bigl((\psi\circ\chi)^{\prime}\bigr)$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)+\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime}).$
Similarly, since $(\varphi\circ\psi)^{\prime}\circ\chi=(\varphi^{\prime}\circ\psi\circ\chi)\cdot(\psi^{\prime}\circ\chi)$,
$\log\bigl((\varphi\circ\psi)^{\prime}\circ\chi\bigr)=\log(\varphi^{\prime}\circ\psi\circ\chi)+\log(\psi^{\prime}\circ\chi),$
hence
$B(\varphi\circ\psi,\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log\bigl((\varphi\circ\psi)^{\prime}\circ\chi\bigr)\,d\log(\chi^{\prime})$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime})+\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime}).$
Substituting these two expressions into $(\delta B)(\varphi,\psi,\chi)$, the terms $\frac{1}{2}\int\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\log(\chi^{\prime})$ cancel, and we obtain
$(\delta B)(\varphi,\psi,\chi)$ $=\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime})-\frac{1}{2}\int_{\mathbb{R}}\log(\psi^{\prime}\circ\chi)\,d\log(\chi^{\prime})$
$\qquad+\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)-\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime})$
$=\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)-\frac{1}{2}\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}).$
Now use the pullback/change-of-variables identity for exact one-forms:
$d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)=\chi^{*}(d\log(\psi^{\prime})),\qquad\int_{\mathbb{R}}F\circ\chi\;\chi^{*}\alpha=\int_{\mathbb{R}}F\,\alpha,$
valid for orientation-preserving $\chi$ and integrable forms. With $F=\log(\varphi^{\prime}\circ\psi)$ and $\alpha=d\log(\psi^{\prime})$, we get
$\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi\circ\chi)\,d\bigl(\log(\psi^{\prime}\circ\chi)\bigr)=\int_{\mathbb{R}}\log(\varphi^{\prime}\circ\psi)\,d\log(\psi^{\prime}),$
so $(\delta B)(\varphi,\psi,\chi)=0$. Hence $B$ is a group $2$-cocycle.
Infinitesimal cocycle. Let $\varphi_{t},\psi_{s}$ be the flows of $u,v\in\mathfrak{g}^{-\infty}$ with $\varphi_{0}=\psi_{0}=\operatorname{id}$. Then
$\varphi_{t}^{\prime}(x)=1+t\,u^{\prime}(x)+O(t^{2}),\qquad\psi_{s}^{\prime}(x)=1+s\,v^{\prime}(x)+O(s^{2}),\qquad\frac{\psi_{s}^{\prime\prime}(x)}{\psi_{s}^{\prime}(x)}=s\,v^{\prime\prime}(x)+O(s^{2}).$
Also $\log(\varphi_{t}^{\prime}\circ\psi_{s})=t\,u^{\prime}(x)+O(t^{2})+O(ts^{2})$, so inserting into the defining integral gives
$B(\varphi_{t},\psi_{s})=\frac{1}{2}\int_{\mathbb{R}}\Bigl(tu^{\prime}(x)+o(t)\Bigr)\Bigl(sv^{\prime\prime}(x)+o(s)\Bigr)\,dx=\frac{ts}{2}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx+o(ts).$
Thus the bilinear form obtained by differentiating $B$ at $(\operatorname{id},\operatorname{id})$ is $\frac{1}{2}\,\omega(u,v)$. Equivalently, the cocycle $2B$ differentiates to $\omega$.
The standard construction of a central extension from a normalized group $2$-cocycle therefore yields a central extension of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ whose Lie algebra is the central extension defined by $\omega$ (using the cocycle $2B$) or by $\frac{1}{2}\omega$ (using $B$). ∎
4.3. $L^{p}$ cocycles and rigidity
The quantity $\varphi^{\prime 1/p}$ is the Jacobian cocycle for the natural action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $p$th-root densities. Variants of this Jacobian cocycle, and Bott-Thurston type constructions for general $p$, appear in the literature on $L^{p}$ diffeomorphism groups and density geometry. The point here is that, in the present decay-controlled setting, one can work in explicit $p$-root coordinates and obtain a clean rigidity statement, namely the induced real-valued Lie algebra $2$-cocycle is always the Gelfand-Fuchs cocycle, up to the explicit scalar factor $p^{-2}$. In particular, there is no new real-valued Lie algebra $2$-cocycle coming from the symmetric pairing $\int u^{\prime}v^{\prime}\,\mathrm{d}x$, which is not alternating and therefore does not define a Chevalley-Eilenberg $2$-cochain.
Definition 4.6 ($p$-root Jacobian cocycle and Bott-Thurston cocycle).
Fix $p\in[1,\infty)$. Define the multiplicative group $1$-cocycle
$\Theta_{p}(\varphi):=(\varphi^{\prime})^{1/p},$
and its logarithm
$\sigma_{p}(\varphi):=\log\Theta_{p}(\varphi)=\frac{1}{p}\log\varphi^{\prime}.$
Define the real-valued group $2$-cochain
$B_{p}(\varphi,\psi):=\int_{\mathbb{R}}\bigl(\sigma_{p}(\varphi)\circ\psi\bigr)\,(\sigma_{p}(\psi))^{\prime}\,dx=\int_{\mathbb{R}}\bigl(\sigma_{p}(\varphi)\circ\psi\bigr)\,d(\sigma_{p}(\psi)).$ (11)
Its infinitesimal cocycle is the alternating bilinear form
$\omega_{p}(u,v):=\frac{1}{p^{2}}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx,\qquad u,v\in\mathfrak{g}^{-\infty}.$ (12)
Theorem 4.7 (Rigidity of the Bott-Thurston class).
Let $p\in[1,\infty)$. Then the following hold.
  • (i)
    The maps $\Theta_{p}$ and $\sigma_{p}$ satisfy the cocycle identities
    $\Theta_{p}(\varphi\circ\psi)=(\Theta_{p}(\varphi)\circ\psi)\,\Theta_{p}(\psi),\qquad\sigma_{p}(\varphi\circ\psi)=\sigma_{p}(\varphi)\circ\psi+\sigma_{p}(\psi).$
  • (ii)
    The cochain $B_{p}$ defined in (11) is a group $2$-cocycle. Its Lie algebra cocycle obtained by differentiating at the identity is $\omega_{p}$ in (12).
  • (iii)
    The form $\omega_{p}$ is a non-trivial Lie algebra $2$-cocycle and
    $\omega_{p}]=\frac{1}{p^{2}}[\omega]\in H^{2}(\mathfrak{g}^{-\infty};\mathbb{R}),\qquad\omega(u,v):=\int_{\mathbb{R}}u^{\prime}v^{\prime\prime}\,dx.$
    In particular the one-dimensional subspace spanned by $\omega_{p}]$ is independent of $p$.
Proof.
Part (i) is the Jacobian cocycle identity already established earlier in the paper (it is an immediate consequence of the chain rule), and the logarithmic identity follows by taking logarithms.
To show that $B_{p}$ is a group $2$-cocycle, write the coboundary as
$(\delta B_{p})(\varphi,\psi,\eta)=B_{p}(\psi,\eta)-B_{p}(\varphi\circ\psi,\eta)+B_{p}(\varphi,\psi\circ\eta)-B_{p}(\varphi,\psi).$
Using $\sigma_{p}(\psi\circ\eta)=\sigma_{p}(\psi)\circ\eta+\sigma_{p}(\eta)$ and bilinearity of the integral,
$B_{p}(\varphi,\psi\circ\eta)$ $=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)+\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\eta)),$
$B_{p}(\varphi\circ\psi,\eta)$ $=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\eta))+\int_{\mathbb{R}}(\sigma_{p}(\psi)\circ\eta)\,d(\sigma_{p}(\eta)).$
Substituting into $\delta B_{p}$, the terms involving $d(\sigma_{p}(\eta))$ cancel, as do $B_{p}(\psi,\eta)$ with the last term above, leaving
$(\delta B_{p})(\varphi,\psi,\eta)=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)-B_{p}(\varphi,\psi).$
Since $d(\sigma_{p}(\psi)\circ\eta)=(\sigma_{p}(\psi)^{\prime}\circ\eta)\eta^{\prime}\,dx$, a change of variables $y=\eta(x)$ gives
$\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi\circ\eta)\,d(\sigma_{p}(\psi)\circ\eta)=\int_{\mathbb{R}}(\sigma_{p}(\varphi)\circ\psi)(y)\,\sigma_{p}(\psi)^{\prime}(y)\,dy=B_{p}(\varphi,\psi),$
so $\delta B_{p}=0$.
For the infinitesimal cocycle, let $\varphi_{t}$ and $\psi_{s}$ be the flows of $u$ and $v$ with $\varphi_{0}=\psi_{0}=\mathrm{Id}$. Then $\varphi_{t}^{\prime}(x)=1+tu^{\prime}(x)+O(t^{2})$ and $\sigma_{p}(\varphi_{t})(x)=\frac{t}{p}u^{\prime}(x)+O(t^{2})$, while $(\sigma_{p}(\psi_{s}))^{\prime}(x)=\frac{s}{p}v^{\prime\prime}(x)+O(s^{2})$. Inserting into (11) yields
$B_{p}(\varphi_{t},\psi_{s})=\frac{ts}{p^{2}}\int_{\mathbb{R}}u^{\prime}(x)v^{\prime\prime}(x)\,dx+o(ts),$
hence $\partial_{t}\partial_{s}|_{0}B_{p}(\varphi_{t},\psi_{s})=\omega_{p}(u,v)$.
Finally, $\omega_{p}$ is alternating, which is immediate from integration by parts with vanishing boundary terms. Since $\omega$ is a non-trivial cocycle, $\omega_{p}=\frac{1}{p^{2}}\omega$ is non-trivial and $\omega_{p}]=\frac{1}{p^{2}}[\omega]$ in cohomology. ∎
Remark 4.8.
The $L^{p}$-Schwarzian does not satisfy a two-term cocycle law. Instead, its composition formula is the three-term identity in Proposition 3.20, and the cross-term is structural, in the sense that it disappears only in the affine case or in the projective limit $p\to\infty$. This highlights a qualitative difference with the metric results of the previous sections in which $L^{p}$-methods linearize the Finsler geometry, but they do not simplify the projective and cohomological structure encoded by Schwarzian-type cocycles.
Remark 4.9.
The Bott-Thurston cocycle admits an expression in $p$-root coordinates and a groupoid coboundary interpretation via a cumulant transgression on the action groupoid $\operatorname{Diff}^{-\infty}(\mathbb{R})\ltimes|\Omega|_{\infty,1}(\mathbb{R})$. This produces a gerbe structure on the density quotient and a canonical absorption of the projective multiplier into Koopman central extensions. These constructions will appear in a companion paper.
5. The real Bers embedding
The Schwarzian cocycle assigns to each $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ a potential $\frac{1}{2}S(\varphi)\in W^{\infty,1}(\mathbb{R})$. This section studies the resulting map $\beta^{-\infty}$, the real Bers map, from three perspectives: (i) as an injective embedding into a Fréchet space of potentials, with the image characterized by the existence of a distinguished positive solution of a Sturm-Liouville equation; (ii) as the curvature correction governing Liouville transport of second-order operators; and (iii) as a local coordinate system on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ near the identity, constructed via the tame inverse function theorem. A canonical inverse reconstruction, energy and factorization identities for the associated Schrödinger operator, and a spectral transport theorem relating Bers potentials to the free Laplacian are also established.
5.1. Definition and injectivity
[…]
Definition 5.1.
The real Bers map is
$\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R}),\qquad\beta^{-\infty}(\varphi)=Q(\varphi)=\frac{1}{2}S(\varphi).$
Theorem 5.2.
[…]
$\widetilde{\operatorname{Diff}}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})\longrightarrow W^{\infty,1}(\mathbb{R}).$
Its image is
$\beta^{-\infty}\bigl(\mathrm{Diff}^{-\infty}(\mathbb{R})\bigr)=\left\{\frac{1}{2}\left(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}\right):f\in W^{\infty,1}(\mathbb{R})\right\}.$
Equivalently, $q$ lies in the image if and only if there exists $f\in W^{\infty,1}(\mathbb{R})$ such that $2q=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$.
Proof.
[…]
To prove injectivity on the quotient, suppose $\beta^{-\infty}(\varphi)=\beta^{-\infty}(\psi)$, i.e. $S(\varphi)=S(\psi)$. Apply the Schwarzian cocycle identity to $\varphi=(\varphi\circ\psi^{-1})\circ\psi$:
$S(\varphi)=\bigl(S(\varphi\circ\psi^{-1})\circ\psi\bigr)(\psi^{\prime})^{2}+S(\psi).$
Since $S(\varphi)=S(\psi)$, we obtain $\bigl(S(\varphi\circ\psi^{-1})\circ\psi\bigr)(\psi^{\prime})^{2}=0$, hence $S(\varphi\circ\psi^{-1})=0$. By the classical characterization of the kernel of the Schwarzian, this implies $\varphi\circ\psi^{-1}\in\mathrm{Aff}(\mathbb{R})$, equivalently $\varphi=A\circ\psi$ for some affine $A$. Thus $\beta^{-\infty}$ is injective on $\widetilde{\operatorname{Diff}}^{-\infty}(\mathbb{R})/\mathrm{Aff}(\mathbb{R})$.
For the image formula, let $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$ and set $f:=\log(\varphi^{\prime})\in W^{\infty,1}(\mathbb{R})$. A direct computation gives
$S(\varphi)=\frac{\varphi^{\prime\prime\prime}}{\varphi^{\prime}}-\frac{3}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}=\bigl(f^{\prime\prime}+(f^{\prime})^{2}\bigr)-\frac{3}{2}(f^{\prime})^{2}=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2},$
hence
$\beta^{-\infty}(\varphi)=\frac{1}{2}S(\varphi)=\frac{1}{2}\left(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}\right),$
so $\beta^{-\infty}(\mathrm{Diff}^{-\infty}(\mathbb{R}))$ is contained in the displayed set.
Conversely, given any $f\in W^{\infty,1}(\mathbb{R})$, let $\varphi:=\Phi_{\infty}^{-1}(f)$ (so that $\log(\varphi^{\prime})=f$). Then the same computation yields $\beta^{-\infty}(\varphi)=\frac{1}{2}(f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2})$, proving that every element of the displayed set is attained. This identifies the image. ∎
Lemma 5.3 (Asymptotic normalization removes the affine gauge).
If $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and $\varphi=A\circ\psi$ for some $A\in\operatorname{Aff}(\mathbb{R})$, then $A=\mathrm{Id}$ and hence $\varphi=\psi$. Consequently, $\beta^{-\infty}$ is injective already on $\operatorname{Diff}^{-\infty}(\mathbb{R})$.
Proof.
Write $A(x)=ax+b$ with $a>0$. Then $(A\circ\psi)^{\prime}=a\,\psi^{\prime}$ so $a=(A\circ\psi)^{\prime}(+\infty)=\varphi^{\prime}(+\infty)$. By the defining asymptotics of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ we have $\varphi^{\prime}(+\infty)=\psi^{\prime}(+\infty)=1$, hence $a=1$. With $a=1$ we have $A(x)=x+b$ and $\varphi(x)-\psi(x)\equiv b$. Since $\varphi-\psi\in W^{\infty,1}(\mathbb{R})$ and $W^{\infty,1}(\mathbb{R})\subset L^{1}(\mathbb{R})$, this forces $b=0$. ∎
Remark 5.4 (Connection to $L^{p}$ linearization).
The proof shows that the Bers map factors as
$\beta^{-\infty}=\frac{1}{2}K_{\infty}\circ\Phi_{\infty},$
where $\Phi_{\infty}(\varphi)=\log(\varphi^{\prime})$ is the $p=\infty$ isometric embedding from Corollary 2.9 and $K_{\infty}(f)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ is the one-dimensional projective curvature operator.
5.2. Affine action on potentials
For $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ define the affine action on potentials
$(\mathcal{A}_{\varphi}q)(x)=q(\varphi(x))\varphi^{\prime}(x)^{2}+Q(\varphi)(x).$
Introduce also the Liouville transport on functions
$(\mathcal{L}_{\varphi}f)(x):=(\varphi^{\prime}(x))^{-1/2}f(\varphi(x)).$ (13)
Proposition 5.5.
For $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and potentials $q$ the following hold.
  1. (i)
    $\mathcal{A}_{\varphi\circ\psi}=\mathcal{A}_{\psi}\circ\mathcal{A}_{\varphi}$.
  2. (ii)
    The Liouville transport intertwines the Sturm-Liouville operators by
    $L_{\mathcal{A}_{\varphi}q}\,\mathcal{L}_{\varphi}=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}\,L_{q},\qquad L_{q}=\partial_{x}^{2}+q.$ (14)
    Equivalently,
    $\mathcal{L}_{\varphi}\,L_{q}\,\mathcal{L}_{\varphi}^{-1}=(\varphi^{\prime})^{-2}\,L_{\mathcal{A}_{\varphi}q}.$ (15)
  3. (iii)
[…]
The orbit of 0 under $\mathcal{A}$ is precisely $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$.
Proof.
(i) Let $\varphi,\psi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and let $q$ be a smooth potential. By definition,
$(\mathcal{A}_{\varphi\circ\psi}q)(x)=q(\varphi(\psi(x)))\,(\varphi\circ\psi)^{\prime}(x)^{2}+Q(\varphi\circ\psi)(x).$
Since $(\varphi\circ\psi)^{\prime}=(\varphi^{\prime}\circ\psi)\,\psi^{\prime}$, the first term equals $q(\varphi(\psi(x)))\,(\varphi^{\prime}(\psi(x))\psi^{\prime}(x))^{2}$. On the other hand,
$(\mathcal{A}_{\psi}(\mathcal{A}_{\varphi}q))(x)=(\mathcal{A}_{\varphi}q)(\psi(x))\,\psi^{\prime}(x)^{2}+Q(\psi)(x),$
and expanding $(\mathcal{A}_{\varphi}q)(\psi(x))$ gives
$(\mathcal{A}_{\psi}(\mathcal{A}_{\varphi}q))(x)=q(\varphi(\psi(x)))\,(\varphi^{\prime}(\psi(x))\psi^{\prime}(x))^{2}+(Q(\varphi)\circ\psi)(x)\,\psi^{\prime}(x)^{2}+Q(\psi)(x).$
Thus $\mathcal{A}_{\varphi\circ\psi}=\mathcal{A}_{\psi}\circ\mathcal{A}_{\varphi}$ is equivalent to
$Q(\varphi\circ\psi)=(Q(\varphi)\circ\psi)\,(\psi^{\prime})^{2}+Q(\psi),$
which holds because $Q=\frac{1}{2}S$ and the Schwarzian cocycle identity gives $S(\varphi\circ\psi)=(S(\varphi)\circ\psi)(\psi^{\prime})^{2}+S(\psi)$.
For (ii), it suffices to verify (14) on $f\in C_{c}^{\infty}(\mathbb{R})$. Write
$g(x):=(\mathcal{L}_{\varphi}f)(x)=(\varphi^{\prime}(x))^{-1/2}f(\varphi(x))=\mathrm{e}^{-u(x)/2}(f\circ\varphi)(x),\qquad u=\log\varphi^{\prime}.$
Then
$\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)g=\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)\bigl(\mathrm{e}^{-u/2}(f\circ\varphi)\bigr)=\mathrm{e}^{-u/2}(f^{\prime}\circ\varphi)\,\varphi^{\prime}=\mathrm{e}^{u/2}(f^{\prime}\circ\varphi),$
and therefore
$\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)g=\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\bigl(\mathrm{e}^{u/2}(f^{\prime}\circ\varphi)\bigr)=\mathrm{e}^{u/2}(f^{\prime\prime}\circ\varphi)\,\varphi^{\prime}=(\varphi^{\prime})^{3/2}(f^{\prime\prime}\circ\varphi).$
Since $Q(\varphi)=\frac{1}{2}S(\varphi)=\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}$, the left-hand side equals $(\partial_{x}^{2}+Q(\varphi))g$. Adding the potential term yields
$L_{\mathcal{A}_{\varphi}q}g=\bigl(\partial_{x}^{2}+q(\varphi)\varphi^{\prime 2}+Q(\varphi)\bigr)g=(\varphi^{\prime})^{3/2}\bigl((f^{\prime\prime}+qf)\circ\varphi\bigr)=(\varphi^{\prime})^{2}\,\mathcal{L}_{\varphi}(L_{q}f),$
which is (14). The conjugation form (15) follows by right-multiplying by $\mathcal{L}_{\varphi}^{-1}$.
[…]
5.3. Miura factorization
Proposition 5.6.
Let $u=\log(\varphi^{\prime})$. Then
$L_{\beta^{-\infty}(\varphi)}=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\left(\partial_{x}+\frac{u^{\prime}}{2}\right).$
The function $y_{2}=\mathrm{e}^{-u/2}=(\varphi^{\prime})^{-1/2}$ is a positive solution of $L_{\beta^{-\infty}(\varphi)}y=0$ annihilated by $(\partial_{x}+\frac{u^{\prime}}{2})$.
Proof.
Set $u=\log(\varphi^{\prime})$, so $\varphi^{\prime}=\mathrm{e}^{u}$. By Theorem 5.2,
$q:=\beta^{-\infty}(\varphi)=\frac{1}{2}S(\varphi)=\frac{1}{2}\Bigl(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\Bigr)=\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}.$
Expanding,
$\Bigl(\partial_{x}-\frac{u^{\prime}}{2}\Bigr)\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)=\partial_{x}^{2}+\frac{u^{\prime\prime}}{2}-\frac{(u^{\prime})^{2}}{4}=\partial_{x}^{2}+q=L_{\beta^{-\infty}(\varphi)}.$
For $y_{2}=\mathrm{e}^{-u/2}$ one computes
$\Bigl(\partial_{x}+\frac{u^{\prime}}{2}\Bigr)y_{2}=\Bigl(-\frac{u^{\prime}}{2}+\frac{u^{\prime}}{2}\Bigr)\mathrm{e}^{-u/2}=0,$
and therefore $L_{\beta^{-\infty}(\varphi)}y_{2}=0$ by the factorization. Since $y_{2}>0$, this is the distinguished positive solution. ∎
5.4. Fréchet structure and tame right inverse
Theorem 5.7.
The map $\beta^{-\infty}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to W^{\infty,1}(\mathbb{R})$ is Fréchet smooth. For $\varphi$ with $u=\log(\varphi^{\prime})$, the differential is
$D_{\varphi}\beta^{-\infty}(\delta\varphi)=\frac{1}{2}(\delta u^{\prime\prime}-u^{\prime}\delta u^{\prime}),\qquad\delta u=\frac{\delta\varphi^{\prime}}{\varphi^{\prime}}.$
The operator $\mathcal{D}_{u}(\delta u)=\delta u^{\prime\prime}-u^{\prime}\delta u^{\prime}$ admits a continuous tame right inverse $R_{u}\colon W^{k,1}(\mathbb{R})\to W^{k+2,1}(\mathbb{R})$ for every $k\geq 0$, depending smoothly on $u$ on bounded subsets.
Proof.
Smoothness of $\beta^{-\infty}$ follows from Proposition 3.2. The differential formula is immediate.
For the right inverse, note $\mathcal{D}_{u}(\delta u)=(\partial_{x}-u^{\prime})\partial_{x}\delta u$. Setting $w=\partial_{x}\delta u$ with $\delta u(-\infty)=0$, we have $(\partial_{x}-u^{\prime})w=v$. Multiplying by $\mathrm{e}^{-u}$ gives $\partial_{x}(\mathrm{e}^{-u}w)=\mathrm{e}^{-u}v$, hence
$w(x)=\mathrm{e}^{u(x)}\int_{-\infty}^{x}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y.$
Define
$(R_{u}v)(x)=\int_{-\infty}^{x}\mathrm{e}^{u(t)}\int_{-\infty}^{t}\mathrm{e}^{-u(y)}v(y)\,\mathrm{d}y\,\mathrm{d}t.$
Then $\partial_{x}(R_{u}v)=w$ and $\mathcal{D}_{u}(R_{u}v)=v$.
For tame bounds, on bounded subsets of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ we have uniform bounds $0<c\leq\mathrm{e}^{u}\leq C$. By Faà di Bruno, $\partial_{x}^{m}(\mathrm{e}^{u(t)-u(y)})$ is a sum of terms $\mathrm{e}^{u(t)-u(y)}P_{m}(u^{\prime},\ldots,u^{(m)})$ with $P_{m}$ polynomial. Since $u^{(j)}\in W^{1,1}\hookrightarrow L^{\infty}$ with $u^{(j)}(\pm\infty)=0$, all coefficients are bounded.
For $m\leq k+2$, the structure of $\partial_{x}^{m}(R_{u}v)$ is
$\partial_{x}^{m}(R_{u}v)=\sum_{j=0}^{m-2}a_{m,j}\partial_{x}^{j}v+b_{m}\partial_{x}(R_{u}v)$
where $a_{m,j},b_{m}$ are bounded by finitely many $\left\|u^{(\ell)}\right\|_{L^{\infty}}$. The term $\partial_{x}(R_{u}v)=w$ satisfies $\left\|w\right\|_{L^{1}}\leq C\left\|v\right\|_{L^{1}}$ by Grönwall. Thus
$\left\|\partial_{x}^{m}(R_{u}v)\right\|_{L^{1}}\leq C_{m}\left(\sum_{j=0}^{m-2}\left\|\partial_{x}^{j}v\right\|_{L^{1}}+\left\|v\right\|_{L^{1}}\right)\leq C_{m}^{\prime}\left\|v\right\|_{W^{k,1}},$
giving $\left\|R_{u}v\right\|_{W^{k+2,1}}\leq C\left\|v\right\|_{W^{k,1}}$.
Smooth dependence on $u$ follows from smoothness of $u\mapsto\mathrm{e}^{\pm u}$ on bounded sets and the linear structure of $R_{u}$. ∎
5.5. Spectral characterization of the image
Theorem 5.8.
A potential $q\in W^{\infty,1}(\mathbb{R})$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if $y^{\prime\prime}+qy=0$ admits linearly independent solutions $y_{1},y_{2}$ with $y_{2}>0$, $W(y_{1},y_{2})\equiv 1$, and $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. The diffeomorphism is $\varphi=y_{1}/y_{2}$.
Proof.
If $q=\beta^{-\infty}(\varphi)$, Proposition 3.25 gives $y_{2}=(\varphi^{\prime})^{-1/2}$, $y_{1}=\varphi(\varphi^{\prime})^{-1/2}$ satisfying the required properties, with $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$.
Conversely, given such $y_{1},y_{2}$, set $\varphi=y_{1}/y_{2}$. Then $\varphi^{\prime}=W(y_{1},y_{2})/y_{2}^{2}=y_{2}^{-2}>0$ and $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. The normalization at -$\infty$ is determined by integrating $\varphi^{\prime}-1$, giving $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. By Proposition 3.25, $S(\varphi)=2q$, so $\beta^{-\infty}(\varphi)=q$. ∎
5.6. Explicit inverse construction
The spectral characterization in Theorem 5.8 becomes a canonical inverse once the asymptotic normalization built into $\operatorname{Diff}^{-\infty}(\mathbb{R})$ is used to remove any residual affine ambiguity.
Lemma 5.9 (Distinguished positive solution).
Let $q\in W^{\infty,1}(\mathbb{R})$ lie in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$. Then there exists a unique positive solution $y_{q}$ of
$y^{\prime\prime}+qy=0$
such that $y_{q}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Moreover $y_{q}(\pm\infty)=1$ and $y_{q}^{\prime}(\pm\infty)=0$.
Proof.
Existence follows from Theorem 5.8 by taking $y_{q}=y_{2}$.
To prove the limits, note that $y_{q}^{-2}-1\in W^{1,1}(\mathbb{R})$ implies $y_{q}^{-2}(x)\to 1$ as $x\to\pm\infty$. Since $y_{q}>0$, this gives $y_{q}(x)\to 1$ as $x\to\pm\infty$. Differentiating $y_{q}^{-2}$ gives $(y_{q}^{-2})^{\prime}=-2y_{q}^{\prime}/y_{q}^{3}\in L^{1}(\mathbb{R})$, and on $W^{1,1}$-bounded subsets we have uniform bounds $0<c\leq y_{q}\leq C$, hence $y_{q}^{\prime}\in L^{1}(\mathbb{R})$ and therefore $y_{q}^{\prime}(\pm\infty)=0$.
For uniqueness, let $y,\tilde{y}>0$ be solutions with $y^{-2}-1,\tilde{y}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Their Wronskian W(y,y~)W($\tilde{y}$) is constant. Using $y(\pm\infty)=\tilde{y}(\pm\infty)=1$ and $y^{\prime}(\pm\infty)=\tilde{y}^{\prime}(\pm\infty)=0$ we obtain $W(y,\tilde{y})=\lim_{x\to\pm\infty}(y\tilde{y}^{\prime}-y^{\prime}\tilde{y})=0$, hence $W\equiv 0$ and $y=c\,\tilde{y}$. Taking $x\to\pm\infty$ yields $c=1$. ∎
Theorem 5.10 (Canonical reconstruction formula).
Let $q\in W^{\infty,1}(\mathbb{R})$. Then $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the distinguished positive solution $y_{q}$ from Lemma 5.9 exists. In that case, the unique $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\beta^{-\infty}(\varphi)=q$ is given by
$\varphi^{\prime}(x)=y_{q}(x)^{-2},\qquad\varphi(x)=x+\int_{-\infty}^{x}\bigl(y_{q}(t)^{-2}-1\bigr)\,dt.$ (16)
Equivalently, with $u=\log\varphi^{\prime}$ one has $u=-2\log y_{q}\in W^{\infty,1}(\mathbb{R})$ and $\beta^{-\infty}(\varphi)=\frac{1}{2}\bigl(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\bigr)=q$.
Proof.
Assume first $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$. Then Lemma 5.9 provides $y_{q}>0$ with $y_{q}^{-2}-1\in W^{\infty,1}(\mathbb{R})$. Define $\varphi$ by (16). Then $\varphi^{\prime}>0$ and $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$, hence $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ by the defining asymptotics. Moreover, setting $u=\log\varphi^{\prime}=-2\log y_{q}$ and using $y_{q}^{\prime\prime}=-qy_{q}$, a direct computation yields
$\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)=\frac{1}{2}\left(-2\frac{y_{q}^{\prime\prime}}{y_{q}}+2\left(\frac{y_{q}^{\prime}}{y_{q}}\right)^{2}-\frac{1}{2}\cdot 4\left(\frac{y_{q}^{\prime}}{y_{q}}\right)^{2}\right)=-\frac{y_{q}^{\prime\prime}}{y_{q}}=q.$
Thus $\beta^{-\infty}(\varphi)=q$.
Conversely, if $y_{q}$ exists, then the construction above produces $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ with $\beta^{-\infty}(\varphi)=q$, hence $q$ lies in the image. Uniqueness of $\varphi$ follows from Lemma 5.3. ∎
Remark 5.11 (Volterra form of the distinguished solution).
If $y_{q}$ is the distinguished solution, then integrating $y_{q}^{\prime\prime}=-qy_{q}$ twice from -$\infty$ gives the Volterra equation
$y_{q}(x)=1-\int_{-\infty}^{x}(x-t)\,q(t)\,y_{q}(t)\,dt,$ (17)
which uniquely characterizes $y_{q}$ among solutions with $y_{q}(-\infty)=1$, $y_{q}^{\prime}(-\infty)=0$.
5.7. Riccati-Miura characterization and energy identity
Proposition 5.12 (Riccati-Miura criterion with integral constraint).
Let $q\in W^{\infty,1}(\mathbb{R})$. The following are equivalent:
  1. (i)
    $q\in\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$.
  2. (ii)
    There exists $w\in W^{\infty,1}(\mathbb{R})$ with $w(\pm\infty)=0$ and
    $\int_{\mathbb{R}}w(x)\,dx=0$ (18)
    such that
    $q=w^{\prime}-w^{2}.$ (19)
  3. (iii)
    There exists a positive solution $y$ of $y^{\prime\prime}+qy=0$ with $y^{-2}-1\in W^{\infty,1}(\mathbb{R})$, in which case $w=-y^{\prime}/y$ satisfies (18) and (19).
Proof.
[…]
$q=\frac{1}{2}\left(u^{\prime\prime}-\frac{1}{2}(u^{\prime})^{2}\right)=\left(\frac{u^{\prime}}{2}\right)^{\prime}-\left(\frac{u^{\prime}}{2}\right)^{2}.$
Set $w=u^{\prime}/2$. Since $u\in W^{1,1}(\mathbb{R})$ we have $u(\pm\infty)=0$ and hence $\int_{\mathbb{R}}w=\frac{1}{2}(u(\infty)-u(-\infty))=0$.
[…]
$y(x):=\exp\left(-\int_{-\infty}^{x}w(t)\,dt\right).$
Then $y>0$, $y^{\prime}=-wy$, and $y^{\prime\prime}=(w^{2}-w^{\prime})y=-qy$, so $y^{\prime\prime}+qy=0$ and $w=-y^{\prime}/y$. Moreover,
$y^{-2}(x)=\exp\left(2\int_{-\infty}^{x}w(t)\,dt\right)=\exp(u(x)),\qquad u(x):=2\int_{-\infty}^{x}w(t)\,dt.$
Since $w\in W^{\infty,1}$, we have $u\in W^{\infty,1}$. The integral constraint (18) implies $u(\infty)=0$, hence $y^{-2}-1=\mathrm{e}^{u}-1\in W^{\infty,1}(\mathbb{R})$.
[…]
Proposition 5.13 (Energy identity and correct factorization).
Let $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $u=\log\varphi^{\prime}$ and $q=\beta^{-\infty}(\varphi)$. Then, with $w=u^{\prime}/2$, one has the energy identity
$\int_{\mathbb{R}}q(x)\,dx=-\int_{\mathbb{R}}w(x)^{2}\,dx=-\frac{1}{4}\int_{\mathbb{R}}(u^{\prime}(x))^{2}\,dx\leq 0.$ (20)
Moreover the Schrödinger operator
$H_{q}:=-\partial_{x}^{2}-q$ (21)
admits the exact nonnegative factorization
$H_{q}=\left(-\partial_{x}+w\right)\left(\partial_{x}+w\right)=A^{\ast}A,\qquad A:=\partial_{x}+w,$ (22)
and therefore $(H_{q}f,f)_{L^{2}}\geq 0$ on $C_{c}^{\infty}(\mathbb{R})$ (and on the natural self-adjoint domain).
Proof.
By Proposition 5.12 we have $q=w^{\prime}-w^{2}$ with $w(\pm\infty)=0$. Integrating gives $\int_{\mathbb{R}}q=\int_{\mathbb{R}}w^{\prime}-\int_{\mathbb{R}}w^{2}=-\int_{\mathbb{R}}w^{2}$, which is (20).
For the factorization, compute for $f\in C_{c}^{\infty}(\mathbb{R})$:
$\left(-\partial_{x}+w\right)\left(\partial_{x}+w\right)f=-f^{\prime\prime}-w^{\prime}f-wf^{\prime}+wf^{\prime}+w^{2}f=-f^{\prime\prime}-(w^{\prime}-w^{2})f=-f^{\prime\prime}-qf.$
Since $A^{\ast}=-\partial_{x}+w$ on $C_{c}^{\infty}(\mathbb{R})$, we obtain $H_{q}=A^{\ast}A$ and hence $(H_{q}f,f)=\|Af\|_{L^{2}}^{2}\geq 0$. ∎
Remark 5.14 (Local coordinates and spectral transport).
Local Bers coordinates centered at arbitrary basepoints can be constructed, with transition maps given by Schwarzian-shifted affine connections. A spectral transport formula conjugates Schrödinger operators along paths in the Bers image. These constructions will appear in a companion paper.
6. Scattering theory and the nonlinear Fourier coordinate
The Sturm-Liouville characterization of the real Bers image in Theorem 5.8 admits a refinement in terms of one-dimensional scattering. For a real potential $q\in W^{\infty,1}(\mathbb{R})$ we consider the Sturm-Liouville operator $\mathcal{L}_{q}=\partial_{x}^{2}+q$ from Proposition 5.5 and the associated Schrödinger operator
$H_{q}:=-\mathcal{L}_{q}=-\partial_{x}^{2}-q$ (23)
on $L^{2}(\mathbb{R})$ with domain $H^{2}(\mathbb{R})$. Since $q\in L^{1}(\mathbb{R})$, the operator $H_{q}$ is short-range: its essential spectrum is $0,\infty)$ and its negative spectrum, if present, consists of finitely many eigenvalues. For Bers-Miura potentials $q=\beta^{-\infty}(\varphi)$, the factorization in Proposition 5.6 implies $H_{q}\geq 0$, hence the scattering data contain no discrete bound state part. In this situation the reflection coefficient provides a nonlinear spectral coordinate on $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
6.1. Jost solutions and scattering data
Definition 6.1 (Jost solutions).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $k\in\mathbb{C}$ with $\Im k\geq 0$. A Jost solution for $H_{q}$ is a solution $y(x,k)$ of
$y^{\prime\prime}(x,k)-q(x)y(x,k)=k^{2}y(x,k)$ (24)
with prescribed plane wave asymptotics at one end. The right and left Jost solutions $f_{\pm}(x,k)$ are uniquely characterized by
[…]
For $k\in\mathbb{R}\setminus\{0\}$ the scattering coefficients $a(k),b(k)$ are defined by the identity
$f_{-}(x,k)=a(k)f_{+}(x,-k)+b(k)f_{+}(x,k),\qquad k\in\mathbb{R}\setminus\{0\}.$ (26)
The transmission and reflection coefficients are $T(k)=a(k)^{-1}$ and $R(k)=b(k)/a(k)$.
Proposition 6.2 (Existence and basic properties of scattering data).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real.
  1. (i)
    For each $k$ with $\Im k\geq 0$ and $k\neq 0$, the Jost solutions exist, are $C^{\infty}$ in $x$, and depend holomorphically on $k$ for $\Im k>0$ and continuously on $k$ for $\Im k\geq 0$. They solve the Volterra equations
    $f_{+}(x,k)=\mathrm{e}^{\mathrm{i}kx}+\int_{x}^{\infty}\frac{\sin(k(t-x))}{k}\,q(t)f_{+}(t,k)\,\mathrm{d}t,\qquad\Im k\geq 0,$ (27)
    and the analogous equation for $f_{-}(x,k)$ with integration from -$\infty$ to $x$.
  2. (ii)
    The coefficients $a(k),b(k)$ are $C^{\infty}$ on $\mathbb{R}\setminus\{0\}$ and satisfy the symmetry
    $a(-k)=\overline{a(k)},\qquad b(-k)=\overline{b(k)},\qquad k\in\mathbb{R}\setminus\{0\}.$ (28)
  3. (iii)
    One has the flux identities
    $|a(k)|^{2}=1+|b(k)|^{2},\qquad|T(k)|^{2}+|R(k)|^{2}=1,\qquad k\in\mathbb{R}\setminus\{0\}.$ (29)
    In particular $|a(k)|\geq 1$, with equality if and only if $b(k)=0$.
  4. (iv)
    For every $N\in\mathbb{N}$ there exists $C_{N}$ such that, for $|k|\geq 1$,
    $|a(k)-1|+|b(k)|\leq C_{N}|k|^{-N}.$
    Consequently a(k)1a(k)-1 and $b(k)$ extend to Schwartz functions of $k$ on $\mathbb{R}$.
Proof.
Fix $k$ with $\Im k\geq 0$ and $k\neq 0$. The Green kernel for $\partial_{x}^{2}-k^{2}$ with outgoing normalization at $\infty$ is $G_{k}(x,t)=\frac{\sin(k(t-x))}{k}\mathbf{1}_{t\geq x}$, so variation of constants gives (27). Let $X=L^{\infty}([x_{0},\infty))$ with norm $\|u\|_{X}=\sup_{t\geq x_{0}}|u(t)|$. For $u\in X$ the integral operator
$(K_{k}u)(x):=\int_{x}^{\infty}\frac{\sin(k(t-x))}{k}\,q(t)u(t)\,\mathrm{d}t$
satisfies, via the bound $|\sin(k(t-x))/k|\leq|t-x|$ valid for all $k\neq 0$,
$\|K_{k}u\|_{X}\leq\left(\int_{x_{0}}^{\infty}(t-x_{0})\,|q(t)|\,\mathrm{d}t\right)\|u\|_{X}.$
Since $q\in W^{\infty,1}(\mathbb{R})$ implies $\int_{\mathbb{R}}(1+|t|)|q(t)|\,\mathrm{d}t<\infty$, choosing $x_{0}$ large enough that $\int_{x_{0}}^{\infty}(t-x_{0})|q(t)|\,\mathrm{d}t<\frac{1}{2}$ makes the map $u\mapsto\mathrm{e}^{\mathrm{i}k\cdot}+K_{k}u$ a contraction on $X$, hence it has a unique fixed point. Patching the resulting solution as $x_{0}\to$-$\infty$ gives a globally defined $f_{+}(x,k)$ satisfying (27) for all $x\in\mathbb{R}$. Differentiating (27) in $x$ shows $f_{+}$ is $C^{\infty}$ in $x$ and solves (24). Holomorphy in $k$ for $\Im k>0$ follows by differentiating the Volterra equation in $k$: each $k$-derivative falls on $\sin(k(t-x))/k$ and produces factors bounded by $|t-x|^{m}$ times $\mathrm{e}^{-(\Im k)(t-x)}$, which remain integrable against $|q(t)|$ for $\Im k>0$. Continuity on $\Im k=0$ follows by dominated convergence, using $|\sin(k(t-x))|\leq\mathrm{e}^{(\Im k)(t-x)}$.
For real $k\neq 0$, the coefficients $a(k),b(k)$ are determined by Wronskians. Since the Wronskian $W(u,v)=uv^{\prime}-u^{\prime}v$ is constant in $x$ for any pair of solutions of (24), applying $W(f_{+}(k),\cdot)$ to (26) yields
$a(k)=\frac{1}{2\mathrm{i}k}W(f_{-}(\cdot,k),f_{+}(\cdot,k)),$ (30)
because $W(f_{+}(\cdot,k),f_{+}(\cdot,-k))=-2\mathrm{i}k$ by the asymptotics (25). Similarly, applying $W(f_{+}(\cdot,-k),\cdot)$ to (26) gives
$b(k)=\frac{1}{2\mathrm{i}k}W(f_{+}(\cdot,-k),f_{-}(\cdot,k)).$ (31)
The dependence asserted in (ii) follows from the smooth dependence of $f_{\pm}$ on $k$ on $\mathbb{R}\setminus\{0\}$ and the identities (30)-(31). The symmetry (28) follows from reality of $q$ and uniqueness of Jost solutions: for $k\in\mathbb{R}\setminus\{0\}$ the function $\overline{f_{+}(x,k)}$ solves (24) and has asymptotic $\mathrm{e}^{-\mathrm{i}kx}$ as $x\to+\infty$, hence $\overline{f_{+}(x,k)}=f_{+}(x,-k)$, and similarly $\overline{f_{-}(x,k)}=f_{-}(x,-k)$.
To prove (29), consider the conserved current for real solutions of the stationary equation: if $y$ solves (24) with real $k$, then
$J[y]:=\frac{1}{2\mathrm{i}}\bigl(\overline{y}\,y^{\prime}-y\,\overline{y}^{\prime}\bigr)$
is constant in $x$ because $J^{\prime}[y]=\Im(\overline{y}y^{\prime\prime})$ and $y^{\prime\prime}=-(k^{2}+q)y$ with $k^{2}+q$ real. Evaluate $J[f_{-}(\cdot,k)]$ as $x\to$-$\infty$ using $f_{-}(x,k)\sim\mathrm{e}^{-\mathrm{i}kx}$ to obtain $J[f_{-}]=-k$. Evaluate the same current as $x\to+\infty$ using (26) and $f_{+}(x,\pm k)\sim\mathrm{e}^{\pm\mathrm{i}kx}$:
$f_{-}(x,k)\sim a(k)\mathrm{e}^{-\mathrm{i}kx}+b(k)\mathrm{e}^{\mathrm{i}kx}\quad\text{as }x\to+\infty,$
hence $J[f_{-}]=-k|a(k)|^{2}+k|b(k)|^{2}$. Equating the two values gives $|a(k)|^{2}=1+|b(k)|^{2}$. Dividing by $|a(k)|^{2}$ yields $|T(k)|^{2}+|R(k)|^{2}=1$ and $|a(k)|\geq 1$, with equality exactly when $b(k)=0$.
For (iv), introduce $m_{+}(x,k):=\mathrm{e}^{-\mathrm{i}kx}f_{+}(x,k)$, so $m_{+}(x,k)\to 1$ as $x\to+\infty$ and $m_{+}$ satisfies the Volterra equation obtained by substituting $f_{+}=\mathrm{e}^{\mathrm{i}k\cdot}m_{+}$ into (27),
$m_{+}(x,k)=1+\int_{x}^{\infty}\frac{\mathrm{e}^{2\mathrm{i}k(t-x)}-1}{2\mathrm{i}k}\,q(t)m_{+}(t,k)\,\mathrm{d}t.$
Iterating this equation shows $m_{+}(x,k)$ is uniformly bounded in $x$ for $|k|\geq 1$ and $\Im k\geq 0$, with bounds controlled by $\|q\|_{L^{1}}$. Using (31) and the asymptotics of $f_{-}$ at -$\infty$ one obtains the standard representation
$b(k)=-\frac{1}{2\mathrm{i}k}\int_{\mathbb{R}}q(t)m_{+}(t,k)\mathrm{e}^{2\mathrm{i}kt}\,\mathrm{d}t,\qquad k\in\mathbb{R}\setminus\{0\},$
and similarly a(k)1a(k)-1 is given by an oscillatory integral with integrand built from $q$ and $m_{+}$. Since $q^{(N)}\in L^{1}(\mathbb{R})$ for every $N$ and $m_{+}(\cdot,k)$ is bounded with bounded $t$-derivatives expressible by the differential equation, repeated integration by parts in $t$ gives $|b(k)|\leq C_{N}|k|^{-N}$ for $|k|\geq 1$. The same argument applies to a(k)1a(k)-1. This proves rapid decay and hence the Schwartz property on $\mathbb{R}$. ∎
6.2. Miura factorization and absence of bound states
Lemma 6.3 (No bound states for Bers-Miura potentials).
Let $q=\beta^{-\infty}(\varphi)$ with $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and write $u=\log(\varphi^{\prime})$. Set $A:=\partial_{x}+\frac{u^{\prime}}{2}$. Then
$H_{q}=A^{*}A\geq 0$ (32)
as a self-adjoint operator on $L^{2}(\mathbb{R})$. In particular $H_{q}$ has no negative eigenvalues, hence no bound states. Moreover $H_{q}y=0$ admits a positive solution $y=\mathrm{e}^{-u/2}=(\varphi^{\prime})^{-1/2}$, which is bounded and satisfies $y(x)\to 1$ as $x\to\pm\infty$.
Proof.
Proposition 5.6 gives
$\mathcal{L}_{q}=\partial_{x}^{2}+q=\left(\partial_{x}-\frac{u^{\prime}}{2}\right)\left(\partial_{x}+\frac{u^{\prime}}{2}\right),$
hence $H_{q}=-\mathcal{L}_{q}=(-\partial_{x}+\frac{u^{\prime}}{2})(\partial_{x}+\frac{u^{\prime}}{2})$. Since the adjoint of $\partial_{x}$ on $L^{2}(\mathbb{R})$ is $\partial_{x}$ and $u^{\prime}/2$ acts by multiplication, this is exactly $A^{*}A$, proving (32). For $f\in H^{1}(\mathbb{R})$ one has $\langle f,H_{q}f\rangle=\|Af\|_{L^{2}}^{2}\geq 0$, so $H_{q}\geq 0$ and there are no negative eigenvalues. Finally $Ay=0$ has the positive solution $y=\mathrm{e}^{-u/2}$ and $u\in W^{\infty,1}(\mathbb{R})$ implies $u(x)\to 0$ as $x\to\pm\infty$, hence $y(x)\to 1$ at both ends. ∎
6.3. Spectral characterization at zero energy
Theorem 6.4 (Spectral characterization of the Bers image).
Let $q\in W^{\infty,1}(\mathbb{R})$ be real and let $f_{+}(x,k)$ be the right Jost solution of $H_{q}$ as in Definition 6.1. Then $q$ lies in $\beta^{-\infty}(\operatorname{Diff}^{-\infty}(\mathbb{R}))$ if and only if the zero-energy Jost solution $f_{+}(x,0)$ satisfies
$f_{+}(x,0)>0\qquad\text{and}\qquad f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R}).$ (33)
In this case the associated diffeomorphism is
$\varphi(x)=x+\int_{-\infty}^{x}\bigl(f_{+}(t,0)^{-2}-1\bigr)\,\mathrm{d}t,$ (34)
so that $\varphi^{\prime}=f_{+}(\cdot,0)^{-2}$ and $q=\beta^{-\infty}(\varphi)$. Moreover, with $w:=-f_{+}^{\prime}(\cdot,0)/f_{+}(\cdot,0)$ one has $w\in W^{\infty,1}(\mathbb{R})$ and the Miura factorization
$\mathcal{L}_{q}=(\partial_{x}-w)(\partial_{x}+w)$
holds.
Proof.
Assume first that $q=\beta^{-\infty}(\varphi)$ for some $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Proposition 5.6 gives a positive solution $y_{2}=(\varphi^{\prime})^{-1/2}$ of $\mathcal{L}_{q}y=0$, equivalently $H_{q}y=0$, and $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. Since $\varphi^{\prime}\to 1$ as $x\to+\infty$, we have $y_{2}(x)\to 1$ as $x\to+\infty$. The Volterra equation (27) at $k=0$ becomes
$y(x)=1+\int_{x}^{\infty}(t-x)q(t)y(t)\,\mathrm{d}t,$
and it has a unique bounded solution with $y(x)\to 1$ as $x\to+\infty$. By the defining property (25) at $k=0$, this solution is exactly $f_{+}(x,0)$. Thus $f_{+}(x,0)=y_{2}(x)>0$ and $f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R})$, proving (33).
Conversely, assume (33). Define $\varphi$ by (34). Since $f_{+}(x,0)^{-2}-1\in W^{\infty,1}(\mathbb{R})$, we have $\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$ and $\varphi^{\prime}>0$, hence $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$. Set $y_{2}:=f_{+}(\cdot,0)$ and $y_{1}:=\varphi y_{2}$. Then $y_{2}$ solves $y_{2}^{\prime\prime}+qy_{2}=0$. Since $\varphi^{\prime}=y_{2}^{-2}$, the quotient rule gives
$\left(\frac{y_{1}}{y_{2}}\right)^{\prime}=\varphi^{\prime}=y_{2}^{-2},$
hence $y_{1}^{\prime}y_{2}-y_{1}y_{2}^{\prime}=\varphi^{\prime}y_{2}^{2}=1$. Thus $y_{1},y_{2}$ are linearly independent solutions with unit Wronskian in the convention of Theorem 5.8. Moreover $y_{2}>0$ and $y_{2}^{-2}-1=\varphi^{\prime}-1\in W^{\infty,1}(\mathbb{R})$. By Theorem 5.8, this implies $q=\beta^{-\infty}(\varphi)$.
Finally, put $w:=-y_{2}^{\prime}/y_{2}$. Since $y_{2}>0$ and $y_{2}^{-2}-1\in W^{\infty,1}(\mathbb{R})$, the function $\log y_{2}$ has derivative w-w in $W^{\infty,1}(\mathbb{R})$, hence $w\in W^{\infty,1}(\mathbb{R})$. Writing $y_{2}=\mathrm{e}^{-\int w}$ and substituting into $y_{2}^{\prime\prime}+qy_{2}=0$ gives $q=w^{\prime}-w^{2}$ and therefore $\mathcal{L}_{q}=(\partial_{x}-w)(\partial_{x}+w)$. ∎
6.4. The scattering transform as a nonlinear Fourier map
For $q\in W^{\infty,1}(\mathbb{R})$ the scattering data consist of the reflection coefficient together with possible discrete bound state data. Lemma 6.3 shows that for $q=\beta^{-\infty}(\varphi)$ no bound states occur, so the reflection coefficient alone captures the scattering data. This yields a nonlinear spectral coordinate on the real Bers image.
Definition 6.5 (Scattering Bers map).
Define
$\mathcal{S}\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to\mathscr{S}(\mathbb{R}),\qquad\mathcal{S}(\varphi):=R_{\beta^{-\infty}(\varphi)},$
where $R_{q}$ is the reflection coefficient of the Schrödinger operator $H_{q}$ in (23).
Theorem 6.6 (Injectivity of the scattering Bers map).
The map $\mathcal{S}$ is constant on left $\operatorname{Aff}(\mathbb{R})$-cosets and descends to an injective map on $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$. Moreover, for every $\varphi\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ the function $R=\mathcal{S}(\varphi)$ satisfies:
  1. (i)
    $\overline{R(-k)}=R(k)$ for all $k\in\mathbb{R}$,
  2. (ii)
    $|R(k)|<1$ for all $k\in\mathbb{R}\setminus\{0\}$,
  3. (iii)
    the boundary modulus determines an outer function $a_{\mathrm{out}}$ on the upper half-plane,
    $a_{\mathrm{out}}(z)=\exp\!\left(-\frac{1}{2\pi\mathrm{i}}\int_{\mathbb{R}}\frac{\log(1-|R(\xi)|^{2})}{\xi-z}\,\mathrm{d}\xi\right),\qquad\Im z>0,$ (35)
    and the transmission denominator $a(z)$ for $H_{\beta^{-\infty}(\varphi)}$ coincides with $a_{\mathrm{out}}(z)$ and has no zeros in $\Im z>0$.
Proof.
If $\psi\in\operatorname{Aff}(\mathbb{R})$ then $\beta^{-\infty}(\varphi\circ\psi)=\beta^{-\infty}(\varphi)$ by Theorem 5.2, hence $\mathcal{S}(\varphi\circ\psi)=\mathcal{S}(\varphi)$, so $\mathcal{S}$ factors through $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
For injectivity, let $\varphi_{1},\varphi_{2}\in\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $q_{j}=\beta^{-\infty}(\varphi_{j})$. Assume $\mathcal{S}(\varphi_{1})=\mathcal{S}(\varphi_{2})$, so $R_{q_{1}}(k)=R_{q_{2}}(k)$ for all $k\in\mathbb{R}\setminus\{0\}$. By Lemma 6.3, the operators $H_{q_{j}}$ have no negative spectrum, hence their scattering data have no discrete part. The Gel’fand-Levitan-Marchenko reconstruction for short-range one-dimensional Schrödinger operators then implies $q_{1}=q_{2}$ from equality of the reflection coefficients in the no-bound-state case, see [7, Theorem 1]. The injectivity of $\beta^{-\infty}$ modulo $\operatorname{Aff}(\mathbb{R})$ from Theorem 5.2 then gives $\varphi_{1}]=[\varphi_{2}]$ in $\operatorname{Diff}^{-\infty}(\mathbb{R})/\operatorname{Aff}(\mathbb{R})$.
For the additional properties, (i) follows from reality of $q$ and the symmetry $b(-k)=\overline{b(k)}$, $a(-k)=\overline{a(k)}$ in (28). For (ii), Proposition 6.2 gives $|T(k)|^{2}+|R(k)|^{2}=1$ for $k\neq 0$, and $T(k)\neq 0$ for real $k\neq 0$ since $T(k)=0$ would produce a nontrivial solution of (24) that is purely outgoing at both ends at energy $k^{2}>0$, which is impossible. Thus $|R(k)|<1$ for $k\neq 0$.
For (iii), the function $a(z)$ extends holomorphically to $\Im z>0$ and has no zeros there exactly because zeros in $\Im z>0$ correspond to bound states of $H_{q}$ and these are excluded by Lemma 6.3. On the real axis one has $|a(k)|^{-2}=|T(k)|^{2}=1-|R(k)|^{2}$ for $k\in\mathbb{R}\setminus\{0\}$ by (29). Since $R\in\mathscr{S}(\mathbb{R})$, the function $\log(1-|R|^{2})$ is integrable and the outer function $a_{\mathrm{out}}$ defined by (35) lies in the Hardy class on $\Im z>0$ and has boundary modulus $|a_{\mathrm{out}}(k)|=|a(k)|$. The quotient $a/a_{\mathrm{out}}$ is then a bounded holomorphic function on $\Im z>0$ with unimodular boundary values, hence constant. The normalization $a(z)\to 1$ as $|z|\to\infty$ forces that constant to be $1$, so $a=a_{\mathrm{out}}$. ∎
Remark 6.7 (Further scattering theory).
The scattering data admit further structure: a relative spectral zeta function whose residue at $z=-\frac{1}{2}$ is a multiple of Fisher information, Bers-Mercer reproducing kernels, and a Marchenko reconstruction providing an explicit inverse for the scattering transform. These developments, along with regularized determinants, will appear in a companion paper.
7. Projective curvature of densities and forced sign-change theorems
7.1. Asymptotically Lebesgue densities
Fix $dx$ as the reference density on $\mathbb{R}$.
Definition 7.1 (Asymptotically Lebesgue densities).
The manifold of asymptotically Lebesgue densities is
$|\Omega|_{\infty,1}(\mathbb{R})=\{\mu=g\,dx\mid g>0,\;g-1\in W^{\infty,1}(\mathbb{R})\}.$
The condition $g-1\in W^{1,1}(\mathbb{R})$ forces $g(x)\to 1$ as $x\to\pm\infty$.
Lemma 7.2 (Regularity, boundedness, and asymptotics).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$. Then $g$ admits a $C^{\infty}$ representative,
$g^{(k)}\in L^{1}(\mathbb{R})\quad\forall k\geq 1,\qquad\lim_{x\to\pm\infty}g^{(k)}(x)=0\quad\forall k\geq 1,\qquad\lim_{x\to\pm\infty}g(x)=1.$
In particular there exist $0<m\leq M<\infty$ with $m\leq g\leq M$ on $\mathbb{R}$, and every derivative $g^{(k)}$ is bounded. Moreover, if $F\in C^{\infty}((0,\infty))$ has bounded derivatives on $[m,M]$, then $F(g)-F(1)\in W^{\infty,1}(\mathbb{R})$.
Proof.
Since $g-1\in W^{k,1}(\mathbb{R})$ for every $k$, the one-dimensional Sobolev embedding yields a representative in $C^{k-1}(\mathbb{R})$ with (k1)(k-1)st derivative absolutely continuous. Varying $k$ gives a $C^{\infty}$ representative and $g^{(k)}\in L^{1}(\mathbb{R})$ for all $k\geq 1$. For $k\geq 1$, we have $g^{(k)}\in W^{1,1}(\mathbb{R})$, hence $g^{(k)}$ has finite limits at $\pm\infty$. Since also $g^{(k)}\in L^{1}(\mathbb{R})$, these limits must be 0. The limit $g(\pm\infty)=1$ follows from $g-1\in W^{1,1}(\mathbb{R})$. Continuity and the asymptotic limit give global upper and lower bounds on $g$. Finally, the boundedness of $g^{(k)}$ follows from $g^{(k)}\in W^{1,1}(\mathbb{R})$, and the last statement follows by iterated chain rule together with the algebra property of $W^{k,1}(\mathbb{R})$ and boundedness of $F^{(j)}$ on $[m,M]$. ∎
7.2. The Jacobian isomorphism
Definition 7.3 (Jacobian map).
The Jacobian map is
$J\colon\operatorname{Diff}^{-\infty}(\mathbb{R})\to|\Omega|_{\infty,1}(\mathbb{R}),\qquad J(\varphi)=\varphi^{\prime}\,dx.$
Theorem 7.4 (Jacobian correspondence).
$J$ is a Fréchet diffeomorphism with inverse
$J^{-1}(g\,dx)(x)=x+\int_{-\infty}^{x}(g(t)-1)\,dt.$
Moreover:
  1. (i)
    (Linearization intertwining) The diagram
    $\operatorname{Diff}^{-\infty}(\mathbb{R})$${\lvert\Omega\rvert}_{\infty,1}(\mathbb{R})$$W^{\infty,1}(\mathbb{R})$$W^{\infty,1}(\mathbb{R})$$\scriptstyle{J}$$\scriptstyle{\Phi_{p}}$$\scriptstyle{\Psi_{p}}$
    commutes, where $\Psi_{p}(g\,dx)=p(g^{1/p}-1)$ for $p<\infty$ and $\Psi_{\infty}(g\,dx)=\log g$.
  2. (ii)
    (Finsler isometry) $J$ is an isometry from $(\operatorname{Diff}^{-\infty}(\mathbb{R}),F_{\dot{W}^{1,p}})$ to $(|\Omega|_{\infty,1}(\mathbb{R}),F_{p}^{\mathrm{FR}})$.
  3. (iii)
    (Equivariance) $J(\varphi\circ\psi)=\psi^{*}J(\varphi)$.
  4. (iv)
    (Schwarzian compatibility) Setting $S(\mu)=f^{\prime\prime}-\frac{1}{2}(f^{\prime})^{2}$ for $f=\log(\mu/dx)$, we have $S(J(\varphi))=S(\varphi)$.
Proof.
Bijectivity follows from the fundamental theorem of calculus: if $\varphi^{\prime}\,dx=g\,dx$ with $g-1\in W^{\infty,1}$ then $\varphi(x)=x+\int_{-\infty}^{x}(g(t)-1)\,dt$, which lies in $\operatorname{Diff}^{-\infty}(\mathbb{R})$, and conversely $g=\varphi^{\prime}$ satisfies $g-1\in W^{\infty,1}$. Smoothness of $J$ follows from smoothness of $\varphi\mapsto\varphi^{\prime}$ and continuity of induced pullback maps in the Fréchet topologies; smoothness of $J^{-1}$ is immediate from the explicit formula since $g\mapsto\int_{-\infty}^{\cdot}(g-1)$ is continuous linear on $W^{\infty,1}$.
For the Finsler isometry, let $\delta\varphi=u\circ\varphi$. Then $D_{\varphi}J(\delta\varphi)=(u^{\prime}\circ\varphi)\varphi^{\prime}\,dx$ and
$F_{p}^{\mathrm{FR}}(J(\varphi),D_{\varphi}J(\delta\varphi))^{p}=\int_{\mathbb{R}}|u^{\prime}\circ\varphi|^{p}\,\varphi^{\prime}\,dx=\int_{\mathbb{R}}|u^{\prime}|^{p}\,dy=F_{\dot{W}^{1,p}}(\varphi,\delta\varphi)^{p}.$
The remaining claims are direct computations. ∎
Corollary 7.5 (Transferred Fisher-Rao geodesics).
All geometric structures on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ transfer to $|\Omega|_{\infty,1}(\mathbb{R})$ via $J$. For $p<\infty$, the $L^{p}$ Fisher-Rao geodesic from $\mu_{0}=g_{0}\,dx$ to $\mu_{1}=g_{1}\,dx$ is
$\mu(t)=\Bigl((1-t)g_{0}^{1/p}+tg_{1}^{1/p}\Bigr)^{p}\,dx.$
7.3. Score functions, projective curvature, and Fisher information
Definition 7.6 (Score).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$, the score function is
$s_{\mu}=\frac{g^{\prime}}{g}=(\log g)^{\prime}.$
Proposition 7.7 (Schwarzian as score curvature).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and $s=s_{\mu}$. Then
$S(\mu)=s^{\prime}-\frac{1}{2}s^{2}.$
Moreover, $S(\mu)=0$ in $|\Omega|_{\infty,1}(\mathbb{R})$ if and only if $\mu=dx$.
Proof.
The identity is the definition of $S(\mu)$ with $f=\log g$ and $s=f^{\prime}$. If $S(\mu)=0$, then $s^{\prime}=\frac{1}{2}s^{2}$. By Lemma 7.2, $s\in C^{0}(\mathbb{R})$ with $s(\pm\infty)=0$. The ODE implies $s$ is nondecreasing. Since the limits at both ends coincide, $s$ must be constant, hence $s\equiv 0$. Thus $g^{\prime}\equiv 0$ and $g(\pm\infty)=1$ gives $g\equiv 1$. ∎
Definition 7.8 (Density Bers potential).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, define
$\beta(\mu):=\frac{1}{2}S(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}.$
Lemma 7.9 (Half-density curvature identity).
With $\psi_{\mu}:=g^{-1/2}$ one has
$\beta(\mu)=-\frac{\psi_{\mu}^{\prime\prime}}{\psi_{\mu}}.$
Proof.
From $\psi=g^{-1/2}$ we have $\psi^{\prime}/\psi=-(1/2)g^{\prime}/g=-(1/2)s$, hence
$\frac{\psi^{\prime\prime}}{\psi}=\left(\frac{\psi^{\prime}}{\psi}\right)^{\prime}+\left(\frac{\psi^{\prime}}{\psi}\right)^{2}=-\frac{1}{2}s^{\prime}+\frac{1}{4}s^{2}.$
Rearranging gives $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}=-\psi^{\prime\prime}/\psi$. ∎
Definition 7.10 (Fisher information).
The Fisher information is
$I(\mu)=\int_{\mathbb{R}}s_{\mu}^{2}\,\mu=\int_{\mathbb{R}}\frac{(g^{\prime})^{2}}{g}\,dx.$
Proposition 7.11 (Integrated Schwarzian and mean negativity of $\beta$).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s$. Then
$\int_{\mathbb{R}}S(\mu)\,\mu=-\frac{3}{2}\,I(\mu),\qquad\int_{\mathbb{R}}\beta(\mu)\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx\leq 0,$
with equality in either identity if and only if $\mu=dx$.
Proof.
Write $\mu=g\,dx$ and $s=g^{\prime}/g$. Then
$\int_{\mathbb{R}}S(\mu)\,\mu=\int_{\mathbb{R}}\Bigl(s^{\prime}-\frac{1}{2}s^{2}\Bigr)g\,dx=\int_{\mathbb{R}}s^{\prime}g\,dx-\frac{1}{2}\int_{\mathbb{R}}s^{2}g\,dx.$
Since $sg=g^{\prime}$ and $g^{\prime},s\to 0$ at $\pm\infty$ by Lemma 7.2,
$\int_{\mathbb{R}}s^{\prime}g\,dx=\int_{\mathbb{R}}(sg)^{\prime}\,dx-\int_{\mathbb{R}}sg^{\prime}\,dx=0-\int_{\mathbb{R}}s^{2}g\,dx=-I(\mu),$
and therefore $\int S(\mu)\,\mu=-I(\mu)-\frac{1}{2}I(\mu)=-\frac{3}{2}I(\mu)$.
For the second identity, use $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ and note $\int_{\mathbb{R}}s^{\prime}\,dx=s(+\infty)-s(-\infty)=0$. Rigidity follows from $I(\mu)\geq 0$ and Proposition 7.7. ∎
Proposition 7.12 (Supersymmetric factorization of the projective Schrödinger operator).
Let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$ and Bers potential $\beta=\beta(\mu)$. Define the first-order operators on $C_{c}^{\infty}(\mathbb{R})$
$\mathcal{D}_{\mu}:=\partial_{x}+\frac{1}{2}s,\qquad\mathcal{D}_{\mu}^{\ast}:=-\partial_{x}+\frac{1}{2}s\quad\text{(adjoint in }L^{2}(dx)\text{)}.$
Then
$\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}=-\partial_{x}^{2}-\beta(\mu).$
Consequently $\partial_{x}^{2}-\beta(\mu)$ is nonnegative as a quadratic form on $L^{2}(dx)$:
$\langle f,(-\partial_{x}^{2}-\beta(\mu))f\rangle_{L^{2}(dx)}=\|\mathcal{D}_{\mu}f\|_{L^{2}(dx)}^{2}=\int_{\mathbb{R}}\Bigl|f^{\prime}(x)+\frac{1}{2}s(x)f(x)\Bigr|^{2}\,dx.$
Proof.
For $f\in C_{c}^{\infty}(\mathbb{R})$ one computes
$\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}f=\Bigl(-\partial_{x}+\frac{1}{2}s\Bigr)\Bigl(f^{\prime}+\frac{1}{2}sf\Bigr)=-f^{\prime\prime}-\frac{1}{2}s^{\prime}f+\frac{1}{4}s^{2}f.$
Using $\beta=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ gives $\mathcal{D}_{\mu}^{\ast}\mathcal{D}_{\mu}f=-f^{\prime\prime}-\beta f$, proving the identity. The quadratic form identity follows by integration by parts, since $\mathcal{D}_{\mu}^{\ast}$ is the $L^{2}(dx)$-adjoint of $\mathcal{D}_{\mu}$ on $C_{c}^{\infty}$. ∎
7.4. Straightening of geodesic dynamics
The only input from the diffeomorphism picture used below is the Riccati evolution of the pulled-back strain along $\dot{W}^{1,p}$-geodesics, which has already been established earlier. The present subsection records the density-side identity that identifies the strain with statistical velocity.
Proposition 7.13 (Statistical velocity as pulled-back strain).
Let $\varphi(t)$ be a smooth curve in $\operatorname{Diff}^{-\infty}(\mathbb{R})$ and set $\mu(t)=J(\varphi(t))=g(t)\,dx$ with $g(t,x)=\varphi^{\prime}(t,x)$. Let $u=\dot{\varphi}\circ\varphi^{-1}$ be the Eulerian velocity. Then the statistical velocity
$v(t,x):=\partial_{t}\log\frac{\mu(t)}{dx}=\partial_{t}\log g(t,x)$
satisfies
$v(t,x)=u_{x}\bigl(t,\varphi(t,x)\bigr).$
Proof.
Differentiate $g=\varphi^{\prime}$ in $t$:
$\dot{g}=\partial_{t}(\varphi^{\prime})=\partial_{x}\dot{\varphi}=\partial_{x}(u\circ\varphi)=(u_{x}\circ\varphi)\varphi^{\prime}.$
Dividing by $g=\varphi^{\prime}$ gives $v=\dot{g}/g=u_{x}\circ\varphi$. ∎
Corollary 7.14 (Riccati law in density variables).
Along any $F_{\dot{W}^{1,p}}$-geodesic $\varphi(t)$ (equivalently, along any $F_{p}^{\mathrm{FR}}$-geodesic $\mu(t)=J(\varphi(t))$), the statistical velocity $v=\partial_{t}\log(\mu/dx)$ satisfies
$\partial_{t}v+\frac{1}{p}v^{2}=0$
pointwise on $\mathbb{R}$.
Proof.
By Proposition 7.13, $v=u_{x}\circ\varphi$ is the pulled-back strain. The Riccati evolution for the pulled-back strain along $\dot{W}^{1,p}$-geodesics was established earlier, hence the conclusion. ∎
7.5. Log-concavity, half-density curvature, and the Schwarzian
Throughout let $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ and write
$f=\log g,\qquad s=s_{\mu}=f^{\prime}=\frac{g^{\prime}}{g},\qquad\kappa_{\mu}=f^{\prime\prime}=s^{\prime},\qquad\beta_{\mu}=\beta(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}.$ (36)
We also set the distinguished half-density ratio $\psi_{\mu}:=g^{-1/2}$.
7.5.1. Curvature decomposition and global impossibilities
Definition 7.15 (Curvature regions).
The log-curvature regions of $\mu$ are
$\Omega_{-}(\mu)=\{x:\kappa_{\mu}(x)<0\},\quad\Omega_{+}(\mu)=\{x:\kappa_{\mu}(x)>0\},\quad\Omega_{0}(\mu)=\{x:\kappa_{\mu}(x)=0\}.$
The projective curvature regions are
$\Pi_{-}(\mu)=\{x:\beta_{\mu}(x)<0\},\quad\Pi_{+}(\mu)=\{x:\beta_{\mu}(x)>0\},\quad\Pi_{0}(\mu)=\{x:\beta_{\mu}(x)=0\}.$
Lemma 7.16 (Decay and zero-mean identities).
For $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$, the functions g1g-1, $g^{\prime}$, $s$, $\psi_{\mu}-1$, and $\psi_{\mu}^{\prime}$ lie in $L^{1}(\mathbb{R})\cap C^{0}(\mathbb{R})$, and
$\lim_{x\to\pm\infty}g(x)=1,\qquad\lim_{x\to\pm\infty}s(x)=0,\qquad\lim_{x\to\pm\infty}\psi_{\mu}(x)=1,\qquad\lim_{x\to\pm\infty}\psi_{\mu}^{\prime}(x)=0.$
Moreover,
$\int_{\mathbb{R}}s(x)\,dx=f(+\infty)-f(-\infty)=0,\qquad\int_{\mathbb{R}}\psi_{\mu}^{\prime}(x)\,dx=\psi_{\mu}(+\infty)-\psi_{\mu}(-\infty)=0.$
Proof.
By Lemma 7.2, $g\in C^{\infty}(\mathbb{R})$, $g(\pm\infty)=1$, and $g$ is bounded above and away from 0. Thus $s=g^{\prime}/g\in L^{1}\cap C^{0}$ and $s(\pm\infty)=0$ since $g^{\prime}(\pm\infty)=0$. Also $\psi_{\mu}=g^{-1/2}$ satisfies $\psi_{\mu}-1\in L^{1}$ and $\psi_{\mu}^{\prime}=-(1/2)g^{-3/2}g^{\prime}\in L^{1}\cap C^{0}$ with $\psi_{\mu}^{\prime}(\pm\infty)=0$.
Finally $f=\log g$ satisfies $f(\pm\infty)=0$ and $f^{\prime}=s\in L^{1}$, hence $\int s=0$; similarly for $\psi_{\mu}^{\prime}$. ∎
Proposition 7.17 (No global log-concavity / log-convexity).
If $\kappa_{\mu}\leq 0$ on $\mathbb{R}$ (globally log-concave), or if $\kappa_{\mu}\geq 0$ on $\mathbb{R}$ (globally log-convex), then $\mu=dx$.
Proof.
Assume $\kappa_{\mu}=s^{\prime}\leq 0$. Then $s$ is non-increasing. By Lemma 7.16, $s(\pm\infty)=0$. A non-increasing function with equal limits at both ends is constant, hence $s\equiv 0$. Therefore $f^{\prime}=0$ and $f(\pm\infty)=0$ gives $f\equiv 0$, so $g\equiv 1$ and $\mu=dx$. The case $\kappa_{\mu}\geq 0$ is identical. ∎
Theorem 7.18 (No global sign for the Bers potential / Schwarzian).
If $\beta_{\mu}\geq 0$ on $\mathbb{R}$ or $\beta_{\mu}\leq 0$ on $\mathbb{R}$, then $\mu=dx$. Equivalently, for every nontrivial $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$, the projective curvature $\beta_{\mu}$ (and hence $S(\mu)=2\beta_{\mu}$) takes both positive and negative values.
Proof.
By Lemma 7.9, $\beta_{\mu}\geq 0$ implies $\psi_{\mu}^{\prime\prime}\leq 0$, so $\psi_{\mu}^{\prime}$ is non-increasing. Lemma 7.16 gives $\psi_{\mu}^{\prime}(\pm\infty)=0$, hence $\psi_{\mu}^{\prime}\equiv 0$ and $\psi_{\mu}\equiv 1$, so $g\equiv 1$ and $\mu=dx$. The case $\beta_{\mu}\leq 0$ gives $\psi_{\mu}^{\prime\prime}\geq 0$ and the same conclusion. ∎
Corollary 7.19 (Mixed curvature is forced).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ be nontrivial. Then:
  • (i)
    $\Omega_{-}(\mu)$ and $\Omega_{+}(\mu)$ are both nonempty open sets.
  • (ii)
    $\Pi_{-}(\mu)$ and $\Pi_{+}(\mu)$ are both nonempty open sets.
In particular, both $\kappa_{\mu}$ and $\beta_{\mu}$ have zeros on $\mathbb{R}$.
Proof.
If $\Omega_{-}(\mu)=\emptyset$, then $\kappa_{\mu}\geq 0$ on $\mathbb{R}$ and Proposition 7.17 forces $\mu=dx$, contradiction. Similarly $\Omega_{+}(\mu)\neq\emptyset$. For $\Pi_{\pm}(\mu)$ apply Theorem 7.18. Openness follows from continuity. ∎
7.5.2. Inflection multiplicity and critical-point sign rules
Proposition 7.20 (At least two log-inflection points and two projective inflection points).
Let $\mu\in|\Omega|_{\infty,1}(\mathbb{R})$ be nontrivial. Then:
  • (i)
    The score $s$ attains a positive maximum and a negative minimum; in particular $\kappa_{\mu}=s^{\prime}$ vanishes at at least two points.
  • (ii)
    The half-density slope $\psi_{\mu}^{\prime}$ attains a positive maximum and a negative minimum; in particular $\beta_{\mu}=-\psi_{\mu}^{\prime\prime}/\psi_{\mu}$ vanishes at at least two points.
Consequently, both $\Omega_{0}(\mu)$ and $\Pi_{0}(\mu)$ contain at least two points.
Proof.
By Lemma 7.16, $s\in C^{0}(\mathbb{R})$, $s(\pm\infty)=0$, and $\int_{\mathbb{R}}s\,dx=0$. If $s\geq 0$ on $\mathbb{R}$ then $\int s=0$ forces $s\equiv 0$, hence $\mu=dx$, contradiction; thus $s$ takes both signs. Since $s(\pm\infty)=0$, it attains a positive maximum and negative minimum; at each extremum, $s^{\prime}=0$.
The same argument applies to $\psi_{\mu}^{\prime}$ using Lemma 7.16. Nontriviality implies $\psi_{\mu}^{\prime}\not\equiv 0$, hence it takes both signs and has extrema where $\psi_{\mu}^{\prime\prime}=0$, i.e. $\beta_{\mu}=0$. ∎
Proposition 7.21 (Sign of $\beta$ at critical points).
Let $x_{0}$ be a critical point of $g$ (equivalently $s(x_{0})=0$). Then
$\beta_{\mu}(x_{0})=\frac{1}{2}\,s^{\prime}(x_{0})=\frac{g^{\prime\prime}(x_{0})}{2g(x_{0})}.$
In particular, $x_{0}$ is a strict local maximum of $g$ if and only if $\beta_{\mu}(x_{0})<0$, and a strict local minimum if and only if $\beta_{\mu}(x_{0})>0$.
Proof.
At $x_{0}$ we have $s=0$, hence $\beta_{\mu}(x_{0})=\frac{1}{2}s^{\prime}(x_{0})$. Also $s^{\prime}=(g^{\prime\prime}/g)-(g^{\prime}/g)^{2}=g^{\prime\prime}/g$ at $x_{0}$. ∎
Proposition 7.22 (Log-concavity forces negative projective curvature).
On $\Omega_{-}(\mu)$ one has $\beta_{\mu}<0$. Equivalently, $\Omega_{-}(\mu)\subset\Pi_{-}(\mu)$. Moreover, $\beta_{\mu}>0$ can occur only on the strict log-convexity region where
$s^{\prime}(x)>\frac{1}{2}s(x)^{2}.$
Proof.
If $s^{\prime}(x)<0$ then $\beta_{\mu}(x)=\frac{1}{2}s^{\prime}(x)-\frac{1}{4}s(x)^{2}<0$. The characterization of $\beta_{\mu}(x)>0$ is a rearrangement. ∎
7.5.3. Fisher information does not control projective curvature
Proposition 7.11 implies
$\int_{\mathbb{R}}\beta_{\mu}\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx,$
so $\beta_{\mu}$ has negative mean with respect to $dx$. Quantitative control of the positive region $\Pi_{+}(\mu)$, or of $\int[\beta_{\mu}]_{+}\,dx$, from Fisher information alone fails.
Lemma 7.23 (Oscillatory averaging).
Let $\phi\in C_{c}^{\infty}(\mathbb{R})$ and let $h$ be $2\pi$-periodic with $h\in L^{\infty}([0,2\pi])$ and mean $\overline{h}=(2\pi)^{-1}\int_{0}^{2\pi}h(t)\,dt$. Then
$\int_{\mathbb{R}}\phi(x)\,h(\lambda x)\,dx\longrightarrow\overline{h}\int_{\mathbb{R}}\phi(x)\,dx\qquad\text{as }\lambda\to\infty.$
Proof.
Expand $h$ in its Fourier series in $L^{2}([0,2\pi])$: $h=\overline{h}+\sum_{k\neq 0}c_{k}e^{ikx}$. Then
$\int_{\mathbb{R}}\phi(x)h(\lambda x)\,dx=\overline{h}\int_{\mathbb{R}}\phi(x)\,dx+\sum_{k\neq 0}c_{k}\int_{\mathbb{R}}\phi(x)e^{ik\lambda x}\,dx,$
and each oscillatory integral tends to 0 by the Riemann-Lebesgue lemma. ∎
Theorem 7.24 (Unbounded positive projective curvature at fixed Fisher information).
Fix $I_{0}>0$. There exists a sequence $\mu_{n}=g_{n}\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ such that
$I(\mu_{n})=I_{0}\,\,\text{for all }n,\quad\int_{\mathbb{R}}[\beta(\mu_{n})]_{+}\,dx\rightarrow\infty,\quad\#\{x:\beta(\mu_{n})(x)=0\}\rightarrow\infty.$
In particular, neither $\int[\beta]_{+}\,dx$ nor the number of sign changes of $\beta$ can be bounded in terms of Fisher information alone.
Proof.
Choose $\phi\in C_{c}^{\infty}(\mathbb{R})$ with $\phi\geq 0$, $\phi\not\equiv 0$, and such that $\phi\equiv 1$ on a nontrivial closed interval $J=[a,b]\subset\mathbb{R}$. For parameters $\lambda\geq 1$ and $\alpha>0$ define
$g_{\lambda,\alpha}(x):=1+\frac{\alpha}{\lambda}\,\phi(x)\sin(\lambda x),\qquad\mu_{\lambda,\alpha}:=g_{\lambda,\alpha}\,dx.$
For every fixed $\alpha$ and all $\lambda\geq 2\alpha\|\phi\|_{L^{\infty}}$ one has $g_{\lambda,\alpha}\geq\frac{1}{2}$ on $\mathbb{R}$. Moreover $g_{\lambda,\alpha}-1\in C_{c}^{\infty}(\mathbb{R})\subset W^{\infty,1}(\mathbb{R})$, hence $\mu_{\lambda,\alpha}\in|\Omega|_{\infty,1}(\mathbb{R})$.
We normalize the Fisher information. Write $I(\mu)=\int(g^{\prime})^{2}/g\,dx$. A direct differentiation gives
$g_{\lambda,\alpha}^{\prime}(x)=\alpha\,\phi(x)\cos(\lambda x)+\frac{\alpha}{\lambda}\,\phi^{\prime}(x)\sin(\lambda x).$
Since $g_{\lambda,\alpha}^{-1}$ is uniformly bounded by $2$ and $g_{\lambda,\alpha}^{-1}=1+O(\lambda^{-1})$ uniformly in $x$ for $\alpha$ restricted to any compact interval,
$I(\mu_{\lambda,\alpha})=\int_{\mathbb{R}}\frac{(g_{\lambda,\alpha}^{\prime})^{2}}{g_{\lambda,\alpha}}\,dx=\int_{\mathbb{R}}(g_{\lambda,\alpha}^{\prime})^{2}\,dx+O(\lambda^{-1}),\qquad\lambda\to\infty,$
where the error is uniform for $\alpha$ in compact sets. Expanding the square and using that $\phi,\phi^{\prime}$ are compactly supported,
$\int_{\mathbb{R}}(g_{\lambda,\alpha}^{\prime})^{2}\,dx$ $=\alpha^{2}\int_{\mathbb{R}}\phi(x)^{2}\cos^{2}(\lambda x)\,dx+\frac{2\alpha^{2}}{\lambda}\int_{\mathbb{R}}\phi(x)\phi^{\prime}(x)\cos(\lambda x)\sin(\lambda x)\,dx$
$\quad+\frac{\alpha^{2}}{\lambda^{2}}\int_{\mathbb{R}}(\phi^{\prime}(x))^{2}\sin^{2}(\lambda x)\,dx.$
The last two terms are $O(\lambda^{-1})$. For the leading term, Lemma 7.23 applied to $h(t)=\cos^{2}t$ yields
$\int_{\mathbb{R}}\phi(x)^{2}\cos^{2}(\lambda x)\,dx\longrightarrow\frac{1}{2}\int_{\mathbb{R}}\phi(x)^{2}\,dx,\qquad\lambda\to\infty.$
Consequently,
$I(\mu_{\lambda,\alpha})=\frac{\alpha^{2}}{2}\int_{\mathbb{R}}\phi^{2}\,dx+o(1),\qquad\lambda\to\infty,$
with $o(1)$ uniform for $\alpha$ in compact sets. Choose $\alpha_{*}>0$ so that
$\frac{\alpha_{*}^{2}}{2}\int_{\mathbb{R}}\phi^{2}\,dx=I_{0}.$
Fix a small $\varepsilon>0$ and consider $\alpha\in[\alpha_{*}-\varepsilon,\alpha_{*}+\varepsilon]$. By uniform convergence, there exists $\lambda_{0}$ such that for all $\lambda\geq\lambda_{0}$,
$I(\mu_{\lambda,\alpha_{*}-\varepsilon})<I_{0}<I(\mu_{\lambda,\alpha_{*}+\varepsilon}).$
For each such $\lambda$, continuity of $\alpha\mapsto I(\mu_{\lambda,\alpha})$ implies the existence of $\alpha(\lambda)$ in this interval with $I(\mu_{\lambda,\alpha(\lambda)})=I_{0}$.
Next, write $s_{\lambda,\alpha}=(\log g_{\lambda,\alpha})^{\prime}$ and use
$\beta(\mu_{\lambda,\alpha})=\frac{1}{2}\,\frac{g_{\lambda,\alpha}^{\prime\prime}}{g_{\lambda,\alpha}}-\frac{3}{4}\left(\frac{g_{\lambda,\alpha}^{\prime}}{g_{\lambda,\alpha}}\right)^{2}.$
A second differentiation gives
$g_{\lambda,\alpha}^{\prime\prime}(x)=-\alpha\lambda\,\phi(x)\sin(\lambda x)+2\alpha\,\phi^{\prime}(x)\cos(\lambda x)+\frac{\alpha}{\lambda}\,\phi^{\prime\prime}(x)\sin(\lambda x).$
Since $g_{\lambda,\alpha}^{-1}$ is uniformly bounded and equals $1+O(\lambda^{-1})$ uniformly, there exists $C_{1}<\infty$ such that
$\beta(\mu_{\lambda,\alpha})(x)=-\frac{\alpha\lambda}{2}\,\phi(x)\sin(\lambda x)+E_{\lambda,\alpha}(x),\qquad\|E_{\lambda,\alpha}\|_{L^{1}(\mathbb{R})}\leq C_{1}.$
Using $A+B]_{+}\geq[A]_{+}-|B|$ and integrating yields
$\int_{\mathbb{R}}[\beta(\mu_{\lambda,\alpha})]_{+}\,dx\geq\frac{\alpha\lambda}{2}\int_{\mathbb{R}}\phi(x)[-\sin(\lambda x)]_{+}\,dx-C_{1}.$
Lemma 7.23 with $h(t)=[-\sin t]_{+}$ gives $\overline{h}=1/\pi$, hence the integral is bounded below by a positive multiple of $\int\phi$ for all $\lambda$ large. Taking $\alpha=\alpha(\lambda)$ and letting $\lambda\to\infty$ implies $\int[\beta(\mu_{\lambda,\alpha(\lambda)})]_{+}\,dx\to\infty$.
Finally, on $J$ we have $\phi\equiv 1$, and the main term oscillates with amplitude $\sim\alpha\lambda$ while the error is uniformly bounded. Thus $\beta$ alternates sign on $J$ at scale $\lambda^{-1}$, producing $\gtrsim\lambda(b-a)$ distinct zeros. Choosing $\lambda_{n}\to\infty$ and setting $\mu_{n}:=\mu_{\lambda_{n},\alpha(\lambda_{n})}$ gives the claim.
Remark 7.25 (Koopman operators and branching diffusions).
The Koopman operator theory on density space, including the statistical Hilbert bundle with unitary half-density representation and the branching diffusion interpretation via Feynman-Kac, will appear in a companion paper. The forced sign-change theorem translates into the statement that every nontrivial density has both a genuine branching region and a genuine killing region.
8. Sharp Functional Inequalities
The Bers image identifies a diffeomorphism with a one-dimensional Schrödinger potential subject to a critical spectral constraint. Namely, for
$q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R}),\qquad H_{q}:=-\partial_{x}^{2}-q,$
the operator $H_{q}$ is nonnegative and has no bound states (Lemma 6.3). We develop the analytic consequences of this constraint.
8.1. Critical Hardy theory and resonance rigidity
For $q=\beta^{-\infty}(\varphi)$ write $u=\log\varphi^{\prime}$ and set
$\psi:=(\varphi^{\prime})^{-1/2}=e^{-u/2}.$ (37)
Since $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$ one has $\psi>0$ and $\psi(\pm\infty)=1$.
Theorem 8.1 (Ground state transform, uniqueness, and criticality).
Let $q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R})$ with $\varphi\in\mathrm{Diff}^{-\infty}(\mathbb{R})$, and let $\psi$ be as in (37). Then:
  • (i)
    $\psi$ is the unique positive solution of $H_{q}\psi=0$ normalized by $\psi(+\infty)=1$.
  • (ii)
    For every $f\in H^{1}(\mathbb{R})$,
    $\mathcal{E}_{q}(f):=\int_{\mathbb{R}}\bigl((f^{\prime})^{2}-qf^{2}\bigr)\,dx=\int_{\mathbb{R}}\psi^{2}\left|\left(\frac{f}{\psi}\right)^{\prime}\right|^{2}dx=\int_{\mathbb{R}}\left|f^{\prime}-\frac{\psi^{\prime}}{\psi}f\right|^{2}dx.$ (38)
    In particular,
    $\int_{\mathbb{R}}q(x)f(x)^{2}\,dx\leq\int_{\mathbb{R}}(f^{\prime}(x))^{2}\,dx,\qquad f\in H^{1}(\mathbb{R}),$ (39)
    and the constant $1$ is sharp.
  • (iii)
    (Criticality / no positive remainder) For every $W\in L^{1}(\mathbb{R})$ with $W\geq 0$ and $W\not\equiv 0$, the strengthened inequality
    $\mathcal{E}_{q}(f)\geq\int_{\mathbb{R}}W(x)f(x)^{2}\,dx\qquad\forall f\in C_{c}^{\infty}(\mathbb{R})$
    fails. Equivalently, $q$ is a critical Hardy weight for the Dirichlet form $\int(f^{\prime})^{2}$.
Proof.
(i) By Lemma 6.3,
$q=\beta^{-\infty}(\varphi)=\frac{1}{2}u^{\prime\prime}-\frac{1}{4}(u^{\prime})^{2},$
hence $\psi=e^{-u/2}$ satisfies
$\psi^{\prime\prime}=\left(\frac{1}{4}(u^{\prime})^{2}-\frac{1}{2}u^{\prime\prime}\right)\psi=-q\psi,$
so $H_{q}\psi=0$. Let $\tilde{\psi}>0$ be another solution of $H_{q}\tilde{\psi}=0$ with $\tilde{\psi}(+\infty)=1$. The Wronskian $W(\tilde{\psi},\psi)=\tilde{\psi}\,\psi^{\prime}-\tilde{\psi}^{\prime}\,\psi$ is constant. Since $u^{\prime}\in L^{1}(\mathbb{R})$, both $\psi^{\prime}$ and $\tilde{\psi}^{\prime}$ tend to 0 along $\infty$, while $\psi,\tilde{\psi}\to 1$, so $W(\tilde{\psi},\psi)=0$. Therefore $\tilde{\psi}/\psi$ is constant; normalization at $\infty$ forces $\tilde{\psi}=\psi$.
(ii) Define $A:=\partial_{x}-\psi^{\prime}/\psi=\partial_{x}+\frac{1}{2}u^{\prime}$. On $C_{c}^{\infty}(\mathbb{R})$ one has
$A^{*}=-\partial_{x}-\frac{\psi^{\prime}}{\psi},\qquad A^{*}A=-\partial_{x}^{2}-\left(\frac{\psi^{\prime}}{\psi}\right)^{\prime}+\left(\frac{\psi^{\prime}}{\psi}\right)^{2}=-\partial_{x}^{2}-q=H_{q}.$
Thus $\mathcal{E}_{q}(f)=\|Af\|_{L^{2}}^{2}$ for $f\in C_{c}^{\infty}(\mathbb{R})$. Since $Af=\psi(f/\psi)^{\prime}$, this is (38); density extends the identity to $H^{1}(\mathbb{R})$. Inequality (39) follows immediately. Sharpness follows by the cutoff construction already used in Theorem 8.1: with $\chi_{R}(x):=\chi(x/R)$ and $f_{R}:=\chi_{R}\psi$, one has $(f_{R}/\psi)^{\prime}=\chi_{R}^{\prime}$, so $\mathcal{E}_{q}(f_{R})=\int\psi^{2}|\chi_{R}^{\prime}|^{2}\lesssim R^{-1}$ while $\|f_{R}\|_{L^{2}}^{2}\sim R$.
(iii) Fix $W\geq 0$, $W\not\equiv 0$ in $L^{1}(\mathbb{R})$. Choose $\chi\in C_{c}^{\infty}(\mathbb{R})$ with $\chi\equiv 1$ on a neighborhood of $\mathrm{supp}(W)$ and set $f_{R}:=\chi\cdot(\chi_{R}\psi)$ with $\chi_{R}$ as above. Then $f_{R}/\psi=\chi\chi_{R}$, so (38) gives
$\mathcal{E}_{q}(f_{R})=\int_{\mathbb{R}}\psi^{2}|(\chi\chi_{R})^{\prime}|^{2}\,dx\leq 2\int\psi^{2}|\chi^{\prime}|^{2}\,dx+2\int\psi^{2}|\chi_{R}^{\prime}|^{2}\,dx=O(1)+O(R^{-1}).$
Since $\chi\equiv 1$ on $\mathrm{supp}(W)$ and $\psi$ is continuous and positive, $\int Wf_{R}^{2}\,dx\to\int W\psi^{2}\,dx>0$ as $R\to\infty$. Hence for $R$ large, $\mathcal{E}_{q}(f_{R})<\int Wf_{R}^{2}$, so no positive remainder term can be added to (39). ∎
Remark 8.2 (Threshold resonance and spectral gap).
The nullfunction $\psi$ is not in $L^{2}(\mathbb{R})$ and encodes a zero-energy resonance: $H_{q}\geq 0$ but 0 is neither an $L^{2}$ eigenvalue nor an isolated point of the spectrum. Consequently, one should not expect a global Poincaré inequality on $\mathbb{R}$ beyond (39). The substitute is a windowed Hardy-Poincaré inequality after modding out the resonant direction; this is provided by Theorem 8.3.
8.2. Hardy-Poincaré coercivity on finite windows
For $R>0$ denote $I_{R}:=(-R,R)$ and define the localized quadratic form
$\mathcal{E}_{q,R}(f):=\int_{I_{R}}\bigl((f^{\prime})^{2}-qf^{2}\bigr)\,dx,\qquad f\in H^{1}(I_{R}).$
The next statement is the precise coercivity principle one can have at a threshold resonance.
Theorem 8.3 (Hardy-Poincaré inequality on a window).
Let $q=\beta^{-\infty}(\varphi)$ and $\psi=(\varphi^{\prime})^{-1/2}$ as above. For every $R>0$ and every $f\in H^{1}(I_{R})$ let
$c_{R}(f):=\frac{\int_{I_{R}}f(x)\psi(x)\,dx}{\int_{I_{R}}\psi(x)^{2}\,dx},\qquad f_{R}^{\perp}:=f-c_{R}(f)\psi.$ (40)
Then
$\mathcal{E}_{q,R}(f)\geq\lambda_{R}(q)\int_{I_{R}}|f_{R}^{\perp}(x)|^{2}\,dx,$ (41)
with an explicit lower bound
$\lambda_{R}(q)\geq\frac{\pi^{2}}{4R^{2}}\cdot\frac{\displaystyle\inf_{I_{R}}\psi^{2}}{\displaystyle\sup_{I_{R}}\psi^{2}}\,.$ (42)
In particular, $\lambda_{R}(q)>0$ for each fixed $R$.
Proof.
By Theorem 8.1(ii), for $f\in H^{1}(I_{R})$ one has
$\mathcal{E}_{q,R}(f)=\int_{I_{R}}\psi^{2}\left|\left(\frac{f}{\psi}\right)^{\prime}\right|^{2}dx.$
Write $g:=f/\psi$. Then $f_{R}^{\perp}=\psi(g-\bar{g}_{R})$ where
$\bar{g}_{R}:=\frac{\int_{I_{R}}g\,\psi^{2}\,dx}{\int_{I_{R}}\psi^{2}\,dx}=c_{R}(f),$
so $\int_{I_{R}}(g-\bar{g}_{R})\psi^{2}\,dx=0$ and
$\int_{I_{R}}|f_{R}^{\perp}|^{2}\,dx=\int_{I_{R}}\psi^{2}|g-\bar{g}_{R}|^{2}\,dx.$
Let $m_{R}:=\inf_{I_{R}}\psi^{2}$ and $M_{R}:=\sup_{I_{R}}\psi^{2}$. Then
$\int_{I_{R}}\psi^{2}|g-\bar{g}_{R}|^{2}\,dx\leq M_{R}\int_{I_{R}}|g-\bar{g}_{R}|^{2}\,dx,\qquad\int_{I_{R}}\psi^{2}|g^{\prime}|^{2}\,dx\geq m_{R}\int_{I_{R}}|g^{\prime}|^{2}\,dx.$
The (sharp) Neumann Poincaré inequality on (R,R)(-R,R) gives
$\int_{I_{R}}|g-\bar{g}_{R}|^{2}\,dx\leq\frac{4R^{2}}{\pi^{2}}\int_{I_{R}}|g^{\prime}|^{2}\,dx,$
hence
$\int_{I_{R}}|f_{R}^{\perp}|^{2}\,dx\leq M_{R}\frac{4R^{2}}{\pi^{2}}\int_{I_{R}}|g^{\prime}|^{2}\,dx\leq\frac{4R^{2}}{\pi^{2}}\frac{M_{R}}{m_{R}}\int_{I_{R}}\psi^{2}|g^{\prime}|^{2}\,dx=\frac{4R^{2}}{\pi^{2}}\frac{M_{R}}{m_{R}}\,\mathcal{E}_{q,R}(f).$
Rearranging yields (41) with (42). ∎
Remark 8.4 (Why no global gap is possible).
The cutoff sequence $f_{R}=\chi_{R}\psi$ from Theorem 8.1 satisfies $\mathcal{E}_{q}(f_{R})\to 0$ while $\|f_{R}\|_{L^{2}}\to\infty$, so no inequality of the form $\mathcal{E}_{q}(f)\geq\lambda\|f\|_{L^{2}}^{2}$ can hold globally on $\mathbb{R}$. Theorem 8.3 isolates the only obstruction, that is, the resonant direction $\psi$, and only on finite windows.
8.3. Trace hierarchy and Fisher energies as spectral entropies
We express the geometric integrals of the Bers potential in spectral variables. The common mechanism is to view $\log a(z)$ as
  1. (i)
    a Cauchy transform of $\log(1-|R_{q}|^{2})$ using outerness (no bound states), and
  2. (ii)
    a large-$|z|$ expansion obtained from the Volterra/Jost representation. Matching coefficients yields trace identities.
Lemma 8.5 (Outer representation and moment expansion).
Let $q\in W^{\infty,1}(\mathbb{R})$ lie in the Bers image and let $R_{q}$ be its reflection coefficient. Let $a(z)$ be the transmission coefficient from (30). Then $a$ has no zeros in $\Im z>0$ and for $\Im z>0$,
$\log a(z)=-\frac{1}{2\pi i}\int_{\mathbb{R}}\frac{\log(1-|R_{q}(\xi)|^{2})}{\xi-z}\,d\xi,$ (43)
where the branch is fixed by $\log a(z)\to 0$ as $|z|\to\infty$ in $\Im z>0$. Moreover, as $|z|\to\infty$ with $\Im z\geq 0$,
$\log a(z)=\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi+\frac{1}{2\pi i}\frac{1}{z^{3}}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi+O(|z|^{-4}).$ (44)
Proof.
By Theorem 6.6 the absence of bound states implies that $a$ has no zeros in $\Im z>0$, hence $a$ is outer and (43) holds with the stated normalization.
By Proposition 6.2, $R_{q}\in\mathcal{S}(\mathbb{R})$, hence $\log(1-|R_{q}|^{2})\in\mathcal{S}(\mathbb{R})$. Expanding the Cauchy kernel gives
$\frac{1}{\xi-z}=-\frac{1}{z}-\frac{\xi}{z^{2}}-\frac{\xi^{2}}{z^{3}}+O(|z|^{-4}),\qquad|z|\to\infty,\ \Im z\geq 0,$
with an error bounded by $C(1+|\xi|^{3})|z|^{-4}$. Since $\log(1-|R_{q}|^{2})$ is Schwartz, multiplying by the kernel expansion is integrable and dominated convergence justifies termwise integration. Finally, for real $q$ one has $R_{q}(-\xi)=\overline{R_{q}(\xi)}$, so $|R_{q}(\xi)|^{2}$ and hence $\log(1-|R_{q}(\xi)|^{2})$ are even. Therefore the $z^{-2}$ coefficient vanishes, that is
$\int_{\mathbb{R}}\xi\,\log(1-|R_{q}(\xi)|^{2})\,d\xi=0,$
which yields (44). ∎
Lemma 8.6 (Volterra/Jost expansion of $\log a(z)$ to order $|z|^{-3}$).
Let $q\in W^{\infty,1}(\mathbb{R})$. Then as $|z|\to\infty$ with $\Im z\geq 0$,
$\log a(z)=\frac{1}{2iz}\int_{\mathbb{R}}q(x)\,dx+\frac{1}{(2iz)^{3}}\int_{\mathbb{R}}q(x)^{2}\,dx+O(|z|^{-4}).$ (45)
Proof.
Let $f_{+}(x,z)$ be the Jost solution and set $m(x,z):=e^{-izx}f_{+}(x,z)$ so that $m(\cdot,z)\to 1$ and $m^{\prime}(\cdot,z)\to 0$ as $x\to+\infty$. Then $m$ satisfies
$m^{\prime\prime}+2iz\,m^{\prime}+q\,m=0.$ (46)
Write $\varepsilon:=(2iz)^{-1}$ and define
$m^{(2)}(x):=1+\varepsilon m_{1}(x)+\varepsilon^{2}m_{2}(x),\qquad m_{1}(x):=\int_{x}^{\infty}q(t)\,dt,\qquad m_{2}(x):=q(x)+\int_{x}^{\infty}q(t)m_{1}(t)\,dt.$ (47)
Using $m_{1}^{\prime}=-q$ and $m_{2}^{\prime}=q^{\prime}-qm_{1}$, a direct substitution into (46) gives
$m^{(2)}{}^{\prime\prime}+2iz\,m^{(2)}{}^{\prime}+q\,m^{(2)}=\varepsilon^{2}\bigl(m_{2}^{\prime\prime}+qm_{2}\bigr).$
Since $q\in W^{\infty,1}$ one has $m_{1}\in W^{\infty,\infty}$, $m_{2}\in W^{\infty,1}\cap W^{\infty,\infty}$ and $m_{2}^{\prime\prime}+qm_{2}\in L^{1}\cap L^{\infty}$. Writing $m=m^{(2)}+r$ and solving (46) for $r$ by the Volterra method (equivalently, variation of constants for $(e^{2izx}r^{\prime})^{\prime}=-e^{2izx}(qr+\varepsilon^{2}(m_{2}^{\prime\prime}+qm_{2}))$ with $r,r^{\prime}\to 0$ at $\infty$) yields the uniform bound
$\sup_{x\in\mathbb{R}}\bigl(|r(x,z)|+|z|^{-1}|r^{\prime}(x,z)|\bigr)\leq C|z|^{-3},\qquad|z|\gg 1,\ \Im z\geq 0,$ (48)
for a constant $C$ depending on finitely many $W^{k,1}$ norms of $q$.
The Wronskian identity defining $a$ gives
$a(z)=1+\varepsilon\int_{\mathbb{R}}q(x)m(x,z)\,dx.$ (49)
Insert $m=m^{(2)}+r$ and use (48) to obtain
$\int_{\mathbb{R}}q\,m\,dx=\int_{\mathbb{R}}q\,dx+\varepsilon\int_{\mathbb{R}}q\,m_{1}\,dx+\varepsilon^{2}\int_{\mathbb{R}}q\,m_{2}\,dx+O(|z|^{-3}).$ (50)
Set $A:=\int_{\mathbb{R}}q$ and $Q(x):=\int_{x}^{\infty}q(t)\,dt=m_{1}(x)$. Then $Q^{\prime}=-q$, $Q(\infty)=0$, $Q(-\infty)=A$, so
$\int_{\mathbb{R}}q\,m_{1}\,dx=\int_{\mathbb{R}}qQ\,dx=\int_{\mathbb{R}}(-Q^{\prime})Q\,dx=\frac{1}{2}A^{2}.$
For the next coefficient,
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx.$
By Fubini and $\int_{-\infty}^{t}q(x)\,dx=A-Q(t)$,
$\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx=\int_{\mathbb{R}}q(t)Q(t)(A-Q(t))\,dt=\frac{1}{2}A^{3}-\int_{\mathbb{R}}qQ^{2}\,dt.$
Finally $\int_{\mathbb{R}}qQ^{2}=\int_{\mathbb{R}}(-Q^{\prime})Q^{2}=-\frac{1}{3}[Q^{3}]_{-\infty}^{\infty}=\frac{1}{3}A^{3}$, hence
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\frac{1}{6}A^{3}.$
Substituting into (50) and then (49) yields
$a(z)=1+\varepsilon A+\varepsilon^{2}\frac{A^{2}}{2}+\varepsilon^{3}\Bigl(\int_{\mathbb{R}}q^{2}\,dx+\frac{A^{3}}{6}\Bigr)+O(|z|^{-4}).$
Expanding $\log(1+x)=x-\frac{1}{2}x^{2}+\frac{1}{3}x^{3}+O(x^{4})$ and using the displayed expansion shows the even powers cancel, and one obtains
$\log a(z)=\varepsilon A+\varepsilon^{3}\int_{\mathbb{R}}q^{2}\,dx+O(|z|^{-4}),$
which is (45). ∎
Theorem 8.7 (First trace identity).
Let $q=\beta^{-\infty}(\varphi)\in W^{\infty,1}(\mathbb{R})$ and let $R_{q}$ be its reflection coefficient. Then
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk,$ (51)
equivalently
$\int_{\mathbb{R}}q(x)\,dx=\frac{1}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}\,dk.$ (52)
If $q=\beta(\mu)$ for $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, then
$\int_{\mathbb{R}}s(x)^{2}\,dx=\frac{4}{\pi}\int_{\mathbb{R}}\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}\,dk.$ (53)
Proof.
By Lemma 8.5 and Lemma 8.6, the coefficient of $z^{-1}$ in the large-$|z|$ expansion of $\log a(z)$ is simultaneously
$\frac{1}{2iz}\int_{\mathbb{R}}q(x)\,dx\qquad\text{and}\quad\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi$
(we provide more explicit analysis for the coefficient of $z^{-3}$ in the proof of Theorem 8.8). Equating these coefficients yields (51), and (52) is its rearrangement.
If $q=\beta(\mu)$, Proposition 7.11 gives $\int_{\mathbb{R}}q\,dx=-\frac{1}{4}\int_{\mathbb{R}}s^{2}\,dx$, and inserting this into (52) gives (53). ∎
Theorem 8.7 converts the linear functional $\int q$ into a spectral entropy integral. The next identity upgrades this to the quadratic energy $\int q^{2}$ and yields, on the density side, a mixed Sobolev/Orlicz functional of the score.
Theorem 8.8 (Second trace identity).
Under the hypotheses of Theorem 8.7,
$\int_{\mathbb{R}}q(x)^{2}\,dx=-\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk=\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}dk.$ (54)
If $q=\beta(\mu)$ for $\mu=g\,dx\in|\Omega|_{\infty,1}(\mathbb{R})$ with score $s=(\log g)^{\prime}$, then
$\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}(x)|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s(x)|^{4}\,dx=\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\,\log\bigl(1-|R_{q}(k)|^{2}\bigr)^{-1}dk.$ (55)
Proof.
By Proposition 6.2 one has $R_{q}\in\mathcal{S}(\mathbb{R})$, hence $\log(1-|R_{q}|^{2})\in\mathcal{S}(\mathbb{R})$ and all moment integrals below are absolutely convergent. Let $a(z)$ be the transmission coefficient from (30). Since $q$ lies in the Bers image, $H_{q}=-\partial_{x}^{2}-q$ has no bound states, so $a$ has no zeros in $\Im z>0$ and is an outer function (Theorem 6.6). Consequently, for $\Im z>0$,
$\log a(z)=-\frac{1}{2\pi i}\int_{\mathbb{R}}\frac{\log(1-|R_{q}(\xi)|^{2})}{\xi-z}\,d\xi,$ (56)
where the branch is fixed by $\log a(z)\to 0$ as $|z|\to\infty$ in $\Im z>0$. Expanding the Cauchy kernel gives, as $|z|\to\infty$ with $\Im z\geq 0$,
$\frac{1}{\xi-z}=-\frac{1}{z}-\frac{\xi}{z^{2}}-\frac{\xi^{2}}{z^{3}}+O(|z|^{-4}),$
with an error dominated by $C(1+|\xi|^{3})|z|^{-4}$. Since $\log(1-|R_{q}|^{2})$ is Schwartz, termwise integration is justified, and because $q$ is real one has $R_{q}(-\xi)=\overline{R_{q}(\xi)}$, hence $\log(1-|R_{q}(\xi)|^{2})$ is even and the $z^{-2}$ moment vanishes. Therefore
$\log a(z)=\frac{1}{2\pi i}\frac{1}{z}\int_{\mathbb{R}}\log(1-|R_{q}(\xi)|^{2})\,d\xi+\frac{1}{2\pi i}\frac{1}{z^{3}}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi+O(|z|^{-4}).$ (57)
We now compute the same large-$|z|$ expansion from the Jost/Volterra side, keeping the first two nontrivial odd orders. Let $f_{+}(x,z)$ be the Jost solution and set $m(x,z):=e^{-izx}f_{+}(x,z)$, so that $m(\cdot,z)\to 1$ and $m^{\prime}(\cdot,z)\to 0$ as $x\to+\infty$. Then $m$ satisfies
$m^{\prime\prime}+2iz\,m^{\prime}+q\,m=0.$ (58)
Write $\varepsilon:=(2iz)^{-1}$. Define
$m^{(2)}(x):=1+\varepsilon m_{1}(x)+\varepsilon^{2}m_{2}(x),\qquad m_{1}(x):=\int_{x}^{\infty}q(t)\,dt,\qquad m_{2}(x):=q(x)+\int_{x}^{\infty}q(t)m_{1}(t)\,dt.$ (59)
Using $m_{1}^{\prime}=-q$ and $m_{2}^{\prime}=q^{\prime}-qm_{1}$, a direct substitution of (59) into (58) yields
$m^{(2)}{}^{\prime\prime}+2iz\,m^{(2)}{}^{\prime}+q\,m^{(2)}=\varepsilon^{2}\bigl(m_{2}^{\prime\prime}+qm_{2}\bigr).$
Because $q\in W^{\infty,1}$, one has $m_{1}\in W^{\infty,\infty}$ and $m_{2}\in W^{\infty,1}\cap W^{\infty,\infty}$, hence $m_{2}^{\prime\prime}+qm_{2}\in L^{1}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$. Writing $m=m^{(2)}+r$, the remainder $r$ solves
$r^{\prime\prime}+2iz\,r^{\prime}+q\,r=-\varepsilon^{2}(m_{2}^{\prime\prime}+qm_{2}),\qquad r,r^{\prime}\to 0\ \text{as }x\to+\infty.$
By the Volterra estimates from Proposition 6.2 (applied to the inhomogeneous equation),
$\sup_{x\in\mathbb{R}}\bigl(|r(x,z)|+|z|^{-1}|r^{\prime}(x,z)|\bigr)\leq C|z|^{-3},\qquad|z|\gg 1,\ \Im z\geq 0.$ (60)
with $C$ depending only on finitely many $W^{k,1}$ norms of $q$.
The Wronskian identity defining $a$ (the same identity used to obtain the $z^{-1}$ coefficient in the first trace formula) gives
$a(z)=1+\varepsilon\int_{\mathbb{R}}q(x)\,m(x,z)\,dx.$ (61)
Insert $m=m^{(2)}+r$ and use (60) to obtain
$\int_{\mathbb{R}}q\,m\,dx=\int_{\mathbb{R}}q\,dx+\varepsilon\int_{\mathbb{R}}q\,m_{1}\,dx+\varepsilon^{2}\int_{\mathbb{R}}q\,m_{2}\,dx+O(|z|^{-3}).$ (62)
Set $A:=\int_{\mathbb{R}}q$ and $Q(x):=\int_{x}^{\infty}q(t)\,dt=m_{1}(x)$. Then $Q^{\prime}=-q$, $Q(\infty)=0$, and $Q(-\infty)=A$. Consequently
$\int_{\mathbb{R}}q\,m_{1}\,dx=\int_{\mathbb{R}}qQ\,dx=\int_{\mathbb{R}}(-Q^{\prime})Q\,dx=\frac{1}{2}A^{2}.$
Moreover,
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx.$
By Fubini, using $\int_{-\infty}^{t}q(x)\,dx=A-Q(t)$,
$\int_{\mathbb{R}}q(x)\int_{x}^{\infty}q(t)Q(t)\,dt\,dx=\int_{\mathbb{R}}q(t)Q(t)\bigl(A-Q(t)\bigr)\,dt=\frac{1}{2}A^{3}-\int_{\mathbb{R}}qQ^{2}\,dt.$
Finally $\int_{\mathbb{R}}qQ^{2}=\int_{\mathbb{R}}(-Q^{\prime})Q^{2}=-\frac{1}{3}[Q^{3}]_{-\infty}^{\infty}=\frac{1}{3}A^{3}$, hence
$\int_{\mathbb{R}}q\,m_{2}\,dx=\int_{\mathbb{R}}q^{2}\,dx+\frac{1}{6}A^{3}.$
Substituting into (62) and then (61) yields
$a(z)=1+\varepsilon A+\varepsilon^{2}\frac{A^{2}}{2}+\varepsilon^{3}\Bigl(\int_{\mathbb{R}}q^{2}\,dx+\frac{A^{3}}{6}\Bigr)+O(|z|^{-4}).$
Expanding $\log(1+x)=x-\frac{1}{2}x^{2}+\frac{1}{3}x^{3}+O(x^{4})$ and inserting the displayed asymptotic shows that the even powers cancel and the cubic $A^{3}$ contributions cancel as well, leaving
$\log a(z)=\varepsilon A+\varepsilon^{3}\int_{\mathbb{R}}q(x)^{2}\,dx+O(|z|^{-4})=\frac{1}{2iz}\int_{\mathbb{R}}q\,dx+\frac{1}{(2iz)^{3}}\int_{\mathbb{R}}q^{2}\,dx+O(|z|^{-4}).$ (63)
Comparing the $z^{-3}$ coefficients in (57) and (63) gives
$\frac{1}{(2i)^{3}}\int_{\mathbb{R}}q^{2}\,dx=\frac{1}{2\pi i}\int_{\mathbb{R}}\xi^{2}\log(1-|R_{q}(\xi)|^{2})\,d\xi,$
hence, since $(2i)^{3}=-8i$,
$\int_{\mathbb{R}}q(x)^{2}\,dx=-\frac{4}{\pi}\int_{\mathbb{R}}k^{2}\log\bigl(1-|R_{q}(k)|^{2}\bigr)\,dk,$
which is the first equality in (54); the second is its reformulation.
If $q=\beta(\mu)$ with $\mu=g\,dx$ and score $s=(\log g)^{\prime}$, then $q=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}$ and $s(\pm\infty)=0$. Expanding $q^{2}$ and integrating by parts eliminates the mixed term:
$\int_{\mathbb{R}}q^{2}\,dx=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}|^{2}\,dx-\frac{1}{4}\int_{\mathbb{R}}s^{\prime}s^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s|^{4}\,dx=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s|^{4}\,dx,$
since $\int s^{\prime}s^{2}=\frac{1}{3}\int(s^{3})^{\prime}=0$. Substituting this into (54) gives (55). ∎
Remark 8.9 (Hierarchy).
The two identities above are the first two nontrivial members of the odd-order large-$|z|$ expansion of $\log a(z)$. Namely, the $z^{-2}$ term is absent by parity, and each subsequent odd coefficient yields a conserved differential polynomial in $q$ equated to a spectral moment of $\log(1-|R_{q}|^{2})$. In this paper, we stop at $n=0,1$ ($n$ indexes the coefficient of $z^{-(2n+1)}$), which already produces the quadratic/superquadratic score functional (55) needed as an analytic input for Part II.
Remark 8.10 (Relation to mKdV and the Miura transform).
The mixed score functional in (55),
$\mathcal{H}(s):=\frac{1}{4}\int_{\mathbb{R}}|s^{\prime}(x)|^{2}\,dx+\frac{1}{16}\int_{\mathbb{R}}|s(x)|^{4}\,dx,$
is not the usual information-geometric energy (compare the Fisher information $\int s^{2}g\,dx$): it is unweighted and mixes an $H^{1}$ term with an $L^{4}$ term. It is nevertheless familiar from one-dimensional integrable PDE. Indeed, setting $v:=\frac{s}{2}$ one has
$\mathcal{H}(s)=\int_{\mathbb{R}}\bigl(|v_{x}|^{2}+|v|^{4}\bigr)\,dx,$
which, up to conventional normalizations of the flow, is the standard Hamiltonian energy for the (defocusing) modified KdV equation [8]. Moreover, the identity
$q=\beta(\mu)=\frac{1}{2}s^{\prime}-\frac{1}{4}s^{2}=v_{x}-v^{2}$
is exactly the Miura transform (up to sign conventions) linking mKdV and KdV variables [15]. In particular, $\int_{\mathbb{R}}q^{2}\,dx=\mathcal{H}(s)$, so (55) can be viewed as a scattering-theoretic representation of a classical integrable Hamiltonian in terms of the reflection coefficient. From the inverse-scattering viewpoint, this is a low-order instance of the trace/sum-rule hierarchy expressing conserved differential polynomials in $q$ as spectral moments of $\log(1-|R_{q}|^{2})$ [7, 5].
9. Background on $L^{p}$ Fisher-Rao geometry
We recall the $L^{p}$ Fisher-Rao framework on compact manifolds, following Bauer-Bruveris-Harms-Michor [4].
Let $M$ be a closed oriented $n$-manifold with density bundle $|\Omega|=|\bigwedge^{n}T^{*}M|$, positive sections $|\Omega|_{+}(M)=\{\mu\in\Gamma(|\Omega|)\mid\mu>0\}$, and probability densities $\operatorname{Prob}(M)=\{\mu\in|\Omega|_{+}(M)\mid\int_{M}\mu=1\}$. Fix a reference density $\mathrm{d}x\in|\Omega|_{+}(M)$.
Definition 9.1 ([4]).
For $p\in[1,\infty)$, the $L^{p}$ Fisher-Rao Finsler structure on $|\Omega|_{+}(M)$ is
$F_{p}(\mu,a)=\left(\int_{M}\left|\frac{a}{\mu}\right|^{p}\mu\right)^{1/p},\qquad a\in T_{\mu}|\Omega|_{+}(M)=\Gamma(|\Omega|).$
For $p=2$ this squares to the classical Fisher-Rao Riemannian metric. The Finsler structure is $\operatorname{Diff}(M)$-invariant: $F_{p}(\varphi^{*}\mu,\varphi^{*}a)=F_{p}(\mu,a)$ for every $\varphi\in\operatorname{Diff}(M)$.
Theorem 9.2 ([4]).
The map $\Phi_{p}(\mu)=p(\mu/\mathrm{d}x)^{1/p}$ is an isometric embedding $(|\Omega|_{+}(M),F_{p})\hookrightarrow(L^{p}(M,\mathrm{d}x),\|\cdot\|_{L^{p}})$ onto the positive cone; on $\operatorname{Prob}(M)$ the image lies in the $L^{p}$ sphere of radius $p$. The unique $F_{p}$-geodesic from $\mu_{0}=\rho_{0}\,\mathrm{d}x$ to $\mu_{1}=\rho_{1}\,\mathrm{d}x$ in $|\Omega|_{+}(M)$ is
$\mu(t)=\bigl((1-t)\rho_{0}^{1/p}+t\rho_{1}^{1/p}\bigr)^{p}\mathrm{d}x.$
The classification theorem of Bauer-Bruveris-Modin [3] shows that in dimension $\dim M\geq 2$, every smooth $\operatorname{Diff}(M)$-invariant weak Riemannian metric on $|\Omega|_{+}(M)$ is a linear combination of the Fisher-Rao metric and the total-mass inner product. On $\operatorname{Prob}(M)$, only the Fisher-Rao metric survives. This classification is intrinsically Hilbertian in the sense that the $L^{p}$ Fisher-Rao structures for $p\neq 2$ are Finsler and fall outside its scope, which is the starting point of the present paper.